A Bibliography of Publications in Byte Magazine: 1995–1998

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

12 January 2024 Version 2.31

Title word cross-reference

000-rpm [Mia97c]. 02 [Row97].

1 [Ano95-55, Ano95-198, Dej95e, Gas95, Lo97, Mul96a, Way97f]. 1-2-3 [Cra96, Ano95-55, Dej95e, Gas95]. 1-800-Internet [Mul96a]. 1-Gb [Lo97]. 1.2 [Ano98-59]. 10-PPM [Ano95-101]. 10.4-inch [Joc95e]. 100 [Ano98a, Bal95g, Cun95]. 100-Mbps [Cun95]. 100-MHz [Ano95-208, Ano98a, Bal95g]. 10ppm [Ano95-101]. 11- [Ano95-101]. 11-to [Ano95-101]. 120 [Ano95a, Ano95-158, Ano96-174, LM95b]. 120- [Ano95c]. 120-and [LM95b]. 120-MB [Ano95a]. 120-MHz [Ano95-158, LM95b]. 1200-dpi [Ano95-118]. 12000e [Law95]. 12X [Emi97]. 133 [Ano95c, Ano95-205, LM95b]. 133-MHz [Ano95c, Ano95-205, LM95b]. 14 [Ano97-41]. 14.4 [FH97, Smi95a]. 14.4-Kbps [FH97]. 14.4-Kpbs [Smi95a]. 15GA [Ano95-204, Ano95-209]. 16-bit [Ano95-93]. 16-to [Ano95-188]. 17- [Ano95-106, HKM96]. 17-inch [Ano95-106]. 17GA [Ano95-204, Ano95-209].

2 [Ano95-111, Ano95-112, Ano96-74, Ano98-55, Gas95, Gio95, KE97, Nan95a, Nan96g, Tam95a, Wes96b, Wes96a]. 2.0 [MC98]. 20- [Ano95-106]. 20-and [Ano95-106]. 200 [Ano96-178, RV96]. 200- [Van97]. 200-MHz [Ano96-178, RV96].

2000 [Ano98s, Ano98-95, Em98]. 202 [Hum97a]. 209 [Ano97z]. 20ppm [Ano95-188]. 21-inch [Ano95-106]. 225 [Van97]. 225-MHz [Van97]. 233 [Yag98a]. 233-MHz [Yag98a]. 24-bit [Van95a]. 24-ppm [HKM96]. 250P [And95a, And95b]. 28.8-Kbps [KH95]. 28A [Ano96-132]. 28A-B [Ano96-132]. 2A [Ano96-132]. 2A-B [Ano96-132]. 2GB [Hum96e].

3 [Ano95e, Ano95f, Ano95-60, Ano95-55, Ano95g, Ano96-48, Ano96-126, Ano96-173, Ano97-110, Ano97-119, Bal95b, Bou95, Dej95e, Dej95c, Dej95b, Die95a, EP97, Flo96a, Fou95a, Fou95b, Gas95, Gre97l, Hal96f, Hal97c, HKM97a, Hum96b, Hum97a, Joh96, Lov95, MD95, Row96a, Row97, Sup97, Tho96a, Tho96r, Yar95]. 3-D [Ano95e, Ano95f, Ano95-60, Ano95g, Ano96-48, Ano96-126, Ano96-173, Ano97-110, Ano97-119, Bal95b, Bou95, Dej95c, Dej95b, Die95a, EP97, Flo96a, Fou95a, Fou95b, Gre97l, Hal96f, Hal97c, HKM97a, Hum96b, Hum97a, Joh96, Lov95, MD95, Row96a, Row97, Tho96a, Tho96r, Yar95]. 3-pound [Ano98-96]. 3.0 [Ano97-114]. 3.5-inch [Ano95h]. 3.x [Ano95-63, Ano95-125, Goo95a]. 30-ppm [Ano95-101, Ano95-101]. 300 [Ano95-198, Ano97b, Ano97-106, Cam98a, Hal96g]. 300-MHz [Ano97b, Ano97-106, Cam98a, Hal96g]. 300-to-1 [Ano95-198]. 32 [Ano95b, LM95b]. 32-bit [Ano95-159, Ano95-157, Api95c, CCS97, Chi96a, Cra96, Moh97b, Tam95a, LM95b]. 32-bitness [Mia95d]. 32-Mb [Ano95b]. 32-ppm [Ano95-189]. 33 [Tho95g]. 33-MHz [Tho95g]. 333 [Tan98]. 333-MHz [Tan98]. 350 [Ano97-118]. 36A [Ano96-133]. 36A-D [Ano96-133]. 386DX4 [Ano95-104]. 386DX4-based [Ano95-104]. 3A [Ano96-178]. 3D [Ano96-173, Ano97-77, Em98, Flo96a].

3Dlabs [MD95]. 3XP [Tam95d].

4 [Ano97-97]. 4-D [Ano97-97]. 4-mm [Ano95-100]. 4-pound [Way97g]. 4.0 [Bla96, Gil96b]. 4.O [Nan95d]. 4.x [Ano95-63]. 400 [Ano97c, Ano98-83]. 400-MHz [Ano97c]. 4000 [Ano98-126, Hum96i]. 4100Z [Van96c]. 440LX [Cam98a]. 486-style [Ano95-159]. 486DX4 [Ano95-104]. 4s [Ano95t].

5 [Ano97-113, Ano98-68, Ano98-115, Gil97a]. 5-pound [Ano98-96, Row96c]. 5.0 [MC98]. 500 [Ano96-161]. 500-MHz [Ano96-161]. 5100 [Ano98-85]. 5300c [Ano95-203]. 533 [Hal96h]. 533-MHz [Hal96h]. 540c [Ano95-104]. 56 [Ano96-45, Ude95d]. 56-Kbps [Ano96-45, Ude95d]. 5K86 [Ano96-170]. 5K86-P75 [Ano96-170]. 5p [Ano97-114].

6- [Ano95-101]. 6-based [Ano97-82]. 6-to [Ano95-101]. 6.0 [Gil98a]. 600 [Ano97e, Ano98-51]. 600-dpi [Hum96i]. 600-MHz [Ano97e]. 6000 [Ano95-186]. 603e [Tho95k]. 6200 [Chi96b]. 6200MX [Hum97c]. 64 [Ano95-151, Ano95-152, Ano95-192, Hal98d]. 64-bit [Ano95-151, Ano95-152, Ano95-159, Ano95-192, Ano97-54, Lac96, Sim96, Ano95-151, Ano95-152, Ano95-192]. 64-hit [Ano97t]. 66 [Ano95-104]. 66-MHz [Ano95-104]. 660Plus [Ano97-120]. 680x0 [Tho95h]. 6bone [Fin98]. 6X [Ano95i, Ano95j]. 6x86 [Ano96-37, Ano96-38].

7 [Ano96-172]. 7000S [Ano98h]. 701C [Joc95e]. 755CV [DeJ95a]. 765 [Ano98-76].

8 [Ano97-27]. 8-mm [Ano95-100]. 8000 [Pou95d]. 810N [Ano95-104].

95 [Ano97-125, Gre96j, Pou97j, Pou98f, VN97a, Won98b]. 95-based [LM95b]. 95/Memphis [Ano97-125]. 95/Office [SK96a]. 950N [Ano95-204, Ano95-209]. 9580A [Ano95-211]. ’97 [Gil96g]. ’98 [Ano98-75, Ano98-83, Gar98].

= [Ano95-159, Pri95, Way96m].

A-B [Ano96-133]. AAAS [Pou95q]. abilities [Way95a]. ability [Ano95-59, Cro95, Goo95b]. able [Gar97, Wat97b]. Abstract [Tho95i, Tho95j]. Abstraction [Tho95j, Nan95g]. accelerate [Ano95-48]. Accelerated [Ano96-48]. Accelerator [Ano96-173]. Accelerators [Ano95-107, Ano98-115, Hum96b, Hum97a]. Accent [Ano98-113]. Acces [Ano98-63]. Access [Ano95x, Ano98-37, Bri97a, Cam98b, Dob97c, Gre97h, KHM^{+}96, Lev98, Lew96, Sal95e, Sal95f, Smi95d, Smi95c, Ano95-27, Ano95-72, Ano95-136, Ano98m, Bal95e, Dob98a, Eng97, Cot95b, Het96d, Hum98a, MC97, Nan95c, Sal95k, Vog95, Way96m, Yag97b, Ano95k, Ano95-64, Ano95-133, Dob95, Dob96]. access-routed [Sal95k]. accessibility [Ano95-136]. accomplish [Fri95a, Joh96, Laz96]. According [ATA96, Pou97g]. accounting [Ano95-77]. accuracy [Ano95-204, Ano95-209, Hil95]. accurate [Cua95b, Dej95g, Sug95]. Achievement [Ano95-79, Ano95-100]. Achievements [Ano95-106, Ano95-104, Ano95-103]. Achilles [Ano95-194]. Acorn [Pal96]. Acquire [Ano95-101]. acquires [Ano95-133]. acquisition [Dej95h]. acres [Pal96]. Acrobat [Ano97-114]. across [Cot95c, Fri97a, Ren98, Sin95a, Won98c, Yag97a]. Act [Ano96-178]. Action [Ano95e, Ano95f, Bal95b]. Actions [Ano98b]. Activator [Gre97q]. active [Ano95-104, Van95a, Gre97q]. active-matrix [Ano95-104, Van95a]. ActiveX [CL97]. Actual [Ano98c]. actually [Ano95-154, Pra95b]. Ada [MH97b]. adapter [Smi95b]. Adapters [Ano95-107, Ano96k]. adaptive [Laz97a]. Add [Ano96z, Ano98-28, Ano95-95, Gre96g]. add-on [Ano95-95]. Adding [Fen95, Tho95c, Tho95d, Ano95-134, Ano95-135, Pou98c]. additional [Nan95c]. Additions [Sta96d, Fri98a]. Address [Sch96h, Way96f, Bri97a, Die95c, Win95, Sch96g]. Addresses [Win95, Sup97]. Addressing [Fri97e, Ano95-116, Ude95d]. Adds [Ano96-95, Ano95-184, Ano97-114, Ano98-93, Bal96d, Dej95e, Mia95c]. Adieu [Pou96h]. Adjustable [Sta97a]. Administration [Hur97a]. admits [Ano95w]. Adobe [Ano97-114, Ano97-116, Bla96, Hum96g, Kay96a, Sha96]. Adoption [Hal98f, Way95e]. ADSL [Cam98b]. advanced [Cha97, Gas95]. Advances [Abr97b, Abr97a, Abr97c, Ano95k, Ano97f, Dob95, Ano95-48]. Advantages [LM95b]. Adventure [Ano95-163, Mon95a, Pou95t]. Adventures [Mon96a, Pou95t, Ude96c, Bal96c, Hal95b]. Advertisers [Ano95-86, Ano95-87, Ano96d, Ano96-79, Ano96-80, Ano96-81, Ano96-82, Ano98-53, Ano98-54]. Advice [Knu96c, Sal95m, Kay97a, Key95, Pou95k, Pou96i]. Aerial [Ano96c]. afar [Won98b]. Affair [Hur97e]. affect [Key95]. Affordable [Hal96f, Hum96b, Tho98a, Hal96h, Nad95b, Wei95]. After [BEM95, Pou97f, Pri95]. AG [Lin95d]. Again [Ano96-86, Ano98-59, MB97, Ude97h, Lin95a]. Against [Ano95-100, GM95a, Hal98a, Lin95d]. Age [Gre97k, Hal96k, Pou97h, Ano97-124, Gre96h]. agency [Cla95a]. agent [Ano95-116, Way95b, Way95a]. Agent-Enhanced [Way95a]. Agents [Ano95l, Ano95-92, Hal96a, Way95b, Ind95, Wat95, Way95f]. agnostic [Ano98-84]. ago [MB97]. AGP [Ano97-77]. agree [Hum96c]. agreements [Ano95-222]. Ahead [Ano95m, Ano96-183, Ano97-79, Nan95d, Het96b]. ahout [Ano96-134]. AI [UT95]. Aiding [Laz96]. Ailments [Nan96c, Nan96a]. aim [Sin95a, VN97b]. aimed [Pou95d]. aims [Los97c]. ain’t [HM95]. Air [Ano95g, Ano97g, Bal95e, Rey96]. Air-tight [Rey96]. Al [Gou95]. alarms [Ano98-81]. algorithm [Ste95, Way95c]. algorithms [Cla95c, Dej95g, Gou95]. alike [Gil95]. Alive [Smi96c, Ano97a]. all-in-one [BHKM96a, Hig95c]. alleged [Ano95w]. Alliances [Ano95-183]. allowed [Ano95-61]. allows [Dob97d, Joh96, MF97c]. Almost [WS96, Bau95]. alone [Smi95b]. along [Ano95c, Hig95c]. Alpha [Ano95m, Ano95-64, Ano97e, Ano97-115, Ano98-48, AG95, Api95c, Ban97, Chi96a, Fau95, Hal96h, Tho96d]. Alpha-based [Fau95]. Alpha-powered [Ano98-48]. alphabet [Ano97g]. Alphabetical [Ano96d, Ano96-79, Ano98-53, Ano98-54]. Alps [Hum96i]. ALR [RV96]. already [Hal95a]. also [Ano95-106, Ano95-104, Bal95h, BD95b, DM98, Gre95e, Hig95d, Hum96i, Krt98, Ric97, Tho96n, Vac95a]. AltaVista [VN97b]. altered [Way96g]. altering [Ano95-197]. Alternate [Ano95n]. Alternative [Ano96e, Ano95-151, Ano95-152, Ano95-195, Ano98n]. Alternatives [Ano98-65]. Am [Ano98-50, Kin96]. Amazing [Kay97b, Abr97b]. Ambiguity [Les95]. AMD [Ano96-170, Goo95b, Hal96b, Hal97c, PH95, Pou98d]. Amending [Pou98a]. America [Ano97-95, Bal95e]. Ami [She95c, She95c]. Amiga [Ano95-112]. among [Ano97-105, Ber95, Cli98c, Smi95b, Tho96e, Yag97a]. amount [Hal96c]. amounts [Fri96a]. Ana [Pou97f]. Analog [Ano98z, Lev98, Sal95e, Sal95f, Die95c, Sal95f]. analysis [Ano95-56, Ano95-103, Bau95, Cua95b, Dej95f, Dej95e, Dej95b, Het97, Wat95, Pot95]. analytical [Het97, Pen98]. analyzer [Ano95-55, Abr97b]. Analyzers [Ano98-78]. analyzes [Sto98]. anarchy [Flo97]. Anatomy [Ano95o]. Animate [Ano96-93]. Animation [EP97, Nad95b]. anniversary [Ano95-111]. announced [Ano95-141]. announcements [Ano95-164]. annoying [Ano97v, Ano98-81, Gre97p]. annual [Pou95q]. anoint [Ano95d]. Answer [Ano95-105, Ano95-221, Nan97c, Gre95b, Gre95c, KH95, Pou96k, Ude98b]. Anti [Ano95-208]. Anti-Virus [Ano95-208]. any [Ano95-61, Api95b, Gre96d]. Anyone [Ano96f, Ess95, Ude98b]. anywhere [Ano95-212, Ano95-213, Bau95, Nee95m]. Apache [Ano98-79, Ude97j]. apart [MFDG98]. APC [Ano97-71]. API [Joh96, Tho97d]. APIs [Ano97-31, Die95a, Ano96-67]. App [Ano95-82, Ano95-83, Gla95, Mel95, Sha98]. appear [Ano95g]. appearing [HM95]. Apple [Ano95q, Ano95r, Ano95s, Ano95-114, Ano95-104, Ano95-120, Ano95-164, Ano95-171, Ano95-203, Ano96g, Ano96-173, Ano97i, Ano97-116, Ano98d, Cot95a, Hal95d, MW95, Rei95e, Sch97a, Tho95c, Tho95d, Tho95e, Tho95f, Tho95h, Tho95j, Tho95l, Tho95m, Tho96g, Tho98a, Tho98c, Tho98e, Van96a]. Apple-sanctioned [Ano95-120]. Apple’s [Way97d]. Applet [Ano96-90, MF97c]. Applets [MF97a, Los97c, Tur97, Way96m]. Appliances [Whi98]. Application [Ano95-166, Ano95-167, Ano97-46, Tho95c, Tra96a, Ano95-74, Ano95-75, Ano95-82, Ber95, Fri97a, Cot95a, Gat98, Gre95e, Kay97c, Mor97, Phi97, Pra95b, Sal96f, SP97, Tho95d, Ude95a, Whi98]. application-development [Gre95e]. Applications [Bur95b, CCS97, Die95b, Fri98b, Gil97c, Sal96d, SP97, Tho95d, Ude97e, Ano95y, Ano95z, Ano95-72, Ano95-128, Ano95-160, Ano95-157, Ano95-162, Ano95-190, Ano95-220, Ano95-201, Ano96-89, Ano97-31, Ano97-81, Ano97-82, Ano97-105, Ano97-121, Ano97-122, Ano97-123, Ano98-92, Arc98, Cir95, DeJ96b, Dej95g, Dej95b, Fou95a, Fou95b, Gil96g, GE95, Gre96b, Gre97h, Gre97n, Hub95, Joh96, Knu96c, Krt98, MG97, Mia95d, Mia95f, Nan95a, Nan95g, QS96, RFD95, Sal95m, Sch96e, Sea96b, She95a, SB96, Sta97b, Sto98, Tho95l, Tho98e, Ude96g, Ude97k, Way98b, Yag97a]. applications-server [Ano97-123]. approach [Cla95c, Pou95g, Ano95-64, Ano96-175]. approaches [VN96b]. appropriateness [Sch97b]. Apps [And97, Ano97-39, Ano97-105, Fom98, Gla96, Gre97n, Nad97, Ren98, Nan96e, Pou97b]. Arbor [Ano95-56, Dej95e, Het97]. Archistrat [Ano95t]. Architecture [Cir95, Mac98b, Pou95i, Ano97-53, Arc98, Cli98a, Die95d, Flo95, Hal97d, Pou95g, Pou97b, Sha96, Tho96n, Tho97c, Die95a]. Architectures [Pou95i, Ano96u]. archive [Ude95b]. Ardent [Gre98b]. areas [Sha95c]. aren’t [Sal95i]. Armada [Ano98-96, Row96c]. Armed [Hur97a]. Armor [Sta96c]. Arrays [Gre95b, Gre95c]. Arrive [Ano95-94, Ano95-158, Ano96-52, Ano96-136, Ano96-141, Die95b, Ano96-89, Ano97-97]. arrived [Ano95-120, Ano95i, Ano95j]. Arrives [Ban97, Pou95u, Pou95v]. Art [Ano96p, Phi97, Ano95-37]. artificial [SK95]. Artisoft [Mia95c]. Ascentia [Ano95-104, Ano95-204, Ano95-209]. Asian [LaP96a]. Aside [Ano97-36, Tho96j]. askSam [Ano95-206]. Aspen [Ano98-48]. asset [Sal95b]. Assets [Ano95u, Bau95, Sal95b]. assignments [Pou96i]. Assistance [Gil97b]. Assistant [Ano95-200]. assisted [Gre97d]. Associate [Gre95e]. associated [Ano95-119]. Association [Ano95-89]. AST [Ano95-104, Ano95-204, Ano95-209, Ano97-117]. AT&T [And95a]. Atari [Bal95b]. Athletes [Sal95i]. Atlanta [Sal95i]. ATM [Bal95c, Coo96c, Coo96b, Fri95c, Fri97d, Fri97e, Joc96c, Kay96a, Loh97, Nan96e, Pri95, Sal95g, Sal96f, Sta97a]. Atomic [Pou97a]. atoms [DeJ98]. Attached [Joc98b, Die95d]. Attachmate [BD95b]. Attack [Ano96-142, Sal96a]. attacking [Sha95c]. Attacks [Kay97a]. attempting [Cob95]. Attention [Cam98b, Wat97a]. Attitude [Ano98q]. attitudes [Ano95-117]. ATX [Ano95v]. Audience [Van95a, van95b]. Audio [Gla96]. Audition [Ano98-67]. augmenting [WS96]. August [Ano95q]. Austin [Ano95-145]. Authentication [Ude98b, Ano97h]. author [Gla95]. authoring [BVN95, Del96a, Die96a, Gio97, Sea96b]. AutoCAD [Ano97-41, Ano98n, Yar95]. Autoconf [Fri97a]. Autodesk [Yar95]. Autoloader [Law95]. automate [Ano96-173]. automated [Law95, Rei95a, Rei95e]. automatically [Ano95-88, KP98]. Automation [BD95b, Gil95]. available [Ano95-50, Ano95-64, Ano95-105, Ano95-194, Nan95f]. Avatar [Hum97d]. Avatars [Hal96a]. avoid [Flo97, Meh95b, Pra95b, TM95]. Avoiding [MF97a]. awaited [Laz97a]. Awards [Ano95-141, Ano95d, Ano96a, Ano97a, Bar95c, Tho96c]. Aware [Sal96d]. Away [Ude95a, Pou95r].

B [Ano96-133, Ano96-132]. Baan [Ano98e]. BabeWatch [Abr97c]. Baby [BDF96, Ind95, Way96f]. Babylon [Har96]. Back [Cam97a, Hal98e, Hur97a, Nad95a, Ano97-127, Ano98-89, MDM95, Pou97i]. Backbone [Sta96b, Ano95-51, Sal95k]. Backbones [Ano97l]. backers [Ano95-192]. backing [Ano95-100, GM95a]. BackOffice [Ano97-102, Joc97a]. Backup [Ano98f, Law95, Mia97a, Mys96a, Ano97-123, Hum96e, Hum98a, KHB96b, KHB96a, Pou96l, VN96b]. Bad [Ano96-88, Ano97j, Joc95d, Ano95-37, Cro95]. baked [Pou97f]. baker [FH97]. balance [MW95]. Ball [Nee96c]. Bandai [Way97d]. bandwagon [Ano95-89, Pri95]. Bandwidth [Ano95-103, Ano96h, Ano96-137, Ano98-66, Hal96c, Mac98a, Way96i, Abr97a, Ano95-102, Ano96-78, Fri96b, Hal98e, Los97c, Sal95g, Ude96d]. Bank [Ano95w]. Banking [Ano96j]. Bar [Esc95b]. Barbie [Ano97f]. Bargain [Hum98b, Ano95-206]. barking [Ano98-81]. Barricading [Kay95a]. Barrier [Hal96c, Ano96-78, Hal96g]. bars [Sal95c]. Base [Ano97-64, Ano95-72, Ano97-121, Tho95e, Bal95e]. baseball [Kor95]. Based [Ano98-89, Ano95-100, Ano97-82, Ano98-47, BEM95, Cro95, Gio95, Gre97q, Gre98c, Krt98, LM95b, Rei95e, Ude97b, Whi95b, Ano95r, Ano95s, Ano95-97, Ano95-104, Fau95, Cot95a, Gre95e, LM95b, MG97, Row97, Ude95a, Way97d]. Bashed [Ano95-42]. Basic [Gil97a, WS96, Ano95-61, Sil97, Ude97h, Gre96h, Ano95-82, Ano97-111, Dob97d, Gil95, Gla95, Gre97c, Gre97r, WS96]. basically [Hal95c]. Basics [Gil97a]. Batlle [MDM95]. Batteries [Ano96-98, Bas95]. Battery [Bas95, And95c, Ano95-74, Ano95-106, Ano95-203, MA98, Tho95k]. battery-juiced [MA98]. Battle [Bal96a, Nan97a, Wat98a]. Battlezone [Ano95f, Bal95b]. Battling [Ric97]. BBS [Ano95x, Esp95]. BBSes [Ano95x]. Be [Bal96b, Way97b, And96b, Ano95-50, Ano95-61, Ano95-63, Ano95-64, Ano95g, Ano95-93, Ano95-137, Ano95-128, Ano95-140, Ano95-154, Ano95-174, Ano95-194, Ano98-92, Api95c, Cox98b, DeJ98, Dej95j, Flo95, Fou95a, Fou95b, Fow97, Gar97, Gre97l, Gre97p, Krt98, Mar95e, MC98, Nad95c, Pou95t, Rei95d, Rei95f, Sal95f, Sch96a, Sch97d, Son95, Tho96f, Wat97b, Way95d]. Be-All [Bal96b]. Be-All-That-You-Can-Be [Way97b].

Beach [Pou96q]. beam [Mon97]. beans [Gre97k]. beastie [Ano95-122]. beat [Bal95e]. Beats [Ano96-86, Tho97c]. beaucoup [Ano95-45]. Beautiful [Ano95-106, Cot96c, Kay96b, SC96, Cot96a, Len95]. because [Bar95b]. become [Ano95-146, Ano95-193, Ano95h, Fin98, Hal97b, Sha95a]. becomes [Ano95-194, Ude96l, Ude97b, Yav95]. becoming [Bon95a, Dej95b, Lin98, Rei95f, Tho95a]. bed [Fin98]. beefs [Ano95-171]. been [Ano95w, Ano95-122, Ano95-141, Sal95i, Tho95j]. Before [Fri97e, Ano95-122, Hal96g, MF97c, Mul96b]. beginning [Ind95, Sha95c]. beginnings [Gre97g]. begins [Ano95-168]. Behind [Hal95e, Hur98c, Nan95a, Ano95-109, Hal95b]. being [Ano95-222, Joc95c, Rei95f]. believe [DFT95b, Hur98c, Joc95e]. belongs [Ano98-100]. benchmark [Gre95a]. benchmarketing [Gra96]. Benchmarks [Gre95a, Gre96c, Hur98c, Yag96c, Ano95-104, LM95b]. benefit [DM98, Ude97g]. Benefits [Wil95, Ano95-182, Fri95a, Ude97d]. Bener [Ano96-99]. Besides [Ano98-60]. Best [Ano95-41, Ano95-107, Ano96w, Ano96s, Ano96l, Ano96k, Ano96-122, Ano96-144, Ano97k, Ano98k, Lin95b, Lin95c, Pou96a, Row96b, You96, Abr97c, Ano95-64, Ano95-95, Ano95-106, Ano95-104, Ano95-100, Ano95-105, Ano95d, Ano95-182, Ano97-32, Ano98-58, BGR95b, BHKM96b, Ber95, Hil95, HKM97b, Pou97h, Rud98, Sch96e, Sea95d, Tho95l, Tho97e, Ude98a, Van97, Way97f, Ano96t]. Bet [Ano98-48]. beta [Pou96l, Way96j]. bets [Gil97b]. Better [And96e, Ano95y, Ano95z, Ano95-27, Ano95-50, Ano95-66, Ano95-193, Ano95-219, Ano95i, Ano95j, Ano96m, Ano96n, Ano96o, Ano96-115, Ano96-137, Ano97r, Ano97l, Ano97m, Ano97n, Ano97-42, Ano97-47, Ano98g, Ano98-126, Ess95, Het96g, MFDG98, Orc98, Pou97k, Sim97, Tho95h, Tho95g, Tur97, VN97a, Way96b, Ano95-93, Ano95-157, Ano96-178, Ano97-116, Ano97-119, Cli98d, Dob97a, Dob97c, Gre96d, Hal98b, Hal98e, Het96f, Hil95, Les95, MC98, Rei95d, Way96l, Wei95]. Betting [Sto98]. between [Ano95-75, Ano95-118, Hal95a, MW95, Pou97a, Way95c]. beware [Hum96h]. Beyond [Ano96-128, Gra96, Gre96b, Hal97c, Hal97d, Joc95a, Vac95a, Ano96-174, Nan95c, Pou96l, Sal95h, Ude97h]. Bible [Die97a]. Big [Aer95, Ano95-70, Ano95-106, Ano95-129, Ano95-190, Col96, DeJ96c, Dej95h, Dej95j, Gio96c, HKM97b, Hum98a, Kay96b, Kha96, Mys96b, Pou98b, Wil96, Ano95-154, Ano98-58, Ano98-60, Ano98-73, Flo95, Gil97b, Hur97b, Len95, Sal95k, Tho97c, Tho97e, VN97b, Yag95c, Nan95a]. Big-Picture [Mys96b]. big-screen [HKM97b]. Bigger [Ano97l, Pep98, Krt98]. biggest [Hal97a, Sha95c]. Bill [ATA96]. billion [Ude95e]. Bills [Ano95-51]. bind [Ude97b]. Bindery [Mia95d]. binding [Dob98a, Gre98b]. Bipolar [Hal96h]. Birth [Gwe96]. bit [Ano95-82, Ano95-93, Ano95-151, Ano95-152, Ano95-159, Ano95-157, Ano95-192, Ano97-54, Api95c, CCS97, Chi96a, Cra96, Gla95, Lac96, Moh97b, Sim96, Tam95a, Van95a, Ano95-151, Ano95-152, Ano95-192, LM95b]. bitness [Mia95d]. Bits [Ano97o, Ano97p, Ano97q]. Blanks [Ano95-68]. Blasts [Ano95-28, Ano95-34, Ano95-29, Ano95-30, Ano95-31, Ano95-32, Ano95-33, Ano95-35, Bar95d, Bri97b]. Blazing [Tho98b]. blazingly [Tho97b]. blew [Ano96-134]. Blind [Ano96q, Ano95-219]. blindly [Pri95]. Blink [Nee96e]. block [Eng97]. Blocks [Ano98-94, Sil97]. Blooming [Pou97i]. Blossoms [EP97]. Blue [Aer95, Ano95-36, Cot96b, Row97, Nan95a]. Blueprint [Boy97]. Blues [Mia95a]. Blunder [Hum98b]. blurs [Hal95a]. board [Ano95-89, Ano96-171, Ano98-120]. Boards [BHKM96a]. Bob [Ano95-115]. bolsters [Gil95]. Bomb [Sch98, Ano95o, Ude95a]. Boning [Fin98]. Bonus [Nan95a]. Book [Die97a, BDF96, Var96, DFT95b, KFT95]. Bookkeeping [She95b, She95a]. Books [Aer95, ATA96, Ano95-37, Ano96p, BDF96, BM95, CFDT95, DFT95a, DFT95b, HCM96, KFT95, MDM95, MG96, RFD95, SK95, TJF95, UT95, UGB95]. Boost [Ano96-90, Hus95, Tho97c]. boosts [Shi98]. Borland [Ano98l, Ano95-72, Ano95-173, Ano95-196, Gre97g]. born [Tho97b]. Bosnia [Bon95a]. Boston [Ano95q]. Both [BD95b, Api95b, Api95c, Fen95, Los97a, Sal95f, Smi95d, Ste95, Tho95k, Tho95l, Ude96b]. Bottlenecks [Mac98a, Tho97c, Ano98-78, Bri97b]. bottom [Key95, Ste98, Wat97a]. bottom-up [Wat97a]. Bounce [Nad95a]. Bouncing [Gre96b]. boundaries [Hal95a, Won98c]. Box [Ano98h, Cot96b, Tho95l, Ano95-158, Col96, Laz97b, Way96h, Way96m, Way97d]. boxes [Ano95-154, Hum97b, RV96]. Brain [Ano98-128, Ano98-128]. Brains [Cla95b, Ano98-92, Tho96f]. Brainy [Bas95]. Branch [Sal95k, Sal95l]. Branch-Office [Sal95k, Sal95l]. brand [Van97]. Brave [Joc96b]. Bravo [Chi96b]. Brawny [Bas95]. breaches [Daw95]. Break [Hal96c, Hus95, Ano95-82, Hal96g]. breakfast [Hur97c]. Breaking [Mac98a]. Breaks [Ano97q]. breakthrough [Hal97d]. breakthroughs [Ano95-36]. breed [Ano95-88, Pou95k]. breeze [Ano96-175]. Brett [Ano95-82]. Bricklin [Ano97-112]. Bridges [Fri97c, Gre98f]. brief [Gwe96]. Briefcase [Ano95k]. Bright [Ano95-36, Ano95-69, Kay96b]. brighter [Ano95-36, Ano97-116, Gre95a]. Bring [Mia95b, Ano95-45, Ano95-60, Cli98d, Cob95, Dej95f, Fau95, Fri95c, Hum96b, Hur97d, Joc95a, Pra95a, Ude96b, Way97a].

Bringing [Dut98, Yag96c, Yar95, Ano95x]. brings [AK98, Ano95-182, Ano98-113, Bal95h, Ban97, Sal95h, Way98a]. British [Pou95a, Pou95h]. Broadband [Ano98i, Mon97]. Broadcasting [Meh95b, Ano95-168, Ano98-66, Los97c]. broader [Sal95h]. Broken [Ano97j]. Brooktrout [Ano95-38]. Brown [Ano96-49]. Browse [Hum96d, BHKM96a, FH97]. Browser [Way96m, Bal96d, Het96d, Smi95a, Way97d]. Browsers [Ano96q, Ano95-61]. Browsing [Ano98-45, Dob97a]. Bruce [Ano96-49]. budget [Hum98a]. Bug [Ano95-42, Ano96-37, Gre96c, Hal95e, Pou97h, Tho95f]. Bug-Free [Gre96c]. Bug-Hunting [Pou97h]. bugs [Ano98-121]. Build [Ano95-82, Ano95-83, Ano98-52, Bry95a, Bry95c, Fri95a, Gio97, Gla95, Joc95d, Mon98a, Ano95-186, Ano97-32, Fen95, Gwe95, Kay97c, Mac95b, Pou97e, Ste95, Tur97, Ude96g]. Builder [Ano98l, Gre97g, Het96g]. Builders [Bal96a]. Building [Ano97r, Ano97-105, Ano98-94, Api95a, Bur95b, Fri96b, Fri97c, Gil97c, Knu96b, Pou97d, Ren98, Tho95h, Tra97, Tur97, Ude98c, Won98c, Ano97s, Cob95, Fri98b, Gre96b, Gre97q, MG97, Sea96b, Sil97, Tho98e, Wat97a, Ano98-87]. Builds [Ano97-47, Sal95h, Gla95, Pou97l, Pou98d]. built [Ano95-40, Ano95-98, Ano95-204, Ano95-209, Cli98a, Hal98g, Ude95a, Way97a, Yag96e]. built-in [Ano95-98, Ano95-204, Ano95-209]. Bulk [Pou96b]. Bulky [Ano95-70]. Bullet [Ano95-151, Ano95-152]. Bulletproof [Lin95d]. Bulletproofing [Fri97d]. Bum [Ano98-106]. bumps [Hig95c]. bunch [Gou95]. Bundle [Pep97]. Bundled [Law95]. Burden [Ano97l]. burglars [Way95g]. Buried [Hal95b]. Burn [Ano95-146]. Burning [Ano97s, Gwe96]. Bus [Ano95-193, Ano97-40, Ano98a, Tan98, Ano95-159, Goo95b, Hal98e]. Business [Ano95-101, Ano95-107, Ano95-148, Ano96-113, Ber95, Ess95, HKM97a, KFT95, Ste98, Ude96m, Ano95-141, Ano95-194, Ano95-203, Ano96-89, Bar95a, GM95b, HR97, Nad95a, Nan95b, Sto98, Ude95a, Ude95c, Vac95a, WS96, Whi97, Yag95c]. business-quality [HR97]. business-to-business [Ude95c]. Businesses [Ano98-110, Ano95-77]. Businesslike [Sea97]. Busting [Sho96]. Butterfly [Nee95l]. Button [Ano96-51]. Buy [Ano96-148, Ano95-125, Mul96b, Van97]. Buyer [Ano95-39, Ano96r, Ano98j, Ano95-39]. Buying [Tri95, Ano95-97, Ano95-122]. buys [BHKM96b]. Bye [Hal97a]. Byte [Ano95-41, Ano95r, Ano95s, Ano95-40, Ano95-41, Ano95-64, Ano95-106, Ano95-100, Ano95-103, Ano95-111, Ano95-141, Ano95d, Ano96w, Ano96s, Ano96t, Ano96u, Ano96v, Ano96-76, Ano96-77, Ano96-78, Ano97t, Ano97u, Ano97-53, Ano97-55, AG95, Bar95c, Die95c, Die95d, Gre95a, Nee96b, Sea95d, Ude95b, Ude95c, Ude95d, Ude96b, Ano98k]. BYTE/NSTL [Ano95-106, Ano95-100, Ano95-103]. BYTEmark [Ano95-42].

C [Ano95-184, Ano95-185, Api95c, Api97, Cot95c, Ano95-47, Ano95-184, Ano95-185, Ano96-34, Ano98l, Api95c, Api97, Chi97, Gre96d, Gre96k, Gre97f, Gre97g, Het96g, KE97, Mil96, Rag97]. C-Cube [Chi97, KE97]. C/C [Ano95-47]. C/S [Het96g]. CA [Gre95e]. CA-Visual [Gre95e]. Cable [Ano98m]. Cables [Ano95-191, Cun95, Pou96p]. cabling [Cun95]. Cache [Ano95-96, Phi97, Dob98a, Mac98a]. Cachet [Ano95-96]. caching [Bri97a]. CAD [Ano98n]. Cadillacs [Sch97b]. CADzooks [Ano98n]. Café [Gre97h]. Cairo [Min96]. CalcPac [Ano95-213]. calculates [Ano95-213]. Calculator [Ano95-156]. Caldera [Ano97-64]. calendar [Ude97b]. Call [Ano95-105, Ano95-172, Ano96x, Ano97-93, Ano98o, Ano98-60, Emm96, Tra95, Way96f, Ano98-87, KH95]. called [Ano95b, Way95g]. Caller [ARH95]. calls [Ano95-38, Ano97v, Kay95a, Yag97a]. camera [Ano96-179]. Cameras [Ano96-42, Ano96-97, Arc98, MW95, Ano95b, Ano96-43, Ano97-117]. Campaign [Ude95e]. camps [Tho95k]. Can [Ano95-197, Ano96-66, Fri95d, Gre97r, Hur98c, Nan95d, Ude95i, Way97b, And96b, Ano95k, Ano95-45, Ano95-56, Ano95-61, Ano95-63, Ano95g, Ano95-88, Ano95-137, Ano95-186, Ano95-199, Ano97d, Ano97-122, Ano98-47, Ano98-123, BEM95, Bar95b, Bau95, Bie95, Bri97a, Bry95e, CCS97, Chi97, Dav95, Daw95, Dej95j, Dob98a, Ess95, Fen95, Fri96a, Fri96d, Fri97d, Goo95a, Gre96b, Gre96g, Gre97h, Gre97p, Hal95d, Hal98g, Hil95, HKM97a, Hum97a, Hus95, Joc95d, Key95, Krt98, Law95, Mac98a, Mac95b, Mar95a, Meh95b, Mul96b, Nan96d, Nan96g, Nan97c, Orc98, Phi97, Pou95t, Ric97, Row96c, Rud98, Sal95d, Sch96e, Sha95b, Sha95c, Sil97, Sim97, Smi95d, Tho95j, Tri95, TM95, Ude95a, Ude95b, Ude96g, Ude97k, Ude98a, Ude98b, Wat95, Way95c, Way95g, Way96b, Yag95c]. Can’t [LM95b, Goo95a, Kor95]. capabilities [Ano95x, Ano95-56, Ano96-178, Ano97-123, Fri97b, She97, Sto98, Ude97h]. capability [Sha95c]. capacities [Ano95-98, Ano95b]. capacity [Ano95-36, Esp95, Way95d]. Capture [Ano95-141, Ber95]. car [Ano98-81, Pot95]. Card [Ano95-105, Ano96-173, Coo96c, Hil95, Hum97a, KH95]. Cards [Ano95-66, Ano95-86, Ano95-90, Ano95-97, Ano96-80, Ano96-83, Ano96-133, Ano96-132, Ano96-129, Ano96-130, Ano96-131, Ano97-83, Ano97-84, Ano97-85, Ano97-86, HKM97a, KM96d, Pou96p, Tri95, Ano95-107, Ano98-92, DL95, KM96b, Yag95c]. Care [Cor97, Ano95-197]. careful [Whi97]. carry [Ano95-51]. Cartridge [Mia97b]. Cascading [Ude97c]. case [Sea97, Pou96a]. Cash [Ano98-118, Sin95a]. Cashing [Ano97-90]. Castanet [Los97c]. Casting [Gre95b, Gre95c]. catalog [Pou95a, Pou95h]. catalogs [Pou95h]. Catch [Ano95-60, Joc96c, Mon97]. Catches [Bal96d]. Category [Ano95-87, Ano96-81, Ano96-82]. cc [Ano96-173]. CD [Aer95, ATA96, Ano95-37, Ano95-36, Ano95-195, Ano95i, Ano95j, Ano96p, Ano96-50, Ano96-121, Bal95h, Bal95i, BDF96, BM95, CFDT95, DFT95a, DFT95b, Emi97, Fro96, HCM96, Hal96d, HM96, KFT95, Kay97a, MDM95, MG96, Nad95b, O’M96, Pou95n, RFD95, SK95, TJF95, UT95, UGB95, Ude96d]. CD- [SK95]. CD-QuickShare [Ano95-195, Bal95i]. CD-ROM [Aer95, Ano95-37, Ano95-36, Ano95-195, Ano95i, Ano95j, Ano96-50, Bal95h, Bal95i, BM95, CFDT95, DFT95b, Emi97, Fro96, KFT95, Kay97a, Nad95b, O’M96, Ude96d]. CD-ROMs [Aer95, ATA96, Ano95-37, Ano96p, Ano96-121, BDF96, BM95, CFDT95, DFT95a, DFT95b, HCM96, Hal96d, KFT95, MDM95, MG96, Pou95n, RFD95, SK95, TJF95, UT95, UGB95, Ude96d]. CDMA [Ano97g]. CDs [Ano95-195, Bal95i, Hal96d]. CE [Ano98-123, MG97, Way97a, Way97e]. CE-based [MG97]. CeBIT [Ano95-41, Ano98k]. CeBlT [Ano96-122]. celebrity [Abr98b]. Cell [Ano97v, Sta97a, Ano98-81]. cell-phone [Ano97v]. cells [Hal95e]. censorship [Meh95b]. cent [Hub95]. Centaur [Ano97w, Hal97c]. Centers [Ano96x]. central [Pou97a]. Centralized [Ano98p]. centric [MC97]. Centura [Het96g]. CEO [Ano97-58, Ano98-120]. certain [Tho98c]. CGI [Cli98d, Ude97k]. Chain [Mia95b]. chairman [Ano97-58, Ano98-120]. Challenge [Ano95-198, Hal97c]. challenged [Mon95b]. challengers [PH95, Ude97g]. Challenges [Ano96-108, Nan98a, Sha95c]. champ [Hal97a]. Change [Ano95l, Ano95-92, Ano96-69, Ano96-165, Ano98-31, Hum96c, Ano95-169, Ano95-194, Ano97-58, Hal95c, Hal98a, Ude97c, Win95]. changed [Bal95b, Way96g]. changers [O’M96]. changes [Dej95h, Dob95]. Changing [DM98, Dej95h]. Channel [And96a, Hur98b, Gla95, Lo97]. Chaos [Pou95j, Pou95v, Pou97f, Pou97k, Pou97l]. Charge [Lac96]. charges [Ano95-74]. Charisma [Yea95b]. Charter [Ano95-210]. Charting [Ano96z]. charts [Hig95d]. chat [Nan95c]. Cheap [Ano96-44, Esp95, Hum96g, Hum98b, Laz97b]. Cheaper [Ano95-49, Ano96-32, Ano97o, Ano97x, Sch98]. cheaply [KFT95]. check [Ano98-58, Cam98b]. Checking [Ano95-173, Orc98]. Checklist [Ano95-43]. CheckPont [Daw95]. checks [Gre96i, MF97c]. cheers [Pou97j]. Cheetah [Mia97c]. Cherokees [Sch97b]. Chickens [Pou96o]. Chinese [Zha96]. Chip [Ano95-44, Ano96-48, Cla95b, Gwe96, HM95, Hal98c, Sim96, Ano95-42, Ano95-145, Cam98a, DL95, Em98, Gwe95, Hal96h, Hal96k, Hal98e, MD95, Phi97, Pou95g, Pou98a, Pou97l, Pou98d, Tho95o]. Chips [Ano95-80, Ano95-131, Ano96-90, Ano97x, Ano97-59, Tho96o, Way96g, Way96l, Abr97b, Ano95-93, Ano95-128, Hal95c, Hal97c, PH95, Pou97a, Pou98b, Way95e]. Choice [Ano95d, Ano96a, Ano97a, Tho96c, Ano95-103, Ano97-30, Ano97-77, Api95a, Gil96d, Goo95b, Way98b]. Choices [Ano95-80, Bry95d, Hal96k]. choose [Ano95-104, Gil97c]. choosing [Ano97-105, AG95, Pou96a, Sal95m]. chores [Way95f]. chose [Rei95e]. CHRP [Ano95-132]. CISC [Hal95a]. Citibank [Ano95w]. Citrix [Qui97]. civility [Sch96b].

Civilization [Ano98-60, Abr98a]. claim [Ano95-64, Ano95-198]. claims [BVN95]. clarify [Ano95-117]. Class [Die97b, Api95c, Ano96-180]. class-library [Api95c]. classes [Ano95-61, Ano97-105, Cli98d, Tur97]. Classroom [DD97]. classy [Gre95e]. clatter [Kay98b, Kay98c]. Claudia [And96d]. Clean [Mil96, Vac95b]. Cleaners [Hur97d]. cleanser [Ano97-117]. cleansing [Kri95, Whi97]. Clear [Ano95-74, Mon95b]. clever [Chi96a, Mon95b]. Click [Abr98a]. Client [Gre97f, Het96c, MC97, OH95, OHE95, UGB95, Yag97a, Ano95-190, Ano97-103, Ano98-68, Ber95, Cli98b, DeJ96b, Dej95f, Dob97a, GE95, Het96g, Mac95b, Pou97b, QS96, Rei95e, Vog95, WS96, You96]. client-side [Dob97a]. Client/Server [Gre97f, Het96c, OH95, OHE95, UGB95, Yag97a, Ano95-190, Ano97-103, Ber95, Cli98b, DeJ96b, Dej95f, GE95, Het96g, Mac95b, Pou97b, QS96, Rei95e, Vog95, WS96, You96]. client/server-based [Rei95e]. Clients [Ano98-102, Sim97, Cam97a, Dob98d, Gil96f, Yag97b]. Clipper [Gre95e]. Clock [Ano96-86, Way95d]. clone [Tho95j, Van97]. cloner [Ano98-114]. Clones [Ano95-120]. close [Ano95r, Ano95s, Het96d]. closer [Bal95h, Mac98a, Way96j]. closest [Het96b]. Clothing [Tho96d]. Cloud [Ano95-109]. CLP [Pou95b]. Clustering [Ano97y, Ano98-114]. Clusters [Ano95-141, Mac98b]. Clutter [Ano96o]. CMOS [Ano97z]. COBOL [Ano95-46, Ano95-45, Bau96]. Code [Ano96-34, Ano96z, Ano96y, Ano96-27, BM95, Gre97l, Knu96a, TJF95, Ano95-45, Ano95-82, Ano95-157, Ano95-214, Ano98-78, Bry95d, CC98, Chi96a, Gla95, Joc95d, Lin95a, Ano95-45, Gre97l]. Codes [Esc95b]. Codetalk [Ano96-28, Ano95-46, Ano95-47, Ano96-31, Ano96-29, Ano96-30]. CodeWarrior [Cot97]. cogent [Jen95].

ColdFire [Cir95, Shi98]. collaborate [DM98]. Collabra [Ano95-75, BD95b]. Collection [Mil96, Tri95, Ude96l]. collectors [Ano95-97]. Collision [Kay95b]. Color [Ano95-48, Ano95-49, Ano95-106, Ano96-44, Ano97-62, Bal95d, Chi95, Hil95, HR97, Hum96d, Hum96i, KH96, Kay97b, Sug95, Wei95, Ano95-101, Ano95-207, Ano95-213, Ano97-118, Ano97-119, Ano98-123, Kay95d, Len95, Van95a, Ano95-107]. Colors [Kay95d]. colour [Tho95b]. column [Ude95c]. COM [Cka97, Gat98, Pou95o]. combinatorial [Pou95b]. Combine [Ano95k, Ude96h]. combined [Ano95-151, Ano95-152, Ano95-165, Gre97l]. combines [Ano95-203, Ano96-174, Ano98-81, Hum96g]. Combining [Sal95g, Yag96d]. combo [Abr97b]. COMDEX [Ano96-144, Ano95-141, Ano96l, Ano97k, Ano98-58, Pou97g]. Come [Ano95-122, Ano95-176, Ano98-52, Ano98-62, Nel95, Ano95f, Ano95-182, Ano95-201, Ano97g]. Comes [Ano95-156, Ano95-168, Ano95-173, Ano98-117, Ano95-128, Cla95b, Laz97a, Pou95v, Way96j]. Coming [Ano95-50, Ano95-127, Ano95-128, Ano96n, Ano96-32, Ano96-33, Ano96-62, Ano97-95, Flo96a, Ano95-70, Chi95, Pou95r]. Comm [Gio96b, Ano95-134, Ano95-135]. Commanding [Wat97b]. Commentaly [Eng95]. Commentary [Bar95b, Bon95a, Emm96, Fra96, Jen95, Key95, Les95, Log96, Mar95e, Meh95a, Meh95b, Nel95, Pri95, Rid96, Sha95a, Web96]. Comments [Ano96-34, Knu96a, Ano95-110]. Commerce [Ano96-146, Ano98-29, Ano96-76, Ano97s, Ano98-114, Smi95a]. Commercial [Ano97-28, Ano95-137, Cla95b, Pou95b, Smi95a]. commercially [Dej95g]. Commodity [Mac98b]. Common [Laz97b, Ano95-74, Gre96h, Tho95a, Ano95-132]. Communicates [Tho96i]. Communication [Smi95a]. Communications [Cli98c, Pou95l, Sar98, Ano95b, Ano95-134, Ano95-135, Ano97-29, Ano98-68, Cli98b, Gro96, KFT95, Ske98b, Way95a, Whi95b]. Communicator [Way95a]. community [Sch97a]. Commuter [Ano96-101]. compact [Coo95]. CompactFlash [Ano95b]. Companies [Ano95-51, Ano95-44, Ano95-154, Ano97-31, Ude95b]. companion [Ano96-40, Bal95d]. Company [Ano96-46, Ano96-47, Ano97-37, Ano97-38, Ano98-32, Ano98-33, Ano98-34, Ano95-206, Ano95-201, Whi95a]. Compaq [Ano95-52, Ano98-77, Ano98-96, Ano98-113, Row96c, RV96, Yag98a]. Compare [Sea96b, Ano95-104, Bry95d, Mar95b]. comparing [AG95]. comparison [Sha95b]. Compass [Bal96e]. Compatibility [Sal95c, Ano95-105, Ano95-184, Ano98-74]. Compatible [Tho95l, Ano95-120]. compelling [Ano95-133, Ano97-118, Mon95b]. compensates [Sta97a]. compete [Ano95-166, Ano95-167, Hum97a, Way95e]. Competing [Mar95b, Sta97b]. competition [Nan95d]. competitor [Het96b]. competitors [Hal97e]. compilation [Cot95c]. Compiler [Ano98-121, Ano95-72, Ano95-196, Cot95c]. Compilers [Ano95-173, Ano95-47, Hal98b, Way96l]. Compiling [Gre96e, Ano95-184]. Complete [Fri96c, Gre95d]. completely [Tho95e]. complex [Ano95-61, Cua95a, Eva98, Gou95]. complexity [Pou96b]. Component [Ano98-78, Cka97, Cli98a, Pou97b, Gil97a, Gre98c, Pou95f, Ude96g]. Componentized [Tho96e]. Components [Ano95-46, Ano95-45, Flo98b, HS96, Joc96d, PM97, Ren98, Ric97, Wat98b, Ano95-50, Ano97-103, Api95a, DeJ98, Tho96e, Way96j, Whi98]. Componentware [Ude96g]. Composes [Way97c]. Compression [Bry95d, Ano95-198]. compressor [KE97]. compressor/decompresser [KE97]. compromise [LM95a, Ano95-104]. Compromises [Ano98-89]. Computer [Laz96, SK95, Way95d, Ano95-60, Ano95-97, Ano97-58, Bou95, Bur95a, Hal95d, Joc95b, MDM95, Pot95, RFD95, Sal95b, Sch97d, She95a, TH96, Whi95b, Ano95-52, Gre95e]. computer-based [Whi95b]. computer-generated [Bou95]. computer-toting [Pot95]. computerize [Tri95]. computerizes [Pou95h]. Computers [Ano96-101, Ano96-162, Ano97-124, Ano98q, Cam97a, Joc96g, Kay98a, Les95, Tho95l, Tho96q, Ano95-77, Ano95-97, Ano95-117, Ano95-165, Ano95-190, Ano97t, Ano97-54, CC97, DFT95b, Esc95b, Hal98g, Joc96a, Joc97b, KP98, Mar95a, Sal95c, Sal95h, Sch98, Sha95c, Tri95, Way96h]. Computing [Abr97b, Abr97a, Abr97c, Alv96, Ano96-69, Ano96-156, Ano97f, Ano97-76, BN95, CE98, OHE95, Sch96a, UGB95, Ano95-110, Ano95-141, Ano96-163, Ano97-58, Arc98, Bas95, Ber95, Eva98, Flo98b, Hum96c, Kay98a, Nel95, OH95, Sal95h, Sha95c]. ComUnity [WS96]. concentrate [Yea95b]. concepts [DeJ96b, Hig95d]. concern [Ude95d]. Concert [Mor97]. concurrent [Ano95-222]. condense [Ano95-88]. Conferencing [BD95b, Ude96b, Ude96j, Ano96-78, BD95a, Sea95d, Sea95b]. confidence [Sta96c]. configuration [Fri97d]. configurations [Ano95-100]. conflict [Ric97]. confuse [Sch97b]. confusion [Ano95-125, MDM95]. congestion [Sta97a]. Connected [Ano96f, Ano96-143, Ano96-168, Sch96c, Cox98b, Joc96f, Ude95e]. connecting [Fri97e]. Connection [Ano95-121, Ano98-47, Ano97d, Fri96c]. Connections [Fri95c, Fri97c, Pou97a, Smi95a]. Connectivity [Hen96, Joc98a, Ano95-196, Ano97-123, Hum96h]. Connectiviy [Ano95-220]. connector [Ano95-193]. connects [Ude95a]. Conquer [Mul96b]. consider [Ano97-31, Fri97e, Pou96g]. Consistent [Sug95]. console [Wat97b]. Consolidate [Ano95-174, Wil95]. Constraint [Pou95b]. Construction [Ano95-57, Ano97-42, Row96c, Tra95, Gil97a]. consultant [Mau96]. Consumer [Tho96f, Tho96b, Tho96m]. consuming [Cox98b, Gre95f]. consumption [Ano95-106, Tho97a]. contact [Ano96-178]. Contender [Ano95-160]. Content [Fow97, Die95c, Abr97b]. Content-enabled [Fow97]. contentious [Ano97-29]. contingencies [Cro95]. Continues [Ano98-92, Pou95w]. continuing [Tho95e]. continuous [Ano98v]. continuously [CCS97]. contrast [Ano95-106]. Control [Ano98o, DeJ96b, DL95, Sal95d, Sal96k, Var95, Way97e, Won98b, Ano95-210, Ano97-105, Cre96, Dob97d, Gou95, Sch96b, Whi97, Api95a]. controllers [Sal95h]. Controls [Api95a, Ano95-45, Ano95-106, Wil95]. Convenience [Ano95-53, Ano98-88, Gre96e, Hum96e]. convention [Ano95-141]. converge [Rei95d]. Convergence [Ano96-35, Ano95-106]. Convert [Gre97c]. converters [Sch96e]. converting [Sea95c]. Converts [Ano95-46, Ano95-45, Ano95-212]. Cookin [Van97]. Cooking [Kay97c]. Cool [Bau96, Row97]. Cooperate [Hyd96]. coordinate [Dob95]. copier [Ano95-207]. Coping [Ano98-31]. Copland [Ano97t, Tho95i, Tho95j, Tho96g, Tho96i]. copy [Ano95-163]. CopyJet [Ano95-207]. Copyrights [Rid96]. CORBA [OH95, OHE97]. Core [Rau97, Ano95-61, Ano96-76, Shi98]. Corel [Ano97-27, Ano97-120, Pep97]. Cornerstones [Wel97]. Corporate [VN96c, Ano95k, Die95c, Fri97c, Moh97a]. corporations [Hed95, Rei95d].

Corrections [Ano95-81]. CorStream [Mia95c]. Cosmic [Gre97l]. Cosmo [Gre97l]. Cost [Ano95-38, Ano95-52, Ano96-126, Ano98-56, Row96b, Ano95-71, Ano95-195, Ano97-120, Ano98n, Ban97, Chi97, Cir95, Flo96a, Goo95b, Hub95, Hum96e, Hum96g, Hum98b, Meh95b, Mia97b, Sha96, Tho95k, Tho95m, Tho97c, Yag95c]. cost-effective [Ano95-195, Mia97b]. cost-effectively [Yag95c]. Costs [Ano98-61, Sal95d, Ano95-169, Ano95-222, Ano98-47, TM95, Whi95b]. Cote [Ano95-72]. could [Abr98a, Ano95-38, Ano95-61, Ano95b, Ano95-128, Ano95-129, Ano95-193, Ano96-55, Dob98d, Flo95, Gwe95, Hal97e, Hur97d, Pou97c, Tho96f]. counterattack [Hur97a]. counts [Cor97]. couple [Ano95-192]. course [Way96g]. Covad [Cam98b]. Cover [Ano95-190, Kay98b, Kay98c, Ano95-113, Ano96-163]. coverage [Ano95-112, Ano96-77]. covering [Ano95-28, Ano95-34, Ano95-29, Ano95-30, Ano95-31, Ano95-32, Ano95-33, Ano95-35]. covers [Way96h]. Cozy [Hes97]. CPU [Ano96-40, Ano96-176, Ano96-170, Ano97-110, Ano98r, Ano98-56, Arc98, PH95, Tho95h, Tho97e, Way95d]. CPUs [Ano95-127, Ano95-128, Ano96-99, Fau95, Gre95a, Joc96e, MA98, Tho97b, Tho98b, Way96l]. CPX [Cot95b]. crabby [Bon95a]. crack [Way95g]. Cracker [Hur98a]. Cracking [Sch98]. cracks [Sta96e]. cramping [Sta97b]. crash [Hal98g, Pot95]. crash-prone [Hal98g]. crashed [Pou96k]. Crashing [Ano95-54]. Crave [Way96c]. Cray [Way96h]. Crazy [Ano95-217]. Create [Win95, Ano96-68, Ano97-122, Cka97, EP97, Pou97b]. creates [Coo95]. Creating [Ano96-36, SK95, Ano95-162, Pou95d, Ude98a]. Credibility [Hac97]. credit [Ano98-92]. Crisis [Hac97, Web96, Joc95d]. Critical [Fom98, CCS97, Ske98b, Vog95]. crooks [Ano95w]. crop [BKM96, Cam97a, Hil95].

Cross

[Mor97, Api95c, Gre95a, Gre97j, Los97b]. cross-development [Api95c]. Cross-Platform [Mor97, Gre95a, Gre97j, Los97b]. Crossroads [Sch96a, Bas95]. Crow [Pou96k]. CRTs [Ano95-70]. CruisePad [Ano95-210]. Cruiser [Ano95-211]. Cruising [Ano96-123]. crunch [Abr97a]. crunching [And95b]. cryptanalysis [Way95g]. Crypto [Sch98, Way95g, Way96d]. cryptographic [Way98b]. Cryptography [Los97a]. Cryptolope [Ano97-48]. CSS [Ude97c]. CTI [Ano95-43, Ano97-28, Ano97-29, Ano97-30, Ano97-31, Ano97-105, Joc97b, Sil97]. CTO [Sto98]. Cube [Chi97, KE97]. Cubicle [Die96b]. Cuff [Sar98]. culture [MDM95]. Cup [Sin96a]. Cure [Ano98-104, Knu96b]. Cures [Gre96d]. Curing [Mia95a]. current [BKM96, Cam97a, Hal96d, Hig95d, Kos96, Tho96g]. currently [Ano95-71, Ude97d]. Custom [Ano97-92, Ano95-107, Arc98, QS96, Sha96]. customer [Ano98-47, Tho95d]. customers [Hed95, Ude95a]. customizable [Ano97-122, Esp95]. Cut [Ano95-51, Ano95-103, Ano95-222, Whi95b]. Cut-Throughs [Ano95-103]. Cuts [Ano95-38, Ano96-126]. Cutting [Cor97]. Cutting-Edge [Cor97]. Cyber [DD97, Die97b, Abr97c]. CyberDavid [Mar95e]. Cyberleaf [Ano95-200]. Cyberspace [Ano95-168, Flo96c]. Cyrix [Ano96-37, Ano96-38, Hal97c, PH95, Pou96k]. cyrus [Pou96h, Pou96j, Pou97f].

D [Ano95e, Ano95f, Ano95-60, Ano95g, Ano96-48, Ano96-126, Ano96-133, Ano96-173, Ano97-97, Ano97-110, Ano97-119, Bal95b, Bou95, Dej95c, Dej95b, Die95a, EP97, Flo96a, Fou95a, Fou95b, Gre97l, Hal96f, Hal97c, HKM97a, Hum96b, Hum97a, Joh96, Lov95, MD95, Pou96m, Row96a, Row97, Tho96a, Tho96r, Yar95]. Daisy [Mia95b]. Daisy-Chain [Mia95b]. Damn [Ude96a]. Dan [Ano97-112]. Dance [Yag97b]. Dangerously [Sch97d]. Dark [Ano95e]. Darkness [BGR95b, BGR95a]. DAT24 [Mia97a]. Data [Ano95-57, Ano95-55, Ano95-56, Ano95-89, Ano95-97, Ano95-100, Ano96x, Ano96-32, Ano96-39, Ano96-62, Ano97u, Ano97-32, Ano97-63, Ano98-63, Bre98, Dej95f, Dej95d, Dej95e, Dej95h, Dob98a, Fry98, Gre97h, Hac97, Het96d, Het97, Hum96e, Hur97d, Kay98b, Kay98c, Kri95, Lin96d, Mys96a, Nan97b, Nan98b, Nan98c, Pen98, Pou97a, Pou96o, Hed95, Sal95i, Sal96i, Tri95, Wat95, Wat97a, Whi97, Ano95-103, Ano95-134, Ano95-135, Ano95-162, Ano97f, Ano98-66, Ano98-100, Ano98-119, Bau95, Ber95, BMW97, DeJ96a, Dob97c, Eva98, Fri96a, Gar97, GM95a, Hal96d, Los97a, Mac98a, Mia95d, Mon97, Sim97, Ske98a, Ske98b, Ude96l, VN96b, Way95g, Wel97, Yag97a, Yav95]. Data-Acces [Ano98-63]. data-analysis [Ano95-56, Bau95]. Data-Mart [Wat97a]. Data-Mining [Kri95, Hed95]. Data-Warehouse [Ano95-57, Fry98]. Database [Ano95k, Ano95-151, Ano95-152, Cua95a, Cua95c, Dob95, Dre98, Gre98c, Joc98a, Nan95b, Ano95-59, Ano95-64, Ano95-72, Ano95-196, Ano96-175, Cox98b, Gre98a, Gre98b, Gre98f, Het96d, Kay97c, Lin95d, Way95h, Wel97, Ano95-95]. database/SGML [Ano95-95]. Databases [Ano95-59, Ano95-64, Ano95-166, Ano96n, Ano98s, Ano98-101, BMW97, Cox98b, Dob97c, Fom98, Kay97c, Nan98a, Ske98a, Wat98a, Ano95k, Ano95-62, Ano95-167, Cua95a, Dob96, Gil96f, Gre97f, Gre98f, Nan97b, Rud98, Sto98, Vog95]. Databses [Ano95-167]. Datapro [Hes97]. date [Ano95-212]. daunting [Yav95]. Dawn [Ano97-33]. Day [PF97, Ano95-146, Ano95-174, Tho97b].

Days [Ano96-183, Ano98t, Sal95a, Ano95-141, Bal95b]. dazzling [Hal95b, Kay95c, MD95]. DB [Ano97-47, Gre98c]. DB2 [Nan95b]. DB2/ [Nan95b]. dBase [Car95]. DBMS [Ano97-47]. DC [Pou95m, Pou95q]. DC/X [Pou95m]. DCE [Cli98a]. DCI [DL95]. DCOM [Cli98a]. Dead [Pou96o, Ano95-135]. Deal [Ano98n, Cro95, Fri96a, Tra95]. Dean [Pot95]. Death [Ano97v, Pou95m]. debates [Ano95-113]. decade [Hum96c]. decades [Ano95-28, Ano95-34, Ano95-29, Ano95-30, Ano95-31, Ano95-32, Ano95-33, Ano95-35, Bar95d]. decent [Ano97-82, Cre96, Gre97p]. decides [Ano95-163]. Decision [Gio96c, Gil96d]. Decisions [Ano95-43, Sim97, Vog95, Wel97, Whi97]. decompresser [KE97]. Decongesting [Ano97-34]. dedicated [Cla95b]. Deep [Flo98b]. defeated [Hal97a]. Defending [Ano97-98]. define [Ano95l]. definitions [Ano95-59]. degrees [Ano97-105]. Deja [MB97]. delayed [Tho95j]. Delete [Abr98a]. Deliver [Ano95-131, Ano96x, Ano96-113, BD95b, Hal96c, Het96d, Ude97k]. delivered [Ano95-197]. delivering [Ano95-107, Ano95-165]. Delivers [Ano95-118, Nad95b, Ano95-157, Ano97z, Ano98-126, Bak95b, Die95d, Gil96f, Gre97e, Hum96e, Tho95m]. Delivery [Ano96-166, Ano97-45, Ano97-101, DeJ97]. Dell [Ano98-96]. DeLorme [Ano97-107]. Delphi [Ano95-58, Ano95-72, Ano95-196, Gre97g, Gre97n]. Delrina [Ano95-207]. delta [Nan95c]. DeltaPoint [Ano96-179]. demand [Ano95-77, Ano95-136, Cob95, EVN96a]. demanded [Ano95-60]. demands [Sup97]. Demon [Hal96h]. demons [Ano95c, AG95]. Demystified [CL97]. Demystifying [Fri97e]. density [Pou98a]. departmental [Cla95c]. departments [Ano97-87]. dependency [Ano95-220]. Depends [Gar98, Ano95-95]. Deploying [Ude98a, Ano95-201, Die95c, Gre97j]. Deployment [Ano98e]. Deschutes [Tan98]. describe [Ano95-116]. describes [Ano95-103]. Design [Ano95k, Ano95-186, Ano96p, Ano96-174, Bie95, Dob95, Mul96b, Ude96k, Ano95-196, Ano97-114, DeJ96c, Fau95, Gwe95, MC97, Smi95b, Tho96h, Vac95b, Ano97-118, Ano97-113]. Designed [Ano98h, Ano95-222, Api95c, Het96d, Ude96l, Wel97]. Designer [Yar95]. designers [Ude97c]. Designing [Fau95, Ano95-73, Mon95a, You96]. Desk [Cla95c, Sea96d, Cla95a]. Deskpro [Ano95-52]. Desks [Ano98-47]. Desktop [Ano95-48, Ano95-59, Ano95-126, Ano96-117, Ano97-67, Ano98-103, Ban97, Bou95, Gar98, Gil98b, Laz97b, Sea95d, Sea96a, Sea96c, Van96a, Van96c, Yag98b, Ano95-60, Ano95-70, Ano95-77, Ano95-137, Ano95-154, Ano95-165, Ano95-166, Ano95-167, Ano96-100, Ano98-51, Ano98-113, Chi97, Cla95b, Die95c, DL95, Fau95, Fri95c, Joc95a, Lac96, Los97a, Pou97c, Pou97h, Pra95a, Sea95b, Way95d]. Desktop-conferencing [Sea95d]. Desktops [Ano98p, Ano98-48, Ano98-96, Bal95h]. Despite [Hal95c, Sal95a]. detail [Ano95-37]. detailed [Hal97d]. Details [Ano98y, PH95]. Detecting [Ano98u]. Detection [Hur98a]. detectives [Pot95]. Detectors [Ano96-140]. develop [Nan95g]. developer [PM97]. Developers [Ano95-60, Ano95-198, Ano97-57, Ano97-111, Ano98-80, Joc98a, Cla95b, Dob98d, Fen95, Gil96f, Gre95e, Gre96g, Gre97d, Hal97d, Tra95, Van96d, Var95, Way95b, Way98a]. Developing [Gre97h, Gre97r, Knu96c, Ano95-72, DeJ96b, Die95c]. Development [Ano95-166, Ano95-167, Ano96-31, Ano96-164, Api95a, Gil97a, Gre96k, Gre97o, Gre98d, MG97, Mor97, Way98b, Ano95l, Ano95-47, Ano95-56, Ano95-72, Ano95-109, Ano97-106, Ano97-105, Ano98-84, Ano98-115, Api95c, Bie95, Cot97, Dej95j, Gre95d, Gre95e, Gre97l, Joc97b, Lin98, Mon95a, Nad97, Sal96b, Sta97b, UGB95]. developments [Ano95-150]. Device [Ano96-27, Ano98-82, Ano95-38, Ano95b, Ano96-177, Die95a, Joh96]. device-driver [Die95a]. Devices [Ano96-122, Ano98c, Lev98, Ano95b, Ano95-199, Coo95, Sal95h, Ude95e, Wat97b, Whi98]. Devilish [Ano98y]. Diagnosing [Nan96c, Nan96a]. Dial [Gil97b, Mul96a, BHKM96a]. Dialects [Ano95-61]. Diamond [Ano96-41]. Dictation [Ano97-69, Ano98v]. did [Ano95-154]. Didn’t [Ano95-199, Pou95p]. differ [Nan95b]. difference [Way95c]. Differences [Ano95-184, Ano95-185, Cot95c, Ano95-118]. Different [Gil98b, Ano95-132, Ano95-157, Bry95e, Cla95c, Dav95, Mia95d, VN96b]. differential [Way95g]. difficult [Pou95b]. difficulties [Flo96a]. difficulty [Ano97-105]. dig [Wat95]. Digital [Ano96-42, Ano96-43, Ano97-48, Ano97-88, Arc98, Lev98, MW95, Pou95o, Pou95n, Ude97d, Ano95b, Ano96-179, Ano97-117, Cla95b, Die95c, DL95, Joc95a, MDM95, Pou98c, Van96c, Yag95a, Ano97-110, Ano98-76, Ano98-86, Chi96a, Fau95, Fen95, Pou97a, Tho96e, Way97d]. Dimension [Ano95-55, Ano95-56, Dej95e, Esc95b]. dimensional [Bal95b]. Dimensions [Dej95f]. diodes [Ano95-36]. direct [Van96c, Vog95, Ano95-145]. Direct3D [Joh96]. Direction [Ano98-73]. Directions [Ano98-63, Ano98-120]. directories [Hur97b]. Directory [Ano98-73, Gil97b, Kor95, Lew96, Los97b]. Dirt [Ano96-44]. Dirt-Cheap [Ano96-44]. Dirty [Hur97d]. disabilities [Laz96]. Disabled [Laz97a, Laz96]. Disappearing [Ano95-71]. Disaster [Ano96-113, Ano96-34, Joc95d]. disasters [Pou97f]. discover [Hed95, Gre97p]. Discovery [Gre97p]. discs [Ano95-36, Hal96d]. discuss [Ano95-117]. discusses [Ano98-120, Gat98, Pou95l]. Disk [Mia97c, Ano98-104, Ano98-113, Kay98b, Kay98c, Pou96j, Pou98c]. Disks [Ano95a, Kay98b, Kay98c, Ano96-178]. disparate [Gar97]. dispatch [Tho95c]. displace [Nan95a]. display [Chi95, DL95]. Displays [Ano95-70, Ano95-36, MD95]. Disposable [Hal98a]. distant [Sal95j]. distinctions [Ano95-61]. distinctly [Cla95c]. distortion [Ano95-106]. distributable [Ano95-184]. distribute [Cot95c]. Distributed [OH95, Cha97, Cli98c, Cua95a, Eva98, Gil95, Gre97h, Gre97n, Knu96c, Los97c, Cli98a]. distributed-object [Gil95]. distribution [MC98]. Ditto [Hum96e, Hum96e]. DiVideD [Ano98-27]. Dividing [Hus95]. diving [Dej95j]. DLT [Ano95-100]. DMV [Whi96]. DNA [Way95h]. Do [Ano98v, BKM96, Chi96a, Ude96d, Ano95y, Ano95-74, Ano95-199, Ano96-179, Ano97d, Ano98-115, Bal95c, BHKM96a, Hal95e, Ind95, Joc95d, Mar95b, Ric97, Smi95b, Ste95, Ude98b, Way95g]. do-it-yourself [Ano98-115]. Dock [Nan96f]. Docks [Ano96-101]. Doctors [Bea96]. Document [Mac95b, Mac95a, Sea95a, Sea95c, BD95b]. documentation [Ano95-106]. Documents [Boy97, Ano95-61, Ano95-88, Ano95-212, Ano97-118, Ano97-112, BVN95, Sea95c]. DocuPrint [Ano97-62]. dodging [Pou95k, Ude97k]. Does [Car95, Gre97g, She97, Tho98b, Ano95-210, Cre96, Hal97e, MD95, Sch97a, Sta96b, Ude95i]. Doesn’t [Joc95d, Bar95b, Gre96d]. dogs [Ano98-81]. Doin [Tra95]. Doing [Ano95-108, Ano98w, Rei95b, Ano95-187, Vac95a]. dollars [Sal95d]. dominant [OHE95]. Domino [Ano98-70, Gil96f]. Done [Ber95]. Don’t [Ano96-162, Nee96e, Pou96i, Pri95, Way96d, Kay97c, Mar95e]. Doom [Abr98a]. Doors [Sal96m]. DOS-based [Gre95e]. DOS-Compatible [Tho95l]. Double [Ano96-45, CC97, Hum96f]. Double-speed [Hum96f]. Doubleplay [Ano95-211]. doublers [Ano95-125, Goo95a]. doubles [Ano95-211, Hum96a]. Douglas [UT95]. Down [Ano95-44, Ano95-62, Ano95-55, Ano95-186, Ano95-211, Gre97r, Hum96b, Hur97d, Pou96k, Tho96f, Wat97a, Way96h]. downloaded [MF97c]. dozen [FH97]. dpi [Ano95-118, Ano97-120, Ano98-126, Hum96i]. Dr. [Ano95-208]. Drafting [Ano95-107]. Drag [Gre96d, Tur97]. Drag-and-drop [Gre96d]. DragNET [Ano95-62, Way95h]. Dragon [Ano97-46]. DragStrip [Ano95-214]. Draw [Ano97-41, KM96d, KM96b]. drawing [Ano96-174, Ano98-51]. dread [Ano97-60]. Dream [Ano95-158, Ano97-119, Ude95g]. Dreaming [Eng95]. Dreams [Ano98-101, Ind95]. drill [Ano95-55]. drill-down [Ano95-55]. Drilling [Lin96d]. Drive [Ano95-44, Ano95-169, Mia97c, Nan97b, Pou97f, Ano95h, Hum96e, Hum97d, Pou95m, Pou96j]. driver [Die95a]. Drivers [Ano96-27, Dob96, Joh96]. Drives [Ano95-139, Ano95-140, Ano96-141, Ano98-43, DeJ96d, Emi97, Mys96a, Ano95-117, Ano95-100, Ano95-195, Bal95i, GM95a, Kay98b, Kay98c, Mia97b, Nad95b, Pal96, Pou96l, Pou98c]. Drop [Ano95-165, Gre96d, Tur97]. DSL [Ano98y, Ano98x]. DSP [Ano95-127, Ano95-128, Bal95h, Dav96]. DSPs [Cla95b]. Dual [Ano98z, Ude96b, Hum96b, Pou97d, Ano97-110]. dual-CPU [Ano97-110]. Dual-Mode [Ude96b]. Dubious [Ano95-79, Ano95-106, Ano95-104, Ano95-100, Ano95-103]. due [Ano95-49, MFDG98]. duel [Gas95]. Duet [Ano95-213]. Duette [Cam97b]. Duplex [Ano96-148]. during [Ano95-82]. dust [Tho97b]. duty [And95b, Ano98n]. DVD [Ano97-35, Ano98-27, Hal96d]. Dylan [Pou96e]. Dynamic [Ano98-28, Dob97a, Dob97b, Dob98a, Way96b, Hum97a].

Dynamite [Kri95].

E-Cash [Ano98-118]. E-Commerce [Ano98-29]. e-mail [Ano97-53, Ano97-97, Ano97-127, Kos96, SY97, Ude96h, VN96a, KFT95, Kos96, Ude96c, Way96e, Hig95d, Vau96, Way95a, Ano97-127, SY97, Sta98]. e-mailers [Abr97a]. E-Pocket [Ano98-118]. Each [Ano95-74]. earlier [Ano95-107]. Early [Ano98m, Ano95-132, Ano95-158, Ano97-48, Tho95j]. Ears [Cla95b]. Earth [Bar95a, RV96, Hum96b, Kin96, Way96g]. Ease [Ano96z, Ano95-106, Car95, Hum96e]. eases [Ano95-123]. Easier [Ano95-49, Ano95-200, BVN95, Ano95-52, Ano95-146, Bla96, Cka97, DeJ96c, Flo95, Fri95a, Joc95a, Joc97b, Sch98, Son95, Way95g]. easily [Ano97-122, Nan97c]. East [Way98a]. Easy [Way98a, Ano96-179, Ano98-68, Api95a, Gil97a, Hig95c, Hum98a]. eat [Hur97c]. ECCO [Mon95b]. ECMAScript [Dob98b]. Economical [Cla95b, Ano97-82]. economically [Sal95g]. economize [Nan95f]. Economizers [Nan95f]. Edge [Ano96-41, Cor97, Ano95l, Hum96h]. EDI [Ano97s]. editable [Ano95-212]. Edited [Ano96-155]. editing [Ano95-165, Ano97-120, BVN95, Chi97, Hum98a, Mar95c]. Edition [Ano95-209, Ano95-111, Ano95-184, Ano95-185, Ano95-204]. Editor [Ano96a, Bee96, Fed97, IF97, MF97b, Mon96a, Nav97, Rag97, Ano95-184, Mar95d, Ude95c, MH97a, MH97b]. Editorial [Ano96-46, Ano96-47, Ano97-37, Ano97-38, Ano97-36, Ano98-30, Ano98-31, Ano98-32, Ano98-33, Ano98-34, Nee95l, Nee95j, Nee95k, Nee95a, Nee95b, Nee95c, Nee95d, Nee95e, Nee95f, Nee95g, Nee95h, Nee95i, Nee96c, Nee96a, Nee96d, Nee96e, Nee96f, Sch97d, Sch97c]. editors [Ano95-41, Ano95-141, Bar95c, Ano97a]. education [Ano95-117]. Effect [Nee95l, Pou97d]. Effective [Ude98a, Ano95-195, BD95b, Mia97b]. effectively [Ind95, Sea97, Yag95c]. effects [Bou95]. efficiency [Pra95a]. efficient [Bry95d, Sal95d, Ste95]. egotists [Bon95a]. Eight [Ano98t, BHKM96b, Cam98a, Hum96c, Hum98c, Sea95c]. Eight-Way [Hum98c]. Electric [Flo96b, Ano96-76]. Electronic [Ano95w, Ano96-55, Mon98a, Ano97s, Ano97-116, Sea95c]. elements [Tho96h]. elevators [Ano95-74]. eliminating [Way95f]. Elsa [Ano96-48]. else [Gre97j]. elsewhere [Ano98-100]. elusive [Ano95-146]. Elwood [Way95c]. Email [Ano97-107]. embarking [Ude95e]. Embeddable [Ano96-119, Fri97b]. Embedded [Ano98-101, Gre96g, Ano98-79, Cir95, Goo95b, Gre95d, Krt98, Lin98, MA98, Sal95h, Shi98, SB96, Whi98, Sal95h]. Embedded-systems [Gre96g]. Embrace [Ano95-134, Ano95-135]. Emerges [Eva98]. emphasis [Gre97n, Hal98e]. employs [Pou95d]. empower [Vog95]. empty [Hal95e]. Emulation [Bal95c, Ano95-135, Fri97d, Loh97]. emulator [Tho95h, Tra97]. enable [GE95, Gre98a, Laz97b, MA98, Smi95a, Tho97d]. enabled [Bal95e, Fow97]. Enables [Laz97b, Cli98b, Cli98c, KP98, KE97, Yag97b]. Enabling [Ske98b]. enclaves [Bon95a]. encoder [Chi97]. encoding [KE97]. encounter [Ano98w]. Encroaches [Ano95-48]. Encryption [Sta97b, Sch98]. End [Ano95-63, Lin95a, Yav95, Ano95-60, Ano95-56, Ano95-64, Ano95-77, Ano95-119, Ano95-151, Ano95-152, Ano98-41, Ano98-60, AG95, EP97, HKM97a, Lin95d, Pou97k, Ude95g, Vog95, Ano95-64, Ano95-77, Ano95-119]. end-user [Ano95-64, Ano95-64]. Endian [Sta95a]. Ending [Ano97-71]. ends [Ano95-55, Ano96-55, Dej95e]. enemies [Ano95-62]. enemy [Tho97b]. Energize [Dob98d]. enforce [Ano95-222]. engaged [Sch97d]. engages [Pou96i]. Engine [BHKM96b, Ude97h, VN97b]. Engines [Tho96r, Ano95-98, Ano95-196, Tho96a]. English [Ano95-211]. enhance [Fou95a, Fou95b]. Enhanced [Pou98d, Ano97-47, Die95d, Gil95, Way95a]. enhancements [Api95c]. Enhances [Ano95-220, Cli98a, Moh98]. Enhancing [TC97, Ano95-201]. Enliven [SP97]. enormous [Ano95-136]. enormously [Bal95b]. Enough [Pou98e, You96, LM95b, Nan95a, Nan95b]. ensues [Ano95-163]. ensure [Ano95-61, Sta96d]. ensures [Ano95-184]. Enter [Gil96c, Tho97b]. entering [Moh97a]. Enterprise [Ano96-74, Ano97-39, Ano97-76, Ano97-100, Ano97-98, Ano97-99, Cha97, Gil95, Gio95, Het96e, Nan95b, Pou97b, Ske98a, Ske98b, SM97, Wat97b, Wat98a, WS96, Ano95-59, Ano95-190, Ano97-29, Fri96d, Het96a, Het97, Hur97b, Lin98, Nan96d, Won98a, Ber95]. Enterprise-ready [SM97]. enterprise-wide [Het97, Nan96d]. Enterprising [Los97b]. Enters [Ano98-68, Hal96j]. entertainment [Dej95b]. entire [Nad95c]. Entomophagous [Ano96-49]. entry [Gre95e, Wil95]. Environment [Ano96-31, Ano95z, Ano95-47, Ano95-72, Ano95-196, Api95c, Cot95b, Gio95, Gre97e, SB96, Yag96d]. environments [Ano95-97, Ano95-100, Pou95f, Way95b]. Envy [Gre96d]. Epic [Way97c]. EpixPro [Ano96-179]. Epson [Kay97b]. Equal [Emi97]. Equals [Pou96m]. equipment [Ano95-137, Mul96b, Rei95a, Rei95e, Ano97-110]. Equium [Ano98h]. equivalent [Pra95b, Ude97k]. Era [Hal96d, Hal96j, Ano95-168]. Eric [Ano98-120]. ERP [Ano98-40]. Error [Ano95-81, Ano95-173, Ano97r, Cob95, Orc98, Pou96j]. Errors [Tho95a]. ESI [Ano95-211]. ESI-9580A [Ano95-211]. especially [Bal95c, Lew96]. Esperant [Lin95d]. Essbase [Ano95-56, Dej95e]. essential [Dej95b, MFDG98]. essentials [Way97g]. establish [Fri97d]. eSuite [Gil98b]. Ethernet [Ano95-27, Ano95-98, Ano95-103, Ano95-191, Ano98-44, Cun95, Fri95c, Lo97, Mia95b, Nan97c, Sal95g]. EtherPhone [Nan97c]. Eudora [VN97a]. Eudorable [VN97a]. Eurocom [And95c]. Europe [KFT95]. evaluate [Ano95-106]. evaluates [Ano95-105, Gio95]. Evaluating [Het96g]. evaluation [GM95b]. Even [Ano98-126, Fou95a, Fou95b, HKM97a, Moh97b]. Event [Fre96, Tho95c]. Events [Tho95c, Tho95d]. eventually [Die95a]. Ever [Kay98b, Kay98c, Ano97-121, Bry95d, Gil98a, Hal98g, Hil95, Lin95a, Wei95]. ever-growing [Ano97-121]. Every [Api97]. Everyman [Ano98x]. Everyone [Ano96b, Chi97, Flo96a, Hur97b, Lin96a]. Everything [DeJ96f, Ano95-137, DeJ96a, Hum96e, Hed95, Way96h]. Everywhere [Ano96-110, Ano97-89, HS96, Sch96a]. evolve [Nad95a]. Evolves [Fow97, Lin98, Wil98]. Evolving [Cla95c, Cla95a]. exacerbating [Abr97a]. exactly [CL97]. examines [Cra96, Fri96d]. eXceed [Bak95b]. excel [Ano97-81, Ano95-55, Ano95-56]. excellent [Ano95-118, Ano96-40, Gre97e]. Exceptions [Orc98]. Exchange [Vau96, VN96a, Yav95, Gar97, Yag97a, Het96b, Mar95b]. execute [MF97c]. execution [Esc95b, Pou95d]. executive [Ude95c]. existing [Tho95e]. Expanding [Joc95e, Ano95-142, Ano95-143, Sal95a]. Expands [Ano95-154]. expect [Ano95-77, Ano95-199, Hum96e, Kay98a, Fri97b]. expected [Ano95q]. expensive [Ano95-44, Ano95-206, Cox98b, Dob96]. experiences [RFD95]. experimental [Abr98b]. expert [Fri96d, Lin95d, Wel97]. experts [Hum96c]. Explained [Dob97a, Dob97b, Dob98a]. explains [Ano95-82, Gil95, Joh96, Ude95f, Ude96b, Ude97h, Wel97]. explanations [Kay97a].

Exploit [Ano95-93]. exploiting [Ber95]. exploration [Ude97b]. explorations [Pou95k]. explores [Pou95n]. explosion [Pot95]. Expo [Ano95q, Ano96w]. Exponential [Hal96h]. exports [Sta97b]. express [Ano95-116]. Expression [Ano96-174, Pou95i]. extend [Arc98, Fri96d, Lew96]. EXTender [Ano98-82]. Extending [Ano97-100, Ano97-98, Ano97-99, Wes96b, Wes96a]. extensible [Ano95-72, Cli98a]. extensive [Ano95-106, Ano95-104]. External [Ano95-105, Hum96e, KH95]. Extranet [Ano97-100]. Extranets [Ano97-99, Ano98-35]. extraordinary [EG96a]. Eye [Ano96-157, Em98, Nee96c, HM95]. eye-popping [HM95]. Eyes [Cla95b, Coy98].

F [Bou95, Pou95f]. F/X [Bou95]. fab [Abr97b]. Face [Ano96-102, Dav95, Dej95g, Yag98a, Hal96k]. Facelift [Hal96l]. facial [Dej95g]. facial-recognition [Dej95g]. facility [Ano97-32, Nan95c]. facing [Win95]. Factoring [Ano95-65]. factory [Dej95h, Esc95b, Pot95, Tri95]. Failing [Ano98-104]. Failure [Pot95]. Faking [Vac95b]. fall [Ano95-140, Pou97f, Ano95-141]. falling [MFDG98]. falls [WS96]. familiar [Dej95e, Mia95c, MB97]. family [Pou97a]. fans [DFT95b]. far [Ano95f]. Fashion [HM95]. Fast [And96a, Ano95-99, Ano95-191, Ano97-59, Ano97-77, Ano98-108, Ano98-124, Cun95, Dob97c, FH97, Gre97f, Hal96f, HR97, Hum96g, Mon98a, Mys96a, Sch96c, Ano95-98, Ano95-158, Ano97-118, Dej95g, Guy95, Hub95, MD95, Mia97b, Nan95a, Sal96f, Smi95d, Smi96c, Tho95m, Tho97b, Tho97e, Won98c]. Faster [And95b, Ano95-49, Ano95-66, Ano96-32, Ano96-50, Ano97-41, Ano97-40, Bri97a, Bri97b, Hal96k, Tho98a, And95a, Ano95-191, Ano95-206, Ano97-116, Ano98a, Hal95c, Joc96e, Sch98, Son95, Tam95a]. Fastest [Ano98-121, AG95, Mia97c, Tho96q, Tho97b, Ano95-44, Ano95-107, Ano95-105, Pou95d, RV96]. Fat [MC97, Sha95a]. fatal [Pot95]. Fatter [Ano96-50, Ano97o]. Fault [CCS97]. favorably [Pou95r]. favorite [Ano95-47, Pou96j]. Fax [BEM95, Mia95a, Won98a, Ano95-38, Ano95-51, Ano95-134, Ano95-135, EVN96a]. fax-on-demand [EVN96a]. faxes [BHKM96a]. Faxing [Ano95-38, Ano97d, BEM95]. FBI [Way95h]. Fear [Tra96b]. Feature [Fil96, Ano95-133, Bal96d, DL95]. Feature-Rich [Fil96]. featured [Row96c]. Features [Ano95-71, Ano95-119, And95c, Ano95k, Ano95-98, Ano97-102, Ano97-114, BD95b, Cha97, Fri97a, Gas95, HKM97a, Hum96g, Hum96h, MFDG98, Nan95c, She95c, Smi95a, Smi95b, Tho98c, Tur97]. features-for-the-price [Smi95b]. February [Ano95-113]. Federated [Gre98c]. feedback [Ude95a]. feeding [Ano95-122]. Feel [Gre97j, Pou97j, Ano95-161, Gre97d]. fees [Sal95d]. feet [Api95c]. fetch [Ude98c]. fever [LaP96a]. few [Ano95-101, BVN95, Pou97g]. Fewer [Ano97-66, Sal96k]. Fiber [Mon97]. Fibre [And96a, Hur98b, Lo97]. field [Gre95e, Rei95a, Rei95e, Tri95]. fields [Wil95]. Fight [Ano96-96, Cra96, Gil96a]. Fighter [Ano96-138]. Fights [Hal98e]. File [Ano95-67, Ano95-123, Ano96-107, Ano97-101, Fri96a, Nan95c, Yag95b, Ano96-170, Ano97-123, Ano98-74, CFDT95]. file-delta [Nan95c]. files [Ano95-157, Car95, Tho95a]. Filling [Ano95-68]. Film [Fil96]. filtering [VN97a]. Final [Wei95]. Finally [Ano96-51, Hur97e, Kay96a, Ano95-107, Ano95-120, Nad95a]. finance [Hac97, Pou98a]. finances [Ano95-203]. Financial [Bau95, Ano97-92, Cla95a, Dej95b]. find [Ano95-107, Ano97-32, Ano98-78, BGR95b, Boy97, Flo95, Ind95, Kor95, MW95, Nad97, Pot95, Pou95n, Sal95j, Ude95b, Van97]. Finding [Ano95-162, Kay95a, RM98]. finds [Pou95r, Pou95s, Pou96k, Pou96l]. Fine [Esc95b, CC98, Esc95a]. Finis [Lin95a]. Finish [Sea96c, Sea96a]. Fire [Ano96-138, Pou97d, Joc96c]. Fires [Ano98e]. Firewall [Bry95c, RM98, Bry95a, Bry95e, Daw95]. FireWall-I [Daw95]. Firewalls [Ano98-39, Bry95b, Bry95e, Cob95]. firm [Whi95b]. Firms [Qui97]. First [Ano96-52, BDF96, Pou98b, Pou95m, Tho97a, Tho98c, Ano95r, Ano95s, Ano95-120, Ano95-158, Ano95-154, Ano95-206, Em98, Emi97, Flo96a, FH97, Hal97d, Pou96l, Tho96e]. Fit [Ano96-118]. Fits [Ano97m, Fri97c, Joc95e, Pal96]. Five [Ano97-42, DeJ97, Hig95c, Hum96j, Ano95-60, Bry95e, Hal95e, Pou95b, Pou98c, Sch96e, Sea97]. five-year-old [Sch96e]. Fiver [Ano95-64]. fixes [Tho95f]. Flaming [Sch96b]. Flannel [Moh97a]. flap [Ano98w]. Flash [Ano95-69, Coo96a, Ano95b, Gla95]. Flashier [And95b, And95a]. FlashPix [Ano96-53]. Flat [Ano95-70, Pep98]. Flattery [Ano95-177]. flavors [Fri97a, Gio95]. Flawed [Key95, Lin95d]. flexibility [Ano97-116, Mia95b]. Flexible [And96a, Arc98, Dob97c, Yag95b, Cli98c, Goo95b, Pou97b, Ude96g]. Flies [Chi96c, Nee96d]. flights [Pou95m]. Float [Ano95g]. floor [Dej95h]. floppy [Ano95h]. Flow [Tra96b, Ano96-127, Cro95, Vac95b]. flowchart [Ess95]. flowcharting [Ano95-210]. Fly [CR96a, Won98c, CR96b]. Flying [Ano96-54]. Focus [Cla95c, Cro95, Dej95j, Nad95c, Pou95h, Rei95e, Sal95k, Sal95l, Way95h, Whi95b, Ano95-46, Ano95-45]. Follow [Ano96-55, Min96]. font [Kay96a]. Fooling [Pou97e]. Footprint [Gre98e, MC97]. force [Nad95c, Pou98a]. Forces [Ano95e, Ano96-165, Pou96j]. forecasting [Ste98]. forecasts [Nel95]. Forefront [Joc97a]. Forever [Ano98l]. Forges [Nan95d]. Forget [Ano95-157, Way95d]. Forging [Die96a]. form [KFT95, Sea95c]. formats [CFDT95]. former [Ano98-114]. formerly [Het96g]. Forms [And96e, Ano96-55, Het96e, Ude98a, Het96a, Ske98a, Ude96l]. forum [KFT95]. Forward [Fou95a, Fou95b, Ano98-89]. Forwards [Ano95-103]. Foul [Pou96r]. found [Nan95a, Sea95c]. Foundation [Tur97]. Four [Ano95-44, Ano96-56, Nad95b, Pou98c, RFD95, Yag98a, Emi97, OHE95, Sea96b]. FoxPro [Ano95-166, Ano95-167]. FP [Hal98e]. FPU [Hal95e, Sha95b]. FPUs [Gre95a]. Fractal [Ano96-174]. Fractals [Ano95-198]. fragmenting [Hal98a]. Fragrant [Pou96r]. Framed [HG96]. framework [Pou95f, Ber95, WS96]. frameworks [Tho98e]. Freaks [Var95]. Free [Ano95-71, Ano95-137, Ano97-43, DL95, Gre96c, Hub95, Meh95a, Meh95b, Ude97j, Way95b, Ano96u, Moh97a, Smi95a]. FreeBSD [Ano95-137, Hub95]. FreeHand [Ano96-172]. Freelance [Yea95b]. freely [Ano95-132]. Freeware [Ude97j, Fri97a]. friend [Abr97c]. friendly [Tam95c]. Frogs [Pou95o]. front [Ano95-55, Ano96-55, Dej95e, Lin95d, Vog95]. front-end [Vog95]. Frontier [Pou97h]. Fronts [Ano98b]. frustration [Ano95-116, TM95]. fuel [Tho97a]. Fuels [Ano95-146]. Fujitsu [Way97g]. fulfill [Ind95]. full [Ano95-157, Ano95-204, Ano95-209, Joc95e, Mia95d, Row96c]. full-featured [Row96c]. full-motion [Ano95-204, Ano95-209]. full-size [Joc95e]. fun [DFT95b]. Function [Ano96-55, Pou95d]. functions [Ano95-127, Ano95-128, Sta96d]. fundamental [Ano97-29]. fundamentally [Ano95-197]. fundamentals [TM95]. funky [Ano98-51]. further [Bou95]. Furthers [Ano95-123, Hal95a]. Fusion [Yag96d]. Future [ATA96, Ano95-36, Ano95-65, Ano95-219, Ano95-221, Ano96-57, Ano97-44, Ano97-125, Ano98-36, Cam97a, Eng95, Hum96c, LaP96b, MD95, Pou96e, Pou97b, Tho96e, UT95, Ude97g, Ano95-61, Ano95-71, Ano96-77, Ano96-78, Ano96-163, Ano97t, Ano98-119, Flo98b, Gat98, Kay98a, Min96, Nel95, Rei95f, Sta96d, Tho95e]. Fuzzy [Ano95-74, Gou95, Hal95a].

G [Ano95-62, Flo95, Way95h]. G-men [Ano95-62, Way95h]. G3 [Ano98-115, Tho97a]. Ga [Ano95-72]. gaining [Cha97]. galaxies [Hed95]. Galaxy [Nan95g]. Gamble [Yag98b]. Gambles [Way96l]. Game [Pou95j, Sim96, Ano95-73]. Games [Ano95-73, Ano96-120, Mon95a, Ano95e, Ano95f, Bal95b, Die95a, Sal95i]. gangs [Ano95-195]. Gantt [Hig95d]. Garbage [Mil96]. GartnerGroup [Ano98-38, Ano98-37, Ano98-40, Ano98-39, Gar98]. Gates [Gat98]. Gateway [Ano95-205, Ano96-40, Ano96-174, Ano96-175, Ano98-85]. Gateways [Bon95b, Daw95]. gauge [Mul96b]. gave [Hur98c]. Gb [Lo97]. Gears [Ano98-44, Nan98d]. gems [Ude97j]. General [Ano95-107, Way96l]. generated [Bou95]. generating [Pou95g]. Generation [Cka97, Ano95-49, Ano95-124, Hal96d, Het96b, Hig95d, Joc95b, PH95, Shi98, Tho95j, Tho97a, Way95b]. generations [Cla95b]. generic [Ste95]. genetic [Cla95c, SK95]. Genius [Bla96]. geography [Sha95c]. geometry [Em98]. Geoprocessing [Ano96-58]. Get [Ano95h, Ano96-59, Ano96-60, Ano97-108, Ano98-41, Gre95d, Hal96l, Kay96a, LS98, Sal95n, Tho98a, Ano96-163, Ano97-87, Ano97-97, Ano97-127, Fri98b, Gre96b, Gre96d, Hig95c, Hil95, Joc95b, LM95b, MF97a, Min96, Pou98b, Sch98, Smi95d, Ude95a, Ude95e, Ude98b, Way95h]. Gets [And96c, Ano95-47, Ano95-89, Ano95-88, Ano97z, Ano97-28, Ano98-67, Ano98-69, Ano98-125, HG96, Los97a, MA98, Sta98, Tho95m, Ano95-163, Bla96, Dob95, Esp95, Gre97g, Pou95v, Pou97k, Sch98, She95c]. Getting [Ano95-74, Ano98-42, Kay95d, Ano95-89, Ano95-146, Cox98b, Joc95a, Joc96e, MDM95, Nee95m, Sch96b, Way95g]. Giant [Ano98-43, Ano98-63, Ano98-113, Mar95e, Mia97c]. Giants [DV96]. Gift [Ano96-63]. Gig [Van96b]. Gigabit [Ano98-44, Lo97]. Gigabyte [Ano95-140, Hal96d, Boy97, Kay98b, Kay98c]. gigantic [Rei95f]. Give [Sal96e, Gre97p, Hal96h, HR97]. gives [Ano95-27, Ano96-178, Cro95, Dob97c, Var95, Wat97a]. giving [Vog95]. Glasgow [Moh98]. Glass [Ano95-82]. glimpse [Min96]. Glint [Em98, MD95, Ano96-48, Em98]. glitches [Hum96h]. Global [Ano96-61, Ano98-123, Gro96, Hur97b, PF97, Ude95b, Fin98, Whi95a]. Globalist [And95b]. Globalization [Fow97]. Globalizing [PF97]. Globalyst [And95a]. Glossary [Ano95-103, Ano95-107]. GMR [Ano98-43]. GMX [Em98]. gnashing [BEM95]. Go [Ano95w, Ano96-97, Bal95d, Chi95, Kay98a, Lov95, Sal96e, Van96b, AK98, Ano95c, BGR95b, Bea96, Nan96g, Pou96l, Tam95d, Ude95i]. goal [WS96]. God [CFDT95]. Goes [ARH95, Ano96-58, Ano96-158, Ano98i, Ano98-123, Api95c, Fri98a, Gre97m, Pou97c, Rei95e, Tho97d, Ano96-174, Rei95a, Tho96b, Tho96m]. going [BGR95b, Hur97c, Sal95j, Smi96c]. Gold [Hed95, Dej95d]. golf [DFT95b]. Goliath [Mar95e]. Good [Ano96-88, Gas95, Hal97a, Kay97a, Pou98e, Sta95b, You96, And95b, Ano96-174, Bal95e, Cam97a, Eva98, Fen95, Goo95b, Hum96g, Nan95a, SB96, Tho95d, Ude97i, VN97b]. Good-Bye [Hal97a]. got [Gre95a, Hac97, Kay97c, Tra95]. GPS [Kin96]. Grade [Cam97b]. Grand [Sha95c, Hum98b, Ude95e]. Graphic [Dej95b, Ano95-66]. Graphics [Ano95-66, Ano95-107, Ano96k, Ano96-53, Ano96-172, Dej95c, DV96, Em98, HKM97a, KM96d, Lov95, Ano95-60, Ano95-204, Ano96-171, Ano98-115, CFDT95, Dej95b, DL95, Die95a, Flo96a, HM95, Hum96b, Hum97a, KM96b, MD95, Pou97h, Pou97e, Pou97j, Pou97k, Sea97, Die95a, Row97, Yea95b, Ano95-209]. graphics-chip [DL95]. grapple [Sha95b]. Gray [Moh97a, Ano95-118]. gray-scale [Ano95-118]. Great [Pou97g, Yag95b, Ano95-157, Ano95-203, Ano95-201, Ano97v, Bal95d, Hum96g, Mia97a, Tho95m]. Greatest [Bar95a, Pou97e]. greatly [Ano95b]. Grehan [Ano95-47]. Grip [LS98]. Ground [Ano96-27, Ano96-182, Api95a]. group [Cob95, Tho95k, Ude97b, Ano95-38]. groups [Hig95d]. Groupware [Ano95-75, Ano98-45, Ano98-49, Rob96, Ude95b, Ano95-142, Ano95-143, BD95a, BD95b, DA95, Gil96d, Het96b, Tra95, Joc95c]. GroupWise [Gil96d, Mar95b]. Grow [Ano95-119, Fra96]. Growing [Ano96-111, Joc98a, Ano97-121, Nan96b, Sta96b]. Grows [Ano98-108, Kos96, Way95c]. growth [EG96a, Sta96d, Tho95e]. GSM [Ano97g, EG96a]. GTE [Dej95j]. guarantee [Mac98a]. Guaranteed [Ano97-45, Joc95b]. guard [Lin95d]. Guerrilla [Ano98i]. GUI [Gas95, Hal97a, Nan95a, Sha95a, Tho98e, Vac95a]. Guide [Ano96r, Ano98j, Ano98s, Eng97, UGB95, Hal96k, TC97, Ano95-39]. Guides [Ano97-42, Gre97r]. GUIs [Hal96l, Sha95a]. gulp [CC97]. guru [You96]. Guy [Var96].

H.32x [Dav95]. Hackers [Ano98u, Ano98-46]. hair [BEM95]. hair-pulling [BEM95]. hall [Abr98b, Die97b]. halls [Ano95-141]. hallway [Ano95-211]. Hand [Ano96-135, Ano98-69, Ano98-125, CE98, MG97, Sup97, Way97f, Ano95-97, Ano96-172, Ano97-60, Ano98-123, Tho96f, Way97a, Whi98, Tri95]. Hand-Held [Ano96-135, MG97, Sup97, Way97f, CE98, Ano95-97, Ano96-172, Ano97-60, Tho96f, Way97a, Whi98, Tri95]. hand-helds [Ano98-123, MG97]. handier [SY97]. handle [BEM95, Cro95, Gre97h, Pou97a, Ske98a]. handlers [Tho95c]. Handles [Yav95]. handling [Ano96-173]. Hands [Law95]. Hands-Off [Law95]. handwriting [Ano95-162]. handwriting-recognition [Ano95-162]. Handy [Ano95-97]. hanging [Ano95-195]. Happen [Sal96h]. Happened [Ano96-169, Dej95j]. Hard [Ano95-139, Ano95-140, Pou97f, Ano95-83, Hur97b, Mia97c, Pal96, Pou95m, Pou96j]. hardly [Pou95j]. Hardware [Ano95-39, Ano95-132, HKM97b, HR97, KHB96b, Kay98b, Ano95d, Ano96-173, Ano96-177, Ano96-178, Ano97-33, Ano97a, Ano97-117, Ano97-119, Ano98-114, Gre95a, Joc97b, Laz97b, Nan95a, Rud98, Way95g, Tho95j]. Hardware/Software [Ano95-39]. Harhor [Tho96e]. Harvard [Yea95b]. hassle [Won98b]. Hats [Ano98-46]. Hauppauge [Ano97-106]. hawing [Ano98-41]. haystack [Boy97]. HDS [Way96m]. Head [Ano96-154]. Head-to-Head [Ano96-154]. Headaches [Ano95-63, Ano97-66, Sal96k]. Headed [Ano98-120]. headings [Ano95-61]. Headroom [Nan98e]. Heads [Ano98-43, Ano98-93, Gre96b, Pou97k]. Health [Ano95-197, Cor97]. Hear [Ano98-50, Ano98v]. Heart [Sup97, Joc95b]. Heartthrobs [Joc95b]. heats [Wat98b]. heaven [PM97]. heavily [Ano95-80]. Heavy [Cam98a, And95b, Ano98n, BEM95]. heavy-duty [And95b, Ano98n]. Heavy-Hitting [Cam98a]. Heels [Ano95-194]. hefty [Ano97-114]. Held [Ano96-135, MG97, Sup97, Way97f, Ano95-97, Ano96-172, Ano97-60, CE98, Tho96f, Way97a, Whi98, Tri95]. helds [Ano98-123, MG97]. Hell [Pou95s]. Hello [Hal97a, Ude95c]. Help [Ano98-47, Cla95c, Sea96d, Ano95-191, Ano96-55, Ano97-32, Ano97-54, Cla95a, CFDT95, Flo97, Gil97c, Goo95a, Ind95, KFT95, MFDG98, Meh95b, Mon95a, Nan96d, Pou97e, Sha95b, Sil97, Sug95, Tho95j, TM95, Ude95c, Way96b, Ano96-170]. Help-Desk [Sea96d]. helped [Vac95b]. Helper [Ano95-217, Tho96n]. Helpers [Sea96d]. Helpful [Ano95-76]. helping [Ano95-97, Gou95, Way95h]. Helps [Laz97a, Ano96-173, Ano97-28, Ano97-92, Ano97-118, Dob95, Pou97a, Whi95b, Wil96, Yag96d]. hemming [Ano98-41]. Herd [Joc95c]. Here [Ano95-77, Ano95-161, Ano95-201, Ano97-127, Bry95e, CCS97, Cla95b, Dej95j, Fri95a, Gwe95, Hal95c, Hal98g, Hil95, Joc95d, Mar95a, MD95, Mor97, Nad97, Pou95g, Pra95b, Hed95, Rei95c, Sha95c, Sta96b, Tho95a, Ude96g, Ude97c, Ano95-187]. Heresy [Jen95]. hesitation [Ano95-163]. Heterogeneous [Gio95]. heuristics [Ano95-110]. Hewlett [Ano95-186, Ano95-208, Ano97-114, Hal97d, Hum96j, Son95, Yag98a]. Hey [Way96f]. Hidden [Mar95a]. hierarchical [Nan95f]. High [Ano95-77, Ano95-151, Ano95-152, Ano95c, Ano96-184, Ano98-48, DL95, Hum96j, LM95b, She95b, She95a, Abr98c, Ano95-60, Ano95-56, Ano95-119, Ano95-169, Ano97l, Ano97z, Ano98m, AG95, Bal95e, Cir95, DM98, EP97, Goo95b, Gre95f, HKM97a, Nad95b, Nan95g, Pou97k, Sha96, Tho95k, Ano95-77]. High-end [Ano95-77, Ano95-151, Ano95-152, Ano95-60, Ano95-56, Ano95-119, AG95, EP97, HKM97a, Pou97k, Ano95-77]. high-level [Nan95g]. High-performance [DL95, Sha96, Tho95k]. High-Quality [Ano96-184, Ano97z]. high-resolution [Gre95f]. High-Speed [Ano98-48, Ano97l, Ano98m, DM98, Nad95b].

High-Tech [She95b, She95a]. Higher [Ano96-27, Ano96-62, Ano95-36, Ano95-77]. higher-capacity [Ano95-36]. Highlight [Ano96-122, Pou95q]. Highlights [Ano95-28, Ano95-34, Ano95-29, Ano95-30, Ano95-31, Ano95-32, Ano95-33, Bar95d, Ano98-59]. highway [Gai96]. hills [Dej95d]. him [Ano98-104, Pou97j]. HiNote [Ano98-76]. hint [Ano95-201]. Hints [Ano95-76]. history [Aer95, Gwe96]. hit [Ano97t, Ano97-115, Rei95c]. Hitachi [Sup97]. Hits [Tho96p, Way95f, Whi96, Cot97, Pou97i]. Hitting [Cam98a]. Hofstadter [UT95]. Hold [Ano95-139, Ano95-140, Ano96-50, Hal96d, Wei95]. holding [Ano95-122, Ano95h]. Holdouts [Ano95-125, Goo95a]. Holes [Smi96a, MB97]. Holiday [Ano96-63]. Hollywood [Bou95, EP97, Joc95a]. Hollywood-quality [EP97]. Hollywood-style [Bou95]. Holograms [DeJ98]. Holographic [Ano96-36]. Holy [Dob98b]. Home [Ano95-171, Ano96-64, MG96, Pou97a, Ano95-137, Fri96c, Gre97d, Ano98-82]. home-desktop [Ano95-137]. Home-EXTender [Ano98-82]. honor [Ano97a]. Honorable [Ano95-78, Ano95-79, Ano95-106, Ano95-104, Ano95-101, Ano95-100, Ano95-103, Ano95-107, Ano95-105, Ano96-65, LM95b]. honored [Hig95d]. hood [Ude95f]. hope [Pou96b, Sch96e]. hopes [Ano95-169]. hopped [Ano97m]. Horizon [Ano95-66, Ano96-124, Lac96]. horror [Ano95-73]. horse [Smi96c]. horsepower [Cla95b]. hosed [Ano97-97]. host [Api95b]. Hostile [MF97a]. Hot [Ano95-80, Ano95-150, Ano97-78, Cir95, Kay97c, Pou96j, Pou97f, Sch97e, Sin96a, Ano95-107, Bal95a, Ber95, Row97, Bal95g]. HotMetal [Ano95-200]. House [Ano97-70]. housekeeping [Way95f]. houses [Ano95-134]. how-to [TC97]. HP [Ano95-207, Ano96-73, Ano98-77, Ano98-115, Ano98-114, Ano98-126, Law95, Mia97a, Pou95c, Pou95d, Sar98, Son95]. HP-UX [Son95]. HPCs [MG97, Way97a]. HR [Ano97-87]. HSC [Mar95d]. HSM [Ano98-100, Nan95f]. hst [Yav95]. HTML [And96e, Ano95-61, Ano95-200, Ano98-28, Ano98-49, BVN95, Dob97a, Dob97b, Dob98a, MFDG98, Sch96e, Ude95f, Ude98a]. HTTP [Cli98b, Sta96b, Ude98b]. Hub [Ano95-103]. Hubs [Ano95-103, Ano95-102, Sal95g]. huge [Ano95-88, Joc95c, Pal96]. Human [Ano97-87, Ano95-88, Nel95]. humming [Ano97-45]. Hummingbird [Bak95b]. hundreds [Ano98-58]. hunger [Tho95o]. hungry [Ano95-122]. Hunting [Kor95, Pou97h]. Huzzahs [Ano95-110]. Hybrid [Ude96d, Ano95-95, Sal95f]. hype [Ano95-157, Nan95a]. Hyper [Flo95]. Hyper-G [Flo95]. hypermedia [Flo95]. hypertext [Ano97-112, Nel95, Ano95-61].

i-publish [Ano97-118]. I/O [Ano95-159, Ano98-97, Bri97b, Tho97c, Way95d]. I2O [Ano97-44, Bri97b, Tho97c]. i960 [Tho96n]. IA [Hal98d]. IA-64 [Hal98d]. IBM [Aer95, Ano95-145, Ano95-164, Ano95-186, Ano96-74, Ano97-47, Ano97-48, Ano97-102, Ano97-103, Ano98-43, Ano98-51, Ano98-113, Bal95h, DeJ95a, Flo98b, Gre97k, Gwe95, Hal97a, Hal98c, Joc95e, Tam95a, Tho95o, Ude95g]. ID [ARH95]. IDBC [Joc98a]. IDE [Gre97l]. idea [Wat98b]. ideas [Ano97-118, Ano98w, Ste98]. identification [Ude97d]. IDs [Ude97d]. If [Ano95y, Ano95-122, Ano95-125, Ano95c, Ano98-52, FH97, Gou95, Kay97c, Mia95c, Tho95l, Bry95d, CC98, Nad95b, Sal95k]. II [Ano97c, Ano98-77, AK98, Ano95-214, Ano97b, Ano98-75, Ano98-76, DeJ98, Dob97b, Hal97d, Hal97e, MDM95, Nan98b, Way98a]. III [Dob98a]. IIOP [Cli98b]. IIS [Nan97a, Tan98]. I’ll [Ano96-75]. I’m [Pou95j]. Image [Ano96-104, Bla96, Pou95g, Sea95c, Ano95-106, Ano97-114, Ano97-117, Hum98a, Mar95d, Mar95c, MW95, Sea95a]. Imagery [Ano96-43, Ano96-173]. Images [Ano95g, Ano96m, Ano95-118, Ano97-97, Hum96i, Kay97b, Pou95g]. IMAP [Kos96, SY97]. impact [Ano97-127, Dob98d]. Implement [TM95, Tho95j]. implementation [Ano97-103, Sil97, Way97d]. implements [Sha96, Ude95b]. impolite [Ano97t]. important [Lin98]. Impossible [DFT95a, DFT95b, Kos96]. Impossible-to-Use [DFT95a, DFT95b]. Impostor [Gil96c]. imprecision [Sha95b]. impressed [Pou95r]. Impression [Ano98-126]. impressive [Ano95-192, Api95c, Hum96j, Mar95d]. Improbable [Ano97-49, Ano97-50]. Improve [Ano95-197, Ano96-74, Ano96-101, Ano96-104, Ano96-116, Kay95c, Ano95b, Ano95-219, Ano98-47, Orc98, Phi97, Ste98, Way96b]. Improved [Sta96d, Tho95f, Ano97-114, Ano97-116, Nan95a, Rei95d]. improvement [Sea95c]. improves [Ano95-194, Ano95-204, Ano95-209]. improving [Ano95-197, Tho96n]. inappropriate [Ano95-222]. Inbox [Ano97-55, Ano97-51, Ano97-52, Gre97b, Ano96-76, Ano96-77, Ano96-78, Ano97-53, Ano97-54]. inch [Ano95-106, Ano95h, Joc95e]. include [Ano95-204, Ano95-209]. includes [Ano95-214, BD95b, Esp95, Nan95c]. including [Ano95-55, Ano95-100, CC97, Gio95, Hal98b, MG97, Nan95c, PH95]. incorporate [Van95a]. Incorporates [Ano96-48]. incorporating [Ano95-136]. increase [Pou98c]. increased [Esp95]. increases [Sal95g]. increasing [Ano95-77]. increasingly [Nan95f]. Incredible [Joc95e]. indeed [Sch97a]. independence [Cli98d]. independent [She97, Way96j]. independently [Laz96]. indepth [Ano96-164]. Index [Ano95-86, Ano95-87, Ano96d, Ano96-46, Ano96-47, Ano96-79, Ano96-80, Ano96-81, Ano96-82, Ano97-37, Ano97-38, Ano98-32, Ano98-33, Ano98-34, Ano98-53, Ano98-54]. Indexing [Ano95-88]. Industrial [BEM95, Ano95-196]. Industrial-Strength [BEM95]. industries [Ano95l]. Industry [Ano95-183, Ano95-60, Ano97-58, Ano97-92, Cob95, Sch97d]. inefficiencies [Ano98-121]. inefficiency [Cor97]. Inexpensive [Ano95-131, Ano97-62, Ano95-137]. inexpensively [Nan97c]. infested [Dej95j]. Infinite [DeJ98]. inflict [Ude97d]. influence [Ano95-141]. Info [Ano96-166, Ano95-55, DeJ97, Moh97b]. Information [Ano95-88, Ano95-114, Ano96-113, EVN96b, RFD95, Whi98, Aer95, Ano95-95, Ano96-172, Ano97-32, Cro95, EVN96a, Gar97, Ind95, Sim97, Ude97i, Wat97a, Ano97-124]. informative [Ude97h]. informed [Vog95]. Informix [Sto98]. Infrared [Ano95-89, Ano95-208, Way95a, Ano95-89]. infrastructure [Ske98b]. ing [Fen95]. initiative [Ano95-192]. Ink [Ano97-88, Ano98-124, KH96, Ano95-101, Kay97b]. Ink-Jet [Ano95-101, Ano98-124, KH96, Kay97b]. ink-jets [Ano95-101]. innovation [Ano97a]. innovations [DL95]. innovative [Ano95-141, Pou95g]. Inprise [Ano98-55]. Input [Ano96e]. Inquiry [Ano95-86, Ano95-90, Ano96-80, Ano96-83, Ano96-133, Ano96-132, Ano96-129, Ano96-130, Ano96-131, Ano97-83, Ano97-84, Ano97-85, Ano97-86]. Ins [TC97, Ric97]. Insanity [Ano98-99]. Insider [Aer95]. insights [RFD95]. Insignia [Ano95z]. install [Ano95-163, Bry95e]. Installation [Pou95s, Ano97-114, Hum96h]. Installations [Bau96, Cor96, Whi96, BEM95]. installing [Pou95s]. installs [Pou97l]. instant [Ess95, Smi95d]. Instead [Ano96-139, Ano95-161]. Instincts [Joc95c]. integrate [Ano95-127]. Integrated [Ano95-91, Ano95-72, Ano95-128, Api95c, Cro95, Ren98]. Integrates [Ano95-169, Ano98-79]. Integrating [Joc97b]. Integration [Ano96-51, DM98, Lin96b, Ano95-157, Ano97-114, Bur95a, Fin98, Hur97b, Mon95b]. Integrator [Tra96a, Ude95g]. integrators [Bau97]. Intel [Ano95-206, AK98, Ano95-92, Ano95-93, Ano95-97, Ano95-115, Ano95-159, Ano96-86, Ano96-76, Ano97m, Ano98-56, Api95c, Hal95a, Hal96b, Hal97c, Hal97d, Hal97e, Hur97a, KB96, PH95, RV96, Sch96a, Tan98, Tho96n]. Intel-based [Ano95-97]. IntelliCAD [Ano98n]. Intelligence [Ano97-118]. Intelligent [Flo97, Mul96b, DeJ96c, Tho96n, Wat95]. Interactive [Ano95-94, Ano96-95, Gai96, Dob96]. InterDev [Ano98-108, Gre97q]. interest [Sch97a]. Interface [Ano98r, Joc96b, Mac98b, Nee96b, Tur97, Wes96b, Goo95b, Hal97a, HKM97a, Hum96h, Mia95c, Wes96a, DL95]. Interfaces [Ano97-124, Bau95, Die95a]. Intergalactic [OHE95]. InterMark [Ano95-144, Gre95a]. Internal [Ano95-38, Pou97f]. international [Ano95-41, Sta97b]. internet [Hig95b, Ano95-51, Ano95-71, Ano95-75, Ano95-82, Ano95-83, Ano95-95, Ano95-137, Ano95-117, Ano95-126, Ano95-121, Ano95-142, Ano95-143, Ano95-136, Ano95-148, Ano95-168, Ano95-174, Ano96-33, Ano96-76, Ano96-78, Ano96-107, Ano96-103, Ano96-157, Ano97f, Ano97s, Ano97-56, Ano97d, Ano97-127, Ano98-64, Ano98-71, Bar95a, Bon95a, Bon95b, Bri97a, Cob95, Daw95, DM98, Dob98a, Eng97, Fin98, Fri96d, Gla95, Hig95c, Hur97e, Jen95, Kay95a, LaP96a, Lew96, Los97a, Mul96a, RFD95, Sal96d, SC96, Sin95a, Smi95a, Smi95d, Smi95c, Sta96b, Sta96c, Sta96d, Tur97, Ude95d, Ude96d, Ude97f, VN97b, Way98b, Won98a, Won98c]. Internet-Aware [Sal96d]. interoperability [Ano95l, Ano95-194, Way95b]. interoperate [Bal95c]. interpreter [Ano98-79]. intervention [Ano95-88]. Interview [Ano97-57, Ano97-58, Knu96a]. interviews [Ano96-87]. intracompany [Ano95-38]. Intranet [Hum97b, Bau97, Gil97b, Bau97]. Intranets [Flo97, Ske98b]. introduce [Ano95q]. introduces [Mia95d, Row97]. introduction [Gre97p]. introductions [Ano95-77]. intruders [Bry95c]. Intrusion [Daw95, Hur98a]. intrusions [Cob95]. Intuit [Ano95-203]. intuitive [Sha95a]. Invades [Ano95-171]. Invasion [Ano98-92, Ano98-91]. invasive [CL97]. inventing [Hal98b]. inventor [Nel95]. inventorying [MC98]. invocation [Cli98c]. involve [Hig95d]. involved [Fri95a]. involves [Gre97f]. Iomega [Hum96e, Pou96l]. IP [Ano95-116, Ano98-68, Cli98b, Dut98, Sch96h, Sch96g, Sta96c, Ude95d, Way96f, Win95]. IP-address [Sch96g]. IPv6 [Fin98]. IR [Nan96f]. Iris [Coy98]. Iris-Scanning [Coy98]. Iron [Ano95-190]. IRQ [Pou96i]. Irresistible [DeJ96d]. ISDN [Ano95-194, Ano96-88, Ano96-125, Ano96-134, Fri95b, Fri96c, Fri96d, FS96, Sal95e, Sal95f, Sal96e, TM95]. ISDN/analog [Sal95f]. Isn’t [Ano96b, SC96, Ano97-75, Gre97k, Lin96a]. Isochronous [Fri95c]. Issue [Bur95c, Bur95a]. Issues [Pou95l, Sta95a, Ano95-113, Ano96u, Die95c, Fri96c, Pou96g].

J [Gil98a, Pou95e]. Jack [Ano97-107]. Jackpot [Ano98-58]. Jake [Way95c]. James [Pot95]. Jasmine [Nan98a]. Java [And96e, And97, And98, Ano96-31, Ano96-66, Ano96-89, Ano96-90, Ano97-43, Ano97-57, Ano97-58, Ano97-65, Ano97-111, BDF96, Cli98c, Cli98d, DD97, Die97b, Dre98, Fed97, Gil97b, Gil98a, Gre96b, Gre97c, Gre97d, Gre97e, Gre97h, Gre97k, Gre97j, Gre97l, Gre97m, Gre97o, Gre97p, Gre98c, Gre98a, Gre98b, Gre98d, Gre98f, Hal97b, Hal98b, Hal98f, Hes97, Joc98a, Kay97a, Lin98, MF97d, MF97c, MH97a, MH97b, Nav97, Orc98, OHE97, Pep97, Rag97, Sch96a, Sch97c, Sho96, Sin96a, Tur97, Ude97k, Ude97b, Way96b, Way96l, Way96m, Way97c, Way98a, Way98b]. Java-based [Gre98c]. Java-enable [Gre98a]. JavaBeans [Gil98b, Moh98, Ren98]. JavaOS [MC97]. JavaScript [Bal96c, Fri98b, Mon96a, Ude97a]. JavaSoft [Ano98-59]. Javatalk [Gre97c, Gre97d, Gre97e, Gre97i, Gre97k, Gre97j, Gre97l]. JBuilder [Ano98-55]. JDK [Ano98-59]. Jerram [Ano96-55]. Jerry [Ano95-163, Ano98-104, Ano98w, Pou95l, Pou95n, Pou95m, Pou95k, Pou95p, Pou95q, Pou95r, Pou95s, Pou95v, Pou95w, Pou96i, Pou96j, Pou96k, Pou96l, Pou97h, Pou97f, Pou97g, Pou97i, Pou97j, Pou97l, Pou98c, Pou98d]. Jet [Ano95-101, Ano98-124, Gre97d, KH96, Kay97b]. Jet-assisted [Gre97d]. JetEye [Ano95-211]. jets [Ano95-101]. JetSend [Sar98]. JFactory [Gre96b]. Jingle [Ano98-118]. Job [Joc95d, Cre96, Joc97b]. Jobs [HKM97b, Cot95c, Ind95, Krt98]. jocks [Sal95i]. join [Ano97-124]. Jon [Ude95c]. Josephson [Ano96-169]. Joy [Pou95e]. JSafe [Way98b]. juiced [MA98]. jump [Api95c, Pri95]. Junctions [Ano96-169]. junk [Abr97a, Ano97-97]. Just [And98, Ano95-137, Ano97-59, Bar95b, Hub95, Row97, Ano95-134, Ano95-135, Ano95-158, Ano96-40, Dob95, Gre95a, Gre97k, Joc96e, Pal96, Way96h, Ano96-114]. Justice [Ano98w].

K6 [Pou98d, Hal96b]. Kaboom [Pot95]. Kai [Ano97-117]. Kaput [Lin95a]. Kayaks [Ano98-114]. Kbps [FH97, KH95, Ano96-45, FH97, Ude95d].

Keep [Ano95-97, Ano96-98, Nan96d, Nee96c, Tri95, Ano97-45, BGR95b, Joc96e, Ude97j, Van95a, Way95g]. keeps [Ano95-212]. kept [Ano95-169]. Kernel [Ano96-119, Var95]. Kerouac [Ano97-107]. Key [Ano95-65, Ano96-56, Abr98a, Ano95-162, Pou96g, Ren98, Won98b]. Keyboard [Joc95e]. Keys [Way96d, Sta95b]. Killer [Joc96d, Mel95, Row97, Ano96-78, Hal97b, Ano95-75]. kind [Hil95]. Kinder [Ano97-60]. kinds [Ano95-66]. king [Ano95i, Ano95j, Chi95]. Kit [Ano95-218, Mon95a, Api95a, Gre97o]. kits [Gil97c]. knife [Cor97]. know [Rei95c]. Knowing [TM95, Way96b]. knowledge [Ano95-82]. known [Het96g]. knows [Ano95-146, Ano98-57, Car95, Smi96c]. Knuth [Knu96a]. Kodak [Ano96-43, MW95]. Kpbs [Smi95a].

Lab [Ano95-117, Ano95-106, Ano95-99, Ano95-98, Ano95-104, Ano95-101, Ano95-100, Ano95-102, Ano95-103, Ano95-107, Ano95-105, HR97, KM96d, KM96c, LM95b, Mys96b, Mys96a, KHB96b, SK96b, Em98]. labs [Flo98b, Em98]. Laces [SM97]. lack [MFDG98, Sch96b]. Lame [MF97b, Kay97a]. Lan [Gio95, Ano95-52, Ano95-100, Ano95-119, Ano95-195, Bal95c, Bal95i, Fri96d, Fri97d, GM95a, HKM97b, Loh97, Mia95b, Nan95d, Nan97c, Row96b, Smi95d, VN97b]. Land [Ano95-109]. landscape [Hal98a]. LANE [Sal96f]. language [Esp95, Fri97b, Fri98b, Hal97b, Orc98, She97, Ano95-61, Ano95-129, Eva98, Flo96c]. languages [RFD95]. Lanier [Ano97-61]. LANs [Ano98-88, Bal95c, Bal95e, Nan95f, Sal95n, Sch96h, Sch96g, UGB95]. LANtastic [Mia95c]. Lap [Ano96-119]. LapLink [Nan95c]. laptops [Bal95h]. Large [Yav95, Ano95-98, BD95b, Fri96a, MC97].

Large-scale [Yav95, MC97]. Laser [Ano95-36, Ano97-62, HKM96, Hum96d, Ano95-49, Ano95-98, Ano95-101, Ano97-118, Hum96g, Sha96, Tho95b]. LaserJet [Ano98-126, Hum96j]. Lasers [Ano95-49, Ano98-124, Ano95-101]. Last [Way97a, Ano95-60]. Lasts [Hum96i]. Late [Ano95-122]. Latest [Hum96j, Ano95f, Ano95c, Ano95-184, Ano95-208, Ano96-178, Ano98-113, Bal95h, Bal96d, Cla95b, Gil97b, HKM96, LM95b, Nan95b, Tho98e, Ude95b, Van97]. Latin [IF97]. Latin-X [IF97]. launch [Ano97-29, Sil97]. launches [Ude95c]. Law [Ano97-53, Pou98a]. layer [DL95, Tho95j]. LC [Ano95-171]. LCD [Ano98-103, Chi95, Joc95e, Van95a]. LCD300 [Col96]. LDAP [Lew96]. Lead [Ano96-73, Ano98m, Ano96-91, Joc95d]. Leaders [Het96c, Yea95b]. leading [Ano95l, Bry95d, Hum96h]. leading-edge [Hum96h]. Leads [Lac96, Ano98-104, Gre95b, Gre95c, Wel97]. leap [Ano98-89, Mia97c]. Leaps [Flo98a]. Learn [Ano95-108, Ano95-175, Ano97-92, Del96b, Del96a, Rei95d, Ude97c]. Learned [Bee96, Ude96e]. Learning [Way95c, RFD95, Rei95d]. Learns [Chi96a]. lecture [Abr98b]. LEDs [Ano95-36]. Legacy [BMW97, Bal95c, Loh97]. Legal [Ano95-109]. legendary [Mia95b]. legibility [Ano95-106]. Less [Coo95, Tam95b, Tam95c, Ano95-63, Ano95-154, Ano95-206, Hum98b, Mon95a, Yav95]. less-scary [Mon95a]. Lesson [Ano96-145, Mon96b, Pou95v]. Lessons [Ano97-63, Bee96, Ude96e, Abr98c, Ude98c]. lest [Ano95-44]. Let [Gre95d, Ude96f, Ano95y, Ano95-132, Ano97-103, BHKM96a, Dav95, Daw95, EP97]. lets [Ano95z, Ano96-179, Ano98-82, Ano98-126, BMW97, Chi96a, Cot95c, MC97, Sal95j, Sta96c, Tur97, Ude96d, VN97b, Won98b].

Letter [Ano95-115, Bee96, Ega95, Fed97, IF97, MF97b, Mon96a, MH97b, Nav97, Rag97]. Letters [Ano95-113, Ano95-110, Ano95-114, Ano95-115, Ano95-117, Ano95-116, Ano95-111, Ano95-112, MH97a]. Level [Api97, Ano95-60, Esc95b, Nad97, Nan95g, Pou98a]. leverage [Ano97-121]. LexMark [Ano95-118]. liability [Ano95-222]. liberate [Ano95-220]. Liberty [Pou95q]. libraries [KHB96b, KHB96a, She97]. Library [Pou95h, Api95c, Tho97d, Pou95a, Pou95h, Ste95]. Libretto [Ano97-106]. licensing [Ano95-222, Sal95d]. Lies [Ude96a]. Life [Ano98-28, Sal96f, SK95, And95c, Ano95-98, Ano95-203, Ano98-81, Tho95k]. LifeBook [Way97g]. Light [Yav95, Way95b]. light-weight [Way95b]. lighting [Em98]. lights [BGR95b]. lightweight [Api95a, Bal95d, Pou95f, Lew96]. Like [Ano96-156, Ano98f, Ano95-125, Ano95-161, Dej95j, Hal95b, Pou95p, Row97, Ude97j, Esp95, Moh97b]. liked [Pou95p, Pou97g]. Limbo [Rau97]. limitations [Fri98b, Ude97k]. Limits [Tho96p, Hur97c]. Limur [Ano95-214]. Line [Ano95u, Ano96-113, Rei95b, Ude96g, Ano95q, Ano95-52, Ano95-82, Ano95-129, Ano95-136, Ano95-171, Ano96j, Ano96-77, Ano96-114, Ano97-54, Bar95b, Bea96, BHKM96a, Dob96, Fri96c, Cot95a, Hal96c, Het97, KFT95, Key95, Lin95a, Nee95m, Pen98, Pou95a, Sal95b, Sal95f, Sch96b, Ste98, Sto98, Whi95a, Ano95-136, Bon95a, Esp95, Pou95h, Pou97a, Whi95b]. line-of-business [Sto98]. Lines [Ano95-161, Ano96-62, Rud98, Sal96g, Ano95-45]. Lingo [Del96b, Del96a]. link [Ano95-142, Ano95-143, Sal95f, Ude95d, Way95a]. linked [Ano96-55]. linking [Way96b]. Links [Ano96-32, Ano96-103, Ano96-116, Ano95-61, Ano97l, Ess95, Loh97, Sal95a, Ude96d]. Linux [Ano97-64, Cox98a, Moh97a, Moh97b, Ude97j, Yag96a]. list [Ano95-192, Ano97-114, Gre96g]. Listen [Bar95b]. Listens [Nan96g]. Little [Ano95-190, Pou96k, Yag95b, Ano95-122, Ano96-175, Ano98-100, Col96, Smi95b, Way97f, Won98c]. little-used [Ano98-100]. Livable [Ano97-64]. Live [Ude95d, Ano98m, BVN95, Ske98a, Ude96d, Mar95d]. Lively [Mar95d, Mar95c]. Lives [Gre97i, Abr97b]. Living [Sch97d]. LN [Tra95]. Loaded [Mia97b, Nan95a]. loads [BEM95]. Local [Ano95x, Bal95e, Ude98d, VN97b]. localization [Fow97]. locally [Ano95-210]. locate [KP98]. Location [KP98]. locations [Pen98, Sal95j]. Locks [Way95g]. Logging [Fre96]. Logic [Pou95b, Ano95-74, Gou95, WS96, Arc98]. Logitech [MW95]. Long [Laz97a, Ano95-98, Cra96, Tho95k, Ude97d]. Long-awaited [Laz97a]. long-battery-life [Tho95k]. long-term [Ude97d]. longer [MA98]. Look [Ano96-103, Bar95b, Tho97a, Tho98c, Aer95, Ano95e, Ano95r, Ano95s, Ano95-158, Ano96-164, Ano96-170, Ano96-171, Ano96-172, Ano97-48, Ano97-67, Ano98-63, Api95a, BD95b, Bry95e, Dob95, Dob98d, Emi97, Fri95a, Fri96c, Fri97c, Gil97c, Gre98a, Hal95b, Hal95d, Hal97d, Het96d, MB97, Nan95f, OHE95, Pou96l, Row97, Sha95c, SK95, Tho96g, UT95]. Looking [Ano95-103, Ano96-114, Flo95]. Looks [Ano95-69, Ano95-64, Ano96-174, Flo98b, Pou97i]. lookup [Hal95e]. lorme [Ano96-40]. Lose [Way96d]. Losing [Ano95-96, Ano96-182]. Loss [Ano95-100, GM95a]. Lost [Nad95c]. lot [Ano95-38, Ano95-199, Gre95a, Won98b]. lots [And95c, Ano96-177, Ano98-115]. Lotus [Ano95-64, Ano95-142, Ano95-143, Ano95-201, Ano96-173, Ano96-175, Ano98-93, Cra96, Dej95e, Gil96f, Gil98b, Het96b, Yea95b]. Loud [Mon95b]. Low [Ano98-56, Row96b, Ano95-60, Ano95-119, Ano97z, Ano97-120, Ano98n, Bal95d, Ban97, Chi97, Cir95, Flo96a, Goo95b, HKM97a, Hum96e, Hum96g, MA98, Sha96, Tho95k, Tho95m, Tho97a, Ano95-119]. Low-Cost [Ano98-56, Row96b, Ano97-120, Ano98n, Ban97, Chi97, Flo96a, Goo95b, Hum96g, Sha96, Tho95k, Tho95m]. low-end [Ano95-60, HKM97a, Ano95-119]. low-power [MA98, Tho95k]. low-voltage [Bal95d]. lowcost [Way96m]. Lower [Ano96-91, Ano95-122, Ano98-47]. LSI [Arc98]. LTX [Api95b]. Lumiere [Ano97-120]. lunch [Ano95-82]. lurk [Ano95-109]. LynxOS [SB96].

M204 [Smi95b]. Mac [Ano95-120, Ano95e, Ano95r, Ano95s, Ano95y, Ano95z, Ano95-27, Ano95-56, Ano95-120, Ano95-165, Ano95-171, Ano95-213, Ano95-214, Ano97-40, Ano97-80, Ano97-106, Ano98-74, Ano98-115, Ano98-114, Cox98a, KFT95, Kay96a, Mar95d, Mar95c, Mor97, Pou95f, Tho95c, Tho95d, Tho95e, Tho95i, Tho95j, Tho95l, Tho95m, Tho96j, Tho97d, Tho98c, Tho98d, Way97d]. Mac-based [Way97d]. Mac/Windows [Ano98-74]. Machine [Ano95-147, Way95c, Ano95-162, Ano96-40, Ano97-110, Ano97-117, Ano98-51, Ano98-115, Bal95h, Chi96a, Esc95b, Pou96k, Pou97k, Pra95b, Row96c, Sim96]. machine-vision [Ano95-162]. Machines [Ano95-187, Les95, Mar95a, Ano95-158, Ano97-82, Fen95, LM95b, Row97, Tho97d]. Macintosh [Ano95-107, Ano95-185, Ano96-125, Ano96-173, Api95c, Cot95b, Ric97, TC97, Tho95g, Tho95j, Tho95l, Tho96e]. Macromedia [Ano96-172]. Macs [Ano95q, Ano95-94, Ano95-132, Tho98a, Van97]. MacWorld [Ano95q]. Made [Ano95-200, BVN95, Mon95b, PM97, Rei95f]. maelstrom [Bau97]. Magic [Ano95-40, Tra96a, Ude95f, Yag97a, Way95a]. Mail [Ano95-39, Ano97-127, Ano98-60, Ano98-109, Cro95, SY97, Ude98e, Ano96-163, Ano97-53, Ano97-97, Ano98-41, BHKM96a, Kos96, Sta98, Ude96h, VN96a, Ano96-173, KFT95, Kos96, Ude96c, VN97a, Way96e, Hig95d, Vau96, Way95a]. mailers [Abr97a]. Mainframe [Rei95f, Ano95-151, Ano95-152, BMW97]. Mainframes [Sal96c, Ano95-190, Nan95f, Rei95f]. Mainstream [Fri98a, Ano95-162, Ano95-169, Ano97-120]. Mainstreaming [Chi96b]. Maintain [BMW97]. Maintaining [DeJ96b]. Maintenance [Phi97, DeJ96c, Gil97c]. Mais [Ano96y]. major [Ano95-164, Die95c, Dob98d, Gio95, KFT95, She95c]. Make [Ano98-47, Ano98-48, Cam97b, Dob96, Hyd96, Joc95d, Lin95d, Lin96c, Pou97j, Sal96h, Tho97b, Ano95-103, Ano95-129, Ano95-164, Ano97-30, Ano97-75, BVN95, Cka97, Cua95b, Die95b, DL95, Goo95b, Hal98g, Hig95c, Hum96h, Mon95a, Pou95b, SY97, Tho95d, Ude97f, Ude97h, VN97a, Vog95, Way96l, Whi97, Yag95c]. Makeover [Tho97e]. maker [Rei95e]. Makers [Ano96-149]. Makes [Ano98-126, CR96a, Fri97a, Yag98b, Ano95-145, Ano96-40, Ano96-179, Ano96-175, Ano98-68, Api95b, Bal95d, Car95, CR96b, Flo95, Gil96d, Mia97c, Nel95, Tam95c]. Making [Ano95-43, Ano95-121, Ano97-64, Gre97a, KE97, Ren98, Vac95b, Way98b, Ano95e, Ano95-52, Ano97-31, Ano97-112, Joc97b, Ude97i]. Manage [Ano97-65, Ano98-126, BMW97, Cro95, Goo95a, Sug95, Wat97b, Wil96, Won98b, Yag96d]. Manageability [Ano98h]. Management [Ano95-130, Ano98p, Hig95d, Mac98a, MC98, Sal95j, Sal96j, Sch96h, Ude96i, Ano95-52, Ano95-106, Ano95-151, Ano95-152, Ano96-178, Ano97-123, BD95b, Eva98, Gio96a, Gre98b, Kay96a, Nan95f, Sal95b, Sch96g, SY97, Wel97, Wil95, Yav95, Hig95a]. manager [Ano95-205, Ano96-172]. Managers [Nan95b, Sea95c, Ano95-182, Hig95d, Sal95j, Sea95a, Vog95, Win95]. Manages [Cro95, Ano95-203, Pou95n]. Managing [Ano98-61, Boy97, Kay96a, MF97d, Pen98, Sta95b, Whi97]. manipulation [Bla96, Gar97]. Manor [Pou95j, Pou95v, Pou97e, Pou97f, Pou97k, Pou97l]. manual [UT95]. Manufacturing [Dej95h, Esc95b, Gou95]. Many [Ano95-147, Ano95-154, Cua95b, Fri95a, Gre97f, Ude97k, Win95]. Map [Ano96-137]. Map’n’Go [Ano96-40]. maps [Gre98f]. marathon [Tho95f, Ano95e]. Marilyn [BM95]. Marimba’s [Los97c]. Mario [Sim96]. market [Ano95-105, Ano95d, Goo95b, Sal95h, Sta97b, Tho95j, Tho95k]. Markov [Mar95a]. Markup [Ano95-61, Ano95-129, Flo96c]. Marries [Gre98b]. Marrying [FS96, Cre96]. Mart [Ano95-39, Wat97a]. Masses [Ano96-94, Ano97-28, Way96a]. Masterpieces [Hal95b]. Masters [Gre97n]. match [PM97]. Matchbook [Ano95b]. matches [Bal96d]. Matching [Kay95d, Fri97b]. Mathematica [Tam95a]. mathematical [Gre96i]. mathematics [Les95]. matrix [Ano95-104, Van95a]. Matrox [Ano96-171, Ano97-33]. Matsushita [Ano95-169]. Matter [Ano97-124]. Matters [Gwe95, Fow97, Yag96a]. Matures [Joc97b]. Maturing [Way98a]. maximize [Cla95c]. Maximum [Mia97a, Ano95-193]. may [Ano95-71, Ano95-161, Ano95-174, Fow97, Hal96g, Ind95, Pou95b, Pou98a, Pra95b, Sal95f, Way95d]. Maybe [Hal97b, Ano95-125, Ano95-146]. MB [Ano95a, Ano95b, Ano95h, Pal96]. MBone [Ano96f]. Mbps [Ano95-193, Cun95]. McNealy [Ano97-58]. MCore [MA98]. MD [Hum96i]. MD-4000 [Hum96i]. Me [Ano98-50, Hal97e, Ude96h, Way96f, Way97g]. Mean [Ano95-101, Ano95-145, Bar95b, Dob97a, GM95b, Hal97d, Hal97e]. means [Ano97-122, CC97, Cir95, Fen95, Gre97q]. measure [Ano95-106, Way95d]. mechanism [Cli98c, MF97c]. mechanisms [OH95]. Media [Gre96f, Ano95-128, Gre96a, Mia97b, Sea96b, Ude95c]. MediaForge [Die96a]. medical [Ano95-197]. Medicine [Cor97]. medium [Ano95-100, Ano95-169, Die96a, GM95a]. Meet [Ano96-138, Rud98]. meeting [Pou95q]. Meets [Ano95-149, Ano95-142, Ano95-143, Ano98-64, Pou96d, Tam95a, Ano95-46]. mega [Kri95, Wil96, Ano95-173]. mega-size [Wil96]. mega-storage [Kri95]. Memories [Ano98-62]. Memory [Ano95-69, Ano95-122, Ano98-104, Coo95, Kay95c, Ano95-66, Ano95b, Ano95-125, Goo95b, Laz97b, MC97, Pal96, Tho95f]. Memphis [Ano97-125]. men [Ano95-62, Way95h]. Menace [CC98]. Mentions [Ano95-78, Ano95-79, Ano95-106, Ano95-104, Ano95-101, Ano95-100, Ano95-103, Ano95-107, Ano95-105, Ano96-65, LM95b]. Mentions/Dubious [Ano95-79]. menus [Esp95]. Merced [Kra98, Nan98d]. merger [Ano95-75]. Merges [Ano96-67]. Merging [Sal95g, Sal96i]. MES [Esc95a, Esc95b]. Mesa [Gas95]. Message [Ano98-42, Gil96d, Gre96f, Gre96a]. MessagePad [Ano95-114, Rei95e, Tho95f, Tho97e]. Messaging [Ano95-212, Cli98b]. MetaCreation [Ano97-119]. Metadata [Gar97]. metal [VN97b]. metaphor [Lew96, Mon95b]. Meter [Ano95-222]. metering [Ano95-222]. method [Cli98c]. methods [Pou96a, Pou98c]. Metrowerks [Cot97, Gre97p]. MFC [Ano95-184]. MHz [Ano95-208, Ano95-104, Ano95-158, Ano95c, Ano95-205, Ano96-100, Ano96-178, Ano96-161, Ano97e, Ano97q, Ano97c, Ano97b, Ano97-106, Ano97-115, Ano98a, Ano98-77, Ano98-83, Bal95a, Bal95g, Cam98a, Chi96c, Hal96g, Hal96h, LM95b, RV96, Tan98, Tho95g, Van97, Yag98a]. Micro [Ano95-46, Ano95-45, Ano95-214, Ano96-170, Gre97r]. Micro-Help [Ano96-170]. microkernel [Krt98]. microkernel-based [Krt98]. Micromirrors [Ano96m]. Microplex [Smi95b]. Micropolis [Api95b]. Microprocessor [Goo95b, Gwe96, Hal97d]. microprocessors [Ano95-127]. Microsoft [Ano95-37, Ano95-64, Ano95-73, Ano95-92, Ano95-115, Ano95-112, Ano95-123, Ano95-166, Ano95-167, Ano95-173, Ano95-219, Ano96-134, Ano96-182, Ano96-171, Ano97-66, Ano98w, Ano98-63, Ano98-74, Ano98-108, Api97, Bal96d, BM95, Cka97, Cli98a, Dob95, Dob96, Dob97c, Cot95b, Gil95, Gil96g, Gil98a, Gre97q, Hal95d, Het96b, Hur97a, Joh96, Lin96a, Mia95d, Mon95a, MC98, Nan97a, Pou95r, SK96b, SK96a, SM97, Wil95, Yea95b]. Microsoft/Justice [Ano98w]. Microsystems [Ano97-58]. Middle [Nan97b, Rym96, Sal96j]. Middleware [Ano98-64, Dej96e, Kad96, Ano97-28, Ano97-92, Nan97b]. midrange [Ano95-100, GM95a, HR97]. might [And96d, Ano95-83, Ano95c, Ano95-187, Ano95-192, Bal95c, MC98]. Mighty [Hum97a, Hum97d]. Migrate [Lin95b, Lin95c]. migrating [Ano95-119]. Migration [Ano95-63, Ano95-218, Het96c, Gre95e]. milestone [Mia97b]. Millenia [Ano97-107]. Millennium [Ano96-123, Ano97-33, Ano98-93, Pri95]. MIME [Ano96-120, Sta98, VN97a]. minds [MDM95]. mine [Ude95c, BM95]. minefields [Kay95a]. Miner [Wat95]. Mines [Ano95-109]. minicomputers [Nan95f]. Mining [Ano97-63, Cua95b, Dej95d, Kri95, Ano98-119, Hed95]. minted [DM98]. Mips [Api95c, Hal96f, Ano95-124, Arc98, AG95, Hal98e]. Mirror [Law96]. miscalculation [Hal95e]. misses [Pou97i]. missing [Tho98c]. Mission [Fom98, CCS97]. Mission-Critical [Fom98, CCS97]. Mistakes [Ano97-58]. Mite [Hum97a, Hum97d]. Mitsubishi [Cam97b]. Mix [Lev98, Ano95-157]. mixed [Ano95-128, Sea96b, Yag96d]. mixed-media [Ano95-128, Sea96b]. mixing [Mia95d]. MkLinux [Cox98a]. mm [Ano95-100]. MMX [Ano96u, Ano96-76, Ano97m, Ano97-59, Ano97-67, Ano98-65, Hal97c, KB96, Pou98d]. MO [Fro96]. Mobile [Ano96-97, Ano97-110, AK98, Ano95-97, Ano95-110, Bas95, Gro96, Nad95a]. MobileAccess [Cam97b]. Mobility [Rei95c]. mod [Ano98-79]. mod-perl [Ano98-79]. Mode [Ude96b, Gre95a]. Model [Cka97, Knu96b, UT95, Ano95-61, Pou96b, Cli98a]. Modeling [Knu96b, Yar95, Ano95-186, Cre96]. Models [Mar95a, Pou95n, Ano97s, BGR95b, Hig95d]. Modem [Ano98-107, Ano95-105, Ano95-161, Esp95, FH97, Smi95a, Way95a]. Modems [Ano95-105, Ano95-161, Ano96-45, Ano97x, Ano98m, FH97, KH95, Pou95l, Sal95f]. Moderate [Ano95-159]. modern [Ano97-107, Ano98-81, Tho98c]. modes [MA98]. modified [Fri96a]. modular [Api95b]. modularity [Ano96-91]. MOLAP [Ano96-87]. molecules [DeJ98]. Mon [She95c]. Money [Ano95-215, Ano95-222, Flo96b, Gre97a, Ano95-38, Ano96-76, Hur97e, Sin95a, TM95]. monitor [Col96, Hil95]. monitoring [Sal95d, Sal95j]. Monitors [Ano95-106, Ano98-103, GE95, Mys96b, Ano95-204, Ano95-209, Ano97-45, Daw95, Hil95, HKM97b, Len95]. Monroe [BM95]. Monster [Ano98-43, Ano98-51]. month [Ano95-117, Ano95-101, Ano96-164, Ano97-75, Pou95q, Pou97f, Tho95c]. monthly [Ude95c]. months [Ano95-50]. Moody [Cla95a, Cla95c]. Moore [Ano97-53, Pou98a]. Morphing [Way97d].

Mosaic [Smi95a, Vac95a]. Moss [Gre97r]. Most [Tho98d, Ano97-29, Ano98-81, Ess95, Hal98g, Pou95b, Sha95c, Tho95a, Wat97a]. Motherboards [Ano95v, LaP96c]. motion [Ano95-204, Ano95-209, Bal95h, Way95f]. motion-video [Bal95h]. Motorola [Ano95-145, Ano97-106, MA98, Shi98, Tho95o, Tho96f]. mountains [Pou95n]. mouse [Abr98c]. Move [Ano95i, Ano95j, Sch97a, Tho96j, Sal95h, Sin95a, Ude97h, Wil95]. Moves [Ano96-117, Ano97-120, Ano98-100, Bal95h, Mia95d, Ske98b]. movie [Hal95b]. movie-like [Hal95b]. Moving [Ano95-97, Ano97-35, Ano98-98, Dej95j, Fom98, Tri95, Lac96]. MPEG [Ano95-198, Ano95i, Ano95j, Ano96-92, Chi97, KE97]. MPEG-2 [KE97]. MR [Ano98-113]. Mrs [Pou96n]. MS [Chi96b]. MS-T [Chi96b]. MSN [Ano95-126]. Much [Het96g, Ano95-47, Ano95-134, Ano95-135, Ano95-169, Rei95d, Sea95c, Sta96b]. much-improved [Rei95d]. much-needed [Sta96b]. Muddle [Rym96, Nan97b]. Multicast [Ano98-66]. multidimensional [Dej95f, Dej95e, Wat95]. multifaceted [Pal96]. Multifunction [BHKM96a, Yag95c, Ano98-115]. Multimedia [Ano95q, Ano95-50, Ano95-109, Ano95-127, Ano95-128, Ano96-93, Ano96-99, Ano96-95, Bal95f, Bal95h, BKM96, DD97, Hal95b, Hal96j, Mul96c, Sal96h, Sea96b, Ano95-80, Ano95-205, Cro95, Del96a, Die96a, Die97b, Rei95d, Row96c, Sal96b]. multimediachips [HM95]. multimegabit [Mon97]. Multiplatform [Api95c, Cot97, Api95b]. multiple [Mac98b, Nan97b, Pou95d, Ren98, Wil95]. Multipoint [Ano97-68]. multiprocessing [Ano97-121, Pra95a, Tho97d]. Multiprocessor [Hum97b, Tho97d, Cro95, Pou97k]. multiprotocol [Smi95b]. multitasking [Tho97d]. multithreaded [Pra95b, Way95b]. multithreading [Pra95a]. Multitools [Sea96b]. MultiWin [Ano98-67]. Muscle [KM96c, KM96a]. Music [Ano95-94]. Must [Tho96a, Tho96r, Ano95-166, Ano95-167, Bal95c, Joc95e, Van95a]. Must-See [Tho96r, Tho96a]. Mustang [Esp95]. muster [Sha95c]. Mutation [Ano95-163, Pou95k]. Mwave [Bal95h]. My [Way96f]. Mystique [Ano96-171].

Naaaaah [Ano95-146]. nab [Way95h].

name [Bri97a, Gre97g, She95c].

name/address [Bri97a]. Names [Gre97i]. nanomechanical [Flo98b]. nasty [Hum96h]. Native [Gre95a, Bak95b]. Natrificial [Ano98-128]. Natural [Ano97-69, Hum97c]. Naturally [Ano97-46]. nature [Ano97-121, MDM95]. Navigating [Bal96e, Ano95-129]. navigator [Ano96-125, Ano97-107, Vac95a, Ano96-95, Bal96d, Bal96f]. NC [Hur97a, Way96h]. NDS [Ano95-63, Ano98-73]. near [Ano95-70, Flo96a]. nearing [Hur97c]. Nearly [Ano97-27, Ske98b]. NEC [Ano97-120, Col96, Hum96g, Hum97c, Yag98a]. necessary [Ske98a]. Need [Ano96-162, Yag96d, Ano95y, Ano95-95, Ano97-32, Cox98b, Hil95, Joc98a, Kay97c, Laz96, Rei95c, Sal95f, VN97b]. Needed [Sha95a, Sta96b]. Needle [Kor95, Boy97]. Needs [Ano96y, Jen95, Ude96m, Ano97-103, Ano97-122, Nad97, Pou97f, Tho98c]. Negroponte [RFD95]. neighbor [Tho95d]. NEST [Sal95h]. Net [Ano95-83, Ano95-129, Ano96-109, Ano96-94, Ano96-149, Ano97-34, Cro95, Ude97e, Vac95a, VN96b, Way96a, EG96b, Ano97d, Ano98-37, Ano98-60, Cli98b, Kay95a, Yag96b]. Netbot [Eng97]. Netcards [Ano97-95]. NetManage [Mon95b]. NetPCs [Hur97a]. Netra [Smi95d]. Netscape [Ano95-75, Ano96-95, Bal96d, Los97b, Nan97a, Smi95a, TC97, Ude97e]. NetWare [Ano95-63, Ano95-123, Ano96-96, Ano98-68, Ano98-120, Gil96a, Gil96c, Gil97b, Gio95, Law95, Law96, Mia95c, Nan95d, Sal95h, Ude95e]. NetWare-to-NT [Ano95-123]. Network [Ano95-40, Ano95-98, Ano95-130, Ano96-97, Ano97l, Ano98f, Cam97a, DM98, Fin98, Gio96d, Hus95, Knu96b, Mul96c, Nan95f, Nan97a, Pou96i, Ren98, Smi96a, Tho96k, TS96b, Ude95c, Win95, Ano95-62, Ano95-151, Ano95-152, Ano95-184, Ano95-182, Ano95-195, Ano95-220, Ano96-68, Ano97g, Ano97t, Ano97-53, Ano97-54, Ano98-68, BGR95b, BEM95, Bal95i, Cot95c, Cro95, Daw95, Fri97c, Fri97e, Gio95, Gio96a, HKM96, Hum96j, KHB96b, KHB96a, KP98, Kor95, Law95, Mia95c, MC97, Mul96b, Pou97d, Rei95f, Sal95g, Ske98a, Smi95b, Sta97a, TS96a, VN96b, Wat97b, Way95d, Way95f, Way96h, Wil95, Yag97a, Way95d]. network-based [BEM95]. network-centric [MC97]. network-distributable [Ano95-184]. Network-Printer-Management [Ano95-130]. Network-Ready [Way95d]. Networked [Ano97-70, Gre97n, Ano95z, Ano95-98, Sal95h]. Networking [Ano95i, Ano95j, Loh97, Nan95e, Cha97, Pou95l, Pou95v, Rei95d, Ude95c]. Networks [Ano95-38, Ano98-69, Bal95e, Cli98c, Daw95, DeJ96c, Gio95, Nan96d, Sta97a, Sta98, Ano95-101, Ano95-100, Bry95a, Bry95c, Cun95, DM98, Fri96d, GM95a, Gou95, Hur97e, Loh97, Sal95k, Wil96]. Neural [Gou95, Smi96a]. Neutrino [Krt98]. Never [Ano97-71, Lin95a]. Never-Ending [Ano97-71]. New.Domain.Names [Ano97-74]. newer [Hig95d]. newest [Ano95-206, Ano96-40, Tho97e]. newly [DM98]. newness [Sch97b]. News [Ano97j, Ude98e, Fen95, Ude96b]. news-style [Ude96b]. Newton [Rei95e, Ano95-114, Rei95a, Rei95e, Tho96l, Tho97e]. NexGen [PH95]. Next [And97, Ano95-129, Ano95-136, Ano96-107, Ano97i, Ano98-71, Ano98-72, Cka97, Cli98b, Kha96, Pou96g, Rei95f, Abr98b, Ano95-48, Ano95-93, Ano95-124, Ano95-122, Ano95-128, Ano95h, Ano97-125, Flo95, HM95, Hal96d, Hal96k, Hal98a, Hum96c, Joc95b, Pou95b, PH95, Tho95j, Tho96f, Tho96g, Tho96p, Ber95, Ude96h]. next-generation [Hal96d, PH95]. NFS [Fri96a]. nice [Way97d]. Nicholas [RFD95]. nifty [Kay97b]. Night [Pou96j, Tho97b]. Niklaus [Pou95f]. Nine [Way98a, Won98a, Ano97-103, PH95]. Nine’s [Ano97-77]. Nineties [Ano98-68]. Ninety [Ano98t]. Ninety-Eight [Ano98t]. Nintendo [Sim96]. nitro [Gwe96]. nitro-burning [Gwe96]. Nixdorf [WS96]. No [Ano96-162, Ano98w, Dej95j, Joc98b, MB97, Nee96a, BVN95, Bou95, Kay95d, LM95a, Pou95j, Tam95a, Ano95-104, Way96h]. no-compromise [LM95a, Ano95-104]. noises [Car95]. nontechnical [Vog95]. North [Bal95e]. Not-So-Crazy [Ano95-217]. Notebook [Ano95-139, Ano95-140, Ano96-57, Ano98g, Ano98-76, Hum97c, Row96c, Tho98d, And95a, Ano95-77, Ano95-204, Ano95-209, Ano95-214, Ano95-208, Ano96-91, Ano96-175, Ano97-106, Ano98-96, Ano98-113]. Notebooks [Ano95-104, Ano95-141, Ano96g, Ano96-108, Ano96-136, Ano97-109, KM96c, Tho95k, Ano95-140, Ano97-115, Ano98-114, Dob96, Hum98a, KM96a, LaP96b, LM95a, Yag98a]. Notes [Ano95-142, Ano95-143, Ano98-70, Ano95-75, Ano95-142, Ano95-143, Ano95-133, Ano95-201, Ano98-45, DA95, Gil96e, Gil96b, Gil96f, Het96b, Mac95b, Mar95b]. Notes-killer [Ano95-75]. nothing [Ano97-60, HM95]. noticeable [Ano95-118]. novel [BM95]. Novell [Ano98-120, Ano98-73, Gil96d, Gil97b, Sal95h, Ude95e]. Novera [Way97c]. NSTL [Ano95-106, Ano95-100, Ano95-103, Ano95-105, Ano95-144, BHKM96a, Gio95, Gre95a, Sea95c, Sea95d, Sea96b]. NT [Ano95-100, Ano95-123, Ano95-141, Ano95-145, Ano96u, Ano96-67, Ano96-77, Ano96-110, Ano96-109, Ano96-103, Ano96-111, Ano96-127, Ano96-134, Ano96-155, Ano96-154, Ano96-159, Ano97-31, Ano97-53, Ano97-76, Ano97-75, Ano97-82, Ano97-121, Ano98-73, Ano98-114, AG95, Bak95a, Bak95b, Cam98a, Cha97, CR96a, Chi96c, CR96b, EP97, Em98, Fen95, Fre96, Gar98, Gio95, Gre96g, Hum96h, Hum98c, Lin96a, Min96, MC98, MB97, Nan95d, Nan96b, Pou96l, Pou97k, Pra95b, Pra95a, Rey96, Row97, SM97, Tan98, UGB95, Ude97f, Way96j, Won98b, Yag96b]. NT-based [Row97]. Nth [Ano95-55, Ano95-56, Dej95e]. Nu [Ano95-173]. Nu-Mega [Ano95-173]. Nuance [Ano98-57]. NuBus [Ano95-107]. nuclear [And96d]. NUI [Hal97a]. number [And95b, Ano97-77]. Numbers [Die95d, Way95h]. NuMega [Ano98-78]. numerous [Ano95-174, Ano95-201].

O [Ano95-159, Ano98-97, Bri97b, Tho97c, Way95d]. Oberon [Pou95f]. Oberon/F [Pou95f]. Object [Ano96-145, Ano98-84, Bie95, Cka97, Cli98a, Cli98c, Gre98b, Ano95-61, Ano95-72, Ano95-184, Ano95-186, Ber95, Car95, Cot95c, Dej95j, Eva98, Gil95, Gre98f, Mon96b, Nan95g, Pou95f, She97, Sto98, Tho98e, Ano95-46, Ano97-113, Eva98, OHE97]. Object-Oriented [Cli98c, Ano95-61, Ano95-72, Ano95-184, Ano95-186, Ber95, Cli98a, Cot95c, Dej95j, Gre98b, Nan95g, Pou95f, Tho98e]. object-relational [Gre98f, Sto98]. ObjectGeode [Cre96]. ObjecTime [Ano95-186, Bie95]. Objectivity [Gre98c]. Objectivity/DB [Gre98c]. Objects [Ano96-112, Ano97-65, Ber95, Gre95e, OH95, She97, Ano95g, Cka97, Dob97a, Dob97d, Dob98a, Ess95, Gil95, Gre98f]. ObjectStore [Ano97-113]. obsolescence [Hal98a]. Obsolete [Ano97-127, Way96l]. obstacles [Sal95a]. OCR [Ano95-146, Ano95-204, Ano95-209, Ano97-119]. ODBC [Ano95-184, Dob96, Dob97c]. ODBCDirect [Dob97c]. Off [Ano95-64, Ano95-77, Cor96, Law95, Nee96c, Sar98, Yag98a, Ano95-195, Ano96-76, Ano97-53, HR97, Pou98a]. Off-the-Cuff [Sar98]. Offer [Ano96-103, Ano95-151, Ano95-152, Ano97-105, Cli98d, Het97, Hur97b, Smi95a, Ude95i]. offering [Rei95d]. offerings [Ano95-64]. Offers [Ano98-66, Fri97b, Ano95-47, Ano95-195, Ano95-204, Ano95-209, Ano95-214, Ano97-75, Ano97-121, Ano98-84, Arc98, Eva98, Fri95a, Mia97a, Mon95b, Pou96b, Pou96i, Pou97h, Tho95l, Tho97a, Tho98c, Tho98e, Way97b]. Office [And97, Ano98-36, BN95, Gil96g, Sal95k, Sal95l, SK96a, Ano95-146, Fri96c, Pou97a, Ano95-157, Gil96g, Mia95d, SK96a]. OfficeMac [Ano98-74]. officers [Mar95a]. offices [Nad95c]. official [Way96h]. offs [Gre97f]. often [Ano95-95, Hal95e, Sea95d]. Oh [Fra96]. OLAP [Het97, Pen98]. old [Ano95-46, Ano98-64, Car95, Sch96e, Smi96c, Wat98b]. OLE [Ano95-45, Ano95-157, Ano95-184, Ano96-115, Api95a, BD95b, Gil95, Mia95d, Mon95b, Van96d]. Olympic [Sal95i]. Olympic-Size [Sal95i]. Omnipresent [Pen98]. on- [BHKM96a]. on-chip [Phi97]. On-Line [Ano96-113, Ude96g, Ano95-136, Ano96j, Ano96-114, Pen98, Ano95-82, Ano95-129, Ano96-77, Ano97-54, Bar95b, Bea96, Dob96, Hal96c, Het97, KFT95, Nee95m, Pou95a, Sch96b, Whi95a, Ano95-136, Bon95a, Esp95, Pou95h, Whi95b]. On-Ramps [Ano96-33]. On-Road [Whi95a, Whi95b]. On-Screen [Hil95]. on-time [Whi95a, Whi95b]. Once [Ano95-221]. Oncoming [Yav95]. One [Ano95p, Ano95-147, Ano95-148, Ano95-217, Pou97k, PF97, Rei95b, Sal95e, Sal95f, Tho95l, Tho96o, Van96b, Ano95-47, Ano95-125, Ano95-146, Ano95c, Ano96-178, Ano97h, Ano97-29, Api95a, BHKM96a, Gil97c, Gio97, Gre97j, Hig95c, Pou95n, Sch96b, Sea97, SP97, Tho97e, Wat97b]. one-step [Ano97h]. ones [Bal95e, Sal95i, Tho95a]. Onions [Pou95p, Pou97i]. Only [Coy98, Ano95-27, Hum96c, Nan95f, Sal95i]. onto [Pou97c, Smi95d]. OO [Ano95-149, Cre96, Pou96a]. OO-modeling [Cre96]. OODBMS [Gre97m, Gre98c]. OOP [Dej95j, She97, Var96]. Oops [Dej95j]. Open [Fen95, Way96j, Ano97-31, Hal95d, Hub95, Moh97a, Sta96c]. OpenDoc [Ano96-115, Cot96c, Cot96a, Tho96e, Van96d, Way96j]. Opening [Sal96m, Sch97a]. OpenLinux [Ano97-64]. OpenMind [BD95b]. Opens [Gil96f, AG95]. OpenVMS [Fen95]. Opera [Pou96j]. operate [Bau95]. operates [Ano95-37]. Operating [Ano95-150, Bal96b, Fri98a, Moh98, Ano95-132, Gio95, Nan95a, Van96d, Way95b]. operation [Gou95]. operational [Cro95]. Operators [Ano95-179]. opinions [Pou97h]. opportunities [Ano97-87, Ude95c]. Optical [Ano95-169, Ano96-32, Flo98b, Hal96d]. Optima [Gre96d, Gre97d]. optimization [Ano98-104]. Optimize [Cua95a, Cua95c, Gre95f, TJF95]. optimized [Ess95, Smi95a]. Optimizing [Cua95a]. optional [Esp95]. Options [Ano95-133, MG97, AG95, Kos96, Sta97b, TM95, Way97b, Win95]. OQL [Eva98]. Oracle [Ano95-133, Ano95-151, Ano95-152, Het97, Nan95b, Way95f]. Oracle8 [Ano97-91]. Orchids [Pou95p, Pou97i]. ORDBMS [Sto98]. Order [Ano95-39, Ano96-79, Ano98-53, Ano98-54, Cob95, Smi95c, Smi95d]. organisations [Sal95d]. Organization [Mon95b]. organizational [Won98c]. organizationally [Mon95b].

Organizations [Ano95-45, Fri95a].

Organizes [Flo95, Ano95-214]. Organizing [Har96]. orientation [Car95]. Oriented [Cli98c, Ano95-61, Ano95-72, Ano95-184, Ano95-186, Ber95, Cli98a, Cli98d, Cot95c, Cro95, Dej95j, Gre98b, Nan95g, Pou95f, Tho98e]. original [CFDT95, Van97]. originally [Ano95-222]. OS-compatible [Ano95-120]. OS/2 [Ano95-111, Ano95-112, Ano96-74, Gas95, Gio95, Nan95a, Nan96g, Tam95a, Wes96b, Wes96a]. OSes [Ano97-54, Gre96g, Lac96, MB97, Pal96]. OT [WS96]. Other [MB97, Nan95g, Ano96-34, Gre96h, Pen98, Sha95c, Tho95d]. Our [Ano95-103, Ano95-107, Ano95-145, Ano98-121, BHKM96b, Bry95d, Cra96, Het96b, Ude95f, Ude97h, Ano95-42, Ano95-117, Ano95-111, Ano95-112, Ano96-77, Ano96-127, Ano96-163, Ano97t, Ano97-54, Ano97-97, Gil95, Gwe95, Hac97, Ind95]. out-of-oder [Pou95d]. out-of-the-box [Laz97b]. Outernet [Ano97-56]. outline [Mon95b]. Outlook [Ano96-171]. Output [Wei95]. Outstanding [Ano95-118, Bar95c]. Ovation [van95b, Van95a]. OverDrive [Ano95-159, Ano95-206, Pou97l]. Overheads [Ano96i]. overlap [Ano96-87]. Overload [RFD95, Ano95-114]. Own [Fri95a, Ano96-179, Ano97-97, Bal96d, Bri97a, Gla95, Gre97l]. Ownership [Ano95-52]. Oxygen [Hum97a].

p [Esc95b, Joc95d, LM95b, Mar95a, Tri95, Way95h]. P1000 [Bal95a]. P6 [Ano95-42, Ano95-93, Ano95-115, Ano95-154, Ano95-155, Hal95a, Hal96b]. P75 [Ano96-170]. PA [Hum98a, Pou95d]. PA-8000 [Pou95d]. PA-RISC [Hum98a]. Pace [Ano95-98]. Pack [Ano95-59, Nan95a]. Package [Sal95e, Sal95f, Ano95-184, Api95c, Tho97e, Way97g]. packages [Ano96-174, Hig95c, Sea97]. Packard [Ano95-186, Ano95-208, Ano97-114, Hal97d, Hum96j, Son95, Yag98a]. Packing [Coo95].

Packs [Way96k]. page [Dob97d]. PageCard [Ano95-212]. PageMaker [Ano97-116]. pager [Way95a]. pagers [Ano95b]. pages [Ano96-179, Ano97-97, Dob96, Dob98a, Los97c, Ude98c]. pain [Knu96b, Mia95f, Pou95s, Ude97d]. painful [Ano95-63, Ude98c]. painfully [Tho95a]. Pains [Ano96-111, Joc98a, Nan96b]. pair [Mia95b]. Palindrome [Law95]. Palm [CE98]. palmtop [Way97a]. Panels [Pep98, Van95a, van95b]. panoramic [Hal95d]. panoramic-video [Hal95d]. Paper [Ano95-146, Ano97s, Ano95-98, Ano96-55, Pou95h, Sea95c]. Paperless [Die96b, Ano95-146]. PaperPort [Ano95-212]. paperwork [Cor97]. paradigm [Ber95, Pou95b]. paradigms [OHE95]. Paradise [Ano95-153]. Paradyne [Ano96h]. paragraphs [Ano95-61]. Parallel [Alv96, Pou96b, Pou97c, Rud98, Kri95]. parallelism [Hal98e]. Part [DeJ98, Dob97a, Dob97b, Dob98a, Fri97d, Hal98b, Kay98c, MDM95, Pou95p, Tho95c, Tho95d, Way98a, Abr97a, Ano95-77, Ano95-95, Nan98b, Wat98a]. particular [Ano97-103, Ano97-122]. partners [Ude95a]. Parts [Gre97e, Hus95, Rei95f, Gre97e]. Party [Ano95-54, Tho95f]. Pascal [Ano95-72, Ano95-196, Gre97g]. Pass [Ano96-64]. Past [Ano95-28, Ano95-34, Ano95-29, Ano95-30, Ano95-31, Ano95-32, Ano95-33, Bar95d, Bar95c]. Pastiche [Pou95w]. patched [MB97]. Patching [Sta96e, Sta96a]. path [Gre95e, Gre97r, Kay95a, Tho95e]. pattern [Cla95b, Cua95b, Fri97b]. pattern-matching [Fri97b]. pattern-recognition [Cla95b]. patterns [Ano95-162]. pave [OH95]. pavement [Sin95a]. Pay [Ano96-66, Ano96-92]. paying [VN97a]. Payoff [Key95, Hur97b, Sal95k]. pays [Pou97l]. PBX [Ano98-82, Sal96m]. PC [Ano96-177, AK98, Ano95-28, Ano95-34, Ano95-29, Ano95-30, Ano95-31, Ano95-32, Ano95-33, Ano95-35, Ano95y, Ano95z, Ano95-27, Ano95-44, Ano95-59, Ano95-77, Ano95g, Ano95-137, Ano95-97, Ano95-105, Ano95-119, Ano95-122, Ano95-128, Ano95-154, Ano95-156, Ano95-165, Ano95-187, Ano95-193, Ano95-211, Ano95-205, Ano96w, Ano96-51, Ano96-60, Ano96-64, Ano96-116, Ano96-134, Ano96-144, Ano96-163, Ano96-179, Ano96-176, Ano97-72, Ano98-61, Bar95d, Chi97, Die95c, Gar98, Hal95c, Hal96e, Hal96k, Hal98a, KH95, Laz97b, Mau96, Nan95f, Nan97c, Rei95f, Tho95g, Tho95l, Tho96d, Tho96f, Tho96m, Tra97, Tri95, VN97b, Way96m, Way97f, Way97e, Yag96d]. pcAnywhere [Way97e]. PCI [Ano95r, Ano95s, Ano95-107, Ano96-171, Ano97-77, Ano97-78, Hum97a, Tho95m]. PCI-based [Ano95r, Ano95s]. PCMCIA [Smi95b]. PCs [AK98, Ano95w, Ano95-52, Ano95-94, Ano95c, Ano95-213, Ano96u, Ano96-52, Ano96-99, Ano96-104, Ano96-180, Ano97-60, Ano97-67, Ano97c, Ano97-79, Ano98-48, Ano98-77, Ban97, CE98, CC97, Hal98a, Hal98g, Hum96b, Hum98b, Hur97a, Joh96, Kay96a, KE97, Laz97b, Rei95f, Row97, Sup97, Way97a, Yar95]. PDA [Ano96-51, Ano96-116, Pep97, Tho95f, Way96k]. PDA-to [Ano96-51]. PDAs [Ano96-168, Ano97-126, Cot97, Nad95a, Tho95k]. Peace [Sch96b]. Peak [Ano96-64, BHKM96b]. peek [Ano98-59]. Peer [Mia95c]. peer-to-peer [Mia95c]. penalty [Way95f]. pension [Ano95-74]. Pentafluge [Pou97l]. pentaflugehello [Pou96h]. Penthouse [Ano95-157, Mia95d]. Pentium [Ano95c, Ano95-205, Ano97c, Ano97b, RV96, Row97, Tan98, Ano98-77, And95c, And96d, AK98, Ano95-104, Ano95-158, Ano95-159, Ano95-206, Ano95-208, Ano96-38, Ano96-40, Ano96-117, Ano96-180, Ano97m, Ano97-67, Ano97-117, Ano98-75, Ano98-76, AG95, Bal95d, Bal95a, Bal95g, BHKM96b, BKM96, CR96a, Chi96b, CR96b, Hal95a, Hal95e, HCM96, Hal96g, Hal97d, Hal97e, Hum96b, Hyd96, LM95b, Pou98b, Pou95m, Pou97d, RV96, Tra97, Yag96d, Yag98a]. Pentium-based [Ano95-104]. Pentium-Class [Ano96-180]. Pentiums [Ano95c, LM95b]. People [Ano95-73, Mon95a, Bar95b, Hig95d, Van95a]. percentage [Ano95-77]. Perfect [Ano96-118, Ano97-27, Ano97-75, Sal95f]. Perfect-Fit [Ano96-118]. PerfectOffice [Ano95-160]. Performance [Ano96-73, Ano96-178, Ano98-78, LM95b, Orc98, Ano95-60, Ano95-66, Ano95-104, Ano95-214, Ano95i, Ano95j, Ano97-81, Bal95g, Cli98d, DL95, Die95d, Goo95b, Hal96h, Hal98e, HR97, Hum96g, Hus95, Mul96b, Phi97, Rud98, Sha96, Shi98, Tho95k, Tho95m, Tho96n, Tho97c, VN97b, Way95d, Way95f, Way96b, Smi95d]. performers [BHKM96b]. performs [MF97c]. Peripherals [Hur97c, Tho95m]. perl [Ano98-79, Ano95-40, Ano98-79, Ano98-80, Cox98b, Ude95f, Ude97g, Ude97j]. Persistence [Gre98a]. Persistent [Ude97b, Ske98a]. Person [Ano95-148]. Personal [KE97, Los97a, Ano95-115, Ano95b, Ano95-190, Ano95-205, She95a]. Perspective [Ude95g]. Perspiration [Lin96b]. PERT [Hig95d]. pervasive [Sal95h]. Pet [Ano98-81]. Peter [Ano96-55]. Peter/Jerram [Ano96-55]. PGP [Sta95b]. Phantasmagoria [Hal95b]. Phar [Ano96-119]. Phase [Ano95-169]. Phaser [Ano97-118]. Philips [Ano98-113]. philosophy [Pou95i]. Phone [Ano95-51, Ano95-161, Ano96-62, Ano95-38, Ano95-205, Ano97v, Nan97c, Rei95a, Rei95e, Sal95f, Way95a, Ano96-177]. phone-testing [Rei95e]. PhoneKits [Ano95-205]. PhoneNet [Mia95b]. Phones [Ano97-72, Cam97b, Ano96-167, Ano98-82, Ano98-81]. Photo [Sea96c, Hal95d, Pou95g, Sea96a]. photo-quality [Hal95d]. photo-realistic [Pou95g]. photographs [Ano95-118]. Photon [Var95]. photos [Ano96-179, Ano98-87]. Photoshop [Bla96, Ano98-115]. PhotoSoap [Ano97-117]. physical [Goo95a, Pou98a]. PI000 [Bal95g]. Pick [Ano95-105, Ano96-176, Ano98-58, Bar95c, Bry95d, Hil95, HKM97b, Sea95d]. Picking [Gre97f, Way95g]. Picks [Ano95-41]. pics [Ano97z]. Picture [Ano97z, Hum97a, Mys96b, Wil96, Col96, Gre95a, Mar95d]. Pictures [Log96, Mar95d, Mar95c]. Piece [Ano97-29, Ano97-30, Ano97-31, Sil97, Mul96b]. pieces [Ano95-199, Ano96-55]. Piecing [Ano95-162]. PII [Ano98-96, Ano98-115, Cam98a]. PIIs [Ano98a]. Pin [HCM96]. Pin-Out [HCM96]. Pipes [Ano97o]. Pippin [Ano96-84, Way97d]. pity [Tra95]. PixelFlow [Pou95g]. pizzazz [Ano95-101]. Place [DeJ96f, DeJ96a, Pou96k]. places [Ano95k]. plagues [Pou97j]. Plain [Ano96-62]. Plan [Fil96]. Planet [Sch96b, Sta97b, Way96g]. Planned [Ano96-161, Ano95-75]. Planning [Ano95-218, Esc95b, Kay95a]. Plans [Hal98f, Ano95-74]. Platform [Ano95-50, Ano97-43, Ano98-98, Mor97, Ano95-113, Ano95-142, Ano95-143, Gil98a, Gre95a, Gre97j, Los97b, Sch97d, She97, Ude97f, Way96h, Way96j, Way98b, Yag97a, Ano95-132]. platform-independent [She97, Way96j]. Platforms [Mar95b, Ano95-165, Ren98]. Platter [Kay98b, Kay98c]. Play [Abr97b, Ano95-199, Ano96-92, Ano97-114, Pou95j, Ano95-199, Ano98w]. playback [Ano97-77]. players [Kor95]. Playing [Ano96-120, Coo96c]. Plays [Ano98d]. Plentiful [Ano95-215, Tho95a]. Plug [Abr97b, Ano97-78, Ano97-114, TC97, Ric97, Ano95-199, Ano98w]. plug-in [TC97]. Plug-Ins [TC97, Ric97]. Plugged [Ano95-199, Ano96-121]. Plugged-In [Ano96-121]. plugins [VN97a]. plunge [Dej95j, Gre97m]. Plus [Ano95-97, Ano98w, Ano98-113, BM95, Dav95, Pou97e, Abr97a, Ano95-37, Ano95-110, Ano95-203, Ano95-205, Ano96-173, Ano96-175, Ano96-170, Ano96-171, Ano96-172, Ano96-178, Ano97-106, Bal96d, VN97a, Gre98c]. PN60 [Ano95-213]. Pocket [Ano98-118, Way96k, Way97a, Way97e]. pod [Ano96-125]. Poet [Gre97m]. poetry [Les95]. Point [Ano98-41, Gre97a, BHKM96b, Fri97c]. Point-of-Sale [Gre97a]. Point-to-Point [Fri97c]. pointers [Moh97b]. pointing [Ano96-177]. points [Bal95e]. Poison [And98]. Polaroid [Ano96-43]. police [Way95h]. policy [Ano96-68]. Politics [Bau97]. Polywell [Ano98-48]. Pool [Sal95i]. POP [Kos96, VN97a]. popping [HM95, Pen98]. popular [Esp95, Gre97g, Nan95f, VN97b]. populated [Bon95a]. Populist [Pou97c]. Pornograph [Nee96a]. Port [Nan96f, Ano95-27, Mor97]. portability [Fri97a]. Portable [Ano96-122, Ano98-82, Bal95f, Bal95g, Fri97a, Pou96o, Ren98, AK98, Bal95a, Hum96e, Hum97d, Kay98a, Pou95f, SB96]. Portables [Ano95-77, Ano96-118, Ano97j, BKM96, Ano95-104]. portend [Ano95-36]. Porting [Ano97-60]. Ports [Pou95o]. position [Tho95f]. positioning [Whi95a]. possibilities [Fau95]. Post [Ano96-123, Fri98a]. Post-Millennium [Ano96-123]. Posta [Ano97-101]. Postal [Cro95]. PostScript [Tho95a]. potent [Ber95]. potential [Kay95a]. Pothole [Whi96]. POTS [Ano95-62]. pound [Ano98-96, Row96c, Way97g]. Pournelle [Ano95-163, Pou95l, Pou95o, Pou95n, Pou95m, Pou95k, Pou95p, Pou95q, Pou95r, Pou95s, Pou95t, Pou95u, Pou95v, Pou95w, Pou96p, Pou96o, Pou96r, Pou96q, Pou96n, Pou96m]. Pours [Ano96-31]. Power [And95c, AK98, Ano95r, Ano95s, Ano95-59, Ano95-131, Ano95-156, Ano97-67, Ano97-71, Ano97-81, Ano97-82, Ano97-108, Bak95b, Lin96d, Mia95c, Pou97a, Pou98b, Tho95l, Van97, Way96k, Yag96d, Ano95-106, Ano95-104, Ano95-203, Ano95-205, Ano96-174, Ano97-116, Arc98, Bak95a, Bal95f, BGR95a, BGR95b, Ban97, Cir95, Cla95b, HM95, Hal97c, MA98, Pou95b, Pra95a, Tho95k, Tho95o, Yag96e, Yea95b, Ano95q, Ano95y, Ano95z, Ano95-27, Ano95-165, Ano95-171, Ano97-80, Cox98a, Het96g, Mar95d, Mar95c, Tho95g, Tho95m, Tho96m, Van97]. Power-Builder [Het96g]. PowerBook [Ano95-104, Ano95-203, Ano97-116]. powered [Ano98-48, Tan98]. Powerful [Ano95-165, Ano95-178, Het97, Laz97b, Tho98d, Yea95a, Yea95b, Ano95k, Ano95-72, Ano95-160, Bla96, Cha97, Cli98c, Fau95, Cot95b, Gil96f, Hub95, Moh97a, Rei95f, Sha95c, Wil95]. Powerhouse [Hal98c, RV96, Bal95d]. PowerJ [Gre97d]. PowerLite [Ano95-214]. PowerMacs [Ano96-105]. PowerPC [Ano95-132, Ano95-145, Ano95-164, Ano95-171, Ano95-185, Ano98-83, Cot95a, Gwe95, Hal96g, Hal96h, TJF95, Tho95k, Tho96f, Tho96b, Tho97a, Tho98b, Tra97, Way97d]. PowerPC-based [Cot95a]. PowerPC-related [Ano95-164]. PowerPCs [Ano96-106, Ano96-124, Ano96-161, Tho95k, Tho95o]. PowerPoint [Yea95b]. Powers [Kha98, Sim96]. PowerSite [Ano98-84]. Powersoft [Gre96d, Gre97d, Gre98c]. ppm [Ano95-101, Ano95-189, HKM96, Ano95-101]. PR [And98]. practical [Ano98-63]. praise [Ano96-76]. Precision [Sha95b]. predictions [RFD95]. preemptive [Tho97d]. prepared [Api95c]. preparing [Ano95-60, Ano95-154]. Present [Sea97]. Presentation [DeJ95a, Dej95i]. Presentations [Yea95b, Hum97c, Van95a, Yea95a]. presents [Cot95a]. Press [Ano95-40, Ude95f, Ude95b]. presto [Ess95]. Pretty [Ano98-85, Sta95b]. Prevails [Ano97-80]. Prevent [Kay97a, Ano95-222]. preview [Ano96-170, Ano96-171, Ano96-172]. Previews [Ano95-207, Ano96-125, Ano96-177, Ano98-86]. Previously [Nan95f]. Price [Ano95-53, Ano95-122, Hen96, Ano97z, Bal95f, Hal96h, MW95, Smi95b]. Prices [Ano95-44, Ano95-165, Ano95-122, Ano96-91, Ano97-82]. pricy [HR97]. Prime [Ano96-58]. Print [Ano95-123, Tho95n, Ano97-123, Ano98-126, Sha96, Smi95b, Tho95b]. Printer [Ano95-130, Ano98-126, Row96b, Ano95-118, Ano95-101, Ano95-207, Ano97-119, Hum96j, Hum96g]. printer/copier [Ano95-207]. Printers [Ano95-98, Ano95-101, Ano95-188, Ano95-189, Ano96-44, Ano96-148, HKM96, HR97, KH96, Ano95-48, Ano95-49, Ano95-208, Ano97-118, GM95b, Sha96, Tho95b, Wei95]. PrintGear [Hum96g, Sha96]. Printing [Hum96g, Ano95-118, Ano95-211]. prints [Ano95-213, Ano97-120, Hum96i, Kay97b]. prioritize [Mac98a]. Privacy [Pou95q, Ano96-76, Sta96c, Sta95b]. Private [Hur97e, KFT95, Ude95d]. Prize [Hum98c]. Prizes [Pou95r]. Pro [Ano95-207, Row97, Ano95-203, Ano95-200, Ano96-117, Ano96-178, Bal95f, BHKM96b, CR96a, Chi96b, CR96b, Gre97h, Mon95b, Pou97d, RV96, She95c]. probably [Ano97-127, Fri95a]. problem [Ano95-163, Ano98-73, Gre95b, Gre95c, Hac97, Pou98b, Pou95n, Ude96d]. Problems [Ano98-87, Sal95k, Sal95l, Ano95-109, Ano96-127, Fow97, Joc95c, Los97c, Meh95b, Pou95b, Pou96b, Sha95c, Tra95]. procedure [Yag97a]. Procedures [CC98]. process [Ano98-126, Ude98c, Ano95-210]. Processes [Ess95, Ber95, Gou95]. Processing [Pou95g, Pou96b, Ano95-203, Ano97-45, Cir95, Dob97a, Dob97c, Het97, Kri95, Pen98, Pou97c, Rud98]. processor [AK98, Ano95-95, Ano95-158, Ano95-163, Ano95-159, Ano95-171, Arc98, Em98, MA98, Sha96, Shi98, Sim96, Tan98, Tho96e, Tho96n]. Processors [Ano97-44, Ano97-73, Ano95-80, Cla95b, Joc95b, Laz97b, Tho95k, Yag95a]. produce [CC98, Dob96]. produces [MD95]. producing [Hil95]. Product [Ano95-87, Ano96-81, Ano96-82, Ano95-59, Ano95-77, Ano98-41, Mac95a]. production [Die95c, Vac95b]. Productive [Gil96g]. Productivity [Gil98b, Ano95-201, Cla95c]. productivity-enhancing [Ano95-201]. Products [Mar95a, Ano95-41, Ano95x, Ano95-142, Ano95-143, Ano95-141, Ano95-187, Ano97-115, Ano98-58, BD95b, Bar95c, Bry95e, Cla95b, Dav95, Flo97, Gre96h, HM95, Los97a, MA98, Nan95f, Ude95b, VN96b]. Professional [Ano95-165, Ano95-204, Ano95-209, SK96a]. Profilers [Ano98-78]. profit [Ano97-92]. profits [Ano97-28]. program [Ano95-186, Gre97f, Het96g, Orc98, Sha95b, Way96b, Yar95]. programmability [Gat98]. Programmer [Ano96y, Abr97c, Joh96]. Programmers [Ano98-111, Dob97c, Cot95b, Sha95b]. Programming [Ano96-94, Dob98b, Fri98b, Gre96i, KB96, Lin95a, Rau97, UGB95, Way96b, Ano95-113, Ano96u, Ano96-34, Ano96-76, Ano97-117, Cot95b, Gre95b, Gre95c, Gre96d, Gre97e, Gre97r, Hal97b, KFT95, Pou95b, Pou96b, SK95, Ude95f, Var95, Way96a, Way98a, Gre97p, Pou95b]. Programs [Ano96z, Die95b, Gio96b, Won98a, Ano95-27, Ano95-134, Ano95-135, Ano95-165, Ano95-190, Ano95-214, Ano95-222, Cli98c, Cli98d, EP97, Het96d, Nan97b, Sea95c, Sea95d, Sea96b, Ude97b, Wat95, Won98c]. Prograph [Cot95b]. Progress [Ano95-182, Ano95l]. Project [Ano95-40, Ano97-75, Ano98-112, Hig95a, Ude97a, Ude97k, Ude97b, Ano95g, Hig95d, Sal96b, Hig95d, Ude95c]. projection [Ano95-36, Van95a, Van96c, van95b]. Projector [Van96c]. projects [Hig95d, Kri95]. Proliferate [Ano95-95]. Promise [Ano97x, Ano98m, BVN95, Los97a, Tho95m, Ude97d, Way96j, Wei95]. promised [Ano95-169]. Promises [Ano96-183, Ano97-66, Ano95-107, Ano95-132, Ano95-220, Ano96-173, Joc98a, Tho97c, Ude97k]. promising [Los97b, Sea95c]. prompts [Ude97b]. prone [Hal98g]. proof [Hum96i]. Proper [Phi97, Fri97d]. properly [Ude96l]. Prophetable [Ste98]. Proponents [Hal98b]. proprietary [Ano97-127, Hal97e]. Pros [RV96, Hum96b, Hyd96]. prospects [DA95]. Prospero [Tra96a]. Protect [Bry95c]. Protection [Daw95, Bry95a]. protocol [Ano95-83, Ano95-220, Cli98b, Fri97c, KP98, Lew96, Sta96d]. protocols [Ano95-82, BMW97]. prototyping [Ano95-72]. prove [Ano97a]. proves [And96b, Ano96-178]. Provide [Ano95x, Ano95-103, Ano95-190, Knu96b, Kos96, Law95, Tho95e, Way95b, Yag97a]. Providers [Ano95-174]. provides [Ano95-210, Ano95-211, Ano97-123, Api95a, CCS97, Cua95b, Gre97q, Mia95d, Pou95k, SB96, Way95a]. Proxima [Van95a, Van96c]. Pruning [Sal95k, Sal95l]. pry [Pou95w]. pseudoprivate [Ano95-51]. Psion [Ano96-172, Ano96-178]. Public [Ano95-65, Ano95-62, Sta95b, Ude95d]. Public-Key [Ano95-65]. publicly [Ano95-141]. Publish [Ano98-64, Ano97-118, Ude95b, Ude96d]. publish/subscribe [Ano98-64]. Publisher [Ega95]. publishers [Ano95-88]. Publishing [Ano95-95, Ano95-200, BVN95, Sea96c, Ano97-116, Sea96a]. Pull [Ano98-64, DeJ97, Ude96h]. pulling [BEM95]. pumping [Joc95b]. Punch [Ano95-196]. Pupils [Joc96g, Joc96a]. pure [Gre98f]. Purely [Cot95b]. Purgatory [Ano98-109]. purpose [Way96l]. Push [Ano96-51, Ano98-117, DeJ97, Ude96h, Ano98-64, Pou97c]. Push-Button [Ano96-51]. Put [Flo96c, Sal96c, SY97, Ano95-44, Ano98-123, Bar95b, Nan96d, Pou97d]. Puts [Em98, Ano95-27, Ano95-159, Ano95-210, Ano95-213, Tho95g, Tho97e]. Puzzle [Zha96]. Puzzles [Ano95-162]. Python [She97].

QIC [Ano95-100]. QNX [Krt98, Var95].

QNX/Neutrino [Krt98]. quad

[Ano95i, Ano95j, Nad95b, RV96, Nad95b]. quad-Pentium [RV96]. quad-speed [Nad95b, Nad95b]. Quality [Ano96-184, Ano98-87, DeJ96b, Dej95i, Dut98, Hum96g, Joc95d, KH96, Knu96c, Ano95-118, Ano95-106, Ano97f, Ano97z, Ano97-114, Bry95d, DeJ95a, Hal95d, HR97, MW95, EP97]. quandaries [Ano95-115]. Quantify [Ano98-78]. Quantum [Flo98a]. Qube [Ano98-88]. Queries [Cua95a, Cua95c, Lin95d, Ske98a]. Query [Eva98, Wat95]. query-and-reporting [Wat95]. querying [Lin95d]. Quest [Gar97, Ano98-104]. question [Ano95-187, Ano98-60]. questioning [Hac97]. Quick [Ano97-62, Hum97d, KM96d, Ano95-72, KM96b]. Quick-Striking [Hum97d]. QuickBasic [Esp95]. QuickBasic-like [Esp95]. QuickBooks [Ano95-203]. QuickDraw [Ano96-173]. quickly [Gre96b, Nan95g, Ude96g]. QuickShare [Ano95-195, Bal95i]. QuickSite [Ano96-179]. QuickTime [Hal95d]. quiet [Ano97-117]. quite [Ano95-135].

R&D [Flo98b]. R10000 [Ano95-124].

R12000 [Hal98e]. R5000 [Ano96-126, Hal96f]. Race [Hum98c, Nav97, SM97, Fri95c, Nel95]. racer [Ano95-161]. RAD [Ano95-111, Ano95-112, Ano98-108, Cre96, Gre97g, Gre97h]. Radar [Ano96-140]. Raders [Ano96-127]. Radical [Ano95-166, Gre95e, Pou95d, Ano95-167]. radically [Hal98a]. Radio [Ano95-168, Sal95a, Meh95a, Meh95b]. RAID [Ano95-67, Ano95-99, Ano95-180, Api95b, Guy95, Pou98c]. Raidion [Api95b]. Rainbow [Tho95n, Tho95b]. rains [Pou97j]. Raising [Sch97c]. Rally [Ano95-192]. RAM [Ano95-125, Ano98-104, Goo95a, Pal96]. Ramps [Ano96-33]. RAMs [Ano95-147]. Ran [Gou95]. range [Ano95-100, Bal95e]. ranging [She97]. Rap [Ano98-106, Hal98b]. Rapid [Ano95-166, Ano95-167, Ano98e, Gre98d, Hal98a, Pou95g]. rapidly [Mor97, Tho98e]. Rapids [Ano98-77]. Rates [Sta97a, Mon97]. rather [Ano95y]. Rational [Ano98-78]. RAVE [Way96c]. raw [Ano96-174]. Ray [Ano97-119]. Raymond [Ano95-72]. razor [Hum96i]. razor-sharp [Hum96i]. Re [Bla96]. Reach [Ano96-128, Ano98-35, Hur97b, Pou96j]. reaches [Mia97b]. react [Ano96-127, Ano97-54]. reactor [And96d]. read [Ano95-61, Car95, DFT95b]. Reader [Ano95d, Ano96-133, Ano96-132, Ano96-129, Ano96-130, Ano96-131, Ano97-83, Ano97-84, Ano97-85, Ano97-86, Ano97-107, Tho96c]. Readers [Ano95-116, Ano95-111, Ano95-112, Ano96-134, Ano97-54, Ano97-55, Ano98-60, Ano95d, Ano96u, Ano96-76, Ano96-77, Ano96-78, Ano97t, Ano97-53, Ude95a]. Ready [Ano95-93, Ano97-28, Ano97-125, Cha97, Gou95, MC98, Coo96b, Hal98a, Joc96c, Mon97, SM97, Way95d, BVN95]. Reaity [Ano95-129]. Real [Ano95-89, Ano96-59, Cre96, Fri98a, Gre96g, MW95, Sal95n, Ske98a, Ske98b, SB96, Way97e, Web96, Ano95-128, Cob95, Cro95, DFT95b, Fri95a, Hur97e, KE97, Pou96k]. Real-Time [Cre96, Ske98a, Ske98b, Ano96-59, Cob95, Cro95, Gre96g, KE97]. realistic [Ano96-173, Pou95g]. Reality [Ano95-129, Ano96-151, Ano96-158, Flo96c, KE97, Pou96d, Ano95-146, Ano95-197, Bau95, Hal95d, Way96j]. Really [Ano98-57, Ano95-157, Fow97, HKM97a, Hur98c]. realtime [Cre96]. reason [Pou96k]. Reasons [Ano96-148]. rebels [Hur97a]. Reborn [Tho96l]. Rebuilding [Ano97-87]. Rebuilt [Gre97e]. Recentralization [Wil95]. Recipes [Way98a]. Recognition [Mar95a, Ano95-162, Ano95-187, Ano98v, Cla95b, Cua95b, Dej95g, Laz97b]. recognize [Ano95-187]. reconfigure [Row96c]. Recovery [Ano96-113, Ano97-123]. Recycling [Ano97-88]. Red [Bal95a, Bal95g]. Red-hot [Bal95a, Bal95g]. redesign [CFDT95]. Redesigned [Gre97n, Tho95e]. Reduce [Ano96o, Bie95, Tho95o]. Reduces [Ano95-52]. redundancy [Fri97d]. Reference [Ano95-132]. regardless [Yag97a]. registries [Won98b]. Registry [Won98b]. Regroups [Hal96g]. Reinvented [DM98]. Reinventing [Gil96h]. related [Ano95-164]. Relational [Dre98, Gre98b, Ano95-95, Ano95-151, Ano95-152, Gre98f, Sto98]. Relations [Nan98b, Nan98c, Yav95]. relatively [Ano95-137]. Release [Ano96-173, Cra96, Gil97b, Lin96a, Gil96d]. released [Gre95a]. Releases [Ano95-151, Ano95-152, Ude95b, Yar95]. reliability [Ano95-105, Ano97-30, Ano97-81, Nan95b]. Reliable [Ano95-99, Bre98, Joc95d, KHM^{+}96, Loh97, Gio95, Guy95, Hal98g, Son95]. reliably [CCS97]. Relief [Ano95-122, Ano96e]. rely [Ano95-80]. remains [Het96b]. remarkable [Ric97]. remarkably [Gre95d]. Remote [Ano98-37, Ano98-89, Cam98b, Cli98c, KHM^{+}96, Lev98, Sal95j, Way97e, Won98b, Yag97a, Ano97-123, Fri96d, Nan95c].

Remote-Control [Won98b]. remotest [Ano95k]. removable [Mia97b]. removable-media [Mia97b]. removes [Won98b]. removing [Mia95f]. renderer [Pou95g]. Rent [Ano96-66]. repackage [Rei95e]. Repeat [Ano97-58]. replace [Ano96-55, Pou97f, Sal95c, Sal95k]. Replacement [Van96a]. replacements [Ano95-77]. Replication [Ano95-133, Ano96n, Ano97-89, Bre98, Dob95, Yav95, Ano95k]. Reply [Ano95-86, Ano95-90, Ano96-80, Ano96-83, Ano96-133, Ano96-132, Ano96-129, Ano96-130, Ano96-131, Ano97-83, Ano97-84, Ano97-85, Ano97-86]. Repoet [LM95b]. Report [Ano95-117, Ano95-106, Ano95-99, Ano95-98, Ano95-104, Ano95-101, Ano95-100, Ano95-103, Ano95-107, Ano95-105, Ano97-100, Ano97-98, Ano97-99, Ano98-38, Ano98-40, Ano98-39, Gar98, Hes97, HKM97b, HR97, KM96d, KM96c, Mys96b, Mys96a, SK96b, Ano95-102, KHB96b]. reporting [Wat95]. reports [Aer95, Hac97]. Repositories [Mac95b, Mac95a]. Repository [Wat98b]. represents [Ano95-105]. reproduction [Sug95]. request [Ano97f]. requiem [Abr98b]. require [Gio95, Pal96, Sha95c]. required [Ano95-199]. requirements [DA95]. requires [Ano97-105, DeJ96b, Hur97b, Sal95h, Whi97]. Research [Ano97-49, Ano97-50, UT95, Dej95g, Ano95-204, Ano95-209]. researcher [DM98]. Researchers [Sha95c, Hed95]. Reseller [Ano97-90, Ano97-91]. Resellers [Ano97-92, Ano97-87]. Resolution [Ano95-118, Gre95f]. resolve [Ano98w]. resource [Ude96l]. Resources [Ano97-87, KP98, Wil95]. respond [Ano97-55, Ano98-60]. responding [Ano95-136]. Response [Lin96c, Ano98-47]. Rest [Ano98o]. restores [Ano98-74]. restrictions [Sta97b]. result [Ano95-75]. results [Ano95-117, Ano95-192, Hur98c]. retain [Ude96d]. Retreats [Abr97b, Abr97a, Abr97c, Ano97f]. Return [Ano96-135]. returns [Pou95n, Pou95k]. reusable [Way96j]. rev [Tho95o]. reveal [Ano98-121]. Revealed [Ano95-155, Joh96]. Reveals [Hal98f, Ano95-158, Die95d, Gre98a]. Revenge [Abr97a]. Review [Gre97o, Kay97a, MF97b, Ano95-200, Ano96-163, Hig95a, KH95, LM95a, Mac95a]. reviewer [Cra96, Gil95]. Reviews [Die97a, AG95]. Revisited [Ano95-170, Shi98, Tho96f, Tho96g, Ude97a]. Revisiting [Ano98-74]. Revive [Cra96]. revived [Sch97a]. revives [Cra96]. Revoir [Ano96i, She95c]. revolt [Hal98a]. Revolution [Ano96h, Ano97-100, Sha95a, Ano95-28, Ano95-34, Ano95-29, Ano95-30, Ano95-31, Ano95-32, Ano95-33, Ano95-35, Bar95d, Hal97a, Ano97-77]. revolutionary [Joc95e]. revved [Tho95h]. Rewritable [Ano95-169]. rewrites [Ano98-52]. RFC [Ano95-170]. Rhapsody [Ano98d, Tho98c, Tho98e]. Rich [Fil96, Tho98e]. Richer [Ano96-136, Ano95-61]. Rick [Ano95-47, Gre96i]. Riddle [Dej96e]. Ride [Ano95-193, Ano98a]. Rides [Ano95-71, Ano98-59]. Right [Ano96-67, Ber95, Lin95a, Ano97-32, Bal95f, Bry95d, Car95, Gil97c, Gre95b, Gre95c, Hal95d, Nad97, Sal95m, Sim97, Way95c, Yav95]. Ring [Nan96c, Ano97-28, Nan96a]. Ringing [Nan97c]. RISC [Ano95-171, Ano95-203, Ano97-81, Hal95a, Hal98e, Hum98a, MA98, Pal96, Pou95c, Pou95d, Pou95i, Row96a, Shi98, Sim96, Tam95d]. risk [Nan96d]. Risks [MF97d]. Risky [KFT95]. Rivals [Ano95-93, Ano97-118]. Road [Ano96-137, Ano97-79, Bal95h, Tam95d, Van96c, Way95f, Whi95a, Ano97-107, Rei95c, Wat98b, Whi95b]. Road-Show [Van96c]. roadmap [Ano98-122, Mon98b]. roam [Ano98-82]. roaming [Bal95e]. roaming-enabled [Bal95e]. Roar [Ano98-50]. Roars [Ano95-145, Ano96-180]. robbed [Ano95w]. Robbers [Ano95w]. Robo [Ano96-138]. RoboCop [Ano96-138]. Robot [Eng97]. Robots [Ano96-54, Ano97t]. robust [Gio95]. robustness [Ano95-105]. rocket [Ano95-159]. rockets [Gwe96]. Rocks [Mar95e]. Rogue [Gre96b]. ROLAP [Ano96-87]. Roll [Ano95-172]. ROM [Aer95, Ano95-37, Ano95-36, Ano95-195, Ano95i, Ano95j, Ano96-50, Bal95h, Bal95i, BM95, CFDT95, DFT95b, Emi97, Fro96, KFT95, Kay97a, Nad95b, O’M96, Pal96, SK95, Ude96d]. Romance [Van96c]. ROMs [Aer95, ATA96, Ano95-37, Ano96p, Ano96-121, BDF96, BM95, CFDT95, DFT95a, DFT95b, HCM96, Hal96d, KFT95, MDM95, MG96, Pou95n, RFD95, SK95, TJF95, UT95, UGB95, Ude96d]. room [Sea95c, Ude95b, Vac95b, Ano95-40, Ude95f]. Ross [Yag96d]. rounds [Wil95]. Roundup [BEM95, BVN95, Gio95, Gio96c, Gio96d, Het96e, Hig95d, Hig95c, Nan95f, Sea95c, Sea95d, Sea96c, Sea96d, Vog95, Yea95b, Hig95b]. Route [Gre97h]. routed [Sal95k]. router [Fri97d]. Routers [Ano98z]. routines [Gre95f]. RP [Tho96n]. RPC [Cli98a]. rpm [Mia97c]. rport [Ano95-103]. RS [Ano95-186]. RS/6000 [Ano95-186]. RSA [Way98b]. Rugged [Ano96-108, Tri95, Ano95-97]. ruggedized [Tri95]. ruminations [Ano96-34]. Run [Ano95-173, Gre97f, Way97a, Ano95y, Ano95z, Ano95-104, Ano95-191, CCS97, Chi96a, Joh96, MF97c, MA98, Pra95b, Tho95l, Tho97b]. Run-Time [Ano95-173, MF97c]. Running [Ano95-222, Ano96-98, Hum96h, Ano95-93, Ano95-137, Ano95-128, Ano97-30, Ano97-81, BGR95b, Mia95c, Sal95i]. runs [Ano95-150, Ano95-186, Gwe95, Hub95, Pal96, Tam95a]. runtime [Orc98]. Rush [Hed95].

S [Het96g, Sta98]. S/MIME [Sta98]. S5 [Ano96-174]. S5-120 [Ano96-174]. Safe [Nan96d, Kay95a, Way95g]. Safeguards [Ano95-100, GM95a]. safer [Ano97-75]. Safety [MF97c]. Salamone [Ano95u]. Sale [Bry95e, Gre97a, Bry95b]. Sales [Ano96-64, Joc98b, Ano95-77, Nad95c]. Salsa [Kay97c]. Salvation [Ano98-66]. Salvatore [Ano95u]. Salvatore/Salamone [Ano95u]. Salvo [BMW97]. Samba [Yag97b]. same [Ude98b]. Sampler [Ano96-63, Dav96]. Samsung [Cam97b]. sanctioned [Ano95-120]. sanity [Abr98a]. Santa [Pou97f]. satellite [Abr98c, Ano97-107, Mon97, Ano98-86, Bal95f]. Satellites [Whi95b]. satire [Ano95-37]. Sauce [Kay97c]. Save [Ano95-103, Abr98a, Ano95-38, Ano95-102, Hur97e, OHE97, Sal95d, TM95]. saved [Abr97b]. Saves [Ano98-100]. Saving [Ano95-222, VN96b]. Saviors [VN96b]. Savvy [UGB95, Nee95m]. Say [Ano97-46, Ano98v, Ano95-75, Ano95-93]. says [Ano95-219, Ude95g, Van96d, You96]. Scalability [Fry98, Ano95-59, Ano97-30, Nan95b]. Scalable [Api95b, Pou95g, Ano98-64, Cli98c, Whi98]. Scale [GE95, Ano95-59, Ano95-118, Ano97-103, MC97, Yag95a, Yav95, Yag95c]. scaled [Krt98]. Scales [Krt98]. Scan [Ano97-119]. Scan-Soft [Ano97-119]. ScanJet [Ano97-114]. scanner [Hum96i]. Scanning [Coy98]. scary [Knu96c, Mon95a]. scenario [Pou96a]. scene [Cot97]. scenes [Hal95b]. scheduling [Ano96-178, Cla95c, Yav95]. scheme [Fri97e]. Schiffer [And96d]. Schmidt [Ano98-120]. School [Ano95-171]. schools [Rei95d]. Science [Gre96i, Gre96i]. scientists [Hed95, Sha95c]. Scitor [Ess95]. SCO [Nan98d]. Scorecard [Bry95d, Kor95]. Scorecards [PH95, Bry95d]. Scott [Ano97-58]. scoured [Ano95-41]. scrambling [Ano95-44]. Scratch [Ano95-181, Api95a]. screaming [Bal95g]. screams [Em98]. screen [Flo96a, Hal95d, HKM97b, Joc95e, Tho97e, Hil95]. Screens [Ano96-57, Ano98g, HKM97b, Ano95-70]. script [Dob97d]. Scripting [Fri97b, Ano98-79, Fri97a, She97]. Scriptlets [Ano98-28, Dob98d, Dob98c]. Scripts [LS98, She97]. SCSI [Api95b, Hum96f, Hur97c]. SDK [Ano95-73]. Seagate [Mia97c]. seamless [Yag97a]. seams [Pou95w]. Search [Ude95h, Ude97h, Ude98c, Ano96-78, VN97b]. search-engine [Ude97h]. searched [Ano95-141]. searches [Ano95-106]. Searching [VN97b]. seasoned [Joh96]. Seasons [Pal96]. second [Ano95-49, Way96j]. Secure [Fri97c, Sta96c, Sta98, TS96b, Gio95, KFT95, Los97a, Sta97b, TS96a, Ude97f, Ude98c, Yag95b]. secured [Way98b]. Securing [Ude98e]. Security [Ano95l, Ano96-102, Ano96-95, Ano97-75, Ano97-93, Ano98-92, Coy98, Hur98a, MF97d, MF97c, Ude97f, Ano95-98, Ano96-68, Ano96-77, Daw95, MB97, Nan95c, Smi95a, Way95b]. See [Hal95d, Tho96r, Wil96, Ano95-192, Cam97a, Joc95e, Tho96a]. seeing [Sch97a]. seen [HM95]. Sees [Ano95-219]. segment [Tho95k]. segments [Loh97]. select [Ano95-101]. selects [Ano95-100]. Self [EVN96b, EVN96a]. Self-Serve [EVN96b, EVN96a]. send [BHKM96a]. Senders [Ano97-97]. sendfax [Way95a]. Sends [Gil96d]. sensor [Ano97z]. Sequel [Het96f]. Serial [Ano95-193, Hur97c]. Series [Ano95-211, Die95a]. serious [Hal97b, Het97, Tra95]. seriously [Key95]. Serve [Ano98-117, EVN96b, Sch97e, Won98a, EVN96a]. Server [Ano95-123, Ano96-74, Ano96-111, Ano96-139, Ano97-81, Ano97-82, Ano97-121, Ano98-89, Ano98-97, Fri95a, Gre97f, Gre97q, Het96c, Los97b, MC98, Nan95b, Nan95d, Nan98e, OH95, OHE95, Smi95b, Tho96k, Tho96n, UGB95, Ude96i, Yag97a, Ano95-100, Ano95-190, Ano96u, Ano96-144, Ano97h, Ano97-103, Ano97-123, Ber95, Cli98b, Cli98d, DeJ96b, Dej95f, Cot95a, Gio95, GE95, Gre98c, Het96f, Het96g, Mac95b, Mia95c, Pou97b, QS96, Smi95a, Tho97c, Ude95a, Vog95, WS96, You96, Ano95-56, Ano97-123, Ano98-69, Ano98-125, Cot95a, Gio95, Nan95d, Nan96b, Wil95, Het96d]. Server-based [Gre97q, Ano95-100, Gio95, Rei95e]. server-oriented [Cli98d]. Servers [Ano95-67, Ano95-119, Ano97-78, Ano97-102, BEM95, Bri97b, Hes97, Hum97b, Hum98c, KHM^{+}96, MM96, Ude98d, Ano95-151, Ano95-152, Ano95-154, Ano97-30, Ano97-45, Ano97-81, Ano97-82, Ano97-122, Ano98-115, Het97, Kay98b, Kay98c, RV96, SY97, SP97, Smi95b, Ude98b, Won98a, Yag97b, Ano97-103]. Serves [Het97]. Service [Ano95-123, Ano95-174, Ano96-133, Ano96-132, Ano96-129, Ano96-131, Ano96-139, Ano97-83, Ano97-84, Ano97-85, Ano97-86, Ano97-99, Dut98, Esp95, KP98, MC97, SP97, Ano95-137, Ano96-130, Ano98-47, Cam98b, Rei95a, Rei95e, Sal95m, Sta97a, TM95, Cro95, KP98]. Services [Ano96-113, Ano98-88, Sha98, Ano95-142, Ano95-143, Ano95-136, Ano97l, Ano97-123, Ano98-79, Bry95e, Cla95a, Kos96, Lew96, MC97, OH95]. Serving [Ano98-90, Het96d]. servlet [Ude97b]. servlet-based [Ude97b]. Servlets [Cli98d, Ude97k]. Set [Ano95-98, Ude97e, Way97d, Cam98a, Die95a, Sal95i, Tho98e, Way96h, Way97b, Whi98]. Set-Top [Way97d, Way96h]. Sets [Ano98-107, Die95c, Gil96g, Van96c, Yag95b]. Setting [Ano97-29, Fri95a, Ano98-120]. setup [Ano95-106]. Seven [Ano95-175, HR97, Ano97-75, AG95, Bry95d, Gil97c, Sea95d]. Several [Ano95-50, Ano95-48, Ano95-199, Pou95m, Pou96j]. Sexes [MDM95]. SGI [Gre97l, Hal98e, Row97]. SGML [Ano95-95, Ano95-117]. SH [Sup97]. SH-3 [Sup97]. Shape [Ano95-176, Ano96-164]. Share [Ano97-94, Van95a, van95b, BD95b]. shared [Bal95i]. shark [Dej95j, Hum97d]. shark-infested [Dej95j]. sharp [Hum96i]. sharpness [Ano95-106]. sheds [Gre97g]. Sheets [Ude97c]. ship [Ano95-154]. Shock [And95c]. shoot’em [Ano95f]. shoot’em-up [Ano95f]. Shooting [Ano98-77]. shop [Cox98b]. Shoppers [Ano96-114, Wat97a]. Shopping [Ano97-90]. Short [Cam98b, Smi95c, And95c, Ude97d, WS96, Smi95d]. Short-Order [Smi95d]. short-term [Ude97d]. Shortage [RFD95]. shortcomings [Nan95g]. shotguns [Ano95w]. Should [Van97, Gas95, Sch97d, Tho95j]. Shoulder [Ano97l]. Show [Bar95a, Sea97, Van96c, Ano95f, Ano95-41, Ano95-141, Ano95-145, Ano95-199, Ano98-58, Pou97e, Yea95b]. Showcase [Ano95-39]. shows [Dob96, Het96b]. shrink [Ano95-163, QS96]. shrink-wrapped [Ano95-163, QS96]. Shrinkwrapper [Ano97-48]. side [Dob97a, Hal96i, Sch97d]. Siemens [Vac95b, WS96]. Siena [Ano96-172]. SIGGRAPH [Pou95n, Pou97e]. Signal [Ano96-99, Cla95b, Yag95a]. Signals [HCM96]. significant [Pou95b]. signs [Min96]. Silicon [Tho96p, Coo95, Way96g, Row97]. silver [Way96g]. Silverstream [Sha98]. Simple [Api95b, Vog95, Ano95-61, Ano98-51, Kay95d, Nan95c, Sha95b, She97]. simplest [Pou95t]. simplicity [Mia95b]. Simplified [Dob98c]. simplifies [Tho96n]. Simplify [QS96]. Simplifying [Sal95j]. Simply [Kay97b]. Simulation [Vac95b, DeJ96c, Gre96i, Knu96b]. simulator [Ess95]. since [Ano95f, Bal95b, Hal97a]. Sincere [Ano95-177]. Sing [Lin96c]. single [Ano95-100, Ano95-220, Em98, GM95a, Hum97a, Mul96b, Pra95b, Tho95k]. single-chip [Em98]. single-medium [Ano95-100, GM95a]. single-slot [Hum97a]. single-threaded [Pra95b]. single-user [Tho95k]. singles [Sal95c]. Sins [Tho95a]. Sir [Nee96a]. Sitara [Ano98-52]. Site [Ano96-85, Ano97-42, Bal96a, Dob98d, Eng97, Gio97, Pou97l, Ano95-111, Pou97e, SP97, Ude96b, Ude96d, Ude96g, Ude97h, Ude98b]. sitebuilding [Gil97c]. Sites [Ano96-150, Nee96e, Ano98-51, Bri97a, Dob97a, Gre97q]. Six [Gio96b, HKM96, Los97a, Mia95f, Pou96g, Sin95a, Hum97b, Yag95c]. size [Joc95e, Wil96, Sal95i]. skill [Nad97]. Sky [Mon97]. sleek [Ano97-114]. slick [Ano95e, Api95c, Gre96d]. Slide [Sea97, Yea95b]. slide-show [Yea95b]. slimmed [Tho96f]. slimmed-down [Tho96f]. Slip [Sha98]. Slot [Smi95b, Hum97a]. slots [Hal97e]. Slow [Ano95-146, Tan98, CC98, Hal98b]. Slow-Bus [Tan98]. slower [Pra95b]. Slowly [Lin95c]. slugs [Ano97-102]. Small [Ano96-113, Ano98-110, BN95, Cot96c, Gre95d, Gre98e, MA98, SC96, Sta97b, Yag95a, Ano95-70, Ano98-88, BD95b, Cot96a, MC97, Yag95c, Yag95c]. Small-Office [BN95]. Small-Scale [Yag95c]. Smaller [Ano97-60, Ano97-87]. Smallest [Flo98b]. smallish [Tho97e]. Smart [Ano95-88, Ano96-60, Ano97-95, Ano98-100, Cam97b, Het96a, Het96e, Sch96c, Ano95l, Ano95-210, Ind95]. Smartcard [Ano98-92, Ano98-91]. smarten [LaP96c]. smarten-up [LaP96c]. Smarter [Ano95-178, Ano96-140, Bri97b, Pou98f, Wel97]. Smarts [Ano97-92]. SmartSuite [Ano96-156, Ano98-93, Gil96h]. smoker [Ano95-128]. Smooth [Ano95-179]. SMP [Ano97-82, Hum97b, Pra95b]. SMS [MC98]. smudge [Hum96i]. smudge-proof [Hum96i]. SNA [Sal95k]. snacks [Abr97b]. snap [Ano95-129, Pou97d]. SNMP [Sta96e, Sta96a]. Socket [Ano97-96, Hal97e, Ano95-159]. Soft [Ano95y, Ano97-119]. Soft-Windows [Ano95y]. Softimage [Ano98-86, Pou97j]. Software [Ano95-39, Ano95-130, Ano95-134, Ano95-180, Ano95i, Ano95j, Ano95-222, Ano96p, Ano96-147, Ano97-103, Ano98f, Ano98n, Ano98-40, Ano98-94, Ano98-128, BEM95, BVN95, DFT95a, DFT95b, Fow97, Fri97a, Gil98b, Gio95, Gio96c, Gio96d, Gre95f, Gre97o, Het96e, Hig95b, Hig95c, Joc95d, Pou95s, PF97, Ric97, Sal95d, Sch96e, Sea95c, Sea95d, SK96b, Sea96c, Sea96d, Ste98, Ude97j, Vog95, Web96, Wil96, Won98c, Yea95b, Abr98b, Ano95l, Ano95x, Ano95-52, Ano95-80, Ano95-93, Ano95-151, Ano95-152, Ano95-150, Ano95d, Ano96-78, Ano96-173, Ano96-179, Ano96-177, Ano96-178, Ano97a, Ano97-117, Ano97-119, Ano97-112, Ano98-51, Ano98-113, Ano98-114, Api95a, Bau95, Daw95, DeJ96b, DL95, Esp95, Gil96g, Gwe95, Hig95a, Ind95, Law95, MG97, Rei95d, Rei95f, Sug95, Tho95e, Tra97, Ude96g, Vac95b, Wat95, WS96, You96]. Software [Ano95-64, Ano97-102, Dej95e, Esp95, Het97, Lin95d, Mar95d, Whi98]. Software-Installation [Pou95s]. software-monitoring [Sal95d]. SoftWindows [Ano95z, Ano95-27, Tho95g]. Sofware [Hig95d]. Solar [Vac95b]. Solaris [Ano97-122, Pra95a]. solid [Ano95g, Ano97-123, Lin98, Yar95]. solid-modeling [Yar95]. Solo [Ano96-175, Ano96-174, Ano98-85]. Solomon [Ano95-208]. Solution [Sal95k, Ano95-137, Eva98, Rei95e, Sal95f, Sea95d]. Solutions [Ano95-180, Cla95c, Cro95, Dej95j, Nad95c, Pou95h, Rei95e, Sal95l, Way95h, Whi95b, Ano95-116, BD95b, Ren98, Way98a, Ano95z]. solve [Ano95-163, Ano98-73, Los97c, Pou98b]. solved [Joc95c]. solves [Fri97a]. Solving [Ano98-95, Zha96, Pou96b]. Some [Flo97, Fri98b, Goo95a, Pou97j, Ude97i, Ano95-73, Ano95-75, Ano95-82, Ano95-107, Ano95-154, Ano95-182, Ano95-187, Ano98w, BEM95, Bal96d, Cot97, Dav95, Flo96a, Gre97h, Hig95c, Hil95, HKM97a, Nan95c, Tho95a, Tho98c, Tra95, Ude97c, Ude97k, Ude98c]. Someday [Gar97]. Something [Ano95-46, Ano97-97, Ano98w, Gre97j]. Sometimes [Gre95b, Gre95c]. Song [Tan98]. Sony [Ano98-96, Way95a]. Soon [Ano96-75, Wat97b, Ano95-49, Ano95-70, DL95, Flo96a, Gre97l, Nad95c]. sophisticated [Ano95-60, Ano95-61, MF97c, She97]. Sort [Gre96h, Ano95-191]. Sorting [Kay95c]. sound [Ano96-76, Ano97-53, Van95a, Yag95c, Ano97z]. Sounds [Ano96-150, Ano98-81]. Soup [Hal98b, Way98a, Ano95-56, Ano97g]. soup-up [Ano95-56]. souped [Ano97-67]. souped-up [Ano97-67]. Sour [Ano97-102, Ano97-103]. source [Ano95-214, Dob98a, Ric97]. sources [Dob97c]. Space [Ano98-100, Coo95, DeJ98, Nad95c, Ano95-70, Ano98-104, Flo96c]. space-coming [Ano95-70]. spacecraft [Pou95m]. Spam [Pou96l]. span [Ano95-190]. Spans [Cam98b]. Sparc [Ano95-214]. SparcBook [Tam95d]. SPARCplug [Yag96d]. SparcStation [Yag96d]. sparks [Ude95d]. Speak [Gio96b, Nan96e]. speakers [Ano95-204, Ano95-209]. Speaking [Ano97-46]. Speaks [MH97b]. spearhead [Ano98-92]. Spec [Ano96-124, Yag96c]. Special [Ano97-100, Ano97-98, Ano97-99, Bou95]. specifies [Cun95]. Spectnum [Ano95-154]. Speech [Laz97b, Mar95a, Ano95-162, Ano95-187, Ano98v]. speech-and [Ano95-162]. Speech-Recognition [Mar95a, Ano95-187]. Speed [Ano96-90, Ano98-48, Hal96h, Nan95a, Nan95e, Ano95-101, Ano95c, Ano95i, Ano95j, Ano97l, Ano97-127, Ano98m, AG95, Bri97a, Bry95d, DM98, Hal98b, HKM97a, Hum96f, Hum96a, LM95b, Mia97a, Mia97c, Nad95b, Pra95a, Ske98b, Tho97a, Way95d, Nad95b]. Speedier [Ano98-97]. speeding [Dob98a]. Speeds [Ano96-62, Pou97a, Bal95i, Ske98a]. Speedy [Ano97-101, Pou95c, Pou95d, Ano98-96, Hal96c]. spell [Hal95e]. Spells [Ano96e]. Spider [Ano95-200]. Spidering [Ude98c]. spiffy [Ano95-161]. Spin [Ano97-77]. Spinning [VN96c, Ude97j]. sports [Nan95a]. Spot [Ano97-27]. Spotlight [Ano98-35]. Spreadsheet [Cra96, Gas95, Ess95]. Spreadsheets [Ano96z, Pen98]. Sprees [Ano97-90]. spring [Ano95-164]. Springtime [Tam95e]. spurred [Cob95]. SQL [Ano95-211, Het96f, Lin95d, Nan95b, Vog95]. SQLWindows [Het96g]. Squared [Ano98-88]. Squeeze [Ano95-70]. squeezes [Way97g]. SSA [Ano96-141]. SSL [Ude98c]. stable [Nan95a]. Stack [Ano96-142, Sal96a]. stacks [Sta96a]. Stage [Ano97h]. stages [Ano95-132]. stand [Nan95d, Smi95b]. stand-alone [Smi95b]. Standard [Ano95-132, Ano98-107, Bur95c, Bur95a, Dob98b, Eva98, Sch97c, Ude97e, Ano95-89, Ano95-118, Ano95-140, Ano95-191, Ano95-193, Ano95h, Ano96-100, Ano96-157, Ano96-174, Cun95, Dav95, DL95, Gil96g, Laz97b, Sta96b, Van96c, Yag95b, Ste95]. Standardize [Gar97]. standards [Ano97-29, Ano97-53, Cli98a, Cre96, Dej95b, Fou95a, Fou95b, Hal98a, Lin98, Sug95]. Standing [Ano98-98]. Stands [Ano98-27, Way97f]. Star [Ano98-48, Mar95a, UT95]. STARE [Ano97-97]. Start [Gre97r]. started [Gre96b]. Starting [Ano95-181, Ano95-51, Api95a, Row97]. Starts [Ano97m]. Starved [Way96i]. State [Moh97b, Ano95-105, Tho96g]. state-of-the-market [Ano95-105]. Station [Ano97-104, Ano95-161, Chi97]. Statistical [Cua95b]. Statistical-pattern [Cua95b].

Statistics [Cua95b, Ste98]. Stay [Ano96-143]. Stays [Ano95m]. steam [Mia95c]. Step [Ano97-36, Fou95a, Ano97h, Ano98-89, Bal95h]. Steps [Fou95b, Ind95]. stereo [Van95a]. Steroids [Nan95c, Vau96, VN96a]. stick [Van97]. Sticker [And95c]. Still [Ano95-182, Ano96-124, Ano95l, Ano95-132, Ano95-201, Chi95, HR97, Joc97b]. Stinking [Ano96-162]. Stop [Ano98-99, Bar95b]. Stopwatch [Gre95f]. Storage [Ano96-36, Ano97p, Ano98-90, Kay98c, Nan95f, Pou98c, Ano95b, Ano95-211, Ano98-104, DeJ96a, DeJ98, Kri95, Ude97b]. Store [Ano95-103, Ano97-113, Mon98a, Gre98a]. Store-and-Forwards [Ano95-103]. Stored [CC98]. Storefront [Sin95b]. stories [Ano96-127, Ano97-54]. Storing [Ano98-100, Dre98]. Storming [Gil95]. Story [Sal96n, Ano96-163, O’M96, Sal96b]. Straddles [Ano98-84]. Strained [Yav95]. Strange [Ano96-183]. Strata [Die96a]. Strategic [Ano95-183, Bar95a]. Strategies [Ano96-35, Rob96, KB96]. Strategy [Ano98-29, Ano98-63, Nan96d, Pou95d, Sal95d, Ude95e]. Stratus [Ano96-144]. street [Daw95, Ano97-92]. Strength [BEM95, Ano95-196]. strengthens [Tho95f]. strengths [Dob98d]. stress [Cam98a, Hum97b]. stress-test [Cam98a, Hum97b]. Strike [Hur97a]. Striking [Hum97d]. Strings [Joc98b, Die95d]. stripped [Way96h]. stripped-down [Way96h]. Strong [Ano95-160, Ano95-141, Nan95b, Nan95d]. StrongARM [Pou96c, Tho97e]. stronger [Hal98e]. struggle [Pou96i]. studio [Die95c, Ano98-86]. study [Die97b]. stuff [Ano95-182]. Stupid [Pou97j]. Style [Alv96, Smi95a, Ude97c, Ano95-159, Ude96b, Bou95, Ude96b]. Stymie [Ano95-161]. subclass [Ano95-61]. subnotebook [Joc95e]. Subscribe [Ano98-64].

Subscriber [Pou97a]. Subsystems [Ano95-99, Ano95-100, Guy95]. succeed [Mar95e]. success [Ude97b]. successful [Sil97]. successfully [Bau97]. Successor [Ano97-96, Hal95a]. such [Ano95-118, Ano95-166, Ano95-167, Daw95, Dej95e, Hig95d, Tur97]. Sudden [BGR95a, BGR95b]. suffering [Mia95f]. suffers [Ano95-201]. Suit [Moh97a]. suitable [Krt98]. Suite [Ano95-144, Ano95-157, Ano96-172, Ano97-27, Ano97-102, Ano97-103, Pep97, Ano95-83, Ano95-160, Esp95, Gre95a, SK96b, SK96a, Yag97b, Mia95d, Whi98]. Suites [Ano95-134, Ano95-135, Gil96g]. SuiteSpot [Nan97a]. Suits [Ano97-76]. summarize [Ano95-88]. summer [Ano95-164]. Sumo [DV96]. Sun [Ano95-186, Ano97-58, Ano97-122, Fri96a, Smi95d, Tam95e, Way96l, Whi98, Yag98b]. SunDisk [Ano95b]. SunSoft [Ano96-145, Mon96b]. Super [Sim96, Kay97b]. SuperCede [Gre97i]. Supercharge [Kri95]. supercomputers [Sha95c]. SuperCow [Pou96q]. Superfast [Hur97c]. SuperScript [Ano97-120, Hum96g]. Superservers [Nan97a]. suppliers [Ude95a]. supplies [BGR95a]. Supply [Ano97-71, BGR95b]. Support [Ano96-74, Ano96-147, Ano95-80, Ano95-157, Ano95-219, Ano97-31, Bal95h, Dej95b, Esp95, Fen95, Fri96c, Laz96, Los97b, Mia95d, Nan95a, Nan95b, Tho95e, Tho95f, Tho95j, Van95a, VN97a]. supporting [Hig95d]. Supports [Ano95-132, Ano95-185, Ano95-193, Gil97a]. sure [Hum96c, MB97]. SureStore [Mia97a, Law95]. Surfing [Vac95a]. surgery [Pou97f]. surprise [Tam95a]. Surprises [Pou95r, Kay95d]. surprising [Ano98-121, Ude97b]. Surround [Hal95d]. Survey [Ano96v, Ano96-146, Ano96-153, Ano97u, Hal98f, Ude95a]. surveyed [Sea95c]. Surveys [Ude96l]. Survival [Ano98s]. suspects [Way95h]. suspended [Ano95g]. sustain [Goo95b]. SVmini [Ano97z]. SVmini-209 [Ano97z]. Swan [Tan98]. Swap [Pou96i]. Sweet [Ano98-101]. Swimming [Hur98b]. Switch [Ano96-96, Fri97d, Gil96a]. switches [Sal95g]. Switching [Ano95-102, Ano95-103, Sal95g]. Swoops [Pou95m]. Sybase [Ano98-84]. Symantec [Ano95-184, Ano95-185, Ano96-31, Ano96-178, Cot95c, Way97e]. Symmetra [Ano97-71]. symmetric [Pra95a]. synchronization [Nan95c]. syntactically [Lin95d]. synthesis [Laz97b]. SyQuest [Mia97b]. Sys [Ano96-178]. System [Ano95-212, Ano96-107, Ano98a, Fri96a, Pou95f, Yag95b, Ano95g, Ano95-186, Ano95-199, Ano97-127, AG95, Api95b, BD95b, Bri97a, Cli98d, Cot95a, Gre95d, Gre97l, Gre98b, Gre98f, Los97c, Nan95a, Nan97c, Pou97d, Pou98d, Way95d, Whi95a, Bak95b]. system-independence [Cli98d]. Systems [Ano95-107, Ano95-150, Ano95-158, Ano95-186, Ano95-178, Ano95-176, Bie95, Fau95, Fri98a, Moh98, MC98, Sal95h, Ano95-60, Ano95-120, Ano95-132, Ano95-145, Ano95-151, Ano95-152, Ano95-154, Ano95-162, Ano97-31, Ano97-67, Ano98-77, Bal95c, Bal96b, Cla95b, Cua95b, Dav95, DL95, Esc95b, Gio95, Gre96g, Mon97, Sin95a, Van96d, Yag95b, Yav95, You96, Wil95].

T [Ano96h, Ano96-177, Chi96b]. T4900CT [Bal95d]. table [Hal95e, Tho95c]. tackle [Pou95b, Sha95c]. Tactics [Pou96c]. Tadpole [Bal95a, Bal95g]. Take [Ano95-64, Ano95-77, Ano96-108, Ano96-149, Ano98m, Fri95d, Gre97o, Hur97d, MG96, Ude95i, Ano95k, Ano95-82, Ano96-164, Ano97-107, Cor97, Het96d, Mia95f, Nan97b, VN96b, Way95f, Ano98-55]. Takedown [MG96]. Takes [Ano96-73, Ano98-124, Cor96, Hal96b, Ano95r, Ano95s, Ano95-73, Ano95-211, Gre97m, Pou95g, Pou96l]. taking [Pou97f]. Tales [Nee95m]. Talk [Ano96z, Ano96y, Ano96-27, Ano96-75, Nan96g, Ude96f, Mar95a, Ano95-45]. Talking [Ano95-187, Mar95a]. Talks [Ano98-76, Ano97-58]. Tame [DeJ96c, Gio96d, Gio96a]. Tangle [MM96]. tangling [Bau97]. Tape [Ano98-116, Ano95-117, Ano95-100, GM95a, KHB96b, KHB96a, Law95]. Taps [Ano95-75]. Target [Ano97-35]. targets [Way98b]. task [Cox98b, Joh96]. tasks [Ano97-29, Laz96]. Taste [Pou96k]. taught [Ano95-197]. TCP [Ano95-116, Ano98-52, Ano98-68, Cli98b]. TCP/IP [Ano95-116, Ano98-68, Cli98b]. TDMA [Ano97g]. te [Ano96-68]. Teach [Joc96g, Nan96e, Joc96a]. teaches [Ude98c]. Team [Ano96-156, Het96g, Won98c, Hal97a, Lo97]. Team-Building [Won98c]. teamware [Won98c]. Teamwork [EG96b]. Tech [Ano96-147, Ano98-87, She95b, She95a]. Tech-Support [Ano96-147]. Technical [Ano96-154, PH95, Cua95b, Flo96a, Joc95c, Pou97c, UT95]. technique [Way95g]. techniques [Hed95]. technological [Ano95-48]. Technologies [Ano95-169, Ano96-56, Bau97, Sch97f, Ano95h, Ano96-164, Bar95c, Chi95, Fri96d, Hal95c, Hal96c, Kay95c, Tho98c, Ude97d, Way95e]. Technologists [Sch96d]. Technology [Ano97p, Ano97y, Dej95c, Dej95b, Hum96c, Pou98a, Sal95i, Abr98c, Aer95, Ano95-116, Ano95-105, Ano97-67, Ano98-120, Ano98-81, Ano98-93, Ano98-113, Api95a, Bas95, Chi96a, DeJ97, DeJ98, Dej95h, Dob95, Dob98d, Gil95, Hal95b, Hal95d, Hal96d, Hal98a, Hum96g, Hur97b, Key95, Laz97a, Nan96g, Sal95m, Wat98b, Way97d, Ano96-178, Sal95h, Smi95d]. teeth [BEM95]. teeth-gnashing [BEM95]. Tektronix [Ano97-118, Ano97-119]. Telecommuter [Fri96c]. Telecommuting [Ano96c]. telephones [Joc97b]. Telephony [Ano95-91, Ano95-193, Ano97-105, BHKM96a, Bur95b, Mel95, Nad97, Tra95, Yag95c, Ano96-157, Ano97-28, Bur95a, Cro95, Yag95a]. Telerights [Rid96]. TeleSpeed [Cam98b]. tell [Kor95, Nan96g, Way95c]. Template [Ste95]. Ten [Ano96-148, Bar95a]. terabytes [Ano97f]. TeraFLOPS [Pou98b]. term [Ude97d]. Terminal [Ano95-135, Ano98-69, Ano98-89, Ano98-125]. terminals [Way96h]. terminate [Ano97v]. Test [Ano95-188, Ano95-189, Ano95-98, Ano95-101, Ano95c, Ano97m, Ano98n, Ano98-77, BGR95b, BEM95, Cam97a, Cam98a, Die95c, Fin98, FH97, HKM96, HKM97a, Hum97b, Hur98c, Kay98b, Kay98c, LM95b, MG97, Nan97b, Pou95m, RV96, Tan98, Van97, Way97a, Won98c, Yag95c]. Tested [Ano95-85, Ano95-106, Ano95-104, Ano95-107, Ano95-105, Ano96-70, Ano96-71, Ano96-72, LM95b, Ano95-84, Ano95-100, Ano95-103]. Testing [QS96, Sal95c, Die95d, Dob96, Gre95a, Het96b, LM95a, Rei95a, Rei95e]. Tests [Ano95-199, Ano97b, Ano95-42, Ano95-106, Ano95-107, Ano95-145, Ano98-121, BHKM96b, BHKM96a, Pou98c, Sea96b]. text [Ano95-118, Ano95-212, Hum96i]. Textbase [Ude97i]. TextBridge [Ano95-204, Ano95-209]. Thanks [Ano95-83, Ano95-128, Hal96d, Nan96g, Ano95-66, Ano95-77, Ano95b, Cla95b, Dav95, Ano96-114]. their [Ano95x, Ano95-59, Ano95-93, Ano95-134, Ano98m, BVN95, Dob98d, Gre96g, Laz96, Mar95a, Tho95o, Van95a]. them [Ano95-71, Ano95-191, MB97, Sha95c, Vog95, Way95g, Way96l]. theology [CFDT95]. theory [Kay97c]. There [CFDT95, Dej95d, MD95, Wil95, Die95d, Joc95d, MF97a]. These [Cli98d, HR97, SY97, Sil97, Ano95l, Ano95-63, Ano95-125, Ano95c, Ano98-92, Gio97, Hum96b, Sea97, SP97, Sha95c]. They’re [MF97a, Cam97a, Dob96, Joc95c, Ric97]. They’ve [Tra95]. thicket [Ano98-104]. Thin [Ano98-85, Ano98-102, Ano98-103, MC97, Qui97, And96c, And96b, Cam97a]. Thing [Ano95-129, Kha96, Flo95, Gre97j, Hum96c, Sch96b]. Things [Ano98-62, Nel95, Pou97j, Ano95-194, Ano95-201, Ano97-107, Ano98-81, Pou95t, Ric97]. think [Ano95-83, Fri95a]. ThinkPad [Ano98-51, DeJ95a, And96c, And96b, Bal95h, Joc95e]. Thinner [Ano96-136]. third [Shi98, Tho95f]. third-generation [Shi98]. third-party [Tho95f]. Those [Ano97-59, MB97, Ano96-55, Dej95d, Hum96h, Hur98c, Knu96c, MD95, Wat97b]. though [HR97]. Thought [Ano95-154]. Thousand [Log96]. Thousands [Ano96-33, Sal95d]. Thread [Pra95a]. threaded [Pra95b]. Threads [Pra95b]. Threat [CC98, Ano97f]. Three [Ano98-48, Ano98-66, Bal95b, EP97, Pou97c, Pou97d, Ano95-64, Ano97-105, Ano98-81, Fau95, Het96d, Nan95f, PH95, Way97a]. Three-dimensional [Bal95b]. thrifty [Kay97b]. throughout [Ano95-105, Ano95-141, Ano95-201, Ano97-29, Bal95e]. throughput [Lo97]. Throughs [Ano95-103]. Throw [Way97g]. thus [Tho95d]. Ticking [Sch98]. Tidal [Ano95-164]. tie [Mac98b]. ties [Gil98a]. tight [Rey96]. Tighter [Hur98a]. tightly [Coo95, Gil98a]. Till [Nee96d]. Tillamooks [Yag98a]. Time [Ano95-173, Ano96-58, Bie95, Cre96, Fri98a, Gre96g, Ske98a, Ske98b, SB96, Ano95-125, Ano96-59, Ano98-89, BVN95, Cob95, Cox98b, Cro95, FH97, Gre95f, Hig95d, KE97, MF97c, MA98, Pou96i, Vog95, Whi95a, Hal95b, Whi95b]. time-consuming [Cox98b, Gre95f]. time-critical [Vog95]. time-honored [Hig95d]. timers [Gre95f]. times [Hal96d]. Tiniest [Ano97-104]. Tiny [Krt98, Ano95-174]. Tip [Nee95m]. Tips [Ano95-57, Ano95-97, Bar95a, Flo97, Hil95, Ude97i, Ude98a]. Tired [Pal96]. Titles [Ano96-50]. Today [Hal97b, Hal98g, Hil95, Ude97d, Ano95-105, Ano95d, Bry95e, Gwe96, Hal95c, Mar95b, Sha95c, Way95g]. Together [Ano95-162, Dob96, Dav95, Mac98b, Tho97b]. Token [Nan96c, Nan96a]. Token-Ring [Nan96c]. Tolerance [CCS97, Abr98c]. Tomb [Gre96k]. Tomorrow [Ano98a, Hal97b, Joc96e, Ano96-164, Ano97g, DeJ98, Mar95b]. too [And96b, Cli98d, Hal98b, Mar95a]. took [Dej95j, Nan95f]. Tool [Ano95-46, Ano95-45, Ano96-149, Gre96d, Ano95-95, Ano98-51, Ano98-108, Bar95a, Die96a, Gre96i, Gre97f, Gre97g, Gre97r, Nad97, Vac95a]. Toolkit [Ano95-208, Way96j, Way98b]. Toolkits [Gil97c, Gla96, Gre97f, Het96g]. Tools [Ano95-95, Ano95-190, Ano96o, Ano96-93, Ano97n, Ano97-105, Dej95f, Gio97, Hes97, Het96g, Lin96d, Nad97, Wat95, Ano95l, Ano95-88, Ano95-136, Ano95-166, Ano95-167, Ano97-106, Ano97-117, Ano97-123, Ano98-115, Ano98-114, BVN95, Ber95, Cot97, DeJ96b, DeJ96c, Del96a, Die95c, Gar97, Gil96f, Gio96a, Hum98a, Joc97b, Lin98, Mac98a, Mul96b, Nad95a, Pou97e, QS96, Sim97, Ude98c, Vog95, WS96, Wil95]. toothpaste [Ano98-87]. Top [Wat97a, Way97d, Ano95-186, Cot95a, Way96h]. Top-down [Wat97a]. top-of-the-line [Cot95a]. TopLink [Gre98f]. Toshiba [Ano97-106, Ano98h, Ano98-86, Bal95d, Yag98a]. Toss [Dej96g]. toting [Pot95]. Touch [Bla96, Hur97b]. Tough [Ano95-80]. tougher [Gil96d]. Tour [Ano96-150]. touting [Ano95-59]. tower [Ano95-205]. Toys [Ano96-151]. TP [GE95]. Track [Yav95, Ano95-62]. Tracking [Hur98a, Sch97f, Hig95d]. tracks [Pou96k]. trade [Gre97f]. trade-offs [Gre97f]. traders [Bau95]. trading [HR97]. Traditional [Nan98a, Ano97-30, Sal95f, Sal95h, Yea95b]. traffic [Ano95-51, Cro95, Mac98a]. traffic-oriented [Cro95]. Train [Yav95]. training [Sal95i]. Transacting [Ano98-105]. Transaction [Ano97-45, Gre98c]. Transfer [Nan95c]. transfers [Nan95c]. Transformations [Whi97]. transforming [DeJ97]. transforms [Tra97]. transistor [Pou98a]. transistors [Hal96h]. Transition [Ano95-123]. translates [Ano95-211]. translation [Ano98-51, Ano98-113]. translations [Bri97a]. transmission [Los97c]. transport [Ano95-220, Pou95i]. Transport-Triggered [Pou95i]. Trap [Ano97r]. traps [TC97]. Travel [Gre98e, KFT95]. traveler [Nee95m]. traveling [Ano96-40, Bal95d]. Trek [Mar95a, UT95]. trend [Ber95, Dej95b]. trends [Ano96-91, Pou97c]. Tri [Ano98-48]. Tri-Star [Ano98-48]. trick [Bal95c, Kay95d]. Tricks [Ude97i, Bal96d, Fri98b, TC97, Ude95f]. tricky [Joc97b, MC98]. tries [Ano98-73]. Triggered [Pou95i]. TriMedia [Kha98]. Trip [Nan95a, Nee95m]. Tripmate [Ano97-107]. Triton [LM95b]. Triton-based [LM95b]. Triumphant [Sch96d]. Trouble [Ano96-45, CC97, Hal97e, OHE97]. troubleshooting [And96d]. Tru [Bal95h]. trucking [Whi95a, Whi95b]. true [Ano95-157, Mia95d, Ano95-107]. True-Color [Ano95-107]. TrueTime [Ano98-78]. Trust [Sta95b]. Trusted [Ano97h]. Truth [Hal95e]. try [Gre97f, Pou97d]. Tsunami [Ano95r, Ano95s]. Tumbleweed [Ano97-101]. tumbling [Hur97d]. Tune [Los97c]. Tuned [Ano96-152, Ano97-122]. Tuning [Fri95b]. Tunnel [Yav95]. Turbocharged [Tho95o, Ano95-161]. Turn [Ano95-58, Los97c, Ano97-118, Chi97, Gou95, Nan97c, Sim97, Ude95a, Ude96g]. turned [Ano95-60]. turning [Ano95-88, Ano97-82]. turns [Ano95-205, Bal96f, Gre98b]. TV [Dej96g, Way97d]. tweaks [Hal98e]. Twin [BHKM96b]. Twin-Engine [BHKM96b].

Twist [Bal95c, Law96]. twisted [Mia95b]. twisted-pair [Mia95b]. Two [Ano95-169, Ano95h, Ano97-107, Gre97o, Gre98f, Pou97k, Tho95l, Tho95o, Tho96o, VN96b, Ano95e, Ano95-28, Ano95-34, Ano95-29, Ano95-30, Ano95-31, Ano95-32, Ano95-33, Ano95-35, Ano95-158, Ano95-194, Ano96-178, BD95b, Bar95d, Daw95, Gio95, Hal95b, Hal98a, Het96b, Het96g, Hus95, MDM95, Nan97b]. two-way [Daw95]. Type [Kay96a, MF97c]. types [Cua95b, Cun95, DeJ97].

U.S. [Cro95, Sta97b]. ubiquitous [CL97]. Udell [Ude95c]. ugliness [Gre97h]. Ugly [Ano96-88, Ano98-87]. UI [Tho98c]. Ultimate [Ano96-85, Kad96, Pou95i]. Ultrafast [Ano96k, Smi96b]. Ultraportables [Ano97-108]. UltraSCSI [Hum96a]. UltraSPARCs [Smi96b]. unattended [Law95]. unauthorized [Daw95, UT95]. Uncertain [Gre97p]. undergo [Hal97a]. underlying [Api95a]. underpinnings [Cua95b]. understand [Ano95-187, Kay97c]. Understanding [Ano97-46, Ano95-162, Phi97]. understood [Ano95-61]. Unearthing [Min96]. unending [Ano98-104]. Unexpected [Pou95t, Pou96k]. unforeseen [Cro95]. Unhinged [Ano97-109]. Unicode [Fow97]. unified [Kay96a]. unify [Die95a]. uninstaller [Mia95f]. uninterruptible [BGR95a, BGR95b]. unit [Hum96e, Smi95b]. unite [Esc95b]. Unites [Lew96, Ano95-196]. units [Pou95d]. Universal [Ano95-193, Gre97b, Wat98b]. Unix [Ano95-137, Ano95-192, Ano95-210, Ano95-214, Ano96-127, Ano96-134, Ano96-153, Ano96-155, Ano96-154, Ano96-160, Ano97-30, Ano97-58, Ano97-81, Ano97-115, Ano98-106, Cor96, Cox98b, DFT95b, Fri97a, Fri97b, Hub95, Lac96, Laz97a, Moh97a, Moh97b, Mor97, Nan98d, Tam95c, Yag95b, Yag96d, Yag96e, Yag97b]. Unix-like [Moh97b]. Unix/Windows [Yag96d]. Unixes [Kra98]. UnixWare [Nan98d]. Unleashed [Sal96g]. unless [Ano95-199]. unlikely [Ano98-87]. unnecessary [Sal95d]. unpleasant [Kay95d]. unplugged [Sal95m]. unprecedented [Hal96c]. Untangling [Ano95-191, Cun95, Sal95m]. Unthinkable [Ano97-98]. until [Ano95-194, Pou97d, Tho95j]. Unveiled [Fed97]. unwanted [Bry95c]. up-to-date [Ano95-212]. upcoming [Ano95-128, Ano98-59]. update [Ano95-42, Ano95-95, Ano95-112, Moh97b]. Updated [Die97a, Nan95a]. Upgrade [Ano95-159, Los97b, Mia95c, Pou96n, Ano95-206, Ano97-125, FH97, Hum96h, She95c]. Upgraded [Het96g, Ano95c, Ano97-33]. Upgrades [Pou95m, Ano98-115]. Upgrading [Ano95-63, MC98, Pou96i]. upon [Pou97j]. UPS [Ano97-71, BGR95b]. upstart [Gas95]. Upwardly [Ano97-110]. urban [Ano96-125]. usability [Yea95b]. USB [Ano98c]. USD4000 [Ano95-154]. Use [Ano97-39, DFT95a, DFT95b, Fri97a, Ano95w, Ano95y, Ano95-51, Ano95-48, Ano95-62, Ano95-61, Ano95-71, Ano95-132, Ano95-222, Ano98-68, Ano98-87, BMW97, Car95, Cka97, DeJ96c, Dob96, Ess95, Hig95d, Hum96e, KP98, Orc98, Phi97, Son95, Sta96c, Ude97c, Ude97i]. used [Ano95-137, Ano95-194, Ano95-222, Ano98-100, Arc98, Cun95, Hal95d]. useful [Fri98b, Nad95a, Nan95a, Sim97, Tur97, Whi97]. Usenet [Meh95a, Meh95b]. user [Ano95-61, Ano95-64, Hal97a, Tho95k, Wes96a, Ano95-64]. Users [Ano95-201, Ano96-156, Ano98f, Ano95-60, Ano95-71, Ano95-104, Ano95-136, Ano95-219, Ano96-178, Ano97t, Ano97-32, Daw95, DA95, Fri96d, Hal97d, Hal97e, Kor95, Laz96, Mac98a, Mac95b, Sal95f, Ude95b, Ude95e, Ude95g, Ude95i, Way97b]. uses [Abr98c, Ano95-142, Ano95-143, Cla95c, Gil98b, Lin95d, MC97, Pou96l, Pou98d]. Using [Ano97z, Ber95, BMW97, Cox98b, Die95c, Ano95-106, Ano95-116, Api95a, Bar95a, Cla95b, Die95a, Fou95a, Fou95b, Fri96c, LM95b, Hed95, Ste98]. utilities [Ano95-63, Esp95, Mia95f, Pou96j, Yag97b]. utility [Ano96-170, CCS97, Fri97a]. utilizes [Dob98a]. UX [Son95].

V.34 [Ano95-105, Ano95-161, Esp95, KH95]. V.90 [Ano98-107]. V120 [Ano96-175]. validate [Mul96b]. valuable [Hum96h, Ude96l]. Value [Dej95g, Ude97j, Tho95d, Wat97a]. variant [Dob97d]. Varieties [Ano97d]. variety [Ano95-55, Ano95-191, Cun95, Dej95g, Joh96, Kay95c, Way98a]. various [Fri97a]. VARs [Ano97s, Bau97]. vary [Ano95-59]. varying [Ano97-105]. VB [Ano95-58, Gre96d]. VBA [Gat98]. Vector [Tho98b, Ano96-174]. vector-drawing [Ano96-174]. Vegas [Ano98-58]. Velo [Ano98-113, Way97f]. Vendor [Ano96-155, Ano95-59, Hur98c]. Vendors [Ano95-193, Ano95-192, Ano96-116, Ano96-157, Ano98-48, Ano95x, Ano95-52, Ano95-59, Ano95-89, Ano95-135, Dav95, Mac98b, Wil95]. Verifier [Ano98-57]. Verilog [Cre96]. Veritas [Yag95b]. Versa [Hum97c]. Version [Ano95-184, Ano95-47, Ano95-61, Ano95-208, Ano96-178, Ban97, Car95, Cli98b, Cra96, Dob95, Fri96a, Nan96g, Pou96l, Tam95a, VN97a]. Versions [Smi95a, Api95c, Het96g, Nan95b]. versus [Ude95d]. Vertical [Rei95e, Rei95a]. Very [Ano98-85, Ano97-97, Lin95c, Sha95a, VN96b]. via [Ano95-184]. viability [Die95c]. viable [Dej95g, Sea95d]. Vibe [Gre97j]. Victory [SK96b, SK96a]. Video [Ano95-107, Ano95-165, Ano96-61, Ano96-184, Chi97, Die95c, DL95, Fri95c, KM96c, KE97, SC96, Ano95-204, Ano95-209, Ano97-77, Ano97-120, Bal95h, Die95a, Hal95d, Hil95, Joc95a, KM96a, Nad95b, Pou98c, Van95a, Hal95d]. video-editing [Chi97]. Videocassett [Ano95-100]. Videoconferencing [Ano95-194, Ano97-68, Dav95, Sea95d]. Videophone [Kha98]. Videophones [Way96i]. Videos [Ano95-94]. VideoTalk [Ano97-106]. View [Aer95, Ano96-155, Hum97c, Gat98, Jen95]. Views [Ano95n]. ViewSonic [Ano95-204, Ano95-209]. Village [Ano96-61]. Virtual [Ano95y, Ano95z, Ano95-27, Ano95-129, Ano95-197, Ano95-195, Ano96-27, Ano96-151, Ano96-158, Bal95i, Hur97e, Mac98b, Sal95n, Sea95d, Sea95b, Sin95b, Tho95h, Tho95g, Tra97, Ano95-211, Bau95, Goo95a, Hal95d, Nad95c, Ude95b, Ano95-40, Ano95-129, Flo96c, Ude95f]. virtual-reality [Bau95, Hal95d]. Virtually [Joc96f, Way95f]. Virus [Ano98-50, Pou95k, Ano95-208]. Viruses [Nan96d]. Visio [Ano98n]. vision [Ano95-61, Ano95-162, Sal95h, Ano97z, Won98b]. visit [Pou95r, Pou97e, Pou97l]. visited [Ano95-40]. Visual [Ano95-47, Ano95-196, Ano98-108, Api95c, Cot95b, Gla96, Gre96i, Gre97k, Gre97f, Gre97e, Ano95-46, Ano95-82, Ano97-111, Ano98-78, Api95c, Api97, Dob97d, Gil95, Gil97a, Gil98a, Gla95, Gre95e, Gre96h, Gre96i, Gre97c, Gre97h, Gre97k, Gre97q, Gre97r, WS96]. visual-programming [Cot95b]. VLIW [Pou96f]. VMS [Fen95, Smi96c]. VO [Tho96n]. Voice [Ano96-39, Ano97-57, Ano98-109, HG96, Lin96c, Sal96i, Ano95-51, Ano95-213, Ano98-41, BHKM96a]. Voice-Mail [Ano98-109, Ano98-41]. VoiceType [Nan96g]. volatility [Ano97-92]. volt [Ano95-104]. voltage [Bal95d]. Vote [Ano96-159, Ano96-160]. VP [Ano98-76, Hac97]. VPNs [Ano98-110]. VR [Ano95-197, Ano97-61, Hal95d, Pou96d]. VRML [Ano95-129]. vs [Ano96-38, Ano96-127, Ano96-155, Ano96-154, Gio96c, Het96b, Nan97a]. Vu [MB97]. VxDs [Gre96j]. vying [Ano95h].

Wabi [Tam95b, Tam95c]. wacky [Sch96f]. wafers [Coo95]. wagon [Ano95-161]. wait [Ano95-122, Die95b]. Waiting [Nee96f]. Wake [Ano97-93, Emm96, Tra95]. Wake-Up [Ano97-93, Emm96, Tra95]. Wall [Ano97-92]. wallet [TH96]. Walls [Ano96-181, Has96]. WAN [Fri96d, Hen96, Sal96l]. WANs [Ano98-64]. Want [Ano95-201, Ano95c, Bar95b, CC98, DA95, Kay95d, MF97a, Nan96g, Tho95l]. Wanted [Ano98-111, Gas95, Sch96e]. War [Ano97g, Cra96, Ano97-102, Sch97d, Smi95b]. ware [Nan97b, Sal96j]. Warehouse [Ano95-57, Ano97u, Ano98-119, Fry98, Wel97, Hur97d]. Warehouses [Ano97-32, Ano97f, Kri95]. Warehousing [Hac97]. warning [Pou97g]. Warp [Nan95a, Nan95e, Ano97-123, Gio96c, Nan96g, Tam95a, Ude95g]. Warped [Ude95g]. warplanes [SK95]. Warrior [Bal95h]. Wars [Ano95-44, Ano98r]. Was [Pou97g, Ano97a]. Washington [Pou95q]. Watch [Ano96-161, Ano97-44]. Watchamacallit [Ano97-112]. Watcom [Ano95-47]. Waterproof [Hum96i]. waters [Dej95j]. WAV [Tho96e]. Wave [Ano95-60, Ano95-136, Ano95-164, Ano95-80, Gre96b]. WaveLAN [Bal95e]. WaveLAN-base [Bal95e]. Wavelets [Ano95-198]. Way [Ano97n, Ano97-91, Cli98d, Coo96a, Hum98c, Ano95-157, Ano95-197, Ano97l, Ano97v, Bri97a, Daw95, Dej95h, DM98, Hig95c, Hur97e, MFDG98, Mor97, Nan95f, OH95, Way96h]. Ways [Ano95-175, Boy97, Hum96c, Pou98c, Rei95d, Ano97-75, Arc98, Dob97c, Fri98b, Hal98b, Sha95c, Ude97c, Ude97b, Way96g]. Weakness [Ano95-93, Ano95-155]. weaknesses [Ano95-201, Dob98d]. Wealth [Ano97-94, Ano95-37]. weapon [Way95h]. Wear [Ano96-60, And96d]. Wearable [And96d]. Weather [Ano97-104, Cro95, Gla95, Sha95c]. Weaving [Gio97, MFDG98, Pra95a]. Web [Ano95-200, Bee96, Ude95h, VN97a, Sta95b, Ano95o, Ano95x, Ano95-61, Ano95-95, Ano95-111, Ano95-134, Ano95-135, Ano95-216, Ano96-55, Ano96-52, Ano96-69, Ano96-77, Ano96-78, Ano96-85, Ano96-93, Ano96-112, Ano96-134, Ano96-146, Ano96-152, Ano96-164, Ano96-165, Ano96-163, Ano96-166, Ano96-179, Ano96-167, Ano97h, Ano97n, Ano97-39, Ano97-45, Ano97-75, Ano97-90, Ano97-114, Ano97-113, Ano98-47, Ano98-51, Ano98-64, Ano98-79, Ano98-84, Ano98-80, Ano98-93, Ano98-105, Ano98-108, Ano98-114, Ano98-112, BVN95, Bal96a, Bal96d, Bal96e, BMW97, Bri97a, Cox98b, Dob96, Dob97a, Dob97d, Dob98a, Dob98d, Eng97, Flo95, Fom98, Fri96a, FH97, Gil96f, Gil97a, Gil97c, Gio97, Gre97n, Gre97q, Hal96e, Hal97b, Het96d, Hum98a, Kha96, Los97c, MFDG98, MM96, Nee96e, Nee96f, OHE97, Pen98, PM97, Pou97e, Pou97l, Sal96c, SP97, Sha98, Sin96b]. Web [Sta95b, Sta96b, Ude95a, Ude95b, Ude96b, Ude96d, Ude96g, Ude96h, Ude96j, Ude96k, Ude96l, Ude96m, Ude97c, Ude97i, Ude97j, Ude97a, Ude97k, Ude97b, Ude98a, Ude98b, Ude98d, Vac95a, VN96c, Way96m, Way97d]. Web-based [Ude95a, Ano98-47]. Web-integration [Ano97-114]. Web-ready [BVN95]. Web-server [Het96d]. Web-style [Ude96b]. Web/CD [Ude96d]. Web/CD-ROM [Ude96d]. WebCD [Ude96d]. Webmaster [Ude96b, Ude97h]. Webmeister [Ude95f]. WebNFS [Ano97-54]. WebObjects [Ude96h]. Webserver [Ano96-163]. Webware [Gil96e, Gil96b]. Weeding [Mia95e, Mia95f]. weeks [Law95]. Weigh [DV96]. weight [Way95b]. Weird [Sch96f]. welcome [Var95]. Well [Ano95-135, Ano96-168, Sch96c, Smi96c, Ano95-117, Ano95-107, Ano96-177, Ano97a, Api95c, Joc96f, VN97a, Wel97]. Well-Connected [Ano96-168]. well-designed [Api95c, Wel97]. went [Pot95]. were [Ano95-222]. We’ve [Hac97, Ano95f]. Whatever [Ano96-169]. Where [Ano98-120, Key95, Kin96, Ano98-100, Nan96g, Tho97b, Ude95b]. Wherefore [Ano98-119]. Whether [Pot95, Ano95-95]. Which [Ano97-121, Ano97-122, Ano97-123, Ano98-121, Hal96i, Sch97d, Ano95-206, Ano95-222, Ano97-105, BGR95b, Cob95, Sha95c, Tho95j, Way96g, Wil95]. While [Ano95-132, Hig95d, Nee96f, And96d, Ano95-122, Ano95-157, Tho95e, Tho96n, Ude97k]. White [Ano98-46]. whiteboarding [Sea95d]. Whiteboards [Sea95d, Sea95b]. Who [Jen95, Var96, Ano95-60, Ano95-97, Ano95-146, Sal95f, Sha95b]. Whole [Nan95g, Gre95a]. Wide [Ano95-95, Ano98-85, Vac95a, Het97, Joh96, Nan96d, Way98a]. widely [Ano95-59, Ano95-194]. Wider [Ano97-31, Hal98e]. Wild [Ano95-217]. Wildcat [Esp95]. Wildfire [Ano95-217]. Will [Ano95h, Ano96-69, Ano96-98, Ano98f, Ano98-52, Ano98-118, Gre98e, Hal97b, Sal95c, Sch96a, Ude97e, Ano95y, Ano95-48, Ano95-61, Ano95-75, Ano95-80, Ano95-93, Ano95-122, Ano95-127, Ano95-140, Ano95-146, Ano95-164, Ano95-191, Ano95-219, Ano96-164, Ano97g, Ano97-58, Ano98-92, Cka97, DeJ98, DM98, DL95, Die95a, Fin98, Flo97, Fou95a, Fou95b, Gar97, Gre97d, Gre97p, HM95, Hal95c, Hal96c, Hal96d, Hal97c, Hal97d, Hal98a, Kos96, Mac98b, Mar95b, Mon95a, Nad95c, Pou97b, Rei95d, Rei95f, Son95, Ude97c, Way96l]. Win [Ano95-50, Ano95-125, Ano95c, Ano96-67, Ano96-103, Ano96-128, Ano96-175, Ano96-180, Gar98, Goo95a, Hum96h, LM95b, Pou96j, Pou97j]. Win32 [Fen95, Mon95a]. Win98 [Ano98-122, Mon98b]. WinChip [Ano97w]. WinDD [Ano95-210]. Window [Bak95b, Way96j, Ano95-219]. windowing [Var95]. Windows [Bak95a, Bak95b, Gil98a, Alv96, Ano95k, Ano95f, Ano95y, Ano95z, Ano95-27, Ano95-46, Ano95-45, Ano95-55, Ano95-64, Ano95-68, Ano95-73, Ano95-104, Ano95-100, Ano95-107, Ano95-112, Ano95-123, Ano95-122, Ano95-125, Ano95-133, Ano95-145, Ano95-134, Ano95-135, Ano95c, Ano95-163, Ano95-157, Ano95-166, Ano95-167, Ano95-184, Ano95-182, Ano95-218, Ano95-210, Ano95-207, Ano95-212, Ano95-213, Ano95-219, Ano96u, Ano96-27, Ano96-135, Ano96-181, Ano97-31, Ano97-53, Ano97-60, Ano97-75, Ano97-125, Ano98-72, Ano98-69, Ano98-74, Ano98-78, Ano98-89, Ano98-102, Ano98-123, Ano98-124, Ano98-125, AG95, Api95a, Bak95b, BD95a, BD95b, Car95, CCS97, Cha97, Chi96a, Cox98b, Cra96, Die95b, Die95c, DL95, Die95a, Die95d, Dob95, Em98, Fre96, Gil98a, Gio95, Gio96b, Gio96c, Goo95a, Gre95a, Gre95e, Gre96g, Gre96j, Hal97b, Has96, Hig95a, Hig95d, Hig95b, Hig95c, HKM97a]. Windows [Hum96g, Kay96a, Laz96, LM95b, Lin95b, Lin95c, Mia95a, Mia95e, Mia95f, Mon95a, Mor97, Nan95a, Nan95c, Pou95f, Pou95n, Pou95k, Pou95u, Pou95v, Pou95w, Pou96i, Pou96l, Pou97k, Pou98f, Pra95b, Pra95a, Qui97, Rey96, Ric97, Sea95d, SM97, Tam95a, Tam95c, UT95, UGB95, Ude95i, Ude97f, Var95, VN97a, Way96j, Way96m, Way97a, Won98b, Yag96d, Yag96e, Yag97b, Yea95a, Yea95b]. Windows-Based [Ano98-89]. Windows-friendly [Tam95c]. WinFax [Ano95-207]. WinFrame [Qui97]. WinHEC [Pou97l]. winning [Smi95b]. Wins [Ano96-144]. WinSock [Ano95-220]. Wintel [Ano98-114, Row97]. wire [Ude95d]. Wired [Sch96b, Sin96b]. Wirehead [Sin95a]. Wireless [Ano95-212, Ano97-126, Sal95a, Sal95m, Ano95-210, Ano95-211, Ano95-208, Ano97g, Bal95e, Way95f]. Wires [Way96e, Ano95-191]. Wirth [Pou95f]. wise [Sch97b]. within [Ano95-141, Ano97-103]. Without [And95c, Tra96b, Way96e, Ano95-88, Ano96-181, Ano97d, Has96, Hig95c, HR97, Kay95d, Kor95, Nan95g, Row96a]. witnesses [Pou95m]. Wizard [Ano95-211]. wizardry [Pou97j]. Wobbly [Tam95b, Tam95c]. Woes [Ano97u]. Womb [Gre96k]. Womb-to-Tomb [Gre96k]. wonder [MB97]. wondered [Ano95-40]. wonderful [Sch96f]. Wonderland [Bau95]. Won’t [Ano95-216, Ano97-58, Ano95-50, Ano95-194, Hub95, Rei95f]. wor [Dav95]. Word [Log96, Pou96f, RFD95, Ano95-95, Ano95-163, Tho96e, She95c]. WordPerfect [Ano96-182, Ano97-27]. Words [Ano97-112, Pou97g]. Work [ARH95, Ano95-182, Ano95-216, Ano96-158, Dob96, Dob97d, Joc95d, MW95, Tra96b, Ano95-97, Ano95-199, Ano95-219, Ano96-175, CFDT95, DM98, Hur97b, KFT95, Nad95b, Pou95j, Tho98c, Ude97c]. Workbench [Pou98d]. workers [Nad95a]. Workgroup [Ano98-126, BD95a, BD95b, Nan98e, BEM95, Hig95d, Hum96j, She95c, Cot95a]. workgroup/network [Hum96j]. Workgroups [HKM96, Nan96f, BD95b, Won98a, Won98c]. working [BM95, Nad95c, BM95]. WorkPerfect [Ano95-160]. Workplace [Gil98b]. Works [Ano95-37, Sha96, Ano95l, Ano95-55, Ano95-81, Ano96-127, Api95b, Dej95h, Pou95g, Sta96b, Way96b]. Workshop [Die95c]. Workstation [Ano96-73, Ano97-110, Ano95-100, Ano95-145, Ano96-178, Ban97, EP97, Fau95, Tam95d, Way96m]. workstation-based [Ano95-100]. Workstations [AG95, BHKM96b, Cam98a, Hum96b, Ano95-60, Ano95-186, Ano95-190, Ano97-115, Ano98-51, Cro95, Kay98b, Kay98c, Tan98]. World [Ano98-78, Ano98-84, Em98, Hal97b, Mia97c, Tho96q, Tho97b, Ude95c, Ano95-45, Dej95f, Fri97e, Hal97a, Hal95d, Mon97, Nee95m, Pou97e, Sta97b, Way96g, Ano95-95, Vac95a]. Worlds [Gre97l, Gre98f, Joc96b, Tho95l]. Worried [Fow97, Ude98b]. worst [Ude97k]. Worth [Ano98-74, Log96, Die95b, Dob96, Hum98b, VN97a]. worthy [Gre97l, Hal95a]. would [Ano95y, Ano95-61, Ano95-141, Ano95-154, Ano95-187]. wrap [Ano95-45]. wraparound [Hal95d]. wrapped [Ano95-163, Mia95c, QS96]. wring [Rud98, Won98a]. Write [Ano95-73, Ano96-27, Gre97f, Gre97l, Mon95a, Ano95-111, Ano95-112, Ano96u, Ano96-77, Ano96-78, Ano96-127, Ano96-134, Ano97t, Lin95a, Tho95c]. Writing [Fri98b, Tho98e, Ano95-190, TC97]. Wrong [Ude95g, DFT95b, Gre95b, Gre95c, Pot95]. Wrote [Var96]. Wuzzatdoonear [Ano95-187]. WWW [Fri95a].

X [Bak95a, Bak95b, Bou95, IF97, Pou95m]. X.500 [Ano95-221]. x8 [Ano97-82]. x86 [Ano95-44, Ano95-127, Ano95-128, Ano97q, Ano97-36, Ano97-73, Chi96a, Gwe95, Hal96j, Hal96k]. Xbase [Gre95e]. xenophobic [Bon95a]. Xerox [Ano95-204, Ano95-209, Ano97-62, Nad95c]. XML [Ano98-127]. XTerminal [Way96m].

yard [CFDT95]. Yeah [Lin95a]. Year [Ano96-183, Ano96-127, Bar95c, HM95, Hal96k, Sch96e, Ude95e, Ano98s, Ano98-95]. years [Ano95-48, Ano95-60, Ano95-192, Hal98a, MB97, Pou95b]. Yes [Hal97b, Pou95n]. yesterday [Hal95c]. yielding [Dej95g]. Yields [Ano95i, Ano95j]. You’ll [Fou95b, Lin95a, Ano95-95, Fou95a, Hal96k, Wat97b]. young [Smi96c]. You’re [Ano95-222, Nee96f, Ano95-95, Flo95]. yours [Hal98g]. Yourself [HM96, Ude96d, Ano98-115]. you’ve [Ano95-122, Ano95c, Kay97c].

zap [Ano96-179]. Zen [Phi97]. Zero [CC97, Hur97a, Tho95k]. zillion [Ano96-55]. zip [Ano96-170, Pou96l]. zippy [Ano96-40, Ano96-175]. Zoo [Gou95]. Zoom [Ano96-184, Lov95].

References

Abrahams:1997:ARCb

[Abr97a] Marc Abrahams. Advances and retreats in computing — revenge on junk e-mailers, part 2; plus, exacerbating the bandwidth crunch. Byte Magazine, 22(7):160–??, July 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Abrahams:1997:ARCa

[Abr97b] Marc Abrahams. Advances and retreats in computing — the amazing new Content Analyzer; the fab combo of chips and snacks; and how plug and play has saved lives. Byte Magazine, 22(5):160–??, May 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Abrahams:1997:ARCc

[Abr97c] Marc Abrahams. Advances and retreats in computing — the programmer’s best friend and Cyber BabeWatch. Byte Magazine, 22(8):168–??, August 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Abrahams:1998:CDD

[Abr98a] Marc Abrahams. The click of doom: A delete key that could save civilization and sanity. Byte Magazine, 23(5):160, May 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Abrahams:1998:ELH

[Abr98b] Marc Abrahams. An experimental lecture hall and software for the next celebrity requiem. Byte Magazine, 23 (6):160–??, June 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Abrahams:1998:MUS

[Abr98c]  Marc Abrahams.  A mouse

that uses satellite technology, and lessons in high tolerance. Byte Magazine, 23(7):144–??, July 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Aertker:1995:BCB

[Aer95] Rowland Aertker. Books and CD-ROMs: Big blue: An insider’s view — an insider’s look at IBM’s history, and information technology reports on CD-ROM. Byte Magazine, 20(8):45– ??, August 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1995:FNW

[AG95] Steve Apiki and Rick Grehan.

Fastest NT workstations: Windows NT 3.5 opens up your options for choosing a high-end Windows system. BYTE reviews seven NT speed demons — comparing Alpha, Mips 4400, and Pentium workstations. Byte Magazine, 20(3):115–120, 122, March 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Andrews:1998:PIG

[AK98] Dave Andrews and Jason Krause. More PC power to go: Intel’s new mobile processor brings Pentium II power to portable PCs. Byte Magazine, 23(5):24, May 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Alves:1996:PCS

[Alv96] Alexandre Alves. Parallel computing — Windows style. Byte Magazine, 21(5):169–??, May 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Andrews:1995:FFA

[And95a] D. Andrews. Not flashier, but faster (AT&T Globalyst 250P notebook). Byte Magazine, 20 (6):173, June 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Andrews:1995:FFG

[And95b] Dave Andrews. Not flashier, but faster: The Globalist 250P — good for heavy-duty number crunching. Byte Magazine, 20(6):175–??, June 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Andrews:1995:PPS

[And95c] Dave Andrews. Pentium power without sticker shock: The eurocom — lots of features but short battery life. Byte Magazine, 20(6):173–??, June 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anderson:1996:FCF

[And96a] Doug Anderson. Fibre channel:

Fast and flexible. Byte Magazine, 21(5):173–??, May 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Andrews:1996:TPY

[And96b] D. Andrews. ThinkPad proves you can be too thin. Byte Magazine, 21(8):47, August 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Andrews:1996:TGT

[And96c] Dave Andrews. ThinkPad gets thin. Byte Magazine, 21(8):47– ??, August 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Andrews:1996:WPW

[And96d] Dave Andrews. Wearable Pentium — what Claudia Schiffer might wear while troubleshooting a nuclear reactor. Byte Magazine, 21(9):228–??, September 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Andrews:1996:JFB

[And96e] David L. Andrews. Java forms: Better than HTML. Byte Magazine, 21(11):38, November 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Discusses the

JetForm (http://www.jetform.
com/) and Caere (http://www.

caere.com/) products for interactive forms applications.

Andrews:1997:WNJ

[And97] David L. Andrews. What’s next for Java office apps. Byte Magazine, 22(11):24–??, November 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Andrews:1998:JPJ

[And98] Dave Andrews. Java poison or just PR? Byte Magazine, 23 (1):22–23, January 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:D

[Ano95a] Anonymous. 120-MB disks. Byte Magazine, 20(10):62–??, October 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:MMS

[Ano95b] Anonymous. 15 MB in a matchbook: The storage capacities of personal communications devices, pagers, and digital cameras could improve greatly in 1995 thanks to a 32-Mb new flash memory device called Compact-Flash from SunDisk. Byte Magazine, 20(1):30–??, January 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:PHW

[Ano95c] Anonymous. 16 Pentiums high on Win 95 — if you’ve upgraded to Windows 95, you might want one of these speed demons to go along with it. we test the latest 120- and 133-MHz Pentium PCs. Byte Magazine, 20(12): 156–??, December 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:RCA

[Ano95d] Anonymous. 1995 readers’ choice awards: BYTE readers anoint the best hardware and software on the market today. Byte Magazine, 20(5):75–??, May 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:ALA

[Ano95e] Anonymous. 3-D action — A look at the making of two slick games for the Mac: Dark forces and Marathon. Byte Magazine, 20(12):123–??, December 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:ALW

[Ano95f] Anonymous. 3-D action — the latest DOS and Windows shoot’em-up games show how far we’ve come since Battlezone. Byte Magazine, 20(12):123–??, December 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:IFA

[Ano95g] Anonymous. 3-D images that float in air — A new system can project images from a PC so that they appear to be solid objects suspended in air. Byte Magazine, 20(10):48–??, October 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:WGY

[Ano95h] Anonymous. 3.5 will get you 100 — two technologies are vying to become the next 3.5-inch floppy drive standard for holding 100 MB or more. Byte Magazine, 20(8):26–??, August 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:YBSa

[Ano95i] Anonymous. 6X yields better software MPEG and networking — move over quad speed, a new CD-ROM performance king has arrived. Byte Magazine, 20(7): 40–??, July 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:YBSb

[Ano95j] Anonymous. 6X yields better software MPEG and networking: Move over quad speed, a new CD-ROM performance king has arrived. Byte Magazine, 20(7):

40–??, July 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:AAD

[Ano95k] Anonymous. Access 95 advances database design — combine the Windows 95 Briefcase with Access 95’s powerful replication features and you can take your corporate databases to the remotest places. Byte Magazine, 20(12): 181–??, December 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:ACA

[Ano95l] Anonymous. Agents of change:

Agents and smart software are still works in progress. security and interoperability define the leading edge of development for these industries software tools. Byte Magazine, 20 (3):94–??, March 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:ASA

[Ano95m] Anonymous. Alpha stays ahead.

Byte Magazine, 20(11):210–??, November 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:AV

[Ano95n] Anonymous. Alternate views of the 615. Byte Magazine, 20(11): 200–??, November 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:AWB

[Ano95o] Anonymous. Anatomy of the Web BOMB. Byte Magazine, 20 (10):106–??, October 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:OA

[Ano95p] Anonymous. And one for all. Byte Magazine, 20(11):131–??, November 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:ANM

[Ano95q] Anonymous. Apple’s new multimedia Macs — at Boston Mac-World Expo this august, Apple is expected to introduce a new line of Power Macs. Byte Magazine, 20(8):26–??, August 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:ATPa

[Ano95r] Anonymous. Apple’s Tsunami: PCI power — BYTE takes a close look at the first PCI-based Mac. Byte Magazine, 20(7):26– ??, July 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:ATPb

[Ano95s] Anonymous. Apple’s Tsunami: PCI power: BYTE takes a close look at the first PCI-based Mac. Byte Magazine, 20(7):26– ??, July 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:A

[Ano95t] Anonymous. Archistrat 4s. Byte Magazine, 20(10):60–??, October 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:ALS

[Ano95u] Anonymous. Assets on the line Salvatore/Salamone. Byte Magazine, 20(9):37–??, September 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:AM

[Ano95v] Anonymous. ATX motherboards. Byte Magazine, 20 (10):54–??, October 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:BRG

[Ano95w] Anonymous. Bank robbers go electronic — Citibank admits it’s been robbed, but the alleged crooks use PCs, not shotguns. Byte Magazine, 20(11): 48–??, November 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:BPL

[Ano95x] Anonymous. BBSes to provide local Web access — BBS software vendors are bringing Web capabilities to their products. Byte Magazine, 20(8):30–??, August 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:BVPa

[Ano95y] Anonymous. The better virtual PC — if you need to run Windows applications but would rather use a Power Mac, Soft-Windows 2.0 will let you do it. Byte Magazine, 20(12):195–??, December 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:BVPb

[Ano95z] Anonymous. The better virtual PC — Insignia Solutions’ Soft-Windows 2.0 lets you run DOS and Windows applications on your Power Mac in a networked PC environment. Byte Magazine, 20(12):195–??, December 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:BVPc

[Ano95-27] Anonymous. The better virtual PC — SoftWindows 2.0 not only puts a 486 inside a Power Mac; it gives DOS and Windows programs access to the Mac’s Ethernet port. Byte Magazine, 20(12): 195–??, December 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:BPHa

[Ano95-28] Anonymous. Blasts from the past — highlights from two decades of covering the PC revolution. Byte Magazine, 20(7): 41–??, July 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:BPHc

[Ano95-29] Anonymous. Blasts from the past: Highlights from two decades of covering the PC revolution. Byte Magazine, 20(1):41– ??, January 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:BPHd

[Ano95-30] Anonymous. Blasts from the past: Highlights from two decades of covering the PC revolution. Byte Magazine, 20(2): 33–??, February 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:BPHe

[Ano95-31] Anonymous. Blasts from the past: Highlights from two decades of covering the PC revolution. Byte Magazine, 20(3):41– ??, March 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:BPHf

[Ano95-32] Anonymous. Blasts from the past: Highlights from two decades of covering the PC revolution. Byte Magazine, 20(4):33– ??, April 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:BPHg

[Ano95-33] Anonymous. Blasts from the past: Highlights from two decades of covering the PC revolution. Byte Magazine, 20(5):

41–??, May 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:BPHb

[Ano95-34] Anonymous. Blasts from the past: Highlights from two decades of covering the PC revolution. Byte Magazine, 20(7): 41–??, July 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:BTD

[Ano95-35] Anonymous. Blasts from two decades of covering the PC revolution. Byte Magazine, 20(6):41– ??, June 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:BLB

[Ano95-36] Anonymous. Blue laser, bright future — recent breakthroughs in blue laser diodes and blue LEDs portend higher-capacity CD-ROM discs and brighter projection displays. Byte Magazine, 20(8):34–??, August 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:BCH

[Ano95-37] Anonymous. Books and CD-ROMs: How Microsoft works — A wealth of detail on how Microsoft operates; plus, a CD-ROM satire on bad art. Byte Magazine, 20(12):41–??, December 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:BCC

[Ano95-38] Anonymous. Brooktrout cuts the cost of internal faxing: A new device from the Brooktrout Networks Group could save you a lot of money on intracompany fax phone calls. Byte Magazine, 20 (2):32–??, February 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:BGM

[Ano95-39] Anonymous. BUYER’s GUIDE — mail order. hardware/software showcase. buyer’s mart. Byte Magazine, 20(10):201–??, October 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:BNP

[Ano95-40] Anonymous. The BYTE network project: Perl magic — have you visited BYTE’s Virtual Press Room and wondered how we built it? Byte Magazine, 20(12):115–??, December 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:BPC

[Ano95-41] Anonymous. Byte picks CeBIT’s best: BYTE editors scoured the CeBIT show for the best in new international products. Byte Magazine, 20(5):34–??, May 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:BBB

[Ano95-42] Anonymous. BYTEmark bug bashed — an update to our tests of the P6 chip. Byte Magazine, 20(12):25–??, December 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:CMC

[Ano95-43] Anonymous. A checklist for making CTI decisions. Byte Magazine, 20(9):206–??, September 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:CWD

[Ano95-44] Anonymous. Chip wars drive down PC prices: Four companies are scrambling to put the fastest and lest expensive x86 chip in you PC. Byte Magazine, 20(6): 24–??, June 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:CTN

[Ano95-45] Anonymous. Code Talk: New Micro Focus tool converts COBOL into components — organizations with beaucoup lines of COBOL code can now wrap it all in OLE controls and bring it into the world of Windows. Byte Magazine, 20(12):36–??, December 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:CNM

[Ano95-46] Anonymous. CodeTalk: New Micro Focus tool converts COBOL into components — something old — COBOL — meets something new — Windows 95–in Micro Focus’s Visual Object. Byte Magazine, 20(12):36–??, December 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:CWC

[Ano95-47] Anonymous. CodeTalk: Watcom C/C++ gets visual — the new version of one of Rick Grehan’s favorite compilers offers a much more visual development environment. Byte Magazine, 20 (11):47–??, November 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:CED

[Ano95-48] Anonymous. Color encroaches on the desktop: Over the next several years, technological advances will accelerate the use of color printers. Byte Magazine, 20(6):28–??, June 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:CLF

[Ano95-49] Anonymous. Color lasers:

Faster, easier, cheaper — A second generation of color laser printers is due soon. Byte Magazine, 20(8):40–??, August 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:CBM

[Ano95-50] Anonymous. Coming A better multimedia platform — several multimedia components for Win 95 won’t be available for months. Byte Magazine, 20 (10):21–??, October 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:CCP

[Ano95-51] Anonymous. Companies cut phone bills: Companies are starting to use the Internet as a pseudoprivate backbone to carry fax and voice traffic. Byte Magazine, 20(5):36–??, May 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:CRC

[Ano95-52] Anonymous. Compaq reduces cost of ownership: Compaq Computer is making it easier for LAN management software vendors with its new line of Deskpro PCs. Byte Magazine, 20(5):30– ??, May 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:CAW

[Ano95-53] Anonymous. Convenience, but at what price? Byte Magazine, 20 (9):229–??, September 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:CP

[Ano95-54] Anonymous. Crashing the party. Byte Magazine, 20(11):137–??, November 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:DNDa

[Ano95-55] Anonymous. Data to the nth dimension — this drill-down info analyzer works with a variety of DOS and Windows front ends, including Excel and 1-2-3. Byte Magazine, 20(12):193–??, December 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:DNDb

[Ano95-56] Anonymous. Data to the nth dimension — you can soup-up Excel for the Mac with the highend data-analysis and development capabilities of Arbor’s Essbase Analysis Server 3.2. Byte Magazine, 20(12):193–??, December 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:DCT

[Ano95-57] Anonymous. Data-warehouse construction tips. Byte Magazine, 20(10):102–??, October 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:DVT

[Ano95-58] Anonymous. Delphi and VB turn 32. Byte Magazine, 20(9):26–??, September 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:DDP

[Ano95-59] Anonymous. Desktop databases pack more power: PC database vendors are touting their products’ ability to scale from the desktop to the enterprise. but vendors’ definitions of scalability vary widely. Byte Magazine, 20(5):22–??, May 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:DCW

[Ano95-60] Anonymous. Developers catch the 3-D wave: For the last five years, computer users who demanded sophisticated 3-D graphics turned to high-end workstations. now the computer industry is preparing to bring this level of performance to low-end desktop systems. Byte Magazine, 20(3):34–??, March 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:DWF

[Ano95-61] Anonymous. Dialects of the Web: A future version of HTML (Hypertext Markup Language) could ensure that simple and complex Web documents can be read by any Web user. in this vision, the Web will use an objectoriented model: the core classes — basic headings, paragraphs, and links — will be understood by all browsers, but richer “subclass” distinctions would be allowed for use by more sophisticated browsers. Byte Magazine, 20(3):30–??, March 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:DGU

[Ano95-62] Anonymous. DragNET — Gmen use a network of databases and POTS to track down public enemies. Byte Magazine, 20(12): 106–??, December 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:EMH

[Ano95-63] Anonymous. End migration headaches — upgrading from NetWare 3.x to NetWare 4.x can be less painful with these NDS migration utilities. Byte Magazine, 20(11):30–??, November 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:EWD

[Ano95-64] Anonymous. End-User Windows databases take off: BYTE looks at three Windows database offerings — Microsoft Access 2.0, Lotus Approach 3.0, and Alpha Software’s Alpha Fiver for Windows — that claim to be the best end-user database available. Byte Magazine, 20(1):28– ??, January 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:FPF

[Ano95-65] Anonymous. Factoring in publickey’s future. Byte Magazine, 20 (10):78–??, October 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:FGC

[Ano95-66] Anonymous. Faster graphics cards on the horizon: Better graphic performance thanks to new kinds of graphics memory. Byte Magazine, 20(4):24– ??, April 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:FSR

[Ano95-67] Anonymous. File servers with RAID. Byte Magazine, 20(9): 252–??, September 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:FWB

[Ano95-68] Anonymous. Filling in Windows blanks. Byte Magazine, 20(11): 103–??, November 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:FML

[Ano95-69] Anonymous. Flash memory looks bright. Byte Magazine, 20(6):188–??, June 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:FDS

[Ano95-70] Anonymous. Flat displays squeeze bulky CRTs — big screens, small space-coming soon to a desktop near you. Byte Magazine, 20(11):38–??, November 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:FRD

[Ano95-71] Anonymous. Free rides are disappearing: Features of the Internet that users currently use for free may cost them in the future. Byte Magazine, 20(4):26– ??, April 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:GCR

[Ano95-72] Anonymous. Ga cote, Raymond:

With its quick object-oriented Pascal compiler, extensible development environment, and integrated database access, Borland’s Delphi is a powerful base for prototyping and developing applications. Byte Magazine, 20(5):133–??, May 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:GPW

[Ano95-73] Anonymous. The games people write — Microsoft’s new Game SDK takes some horror out of designing for Windows 95. Byte Magazine, 20(12):135–??, December 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:GCA

[Ano95-74] Anonymous. Getting clear about fuzzy: What do pension plans, elevators, and battery charges have in common? each is an application of fuzzy logic. Byte Magazine, 20(6):34– ??, June 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:GTI

[Ano95-75] Anonymous. Groupware taps the Internet — the planned merger between Netscape and Collabra will result in a Notes-killer of an application, some say. Byte Magazine, 20(12):24–??, December 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:HH

[Ano95-76] Anonymous. Helpful hints. Byte Magazine, 20(9):259–??, September 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:HPT

[Ano95-77] Anonymous. High-End portables take off: High-end notebook computers are accounting for an increasing percentage of portables sales thanks in part to a higher demand by businesses for desktop PC replacements. here’s what to expect from new product introductions in 1995. Byte Magazine, 20(2):26–??, February 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:HM

[Ano95-78] Anonymous. Honorable mentions. Byte Magazine, 20(9): 259–??, September 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:HMD

[Ano95-79] Anonymous. Honorable mentions/dubious achievement. Byte Magazine, 20(11):244–??, November 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:HCT

[Ano95-80] Anonymous. Hot chips, tough choices — A new wave of multimedia processors will rely heavily on software support. Byte Magazine, 20(8):25–??, August 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:HEC

[Ano95-81] Anonymous. How error corrections works. Byte Magazine, 20 (9):250–??, September 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:HBIa

[Ano95-82] Anonymous. How to build an Internet app — Brett Glass explains how to take a bit of Visual Basic code, some knowledge of Internet protocols, and build an on-line application — during your lunch break. Byte Magazine, 20(12):211–??, December 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:HBIb

[Ano95-83] Anonymous. How to build an Internet app — thanks to the net’s protocol suite, it’s not as hard as you might think. Byte Magazine, 20(12):211–??, December 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:HWTb

[Ano95-84] Anonymous. How we tested. Byte Magazine, 20(9):254–??, September 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:HWTa

[Ano95-85] Anonymous. How we tested. Byte Magazine, 20(11):245–??, November 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:IAI

[Ano95-86] Anonymous. Index to advertisers by inquiry reply cards. Byte Magazine, 20(10):96A–??, October 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:IAP

[Ano95-87] Anonymous. Index to advertisers by product category. Byte Magazine, 20(10):246–??, October 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:IGS

[Ano95-88] Anonymous. Indexing gets smart: Information publishers are turning to a new breed of smart indexing tools that can automatically summarize and condense huge documents without human intervention. Byte Magazine, 20(1):36–??, January 1995.

CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:IGR

[Ano95-89] Anonymous. Infrared gets real:

The Infrared Data Association’s new standard has vendors getting on board the infrared bandwagon. Byte Magazine, 20(1):30– ??, January 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:IRC

[Ano95-90] Anonymous. Inquiry reply cards.

Byte Magazine, 20(11):122–??, November 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:IT

[Ano95-91] Anonymous. Integrated telephony. Byte Magazine, 20(10): 64–??, October 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:IMA

[Ano95-92] Anonymous. Intel and Microsoft:

The agents of change. Byte Magazine, 20(10):58–??, October 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:IRR

[Ano95-93] Anonymous. Intel’s rivals ready to exploit P6 weakness — Intel’s rivals say their next chips will be better than the P6 running 16bit software. Byte Magazine, 20 (10):26–??, October 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:IMV

[Ano95-94] Anonymous. Interactive music videos arrive for Macs and PCs. Byte Magazine, 20(9): 32–??, September 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:IPT

[Ano95-95] Anonymous. Internet publishing tools proliferate: The best World Wide Web publishing tool for you, whether it’s a word processor add-on or a relational database/SGML hybrid publishing tool, depends in part on how often you’ll need to update the information that you’re publishing on the Internet. Byte Magazine, 20(3):30–??, March 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:ICL

[Ano95-96] Anonymous. Is cache losing its cachet? Byte Magazine, 20 (6):191–??, June 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:KDM

[Ano95-97] Anonymous. Keep the data moving — handy Intel-based computers and PC cards are helping mobile data collectors who work in rugged environments. plus, tips on buying a hand-held computer. Byte Magazine, 20 (12):87–??, December 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:LRNa

[Ano95-98] Anonymous. Lab report: 12 network printers set the pace — we test networked laser printers with fast engines, built-in Ethernet, large paper capacities, long life, and security features. Byte Magazine, 20(11):238–??, November 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:LRF

[Ano95-99] Anonymous. Lab report: 16 fast, reliable RAID subsystems. Byte Magazine, 20(9):248–??, September 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:LRSa

[Ano95-100] Anonymous. Lab report: 26 safeguards against LAN data loss: BYTE/NSTL selects the best single-medium tape drives for backing up midrange networks. we tested tape subsystems in a range of environments, including server-based, workstation-based, and Windows NT configurations. 4-mm DAT drives; QIC drives; 8-mm videocassett drives; DLT drives; honorable mentions; dubious achievement. Byte Magazine, 20(3):144–??, March 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:LRP

[Ano95-101] Anonymous. Lab report: 29 printers that mean business: Acquire speed and pizzazz with printer from this month’s lab report. we test color ink-jets, 6-to 10ppm laser printers, 11-to 30ppm lasers for networks, and a few select color lasers. 11- to 30-PPM laser printers; honorable mentions; 6- to 10-PPM laser printers; color ink-jet printers; color laser printers. Byte Magazine, 20(5):154–??, May 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:LRSb

[Ano95-102] Anonymous. Lab report: 29 switching hubs save the bandwidth. Byte Magazine, 20 (7):162–??, July 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:LRSc

[Ano95-103] Anonymous. Lab report: 29 switching hubs save the bandwidth: Looking for more bandwidth? our BYTE/NSTL lab rport describes how we tested 29 Ethernet switching hubs. we provide the data and the analysis — you make the choice. storeand-forwards; cut-throughs; how we tested; hub glossary; honorable mentions; dubious achievements. Byte Magazine, 20 (7):162–??, July 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:LRNb

[Ano95-104] Anonymous. Lab report: 30 No-Compromise notebooks: We run 30 active-matrix, 3.3 volt 486DX4 and Pentium notebooks through extensive benchmarks and choose the best Windows and DOS portables for power users. we also report on Apple’s PowerBook 540c and compare the performance of AST’s 66-MHz Ascentia 810N to the 75 MHz 486DX4 notebooks. best 386DX4-based notebooks; how we tested; best Pentium-based notebooks. honorable mentions; dubious achievements. Byte Magazine, 20(4):194–??, April 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:LRV

[Ano95-105] Anonymous. Lab report: 32 V.34 modems answer the call: V.34 represents the best and fastest modem technology available today. NSTL evaluates 32 state-of-the-market modems for throughout, reliability, robustness, and compatibility. external V.34 modems; how we tested; pick a PC card; honorable mentions. Byte Magazine, 20(6):224– ??, June 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:LRB

[Ano95-106] Anonymous. Lab report: 62 big, beautiful color monitors: BYTE/NSTL searches for the best 17-, 20-, and 21-inch monitors, using an extensive battery of tests to measure image quality, sharpness, contrast, convergence, legibility, power consumption, and distortion. we also evaluate ease of setup, controls, power management, and documentation. the best 17-inch monitors; how we tested; the best 20-and 21-inch monitors; honorable mentions; dubious achievements. Byte Magazine, 20(1): 204–??, January 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:LRT

[Ano95-107] Anonymous. Lab report: True-Color graphics accelerators: PCI cards are finally delivering on earlier promises. our custom tests find the fastest PCI graphics accelerators under Windows, as well as some hot new Macintosh cards. the best graphics accelerators for general business; graphics glossary; video tests; the best PCI graphics adapters for drafting; how we tested; the best graphics accelerators for Macintosh NuBus systems; video glossary; honorable mentions. Byte Magazine, 20(2): 136–??, February 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:LD

[Ano95-108] Anonymous. Learn by doing. Byte Magazine, 20(11): 72–??, November 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:LLM

[Ano95-109] Anonymous. Legal land mines cloud multimedia: Legal problems lurk behind multimedia development. Byte Magazine, 20 (4):31–??, April 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:LHH

[Ano95-110] Anonymous. Letters — huzzahs for heuristics, plus comments on mobile computing. Byte Magazine, 20(8):18–??, August 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:LRWa

[Ano95-111] Anonymous. Letters — readers write about our anniversary edition, OS/2, RAD, the BYTE Web site, and more. Byte Magazine, 20(11):20–??, November 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:LRWb

[Ano95-112] Anonymous. Letters — readers write about our coverage of Windows 95, Microsoft, OS/2, and RAD; an Amiga update; and more. Byte Magazine, 20(10):20– ??, October 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:LFL

[Ano95-113] Anonymous. Letters: February letters cover platform debates and programming issues. Byte Magazine, 20(2):18–??, February 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:LIO

[Ano95-114] Anonymous. Letters: Information overload and Apple’s Newton MessagePad. Byte Magazine, 20(4):18–??, April 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:LLI

[Ano95-115] Anonymous. Letters: Letter on Intel’s P6, Microsoft’s Bob, and personal quandaries. Byte Magazine, 20(6):18–??, June 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:LRD

[Ano95-116] Anonymous. Letters: Readers describe solutions using agent technology and express frustration with TCP/IP addressing. Byte Magazine, 20(1):18–??, January 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:LML

[Ano95-117] Anonymous. Letters: This month’s letters discuss computers in education, Internet attitudes, and SGML, as well as clarify results in our lab report on tape drives. Byte Magazine, 20(5):17–??, May 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:LDO

[Ano95-118] Anonymous. LexMark delivers outstanding resolution: Lexmark’s new 1200-dpi printer delivers excellent quality when printing gray-scale images, such as photographs. but when printing standard text, the differences between 1200 and 600 dpi are not as noticeable. Byte Magazine, 20 (2):23–??, February 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:LSG

[Ano95-119] Anonymous. Low-End servers grow up: Features associated with high-end PC LAN servers are migrating to the low end. Byte Magazine, 20(4):23–??, April 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:MCF

[Ano95-120] Anonymous. Mac clones — finally: The first Applesanctioned Mac OS-compatible systems have finally arrived. Byte Magazine, 20(4):22–??, April 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:MIC

[Ano95-121] Anonymous. Making the Internet connection. Byte Magazine, 20(1):126–??, January 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:MPR

[Ano95-122] Anonymous. Memory price relief to come in late 1996 — if you’ve been holding out for lower prices before buying more memory, that hungry little beastie inside your Windows PC will have to wait a while for its next feeding. Byte Magazine, 20(12): 34–??, December 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:MFN

[Ano95-123] Anonymous. Microsoft furthers NetWare-to-NT transition: Microsoft’s file and print service for NetWare eases the transition to Windows NT server. Byte Magazine, 20(4):26–??, April 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:MNG

[Ano95-124] Anonymous. MIPS’ next generation: R10000. Byte Magazine, 20(11):190, November 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:MRW

[Ano95-125] Anonymous. More RAM for Win 3 holdouts — if you like Windows 3.x but not its memory confusion, maybe it’s time to buy one of these RAM doublers. Byte Magazine, 20(12):

191–??, December 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:MDI

[Ano95-126] Anonymous. MSN: Desktop Internet. Byte Magazine, 20(9): 231–??, September 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:MXCa

[Ano95-127] Anonymous. Multimedia x86 CPUs coming in 1996 — new x86 microprocessors will integrate DSP functions. Byte Magazine, 20(12):30–??, December 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:MXCb

[Ano95-128] Anonymous. Multimedia x86 CPUs coming in 1996 — thanks to upcoming chips with integrated DSP functions, your next PC could be a real smoker when it comes to running mixed-media applications. Byte Magazine, 20 (12):30–??, December 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:NNB

[Ano95-129] Anonymous. The net’s next big thing: Virtual reality: VRML (Virtual Reaity Markup Language) could make navigating on-line a snap. Byte Magazine, 20(4):28–??, April 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:NS

[Ano95-130] Anonymous. Network-printer-management software. Byte Magazine, 20(11):244–??, November 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:NCD

[Ano95-131] Anonymous. New 486 chips deliver inexpensive power. Byte Magazine, 20(9):30–??, September 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:NPS

[Ano95-132] Anonymous. New PowerPC standard supports Macs: While still in its early stages, the new CHRP (Common Hardware Reference Platform) standard promises to let PowerPC systems freely use different operating systems. Byte Magazine, 20(3):24–??, March 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:NROa

[Ano95-133] Anonymous. New replication options in Access, Oracle, and Notes — Access for Windows 95 acquires a compelling feature. Byte Magazine, 20(10):30– ??, October 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:NSEa

[Ano95-134] Anonymous. New suites embrace the Web — software houses are adding much more to their Windows communications programs than just data comm and fax. Byte Magazine, 20(12): 25–??, December 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:NSEb

[Ano95-135] Anonymous. New suites embrace the Web — terminal emulation is dead. well, not quite. but vendors are adding much more to Windows communications programs than just data comm and fax. Byte Magazine, 20(12): 25–??, December 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:NOW

[Ano95-136] Anonymous. The next On-Line wave: On-line services are responding to the enormous demand of users for Internet accessibility by incorporating Internet access and tools. Byte Magazine, 20(1):40–??, January 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:JAF

[Ano95-137] Anonymous. Not just another free Unix — FreeBSD can be used for everything from commercial Internet service to a home-desktop solution, running on relatively inexpensive PC equipment. Byte Magazine, 20(12):207–??, December 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:

[Ano95-138] Anonymous. Not that DOS. Byte Magazine, 20(11):113–??, November 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:NHDa

[Ano95-139] Anonymous. Notebook hard drives hold more. Byte Magazine, 20(7):25–??, July 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:NHDb

[Ano95-140] Anonymous. Notebook hard drives hold more: Gigabyte drives for notebooks will be standard this fall. Byte Magazine, 20(7):25–??, July 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:NNC

[Ano95-141] Anonymous. Notebooks, NT clusters capture awards: BYTE editors searched throughout the convention halls of Fall Comdex 1994 for innovative new products that had been publicly announced within 30 days of the show and that would have a strong influence on business computing. Byte Magazine, 20(2): 24–??, February 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:NMIa

[Ano95-142] Anonymous. Notes meets the Internet — new products and services that link Lotus Notes to the Internet are expanding the uses of Lotus’s groupware platform. Byte Magazine, 20(7):26– ??, July 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:NMIb

[Ano95-143] Anonymous. Notes meets the Internet: New products and services that link Lotus Notes to the Internet are expanding the uses of Lotus’s groupware platform. Byte Magazine, 20(7):26– ??, July 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:NNI

[Ano95-144] Anonymous. NSTL’s new InterMark suite. Byte Magazine, 20(3):80–??, March 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:NROb

[Ano95-145] Anonymous. NT roars on the 604 — our tests of new systems from Austin Direct, IBM, and Motorola show that the PowerPC 604 chip makes for a mean Windows NT workstation. Byte Magazine, 20(11):209–??, November 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:OFS

[Ano95-146] Anonymous. OCR fuels slow paper burn: OCR is getting easier, and who knows, maybe the elusive paperless office will one day become reality. naaaaah.

Byte Magazine, 20(5):40–??, May 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:OMM

[Ano95-147] Anonymous. One machine, many RAMs. Byte Magazine, 20 (6):192–??, June 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:OPI

[Ano95-148] Anonymous. One person’s Internet business. Byte Magazine, 20 (1):132–??, January 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:M

[Ano95-149] Anonymous. OO meets OS. Byte Magazine, 20(11):123–??, November 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:OSH

[Ano95-150] Anonymous. Operating systems — hot new developments in the software that runs the software. Byte Magazine, 20(11): 77–??, November 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:ORDa

[Ano95-151] Anonymous. Oracle releases 64-Bit database bullet — high-end network servers offer an alternative to mainframe systems when combined with new 64-bit relational database management software from Oracle. Byte Magazine, 20(7):30–??, July 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:ORDb

[Ano95-152] Anonymous. Oracle releases 64-Bit database bullet: High-end network servers offer an alternative to mainframe systems when combined with new 64-bit relational database management software from Oracle. Byte Magazine, 20(7):30–??, July 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:P

[Ano95-153] Anonymous. OS paradise. Byte Magazine, 20(11):81–??, November 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:PEP

[Ano95-154] Anonymous. P6 expands PC spectnum — thought the first P6 systems would be big servers, did you? many companies are actually preparing to ship desktop boxes first, some for less than USD4000. Byte Magazine, 20 (11):28–??, November 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:PWR

[Ano95-155] Anonymous. P6 weakness revealed. Byte Magazine, 20(9): 24–??, September 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:PPC

[Ano95-156] Anonymous. PC power comes to the calculator. Byte Magazine, 20(9):25–??, September 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:PSF

[Ano95-157] Anonymous. The penthouse suite — forget the Windows 95 hype for a while Office 95 really delivers: true 32-bit code, full OLE 2 support, better integration, and a great way to mix files from different applications. Byte Magazine, 20(12): 179–??, December 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:PDS

[Ano95-158] Anonymous. Pentium dream systems arrive: An early look at two of the first 120-MHz Pentium machines reveals more than just a fast processor inside the box. Byte Magazine, 20(5):23– ??, May 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:POM

[Ano95-159] Anonymous. Pentium Over-Drive = moderate upgrade: Intel’s Pentium OverDrive processor puts a rocket in your socket, but the upgrade has a 32-bit 486style I/O bus — not the 64-bit bus of a Pentium. Byte Magazine, 20(6):30–??, June 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:PSC

[Ano95-160] Anonymous. PerfectOffice a strong contender: WorkPerfect’s new applications suite is powerful. Byte Magazine, 20(4):30– ??, April 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:PLS

[Ano95-161] Anonymous. Phone lines stymie V.34 modems — that spiffy new V.34 modem may feel like a station wagon instead of a turbocharged racer. here’s why. Byte Magazine, 20(11):40–??, November 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:PTP

[Ano95-162] Anonymous. Piecing together puzzles: Finding and understanding patterns in data are key to creating mainstream applications for speechand handwriting-recognition or machine-vision systems. Byte Magazine, 20(2):82–??, February 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:PNM

[Ano95-163] Anonymous. Pournelle: A new mutation — Jerry gets a hesitation problem in his DOS word processor. to solve it, he decides to install his shrink-wrapped copy of Windows 95. adventure ensues. Byte Magazine, 20(12): 197–??, December 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:PTW

[Ano95-164] Anonymous. PowerPC tidal wave: Apple and IBM will make major PowerPC-related announcements this spring and summer. Byte Magazine, 20(6): 25–??, June 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:PVP

[Ano95-165] Anonymous. Professional video prices drop: Powerful desktop computers combined with new video editing programs are delivering professional video editing to the PC, Mac, and Power Mac platforms. Byte Magazine, 20(1):34–??, January 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:RDWa

[Ano95-166] Anonymous. Radical databases — Windows desktop databases, such as Microsoft’s FoxPro, must compete with rapid application development tools. Byte Magazine, 20(7):24–??, July 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:RDWb

[Ano95-167] Anonymous. RADical databses: Windows desktop databases, such as Microsoft’s FoxPro, must compete with rapid application development tools. Byte Magazine, 20(7):24–??, July 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:RCC

[Ano95-168] Anonymous. Radio comes to cyberspace — A new era in broadcasting begins on the Internet. Byte Magazine, 20(10):46– ??, October 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:RDI

[Ano95-169] Anonymous. Rewritable drive integrates two optical technologies: Phase change promised much as a rewritable medium, but high costs have kept it out of the mainstream. Matsushita hopes to change that. Byte Magazine, 20(2):28–??, February 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:RR

[Ano95-170] Anonymous. RFC 1597 revisited. Byte Magazine, 20(11): 203–??, November 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:RIS

[Ano95-171] Anonymous. RISC invades the school and home: Apple’s Power Mac 5200 beefs up the LC line for the school and home with a new PowerPC processor. Byte Magazine, 20(5):30–??, May 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:RC

[Ano95-172] Anonymous. Roll call. Byte Magazine, 20(11):246–??, November 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:REC

[Ano95-173] Anonymous. Run-time error checking comes to compilers — Borland, Microsoft, and Nu-Mega are. Byte Magazine, 20 (10):34–??, October 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:SPC

[Ano95-174] Anonymous. Service providers consolidate: The day of numerous, tiny Internet service providers may be over. Byte Magazine, 20(6):28–??, June 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:SNW

[Ano95-175] Anonymous. Seven new ways to learn. Byte Magazine, 20(3):54– ??, March 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:SSC

[Ano95-176] Anonymous. The shape of systems to come. Byte Magazine, 20(4):52–??, April 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:SF

[Ano95-177] Anonymous. Sincere OS flattery. Byte Magazine, 20(11): 119–??, November 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:SMP

[Ano95-178] Anonymous. Smarter, more powerful systems. Byte Magazine, 20(4):56–??, April 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:SO

[Ano95-179] Anonymous. Smooth operators. Byte Magazine, 20(11): 77–??, November 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:SRS

[Ano95-180] Anonymous. Software RAID solutions. Byte Magazine, 20(9): 256–??, September 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:SS

[Ano95-181] Anonymous. Starting from scratch. Byte Magazine, 20 (3):62–??, March 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:SWP

[Ano95-182] Anonymous. Still a work in progress — Windows 95 brings benefits to network managers, but some of the best stuff is yet to come. Byte Magazine, 20 (11):34–??, November 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:SIA

[Ano95-183] Anonymous. Strategic industry alliances. Byte Magazine, 20(9): 203–??, September 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:SCDa

[Ano95-184] Anonymous. Symantec C++ differences — the latest edition of Symantec’s C++ package adds network-distributable compiling and an object-oriented editor. version 7.2 ensures Windows compatibility via MFC 3.0, ODBC, and OLE 2. Byte Magazine, 20(12):183–??, December 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:SCDb

[Ano95-185] Anonymous. Symantec C++ differences — the new edition of Symantec’s C++ for the Macintosh supports the PowerPC. Byte Magazine, 20(12):183–??, December 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:SDO

[Ano95-186] Anonymous. Systems design in ObjecTime — with this modeling program, you can build an object-oriented system from the top down. ObjecTime runs on workstations from Sun and Hewlett–Packard and the IBM RS/6000. Byte Magazine, 20 (12):189–??, December 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:TMW

[Ano95-187] Anonymous. Talking to machines — “wuzzatdoonear?” your PC might not understand that question, but some speechrecognition products would recognize “what’s that doing here?”. Byte Magazine, 20(12): 97–??, December 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:TPa

[Ano95-188] Anonymous. Test: 16-to 20ppm printers. Byte Magazine, 20(11): 241–??, November 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:TPb

[Ano95-189] Anonymous. Test: 30 to 32-ppm printers. Byte Magazine, 20(11): 243–??, November 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:TCB

[Ano95-190] Anonymous. Tools cover big and little iron — new enterprise client/server programs provide tools for writing applications that span mainframes, workstations, and personal computers. Byte Magazine, 20(12): 26–??, December 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:UFE

[Ano95-191] Anonymous. Untangling fast Ethernet cables — this faster standard will run over a variety of wires. we help you sort them out. Byte Magazine, 20(12): 213–??, December 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:VRU

[Ano95-192] Anonymous. Vendors rally for 64-Bit Unix — the list of backers is impressive, but we might not see the results of a new 64bit Unix initiative for a couple years. Byte Magazine, 20(11): 47–??, November 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:VRB

[Ano95-193] Anonymous. Vendors ride bus to better telephony: A new universal serial bus that supports a maximum 12 mbps could become standard PC connector. Byte Magazine, 20(6):30– ??, June 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:VAH

[Ano95-194] Anonymous. Videoconferencing’s Achilles’ heels — videoconferencing won’t be used widely in business until two things change: interoperability improves and ISDN becomes more available. Byte Magazine, 20(8):24–??, August 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:VCL

[Ano95-195] Anonymous. Virtual CDs on the LAN — CD-QuickShare offers a cost-effective alternative to hanging gangs of CD-ROM drives off your network. Byte Magazine, 20(12):153–??, December 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:VPP

[Ano95-196] Anonymous. Visual Pascal with a punch: Borland’s Delphi unites a visual design environment with the industrial strength of Borland’s Pascal compiler and database connectivity engines. Byte Magazine, 20(2):22–??, February 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:VCI

[Ano95-197] Anonymous. VR can improve your health: Virtual reality is fundamentally altering and improving the way medical care is taught and delivered. Byte Magazine, 20(5):26–??, May 1995.

CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:WCM

[Ano95-198] Anonymous. Wavelets challenge MPEG, fractals — developers claim 300-to-1 compression. Byte Magazine, 20(12):34–??, December 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:WPT

[Ano95-199] Anonymous. We plugged, but they didn’t all play — tests of several new Plug and Play devices show that unless you have all the required pieces in your system, you can expect to do a lot of work. Byte Magazine, 20 (10):36–??, October 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:WPM

[Ano95-200] Anonymous. Web publishing made easier — A review of Cyberleaf, HotMetal Pro, HTML Assistant Pro, and Spider. Byte Magazine, 20(12):170–??, December 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:WNU

[Ano95-201] Anonymous. What Notes users want: As great as Lotus Notes is for deploying productivityenhancing applications throughout a company, it still suffers from numerous weaknesses. here’s what users want and a hint of things to come from Lotus. Byte Magazine, 20(1):24–??, January 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:WN

[Ano95-202] Anonymous. What’s new. Byte Magazine, 20(9):286–??, September 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:WNAa

[Ano95-203] Anonymous. What’s new — Apple’s PowerBook 5300c combines RISC processing power with great battery life; Intuit’s QuickBooks Pro manages your business finances; plus more. Byte Magazine, 20(10):190–??, October 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:WNAb

[Ano95-204] Anonymous. What’s new — AST Research’s Ascentia 950N notebook offers full-motion video and graphics; Xerox improves the OCR accuracy in TextBridge Professional Edition 3.0; View-Sonic’s 15GA and 17GA monitors include built-in speakers; and more. Byte Magazine, 20 (7):192–??, July 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:WNG

[Ano95-205] Anonymous. What’s new — Gateway 2000’s 133-MHz Pentium PC is a tower of multimedia power; PhoneKits turns your PC into a personal phone manager; plus more. Byte Magazine, 20(8):168–??, August 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:WNI

[Ano95-206] Anonymous. What’s new — intel’s newest OverDrive, which is faster and less expensive than the company’s first Pentium Over-Drive, is a bargain upgrade; askSam is. Byte Magazine, 20(11):262–??, November 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:WNPa

[Ano95-207] Anonymous. What’s new — previews of HP’s CopyJet color printer/copier and Delrina’s WinFax Pro for Windows 95. Byte Magazine, 20(12):216–??, December 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:WNPe

[Ano95-208] Anonymous. What’s new: A 100-MHz Pentium notebook; wireless infrared printers from Hewlett–Packard; the latest version of Dr. Solomon’s Anti-Virus toolkit; and more. Byte Magazine, 20(6):258–??, June 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:WNAc

[Ano95-209] Anonymous. What’s new: AST Research’s Ascentia 950N notebook offers full-motion video and graphics. Xerox improves the OCR accuracy in TextBridge professional edition 3.0; View-Sonic’s 15GA and 17GA monitors include built-in speakers; and more. Byte Magazine, 20 (7):192–??, July 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:WNC

[Ano95-210] Anonymous. What’s new: The CruisePad provides wireless control locally; Process Charter for Windows does smart flowcharting; WinDD puts Windows on Unix; and more. Byte Magazine, 20(2):180–??, February 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:WND

[Ano95-211] Anonymous. What’s new: The Doubleplay Series I doubles your PC storage; the JetEye ESI-9580A provides wireless printing; English Wizard translates English into SQL; Cruiser takes you down a virtual hallway; and more. Byte Magazine, 20(3):186– ??, March 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:WNPb

[Ano95-212] Anonymous. What’s new: The PageCard Wireless Messaging System keeps you up-to-date anywhere; PaperPort 2.0 for Windows converts documents into editable text; and more. Byte Magazine, 20(5):202–??, May 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:WNPc

[Ano95-213] Anonymous. What’s new: The PN60 prints color anywhere; CalcPac calculates in Windows; Duet puts a Mac voice in PCs; and more. Byte Magazine, 20 (1):240–??, January 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:WNPd

[Ano95-214] Anonymous. What’s new: The PowerLite 85 notebook offers Micro Sparc II performance, Limur includes Unix source code; DragStrip organizes your Mac programs; and more. Byte Magazine, 20(4):232–??, April 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:WMI

[Ano95-215] Anonymous. When money is plentiful. Byte Magazine, 20 (3):66–??, March 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:WWW

[Ano95-216] Anonymous. Why the Web won’t work. Byte Magazine, 20(11): 60–??, November 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:WOW

[Ano95-217] Anonymous. Wildfire: One wild and not-so-crazy helper. Byte Magazine, 20(9):216–??, September 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:WMP

[Ano95-218] Anonymous. Windows 95 migration planning kit. Byte Magazine, 20(7):52–??, July 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:WSB

[Ano95-219] Anonymous. Windows 95 sees better future: Microsoft says it will work to improve Window’s support for blind users. Byte Magazine, 20(6):36–??, June 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:WEC

[Ano95-220] Anonymous. WinSock 2 enhances connectiviy — the new WinSock promises to liberate network applications from dependency on a single transport protocol. Byte Magazine, 20 (8):30–??, August 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:XOF

[Ano95-221] Anonymous. X.500: The once and future answer. Byte Magazine, 20(11):52–??, November 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:YSM

[Ano95-222] Anonymous. You’re saving money when the meter’s running: Software metering programs, which were originally designed to enforce concurrent licensing agreements and prevent liability for inappropriate use of software, are now being used to cut software costs. Byte Magazine, 20(3):26–??, March 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:ECA

[Ano96a] Anonymous. 1995 editors’ choice awards. Byte Magazine, 21(2): 43–??, February 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:IE

[Ano96b] Anonymous. 4.0 isn’t for everyone. Byte Magazine, 21(7):121– ??, July 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:AT

[Ano96c] Anonymous. Aerial telecommuting. Byte Magazine, 21(10):30– ??, October 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:AIA

[Ano96d] Anonymous. Alphabetical index to advertisers. Byte Magazine, 21(5):222–??, May 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:AIS

[Ano96e] Anonymous. Alternative input spells relief. Byte Magazine, 21 (2):26–??, February 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:ACM

[Ano96f] Anonymous. Anyone connected to the MBone? Byte Magazine, 21(2):60–??, February 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:ANN

[Ano96g] Anonymous. Apple’s new notebooks. Byte Magazine, 21(12): 34–??, December 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:APB

[Ano96h] Anonymous. AT and T Paradyne’s bandwidth revolution. Byte Magazine, 21(2):60–??, February 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:ARO

[Ano96i] Anonymous. Au revoir, overheads. Byte Magazine, 21(2): 34–??, February 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:BO

[Ano96j] Anonymous. Banking on-line. Byte Magazine, 21(12):38–??, December 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:BUG

[Ano96k] Anonymous. The best in ultrafast graphics adapters. Byte Magazine, 21(2):144–??, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:BC

[Ano96l]  Anonymous.  The best of

Comdex. Byte Magazine, 21 (1):40–??, January 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:BIT

[Ano96m] Anonymous. Better images through micromirrors. Byte Magazine, 21(7):28–??, July 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:BRC

[Ano96n] Anonymous. Better replication coming for databases. Byte Magazine, 21(5):36–??, May 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:BTR

[Ano96o] Anonymous. Better tools reduce clutter. Byte Magazine, 21 (1):26–??, January 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:BCA

[Ano96p] Anonymous. Books and CD-ROMs: The art of software design. Byte Magazine, 21(5):41– ??, May 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:BBc

[Ano96q] Anonymous. Browsers for the blind. Byte Magazine, 21(12): 36–??, December 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:BG

[Ano96r]  Anonymous.  Buyer’s guide.

Byte Magazine, 21(5):183–??, May 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:BBa

[Ano96s] Anonymous. BYTE best. Byte Magazine, 21(3):125–??, March 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:BBb

[Ano96t]  Anonymous.  BYTE BEST.

Byte Magazine, 21(4):159–??, April 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:BRW

[Ano96u] Anonymous. BYTE readers write about new server architectures, Windows NT issues, MMX programming, and “free PCs”. Byte Magazine, 21(10): 19–??, October 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:BS

[Ano96v] Anonymous. BYTE survey. Byte Magazine, 21(12):32–??, December 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:BBP

[Ano96w] Anonymous. BYTE’s best of PC Expo. Byte Magazine, 21 (9):30–??, September 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:CCD

[Ano96x] Anonymous. Call centers deliver data. Byte Magazine, 21 (12):28–??, December 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:CTP

[Ano96y] Anonymous. Code talk: A programmer needs a Mais. Byte Magazine, 21(1):38–??, January 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:CTA

[Ano96z] Anonymous. Code talk: Add spreadsheets, charting to programs with ease. Byte Magazine, 21(2):38–??, February 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:CTW

[Ano96-27] Anonymous. Code talk: Write Windows 95 virtual device drivers on higher ground. Byte Magazine, 21(5):38–??, May 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:Ca

[Ano96-28] Anonymous. Codetalk. Byte Magazine, 21(6):36–??, June 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:Cc

[Ano96-29] Anonymous. Codetalk. Byte Magazine, 21(7):32–??, July 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:Cd

[Ano96-30] Anonymous. Codetalk. Byte Magazine, 21(8):32–??, August 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:CSP

[Ano96-31] Anonymous. Codetalk: Symantec pours Java into its development environment. Byte Magazine, 21(3):38–??, March 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:CFC

[Ano96-32] Anonymous. Coming faster, cheaper optical data links. Byte Magazine, 21(2):40–??, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:CTI

[Ano96-33] Anonymous. Coming thousands of Internet on-ramps. Byte Magazine, 21(1):36–??, January 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:CCW

[Ano96-34] Anonymous. Comments on code — why C is a disaster, and other ruminations on programming. Byte Magazine, 21(7):163– ??, July 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:CS

[Ano96-35] Anonymous. Convergence strategies. Byte Magazine, 21(9): 123–??, September 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:CHS

[Ano96-36] Anonymous. Creating holographic storage. Byte Magazine, 21(4):48–??, April 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:CB

[Ano96-37] Anonymous. Cyrix 6x86 bug. Byte Magazine, 21(11):30–??, November 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:CVP

[Ano96-38] Anonymous. Cyrix 6x86 vs. Pentium. Byte Magazine, 21(9): 26–??, September 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:DNV

[Ano96-39] Anonymous. Data’s new voice.

Byte Magazine, 21(9):129–??, September 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:LMM

[Ano96-40] Anonymous. de lorme’s Map’n’Go makes an excellent traveling companion; Gateway’s newest Pentium machine has more than just a zippy CPU. Byte Magazine, 21(9):220–??, September 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:DNE

[Ano96-41] Anonymous. Diamond’s new edge. Byte Magazine, 21(2): 147–??, February 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:DC

[Ano96-42] Anonymous. Digital cameras.

Byte Magazine, 21(12):32–??, December 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:DIN

[Ano96-43] Anonymous. Digital imagery — new digital cameras from Kodak and Polaroid. Byte Magazine, 21(7):220–??, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:DCP

[Ano96-44] Anonymous. Dirt-cheap color printers. Byte Magazine, 21 (10):36–??, October 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:DTM

[Ano96-45] Anonymous. Double trouble for 56-Kbps modems. Byte Magazine, 21(12):27–??, December 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:EICa

[Ano96-46] Anonymous. Editorial index by company. Byte Magazine, 21 (5):226–??, May 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:EICb

[Ano96-47] Anonymous. Editorial index by company. Byte Magazine, 21 (7):212–??, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:EIG

[Ano96-48] Anonymous. Elsa incorporates Glint chip for accelerated 3-D. Byte Magazine, 21(2):147–??, February 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:EBB

[Ano96-49] Anonymous. Entomophagous Bruce Brown. Byte Magazine, 21 (10):48–??, October 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:FFC

[Ano96-50] Anonymous. Faster, fatter CD-ROM titles on hold. Byte Magazine, 21(6):46–??, June 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:FPP

[Ano96-51] Anonymous. Finally,pushbutton PDA-to PC integration. Byte Magazine, 21(3):40– ??, March 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:FWP

[Ano96-52] Anonymous. First Web PCs arrive. Byte Magazine, 21(4):24– ??, April 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:FG

[Ano96-53] Anonymous. FlashPix graphics. Byte Magazine, 21(11): 36–??, November 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:FR

[Ano96-54] Anonymous. Flying robots. Byte Magazine, 21(11):48–??, November 1996. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:FFF

[Ano96-55] Anonymous. Forms follow function — Peter/Jerram electronic forms linked to Web front ends could help replace those zillion pieces of paper. Byte Magazine, 21(9):153–??, September 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:FKT

[Ano96-56] Anonymous. Four key technologies. Byte Magazine, 21(3):46– ??, March 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:FNS

[Ano96-57] Anonymous. Future notebook screens. Byte Magazine, 21(11): 42–??, November 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:GGP

[Ano96-58] Anonymous. Geoprocessing goes prime time. Byte Magazine, 21(7):27–??, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:GR

[Ano96-59] Anonymous. Get real-time. Byte Magazine, 21(2):53–??, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:GSW

[Ano96-60] Anonymous. Get smart — wear a PC. Byte Magazine, 21(3):36– ??, March 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:GVV

[Ano96-61] Anonymous. Global video village. Byte Magazine, 21(9): 137–??, September 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:HDS

[Ano96-62] Anonymous. Higher data speeds coming for plain phone lines. Byte Magazine, 21(1):30–??, January 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:HGS

[Ano96-63] Anonymous. Holiday gift sampler. Byte Magazine, 21(12): 30–??, December 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:HPS

[Ano96-64] Anonymous. Home PC sales pass peak. Byte Magazine, 21 (2):36–??, February 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:HM

[Ano96-65] Anonymous. Honorable mentions. Byte Magazine, 21(3):128– ??, March 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:HJC

[Ano96-66] Anonymous. How Java can pay the rent. Byte Magazine, 21(6): 27–??, June 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:HNM

[Ano96-67] Anonymous. How NT 4.0 merges Win APIs… right! Byte Magazine, 21(7):36–??, July 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:HTC

[Ano96-68] Anonymous. How te create a network security policy. Byte Magazine, 21(8):36–??, August 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:HWW

[Ano96-69] Anonymous. How the Web will change computing. Byte Magazine, 21(1):24–??, January 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:HWTa

[Ano96-70] Anonymous. How we tested. Byte Magazine, 21(2):148–??, February 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:HWTb

[Ano96-71] Anonymous. How we tested. Byte Magazine, 21(3):130–??, March 1996. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:HWTc

[Ano96-72] Anonymous. How we tested.

Byte Magazine, 21(4):164–??, April 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:HTL

[Ano96-73] Anonymous. HP takes the lead in workstation performance. Byte Magazine, 21(7):36–??, July 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:IIE

[Ano96-74] Anonymous. IBM to improve enterprise support in OS/2 server. Byte Magazine, 21(6):27–??, June 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:ITY

[Ano96-75] Anonymous. I’ll talk to you soon. Byte Magazine, 21(3):30– ??, March 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:IBRa

[Ano96-76] Anonymous. INBOX — BYTE readers sound off on electric money, Internet commerce and privacy, programming for Intel’s MMX, and praise the core of BYTE. Byte Magazine, 21(9): 19–??, September 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:IBRb

[Ano96-77] Anonymous. INBOX — BYTE readers write about our Web coverage, NT security, the online future, and more. Byte Magazine, 21(11):19–??, November 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:IBRc

[Ano96-78] Anonymous. INBOX — BYTE readers write about that bandwidth barrier and the future of the Internet, Web conferencing, and the search for killer software. Byte Magazine, 21(12): 19–??, December 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:IAA

[Ano96-79] Anonymous. Index to advertisers by alphabetical order. Byte Magazine, 21(7):208–??, July 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:IAI

[Ano96-80] Anonymous. Index to advertisers by inquiry reply cards. Byte Magazine, 21(5):64–??, May 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:IAPa

[Ano96-81] Anonymous. Index to advertisers by product category. Byte Magazine, 21(5):224–??, May 1996.

CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:IAPb

[Ano96-82] Anonymous. Index to advertisers by product category. Byte Magazine, 21(7):210–??, July 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:IRC

[Ano96-83] Anonymous. Inquiry reply cards.

Byte Magazine, 21(4):144–??, April 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:IP

[Ano96-84] Anonymous. Inside the Pippin. Byte Magazine, 21(3):54– ??, March 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:IUW

[Ano96-85] Anonymous. Inside the ultimate Web site. Byte Magazine, 21(5): 26–??, May 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:IBC

[Ano96-86] Anonymous. Intel beats the clock… again. Byte Magazine, 21(1):25–??, January 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:IMO

[Ano96-87] Anonymous. The interviews:

MOLAP,ROLAP, overlap. Byte Magazine, 21(8):40–??, August 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:IGB

[Ano96-88] Anonymous. ISDN: The good, bad, and ugly. Byte Magazine, 21(7):32–??, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:JBA

[Ano96-89] Anonymous. Java business applications arrive. Byte Magazine, 21(8):30–??, August 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:JCB

[Ano96-90] Anonymous. Java chips boost applet speed. Byte Magazine, 21(4):25–??, April 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:LPM

[Ano96-91] Anonymous. Lower prices, modularity lead notebook trends. Byte Magazine, 21(8):38–??, August 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:MPP

[Ano96-92] Anonymous. MPEG: Pay for play. Byte Magazine, 21(2): 144–??, February 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:MTA

[Ano96-93] Anonymous. Multimedia tools animate the Web. Byte Magazine, 21(3):28–??, March 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:NPM

[Ano96-94] Anonymous. Net programming for the masses. Byte Magazine, 21(2):101–??, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:NNA

[Ano96-95] Anonymous. Netscape’s Navigator 2.0 adds? security and interactive multimedia. Byte Magazine, 21(3):126–??, March 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:NFSa

[Ano96-96] Anonymous. NetWare: Fight or switch? Byte Magazine, 21(7):127–??, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:NCG

[Ano96-97] Anonymous. Network cameras go mobile. Byte Magazine, 21 (10):44–??, October 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:NBW

[Ano96-98] Anonymous. New batteries will keep you running. Byte Magazine, 21(3):27–??, March 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:NCS

[Ano96-99] Anonymous. New CPUs signal bener multimedia PCs. Byte Magazine, 21(4):30–??, April 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:NDS

[Ano96-100] Anonymous. New desktop standard: 200 MHz. Byte Magazine, 21(8):26–??, August 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:NDI

[Ano96-101] Anonymous. New docks improve commuter computers. Byte Magazine, 21(2):38–??, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:NFSb

[Ano96-102] Anonymous. New face of security. Byte Magazine, 21(10):30– ??, October 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:NNO

[Ano96-103] Anonymous. New NT to offer Win 95 look, Internet links. Byte Magazine, 21(4):34–??, April 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:NPI

[Ano96-104] Anonymous. New PCs improve your image. Byte Magazine, 21(5):25–??, May 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:NPa

[Ano96-105] Anonymous. New PowerMacs.

Byte Magazine, 21(9):34–??, September 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:NPb

[Ano96-106] Anonymous. New PowerPCs.

Byte Magazine, 21(11):44–??, November 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:NIF

[Ano96-107] Anonymous. The next Internet file system. Byte Magazine, 21 (10):32–??, October 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:NTR

[Ano96-108] Anonymous. Notebooks take on rugged challenges. Byte Magazine, 21(2):30–??, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:NN

[Ano96-109] Anonymous. NT and the net.

Byte Magazine, 21(7):133–??, July 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:NE

[Ano96-110] Anonymous. NT everywhere. Byte Magazine, 21(7):119–??, July 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:NSG

[Ano96-111] Anonymous. NT server’s growing pains. Byte Magazine, 21 (7):137–??, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:OW

[Ano96-112] Anonymous. Objects and the Web. Byte Magazine, 21(9): 36–??, September 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:OIS

[Ano96-113] Anonymous. On-line information services deliver disaster recovery for small business. Byte Magazine, 21(5):30–??, May 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:OSJ

[Ano96-114] Anonymous. On-line shoppers: “Just Looking, Thanks”. Byte Magazine, 21(3):34–??, March 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:OBO

[Ano96-115] Anonymous. OpenDoc: A better OLE than OLE? Byte Magazine, 21(7):30–??, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:PVI

[Ano96-116] Anonymous. PDA vendors improve PC links. Byte Magazine, 21(1):34–??, January 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:PPM

[Ano96-117] Anonymous. Pentium Pro moves to the desktop. Byte Magazine, 21(6):26–??, June 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:PP

[Ano96-118] Anonymous. Perfect-fit portables. Byte Magazine, 21(5):83– ??, May 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:PLE

[Ano96-119] Anonymous. Phar Lap’s embeddable kernel. Byte Magazine, 21(4):38–??, April 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:PMG

[Ano96-120] Anonymous. Playing MIME games. Byte Magazine, 21(2): 56–??, February 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:PCb

[Ano96-121] Anonymous. Plugged-in CD-ROMs. Byte Magazine, 21(11):

32–??, November 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:PDH

[Ano96-122] Anonymous. Portable devices highlight best of CeBlT. Byte Magazine, 21(6):32–??, June 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:PCa

[Ano96-123] Anonymous. Post-millennium cruising. Byte Magazine, 21(11): 30–??, November 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:PSS

[Ano96-124] Anonymous. PowerPCs still a spec on the horizon. Byte Magazine, 21(2):25–??, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:PUN

[Ano96-125] Anonymous. Previews of an urban navigator and a Macintosh ISDN pod. Byte Magazine, 21(7):214–??, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:RCC

[Ano96-126] Anonymous. R5000 cuts 3-D cost. Byte Magazine, 21(4):26– ??, April 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:RRO

[Ano96-127] Anonymous. Raders react to our stories on Unix vs. NT and works flow, and write about problems with the year 2000. Byte Magazine, 21(8):19–??, August 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:RBW

[Ano96-128] Anonymous. Reach beyond Win 95. Byte Magazine, 21(5):91– ??, May 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:RSIc

[Ano96-129] Anonymous. Reader service — inquiry reply cards. Byte Magazine, 21(10):44A–44B, 200A– 200B, October 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:RSId

[Ano96-130] Anonymous. Reader service — inquiry reply cards. Byte Magazine, 21(11):104A–B, 232A–B, November 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:RSIe

[Ano96-131] Anonymous. Reader service — inquiry reply cards. Byte Magazine, 21(12):152C–D, 192A–B, December 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:RSIb

[Ano96-132] Anonymous. Reader service — inquiry reply cards 1 28A-B, 21 2A-B. Byte Magazine, 21(9):??, September 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:RSIa

[Ano96-133] Anonymous. Reader service — inquiry reply cards 1 36A-D, 208 A-B. Byte Magazine, 21(7):??, July 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:RWA

[Ano96-134] Anonymous. Readers write ahout Unix and NT, the Web PC, and how Microsoft blew it with ISDN. Byte Magazine, 21(7):19–??, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:RHW

[Ano96-135] Anonymous. The return of handheld Windows. Byte Magazine, 21(11):26–??, November 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:RTN

[Ano96-136] Anonymous. Richer, thinner notebooks arrive. Byte Magazine, 21(7):26–??, July 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:RMB

[Ano96-137] Anonymous. A road map to better bandwidth. Byte Magazine, 21(6):27–??, June 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:RMR

[Ano96-138] Anonymous. RoboCop, meet robo fire fighter. Byte Magazine, 21(7):34–??, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:SIS

[Ano96-139] Anonymous. Service instead of server? Byte Magazine, 21 (3):128–??, March 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:SRD

[Ano96-140] Anonymous. Smarter radar detectors. Byte Magazine, 21(9): 28–??, September 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:SDA

[Ano96-141] Anonymous. SSA drives arrive. Byte Magazine, 21(9): 28–??, September 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:SA

[Ano96-142] Anonymous. Stack attack. Byte Magazine, 21(2):105–??, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:SC

[Ano96-143] Anonymous. Stay connected. Byte Magazine, 21(5):95–??, May 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:SPS

[Ano96-144] Anonymous. Stratus PC server wins best of COMDEX. Byte Magazine, 21(8):28–??, August 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:SOL

[Ano96-145] Anonymous. SunSoft’s object lesson. Byte Magazine, 21(2): 117–??, February 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:SWC

[Ano96-146] Anonymous. Survey: Web commerce. Byte Magazine, 21(11): 36–??, November 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:TS

[Ano96-147] Anonymous. Tech-support software. Byte Magazine, 21(10):26– ??, October 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:TRB

[Ano96-148] Anonymous. Ten reasons to buy duplex printers. Byte Magazine, 21(7):32–??, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:TMT

[Ano96-149] Anonymous. Tool makers take it to the net. Byte Magazine, 21(6): 31–??, June 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:TSS

[Ano96-150] Anonymous. Tour the sites and sounds. Byte Magazine, 21(2): 62–??, February 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:TVR

[Ano96-151] Anonymous. Toys — virtual reality. Byte Magazine, 21(6):224– ??, June 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:TW

[Ano96-152] Anonymous. Tuned into the Web. Byte Magazine, 21(12): 40–??, December 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:US

[Ano96-153] Anonymous. Unix survey. Byte Magazine, 21(10):42–??, October 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:UVNb

[Ano96-154] Anonymous. Unix vs. NT: Technical head-to-head. Byte Magazine, 21(5):50–??, May 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:UVNa

[Ano96-155] Anonymous. Unix vs. NT: The (edited) vendors’ view. Byte Magazine, 21(5):44–??, May 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:ULS

[Ano96-156] Anonymous. Users like Smart-Suite’s team computing. Byte Magazine, 21(3):38–??, March 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:VEI

[Ano96-157] Anonymous. Vendors eye Internet telephony standard. Byte Magazine, 21(8):27–??, August 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:VRG

[Ano96-158] Anonymous. Virtual reality goes to work. Byte Magazine, 21 (3):26–??, March 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:VN

[Ano96-159] Anonymous. A vote for NT. Byte Magazine, 21(5):48–??, May 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:VU

[Ano96-160] Anonymous. A vote for Unix. Byte Magazine, 21(5):46–??, May 1996. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:WPP

[Ano96-161] Anonymous. Watch out: 500-MHz PowerPCs planned for 1997. Byte Magazine, 21(5):40– ??, May 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:WDN

[Ano96-162] Anonymous. We don’t need no stinking computers. Byte Magazine, 21(2):54–??, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:WGM

[Ano96-163] Anonymous. We get mail about our Webserver review, the future of computing, and our cover story on the Web PC. Byte Magazine, 21(6):19–??, June 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:WDM

[Ano96-164] Anonymous. Web development — this month we take an indepth look at the technologies that will shape the Web of tomorrow. Byte Magazine, 21(12): 118–??, December 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:WFC

[Ano96-165] Anonymous. Web forces change.

Byte Magazine, 21(2):25–??, February 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:WID

[Ano96-166] Anonymous. Web info delivery. Byte Magazine, 21(11): 27–??, November 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:WPb

[Ano96-167] Anonymous. Web phones. Byte Magazine, 21(8):188–??, August 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:WPa

[Ano96-168] Anonymous. Well-connected PDAs. Byte Magazine, 21 (5):103–??, May 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:WHJ

[Ano96-169] Anonymous. Whatever happened to Josephson junctions? Byte Magazine, 21(4):54–??, April 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:WNPc

[Ano96-170] Anonymous. What’s new — A preview of AMD’s 5K86-P75 CPU, plus a look at Micro-Help’s zip file utility. Byte Magazine, 21(8):182–??, August 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:WNPd

[Ano96-171] Anonymous. What’s new — A preview of Matrox’s Mystique PCI graphics board, plus a look at Microsoft’s Outlook 97. Byte Magazine, 21(10):208–??, October 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:WNPe

[Ano96-172] Anonymous. WHAT’S NEW — A preview of Psion’s Siena handheld information manager, plus a look at Macromedia’s Free-Hand Graphics Suite 7. Byte Magazine, 21(12):200–??, December 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:WNA

[Ano96-173] Anonymous. What’s new — Apple’s QuickDraw 3D accelerator card promises more realistic 3-D imagery: Lotus cc:Mail for Macintosh Release 3 helps you automate handling; plus more new hardware and software. Byte Magazine, 21(3):154– ??, March 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:WNF

[Ano96-174] Anonymous. WHAT’S NEW — fractal design’s expression goes beyond standard vector-drawing packages; Gateway’s Solo 2100 S5-120 combines raw power and good looks. Byte Magazine, 21(11):240–??, November 1996.

CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:WNG

[Ano96-175] Anonymous. What’s new — Gateway’s Solo V120 is a zippy little Win 95 notebook: Lotus Approach 96 makes database work a breeze; plus more. Byte Magazine, 21(1):168–??, January 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:WNPb

[Ano96-176] Anonymous. What’s new — pick your CPU for this PC. Byte Magazine, 21(6):216–??, June 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:WNPa

[Ano96-177] Anonymous. What’s new — previews of a pointing device and AT and T’s PC Phone, as well as lots of new hardware and software. Byte Magazine, 21(2): 202–??, February 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:WNS

[Ano96-178] Anonymous. What’s new — Sys Technology’s 200-MHz Performance Pro 200 workstation proves that two disks are better than one: the latest version of Symantec’s Act! gives Psion 3A users contact management and scheduling capabilities: plus more new hardware and software. Byte Magazine, 21(4):187– ??, April 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:WNE

[Ano96-179] Anonymous. What’s new — the EpixPro digital camera makes it easy to zap photos into your PC; DeltaPoint’s QuickSite software lets you do your own Web pages. Byte Magazine, 21(5):176–??, May 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:WRP

[Ano96-180] Anonymous. Win 95 roars with Pentium-Class PCs. Byte Magazine, 21(2):24–??, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:WW

[Ano96-181] Anonymous. Windows without walls. Byte Magazine, 21(2): 109–??, February 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:WLG

[Ano96-182] Anonymous. WordPerfect losing ground to Microsoft. Byte Magazine, 21(5):38–??, May 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:YPS

[Ano96-183] Anonymous. Year 2000 promises strange days ahead. Byte Magazine, 21(2):36–??, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:ZHV

[Ano96-184] Anonymous. Zoom in on highquality video. Byte Magazine, 21(4):162–??, April 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:ECA

[Ano97a] Anonymous. 1997 Editors’ choice awards — we honor software and hardware that prove innovation was alive and well in 1997. Byte Magazine, 22(12):103–??, December 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:PIT

[Ano97b] Anonymous. 300-MHz Pentium II tests. Byte Magazine, 22(7): 22–??, July 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:PIP

[Ano97c] Anonymous. 400-MHz Pentium II PCs. Byte Magazine, 22(11): 22–??, November 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:VIF

[Ano97d] Anonymous. 57 varieties of Internet faxing — now you can do it without a Net connection. Byte Magazine, 22(2):34–??, February 1997. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:A

[Ano97e] Anonymous. 600-MHz Alpha.

Byte Magazine, 22(11):18–??, November 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:ARC

[Ano97f] Anonymous. Advances and retreats in computing — Internet Barbie, a new threat to data warehouses, and a request for terabytes of quality. Byte Magazine, 22(4):188–??, April 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:AWT

[Ano97g] Anonymous. Air war — TDMA, CDMA, GSM — from this alphabet soup will come tomorrow’s wireless network. Byte Magazine, 22(8):93–??, August 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:AWS

[Ano97h] Anonymous. All the Web’s a stage — trusted Web, a one-step authentication server. Byte Magazine, 22(8):135–??, August 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:AN

[Ano97i] Anonymous. Apple’s next OS.

Byte Magazine, 22(4):26–??, April 1997. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:BNB

[Ano97j] Anonymous. Bad news for broken portables. Byte Magazine, 22(2):32–??, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:BC

[Ano97k] Anonymous. Best of Comdex. Byte Magazine, 22(2):27–??, February 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:BBS

[Ano97l] Anonymous. Better backbones to shoulder bigger network burden — new high-speed links and services are on the way. Byte Magazine, 22(2):36–??, February 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:BFS

[Ano97m] Anonymous. Better in fits and starts — we test a new Pentium hopped up with Intel’s MMX. Byte Magazine, 22(2): 26–??, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:BWT

[Ano97n] Anonymous. Better Web tools on the way. Byte Magazine, 22 (1):26–??, January 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:BFC

[Ano97o] Anonymous. Bits — fatter, cheaper pipes. Byte Magazine, 22(5):32–??, May 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:BNS

[Ano97p] Anonymous. Bits — new storage technology. Byte Magazine, 22(5):28–??, May 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:BXB

[Ano97q] Anonymous. Bits — x86 breaks 200 MHz. Byte Magazine, 22(5): 26–??, May 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:BBE

[Ano97r] Anonymous. Building a better error trap. Byte Magazine, 22(3):32–??, March 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:BPI

[Ano97s] Anonymous. Burning paper with Internet EDI — VARs are building new models for electronic commerce. Byte Magazine, 22 (8):116–??, August 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:BRW

[Ano97t] Anonymous. BYTE readers write about Copland, network computers, impolite robots — and users, and our 64-hit future. Byte Magazine, 22(2): 19–??, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:BSD

[Ano97u] Anonymous. BYTE survey: Data warehouse woes. Byte Magazine, 22(1):32–??, January 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:CDG

[Ano97v] Anonymous. Cell death — A great new way to terminate annoying cell-phone calls. Byte Magazine, 22(12):168–??, December 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:CNW

[Ano97w] Anonymous. Centaur’s new WinChip. Byte Magazine, 22 (12):24–??, December 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:CPC

[Ano97x] Anonymous. Chips promise cheaper modems. Byte Magazine, 22(12):28–??, December 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:CT

[Ano97y] Anonymous. Clustering technology. Byte Magazine, 22(7):26– ??, July 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:CGP

[Ano97z] Anonymous. CMOS gets the picture — using a CMOS sensor, Sound Vision’s SVmini-209 delivers high-quality pics at a low price. Byte Magazine, 22 (11):34–??, November 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:CNP

[Ano97-27] Anonymous. Corel’s nearly perfect suite spot — WordPerfect 8. Byte Magazine, 22(7):36–??, July 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:CGR

[Ano97-28] Anonymous. CTI gets ready for the masses — commercial middleware helps ring up profits with telephony. Byte Magazine, 22 (8):116–??, August 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:CPPa

[Ano97-29] Anonymous. CTI, piece by piece — setting communications standards throughout an enterprise is one of the most, fundamental yet contentious tasks in a CTI launch. Byte Magazine, 22 (2):85–??, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:CPPb

[Ano97-30] Anonymous. CTI, piece by piece — the scalability and reliability of Unix make it the traditional choice for running CTI servers. Byte Magazine, 22(2): 85–??, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:CPPc

[Ano97-31] Anonymous. CTI, piece by piece — wider applications support and open APIs are making more companies consider Windows NT for CTI systems. Byte Magazine, 22(2):85–??, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:DWH

[Ano97-32] Anonymous. Data warehouses — how to build the best facility for your information, and how to help users find the data they need right now. Byte Magazine, 22(1):82–??, January 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:DNM

[Ano97-33] Anonymous. Dawn of the new millennium — Matrox’s upgraded hardware. Byte Magazine, 22(7):121–??, July 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:DN

[Ano97-34] Anonymous. Decongesting the net. Byte Magazine, 22(4):40– ??, April 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:DMT

[Ano97-35] Anonymous. DVD: A moving target. Byte Magazine, 22(12): 30–??, December 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:ESA

[Ano97-36] Anonymous. Editorial — step aside, x86. Byte Magazine, 22 (12):14–??, December 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:EICa

[Ano97-37] Anonymous. Editorial index by company. Byte Magazine, 22 (7):154–??, July 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:EICb

[Ano97-38] Anonymous. Editorial index by company. Byte Magazine, 22(11):158–??, November 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:EAU

[Ano97-39] Anonymous. Enterprise apps use Web. Byte Magazine, 22(8):26– ??, August 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:FMB

[Ano97-40] Anonymous. Faster Mac bus.

Byte Magazine, 22(8):24–??, August 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:FDA

[Ano97-41] Anonymous. Faster on the draw — AutoCAD 14. Byte Magazine, 22(7):119–??, July 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:FGB

[Ano97-42] Anonymous. Five guides to better site construction. Byte Magazine, 22(3):34–??, March 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:FJP

[Ano97-43] Anonymous. A free Java platform. Byte Magazine, 22(7):158, July 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:FWI

[Ano97-44] Anonymous. Future watch: I2O processors. Byte Magazine, 22 (12):28–??, December 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:GDT

[Ano97-45] Anonymous. Guaranteed delivery — transaction processing monitors keep Web servers humming. Byte Magazine, 22(8):77– ??, August 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:SUA

[Ano97-46] Anonymous. I say! an understanding application — dragon’s naturally speaking. Byte Magazine, 22(8):33–??, August 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:IBB

[Ano97-47] Anonymous. IBM builds a better DBMS — IBM’s enhanced DB 2 5.0. Byte Magazine, 22(8):133– ??, August 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:IDS

[Ano97-48] Anonymous. IBM’s digital shrinkwrapper — an early look at Cryptolope. Byte Magazine, 22(8):138–??, August 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:IRa

[Ano97-49] Anonymous. Improbable research. Byte Magazine, 22(2): 38–??, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:IRb

[Ano97-50] Anonymous. Improbable research. Byte Magazine, 22(3):38– ??, March 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:Ia

[Ano97-51] Anonymous.  Inbox.  Byte

Magazine, 22(11):13–??, November 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:Ib

[Ano97-52] Anonymous.  Inbox.  Byte

Magazine, 22(12):17–??, December 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:IBR

[Ano97-53] Anonymous. INBOX — BYTE readers sound off on Moore’s law, Windows NT network architecture, e-mail standards, and more. Byte Magazine, 22(3):19– ??, March 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:IRRa

[Ano97-54] Anonymous. INBOX — readers react to our stories on 64bit OSes, WebNFS, network computers, on-line help, and more. Byte Magazine, 22(1):19– ??, January 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:IRRb

[Ano97-55] Anonymous. Inbox — readers respond to BYTE. Byte Magazine, 22(4):19–??, April 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:IO

[Ano97-56] Anonymous. Internet and outernet. Byte Magazine, 22(1):34– ??, January 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:IVJ

[Ano97-57] Anonymous. Interview: A voice for Java developers. Byte Magazine, 22(12):44, December 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:IWJ

[Ano97-58] Anonymous. Interview: Why Java won’t repeat the mistakes of Unix. Scott McNealy, chairman and CEO of Sun Microsystems, talks about how Java will change computing and the computer industry. Byte Magazine, 22(1):40, January 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Some McNealy quotes: “And I’m urging every CEO to make sure it [Windows] runs really slowly, so people stop using that stuff. The right way is to ban PowerPoint from your company.” “I banned PowerPoint from my company and we have had the best two quarters we’ve ever had… ” “I want to give everybody plastic Mylar sheets and all the pens they need to scribble on them. And I said use what I call the [Sun cofounder] Bill Joy font. You can see where he licked his thumb and erases; it’s so much faster.”.

Anonymous:1997:JHF

[Ano97-59] Anonymous. Just how fast are those new MMX chips? Byte Magazine, 22(2):26–??, February 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:KSW

[Ano97-60] Anonymous. A kinder, smaller Windows — porting to handheld PCs is nothing to dread. Byte Magazine, 22(8):103–??, August 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:LV

[Ano97-61] Anonymous. Lanier on VR. Byte Magazine, 22(8):32–??, August 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:LCI

[Ano97-62] Anonymous. Laser color, inexpensive and quick — Xerox’s DocuPrint. Byte Magazine, 22 (8):37–??, August 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:LDM

[Ano97-63] Anonymous. Lessons in data mining. Byte Magazine, 22(2): 40–??, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:MLM

[Ano97-64] Anonymous. Making Linux more livable — Caldera’s OpenLinux base. Byte Magazine, 22(7):124– ??, July 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:MOJ

[Ano97-65] Anonymous. Manage objects in Java. Byte Magazine, 22(7):158, July 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:MPF

[Ano97-66] Anonymous. Microsoft promises fewer headaches. Byte Magazine, 22(3):40–??, March 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:MPD

[Ano97-67] Anonymous. MMX power for desktop PCs: We look at 10 Pentium systems souped-up with MMX technology. Byte Magazine, 22(7):106–??, July 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:MV

[Ano97-68] Anonymous. Multipoint videoconferencing. Byte Magazine, 22(4):32–??, April 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:ND

[Ano97-69] Anonymous. Natural dictation. Byte Magazine, 22(7):22– ??, July 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:NH

[Ano97-70] Anonymous. The networked house. Byte Magazine, 22(2):

28–??, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:NPS

[Ano97-71] Anonymous. Never-ending power supply — APC’s Symmetra UPS. Byte Magazine, 22 (8):38–??, August 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:NPP

[Ano97-72] Anonymous. New PC phones. Byte Magazine, 22(3):28–??, March 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:NXP

[Ano97-73] Anonymous. New x86 processors. Byte Magazine, 22(8):20– ??, August 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:N

[Ano97-74] Anonymous. New.domain.names. Byte Magazine, 22(7):25–??, July 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:NSW

[Ano97-75] Anonymous. NT security — Windows NT security isn’t perfect. this month’s Web project offers seven ways to make NT safer. Byte Magazine, 22(2): 117–??, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:NSE

[Ano97-76] Anonymous. NT suits up for enterprise computing. Byte Magazine, 22(3):36–??, March 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:NNN

[Ano97-77] Anonymous. Number Nine’s new spin: Revolution 3D — fast video playback and your choice of AGP or PCI. Byte Magazine, 22(11): 33–??, November 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:PHP

[Ano97-78] Anonymous. PCI hot plug for servers. Byte Magazine, 22(7): 23–??, July 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:PRA

[Ano97-79] Anonymous. PCs: The road ahead. Byte Magazine, 22(3):26– ??, March 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:PMP

[Ano97-80] Anonymous. Power Mac prevails. Byte Magazine, 22(11): 22–??, November 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:PSP

[Ano97-81] Anonymous. Power to the server — for performance and reliability, RISC servers running Unix excel as applications servers. Byte Magazine, 22(2): 133–??, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:PSS

[Ano97-82] Anonymous. Power to the server — SMP and economical prices are turning NT and x8 6-based machines into decent applications servers. Byte Magazine, 22 (2):133–??, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:RSIa

[Ano97-83] Anonymous. Reader service — inquiry reply cards. Byte Magazine, 22(1):48A–B, 176A–B, January 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:RSIb

[Ano97-84] Anonymous. Reader service — inquiry reply cards. Byte Magazine, 22(2):144A–B, 144C–D, 176A–B, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:RSIc

[Ano97-85] Anonymous. Reader service — inquiry reply cards. Byte Magazine, 22(7):152A–??, July 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:RSId

[Ano97-86] Anonymous. Reader service — inquiry reply cards. Byte Magazine, 22(11):156A–??, November 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:RHR

[Ano97-87] Anonymous. Rebuilding human resources — as HR departments get smaller, resellers have more opportunities. Byte Magazine, 22(7):104K–??, July 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:RDI

[Ano97-88] Anonymous. Recycling digital ink. Byte Magazine, 22(8):22– ??, August 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:RE

[Ano97-89] Anonymous. Replication everywhere. Byte Magazine, 22(4):34– ??, April 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:RCW

[Ano97-90] Anonymous. Reseller — cashing in on Web shopping sprees. Byte Magazine, 22(12):116C–??, December 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:ROI

[Ano97-91] Anonymous. Reseller — Oracle8: Have it your way. Byte Magazine, 22(12):116K–??, December 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:RLW

[Ano97-92] Anonymous. Resellers learn Wall Street smarts — custom middleware helps resellers profit from the financial industry’s volatility. Byte Magazine, 22(7):104C– ??, July 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:SWC

[Ano97-93] Anonymous. Security wake-up call. Byte Magazine, 22(1):36– ??, January 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:SW

[Ano97-94] Anonymous. Share the wealth. Byte Magazine, 22(2):123–??, February 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:SNC

[Ano97-95] Anonymous. Smart netcards coming to America. Byte Magazine, 22(1):28–??, January 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:SS

[Ano97-96] Anonymous. Socket 7 successor? Byte Magazine, 22(11): 19–??, November 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:SSA

[Ano97-97] Anonymous. Something to STARE at — 4-D images arrive (on our very own pages)! senders of junk e-mail get hosed. Byte Magazine, 22(11):164–??, November 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:SREb

[Ano97-98] Anonymous. Special report — extending the enterprise — defending from the unthinkable. Byte Magazine, 22(12):67–??, December 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:SREc

[Ano97-99] Anonymous. Special report — extending the enterprise — extranets at your service. Byte Magazine, 22(12):75–??, December 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:SREa

[Ano97-100] Anonymous. Special report — extending the enterprise — the extranet revolution. Byte Magazine, 22(12):65–??, December 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:SFD

[Ano97-101] Anonymous. Speedy file delivery — Tumbleweed’s Posta. Byte Magazine, 22(8):137–??, August 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:SSB

[Ano97-102] Anonymous. Suite and sour — BackOffice slugs it out in the features war with IBM Software servers. Byte Magazine, 22(2): 131–??, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:SSN

[Ano97-103] Anonymous. Suite and sour — the nine server components within IBM Software Servers let you scale a client/server implementation to your particular needs. Byte Magazine, 22(2): 131–??, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:TWS

[Ano97-104] Anonymous. Tiniest weather station. Byte Magazine, 22(4):28– ??, April 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:TTA

[Ano97-105] Anonymous. Tools for telephony apps — building CTI applications requires choosing among three classes of CTI development tools, which offer varying degrees of control and difficulty. Byte Magazine, 22(2):91–??, February 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:TLN

[Ano97-106] Anonymous. Toshiba’s Libretto notebook, Motorola’s 300-MHz Mac, Hauppauge’s VideoTalk, plus new development tools. Byte Magazine, 22(7):155–??, July 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:TTM

[Ano97-107] Anonymous. Two things for the modern Jack Kerouac to take on the road: DeLorme’s Tripmate satellite navigator and Millenia’s Email Reader. Byte Magazine, 22(2):182–??, February 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:UGM

[Ano97-108] Anonymous. Ultraportables get more power. Byte Magazine, 22 (12):25–??, December 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:UN

[Ano97-109] Anonymous. Unhinged notebooks. Byte Magazine, 22(2): 30–??, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:UMW

[Ano97-110] Anonymous. Upwardly mobile 3-D workstation — Digital Equipment’s dual-CPU machine. Byte Magazine, 22(7):35– ??, July 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:VJV

[Ano97-111] Anonymous. Visual Java for Visual Basic developers. Byte Magazine, 22(9):145–146, September 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:WWD

[Ano97-112] Anonymous. A watchamacallit for words — Dan Bricklin’s software for making hypertext documents. Byte Magazine, 22(11): 37–??, November 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:WSW

[Ano97-113] Anonymous. What’s in store for the Web — Object Design’s ObjectStore 5. Byte Magazine, 22 (8):34–??, August 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:WNAa

[Ano97-114] Anonymous. What’s new — Adobe’s Acrobat 3.0 adds a hefty list of Web-integration features. Hewlett–Packard’s ScanJet 5p adds improved image quality, a sleek design, and plug and play installation. Byte Magazine, 22 (3):178–??, March 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:WNAd

[Ano97-115] Anonymous. What’s new — Alpha workstations hit 600 MHz; Unix notebooks; and more new products. Byte Magazine, 22 (11):159–??, November 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:WNAb

[Ano97-116] Anonymous. What’s new — Apple’s PowerBook 3400 is faster, brighter, and better; Adobe’s PageMaker 6.5 has improved power, flexibility, and electronic publishing. Byte Magazine, 22 (4):178–??, April 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:WNAc

[Ano97-117] Anonymous. What’s New — AST’s quiet Pentium machine, Kai’s PhotoSoap image cleanser, digital cameras, programming tools, and more new hardware and software. Byte Magazine, 22 (8):163–??, August 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:WND

[Ano97-118] Anonymous. What’s new — Design Intelligence’s i-publish helps you turn ideas into compelling documents; the Tektronix Phaser 350 rivals fast color laser printers. Byte Magazine, 22 (1):184–??, January 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:WNM

[Ano97-119] Anonymous. What’s new — MetaCreations’ Ray Dream 3-D, Tektronix’s new color printer, better OCR from Scan-Soft, and more hardware and software. Byte Magazine, 22(12):163–??, December 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:WNN

[Ano97-120] Anonymous. What’s new — NEC’s low-cost SuperScript 660Plus prints 600 dpi; Corel’s Lumiere moves video editing into the mainstream. Byte Magazine, 22(5):155–??, May 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:WNS

[Ano97-121] Anonymous. Which OS? — NT server offers an ever-growing base of applications that leverage the multiprocessing nature of the OS. Byte Magazine, 22 (2):125–??, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:WSS

[Ano97-122] Anonymous. Which OS? — Sun’s Solaris is easily customizable, which means you can create applications servers that are tuned to your particular needs. Byte Magazine, 22(2): 125–??, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:WWS

[Ano97-123] Anonymous. Which OS? — Warp Server provides solid applications-server capabilities, file and print services, management tools, remote connectivity, and backup and recovery. Byte Magazine, 22(2):125–??, February 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:WIM

[Ano97-124] Anonymous. Why interfaces matter — computers are about to join the Information Age. Byte Magazine, 22(7):12–??, July 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:WMR

[Ano97-125] Anonymous. Windows 95/ Memphis: Ready for the future — the next Windows upgrade. Byte Magazine, 22(7):33– ??, July 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:WP

[Ano97-126] Anonymous. Wireless PDAs.

Byte Magazine, 22(1):30–??, January 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:YEO

[Ano97-127] Anonymous. Your E-mail is obsolete — the impact of the Internet on your proprietary e-mail system? it’s probably obsolete now. here’s how to get back up to speed. Byte Magazine, 22 (2):66–??, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:MSB

[Ano98a] Anonymous. 100-MHz: A system bus for tomorrow: PIIs ride a faster bus. Byte Magazine, 23(6):125–??, June 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:AAF

[Ano98b] Anonymous. Actions on all fronts. Byte Magazine, 23(1):88– ??, January 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:AUD

[Ano98c] Anonymous. Actual USB devices. Byte Magazine, 23(2): 32–??, February 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:APR

[Ano98d] Anonymous. Apple plays a Rhapsody. Byte Magazine, 23 (1):70–??, January 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:BFR

[Ano98e] Anonymous. Baan fires up rapid deployment. Byte Magazine, 23 (1):96C–??, January 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:BSN

[Ano98f] Anonymous. Backup software: Network backup that users will like. Byte Magazine, 23(7):38– ??, July 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:BNS

[Ano98g] Anonymous. Better notebook screens. Byte Magazine, 23(7): 20–??, July 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:BDM

[Ano98h] Anonymous. A “box” designed for manageability: Toshiba’s Equium 7000S. Byte Magazine, 23(6):128–??, June 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:BGG

[Ano98i] Anonymous. Broadband goes guerrilla. Byte Magazine, 23 (1):78–??, January 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:BG

[Ano98j]  Anonymous.  Buyer’s guide.

Byte Magazine, 23(6):112K–??, June 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:BBC

[Ano98k] Anonymous. Byte’s best of CeBIT. Byte Magazine, 23(6): 30–??, June 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:CFB

[Ano98l] Anonymous. C++ forever!: Borland’s C++ builder 3. Byte Magazine, 23(6):124–??, June 1998.

CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:CMT

[Ano98m] Anonymous. Cable modems take the early lead — cable modems live up to their promise of highspeed access. Byte Magazine, 23 (1):119–??, January 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:CSC

[Ano98n] Anonymous. CAD software — CADzooks! what a deal — we test Visio’s IntelliCAD, a heavyduty but low-cost alternative to AutoCAD. Byte Magazine, 23 (2):130–??, February 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:CCR

[Ano98o] Anonymous. Call control for the rest of us. Byte Magazine, 23 (1):86–??, January 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:CMD

[Ano98p] Anonymous. Centralized management for desktops. Byte Magazine, 23(1):63–??, January 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:CA

[Ano98q] Anonymous. Computers with attitude. Byte Magazine, 23(7):87– ??, July 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:CIW

[Ano98r]  Anonymous.  CPU interface

wars. Byte Magazine, 23(1):58– ??, January 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:DYS

[Ano98s] Anonymous. Databases Year 2000 survival guide. Byte Magazine, 23(7):52–??, July 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:DNE

[Ano98t] Anonymous. The days of ninetyeight. Byte Magazine, 23(7):111– ??, July 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:DH

[Ano98u] Anonymous. Detecting hackers. Byte Magazine, 23(6):40– ??, June 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:DYH

[Ano98v] Anonymous. Do you hear what I say? dictation with continuous speech recognition. Byte Magazine, 23(1):115–??, January 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:DSA

[Ano98w] Anonymous. Doing something about Microsoft — Jerry has some ideas about how to resolve the Microsoft/Justice flap. plus, another encounter with Plug and No Play. Byte Magazine, 23 (3):129–??, March 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:DE

[Ano98x] Anonymous. DSL for everyman. Byte Magazine, 23(2): 34–??, February 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:DDD

[Ano98y] Anonymous. DSL’s devilish details. Byte Magazine, 23(7):72– ??, July 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:DAR

[Ano98z] Anonymous. Dual analog routers. Byte Magazine, 23 (3):32–??, March 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:DSD

[Ano98-27] Anonymous. DVD stands for DiVideD. Byte Magazine, 23 (1):77–??, January 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:DHS

[Ano98-28] Anonymous. Dynamic HTML and scriptlets add life. Byte Magazine, 23(1):79–??, January 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:ES

[Ano98-29] Anonymous. E-commerce strategy. Byte Magazine, 23(4):40– ??, April 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:E

[Ano98-30] Anonymous. Editorial. Byte Magazine, 23(7):12–??, July 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:ECC

[Ano98-31] Anonymous. Editorial — coping with change. Byte Magazine, 23 (2):14–??, February 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:EICa

[Ano98-32] Anonymous. Editorial index by company. Byte Magazine, 23 (4):162–??, April 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:EICb

[Ano98-33] Anonymous. Editorial index by company. Byte Magazine, 23 (5):154–??, May 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:EICc

[Ano98-34] Anonymous. Editorial index (by company). Byte Magazine, 23 (7):138–??, July 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:ERS

[Ano98-35] Anonymous. Extranets reach the spotlight. Byte Magazine, 23 (1):71–??, January 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:FO

[Ano98-36] Anonymous. Future office. Byte Magazine, 23(7):32–??, July 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:GRA

[Ano98-37] Anonymous. GartnerGroup on remote access and the Net. Byte Magazine, 23(4):34–??, April 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:GR

[Ano98-38] Anonymous. GartnerGroup report. Byte Magazine, 23(7):28– ??, July 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:GRF

[Ano98-39] Anonymous. GartnerGroup report: Firewalls. Byte Magazine, 23(6):34–??, June 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:GRE

[Ano98-40] Anonymous. GartnerGroup report on ERP software. Byte Magazine, 23(3):34–??, March 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:GPP

[Ano98-41] Anonymous. Get to the point — A product to end voice-mail hemming and hawing. Byte Magazine, 23(3):158–??, March 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:GM

[Ano98-42] Anonymous. Getting the message. Byte Magazine, 23(1):87– ??, January 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:GHM

[Ano98-43] Anonymous. Giant heads, monster drives — IBM’s GMR drives. Byte Magazine, 23(4):42– ??, April 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:GEG

[Ano98-44] Anonymous. Gigabit Ethernet gears up. Byte Magazine, 23 (1):84–??, January 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:GBT

[Ano98-45] Anonymous. Groupware: Browsing through Notes. Byte Magazine, 23(7):33–??, July 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:HWH

[Ano98-46] Anonymous. Hackers in white hats. Byte Magazine, 23(6): 112O–??, June 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:HDM

[Ano98-47] Anonymous. Help desks make the Web connection — Webbased customer service can lower costs and improve response. Byte Magazine, 23(4):112K–??, April 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:HDT

[Ano98-48] Anonymous. High-speed desktops — three vendors make an Alpha bet — Alpha-powered PCs from Aspen, Polywell, and Tri-Star. Byte Magazine, 23 (2):43–??, February 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:HG

[Ano98-49] Anonymous. HTML groupware. Byte Magazine, 23(1):73– ??, January 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:AVH

[Ano98-50] Anonymous. I am virus: Hear me roar. Byte Magazine, 23(1):83– ??, January 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:ITM

[Ano98-51] Anonymous. IBM’s ThinkPad 600, monster workstations, a funky desktop machine, simple drawing software, and a translation tool for Web sites. Byte Magazine, 23(7):139–??, July 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:IYB

[Ano98-52] Anonymous. If you build it, will they come? — Sitara rewrites TCP. Byte Magazine, 23(3):125– ??, March 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:IAAa

[Ano98-53] Anonymous. Index to advertisers — alphabetical order. Byte Magazine, 23(1):148–??, January 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:IAAb

[Ano98-54] Anonymous. Index to advertisers: Alphabetical order. Byte Magazine, 23(5):152–??, May 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:IJT

[Ano98-55] Anonymous. Inprise’s JBuilder:

Take 2. Byte Magazine, 23(7): 85–??, July 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:ILC

[Ano98-56] Anonymous. Intel’s low-cost CPU. Byte Magazine, 23(6):24– ??, June 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:RYN

[Ano98-57] Anonymous. Is that really you?: Nuance Verifier knows. Byte Magazine, 23(6):123–??, June 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:JAC

[Ano98-58] Anonymous. Jackpot at Comdex — we check out the hundreds of new products at the big show in Vegas and pick the best. Byte Magazine, 23(2):77–??, February 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:JRA

[Ano98-59] Anonymous. JavaSoft rides again — A peek at the highlights of the upcoming JDK 1.2. Byte Magazine, 23(4):46–??, April 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:MCR

[Ano98-60] Anonymous. Mail call — readers respond to the big question: Besides civilization, what is the Net the end of? Byte Magazine, 23 (2):160–??, February 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:MPC

[Ano98-61] Anonymous. Managing PC costs. Byte Magazine, 23(7):65– ??, July 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:MTC

[Ano98-62] Anonymous. Memories of things to come. Byte Magazine, 23 (1):74–??, January 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:MDD

[Ano98-63] Anonymous. Microsoft’s dataacces directions — A practical look at the giant’s strategy. Byte Magazine, 23(3):104K–??, March 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:MPS

[Ano98-64] Anonymous. Middleware — publish and subscribe meets the Internet — pull out the old and push in the new with scalable publish/subscribe middleware for WANs and the Web. Byte Magazine, 23(2):125–??, February 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:MA

[Ano98-65] Anonymous. MMX alternatives. Byte Magazine, 23(7):26– ??, July 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:MOB

[Ano98-66] Anonymous. Multicast offers bandwidth salvation — three for data broadcasting. Byte Magazine, 23(3):121–??, March 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:MGA

[Ano98-67] Anonymous. MultiWin gets its audition. Byte Magazine, 23 (1):66–??, January 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:NEN

[Ano98-68] Anonymous. NetWare enters the nineties — Netware 5 makes it easy to use TCP/IP for all network client communications. Byte Magazine, 23(3):44– ??, March 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:NWT

[Ano98-69] Anonymous. Networks: Windows Terminal Server gets a hand. Byte Magazine, 23(7):30– ??, July 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:NNDb

[Ano98-70] Anonymous. New notes and domino. Byte Magazine, 23 (4):24–??, April 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:NI

[Ano98-71] Anonymous. The next Internet. Byte Magazine, 23(1):82– ??, January 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:NW

[Ano98-72] Anonymous. The next Windows. Byte Magazine, 23(1):56– ??, January 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:NNDa

[Ano98-73] Anonymous. Novell’s new directory direction — with NDS for NT, Novell tries to solve a big problem. Byte Magazine, 23 (3):126–??, March 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:OWR

[Ano98-74] Anonymous. OfficeMac: Worth revisiting in 98 — Microsoft restores Mac/Windows file compatibility. Byte Magazine, 23 (3):128–??, March 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:PI

[Ano98-75] Anonymous. Pentium II in ’98. Byte Magazine, 23(3):24– ??, March 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:PIN

[Ano98-76] Anonymous. A Pentium II notebook that talks to you: Digital’s HiNote VP 765. Byte Magazine, 23(6):122–??, June 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:PIP

[Ano98-77] Anonymous. PENTIUM II PCs — shooting the rapids at 333 MHz — we test new systems from Compaq and HP. Byte Magazine, 23(3):41–??, March 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:PAP

[Ano98-78] Anonymous. Performance analyzers — profilers for a component world — NuMega’s True-Time and Rational’s Visual Quantify find bottlenecks in Windows code. Byte Magazine, 23(2):46–??, February 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:PAA

[Ano98-79] Anonymous. Perl and Apache — Apache’s embedded Perl interpreter, mod-perl, integrates Web services with scripting. Byte Magazine, 23(3):97–??, March 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:PWD

[Ano98-80] Anonymous. A Perl for Web developers. Byte Magazine, 23(7): 30–??, July 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:PSN

[Ano98-81] Anonymous. Pet sounds — new technology combines three of the most annoying things about modern life: car alarms, barking dogs, and cell phones. Byte Magazine, 23(1):156–??, January 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:PDP

[Ano98-82] Anonymous. A portable device for PBX(a)Home-EXTender lets phones roam. Byte Magazine, 23(6):48–??, June 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:PM

[Ano98-83] Anonymous. PowerPC: 400 MHz in ’98. Byte Magazine, 23(4):30– ??, April 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:PSO

[Ano98-84] Anonymous. PowerSite straddles the object world — Sybase PowerSite offers agnostic Web development. Byte Magazine, 23 (1):39–??, January 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:PTV

[Ano98-85] Anonymous. Pretty thin, very wide — Gateway’s Solo 5100. Byte Magazine, 23(4):41–??, April 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:PNT

[Ano98-86] Anonymous. Previews of a new Toshiba Satellite and Softimage’s Digital Studio. Byte Magazine, 23(3):153–??, March 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:QPU

[Ano98-87] Anonymous. Quality problems — an unlikely new use for toothpaste and a call for Ugly Tech Building photos. Byte Magazine, 23(4):168–??, April 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:QCS

[Ano98-88] Anonymous. Qube is convenience squared — services for small LANs. Byte Magazine, 23(6):47–??, June 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:RTS

[Ano98-89] Anonymous. Remote terminal server — terminal compromises — Windows-based terminal server: a leap forward or a step back in time? Byte Magazine, 23(3):43–??, March 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:SS

[Ano98-90] Anonymous. Serving up storage. Byte Magazine, 23(1):72– ??, January 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:SIb

[Ano98-91] Anonymous. The Smartcard invasion. Byte Magazine, 23(1):76– ??, January 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:SIC

[Ano98-92] Anonymous. Smartcard invasion continues — security applications will be the spearhead for these “credit cards with brains.”. Byte Magazine, 23(4):112C–??, April 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:SHM

[Ano98-93] Anonymous. SmartSuite heads for the millennium — Lotus adds Web technology. Byte Magazine, 23(6):42–??, June 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:SBB

[Ano98-94] Anonymous. Software building blocks. Byte Magazine, 23 (2):40–??, February 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:SY

[Ano98-95] Anonymous. Solving for the Year 2000. Byte Magazine, 23(1):75– ??, January 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:SND

[Ano98-96] Anonymous. Sony’s 3-pound notebook, Dell’s speedy PII desktops, and Compaq’s 5-pound Armada. Byte Magazine, 23 (4):163–??, April 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:SSI

[Ano98-97] Anonymous. Speedier server I/O. Byte Magazine, 23(2): 28–??, February 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:SMP

[Ano98-98] Anonymous. Standing on a moving platform. Byte Magazine, 23 (1):54–??, January 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:SIa

[Ano98-99] Anonymous. Stop the insanity. Byte Magazine, 23(1):64–??, January 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:SSS

[Ano98-100] Anonymous. Storing smart saves space — HSM moves little-used data where it belongs: elsewhere. Byte Magazine, 23(3):104C–??, March 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:SDE

[Ano98-101] Anonymous. Sweet dreams with embedded databases. Byte Magazine, 23(6):112C–??, June 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:TCW

[Ano98-102] Anonymous. Thin clients for Windows. Byte Magazine, 23 (2):24–??, February 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:TLM

[Ano98-103] Anonymous. Thin’s in: 20 LCD monitors for your desktop. Byte Magazine, 23(7):90– ??, July 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:CFM

[Ano98-104] Anonymous. To cure a failing memory — Jerry’s unending quest for more storage space and RAM leads him into the thicket of disk and memory optimization. Byte Magazine, 23 (2):131–??, February 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:TW

[Ano98-105] Anonymous. Transacting on the Web. Byte Magazine, 23(1):81– ??, January 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:UBR

[Ano98-106] Anonymous. Unix’s bum rap.

Byte Magazine, 23(1):96I–??, January 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:VMS

[Ano98-107] Anonymous. V.90 modem sets a standard. Byte Magazine, 23(7): 37–??, July 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:VIG

[Ano98-108] Anonymous. Visual InterDev grows up fast — Microsoft’s RAD Web tool. Byte Magazine, 23(6):41–??, June 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:VMP

[Ano98-109] Anonymous. Voice-mail purgatory. Byte Magazine, 23(6):34– ??, June 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:VSB

[Ano98-110] Anonymous. VPNs for small businesses. Byte Magazine, 23 (7):21–??, July 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:WPa

[Ano98-111] Anonymous. Wanted: Programmers. Byte Magazine, 23(3):40– ??, March 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:WPb

[Ano98-112] Anonymous. Web project. Byte Magazine, 23(7):81–??, July 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:WNC

[Ano98-113] Anonymous. What’s new — Compaq’s notebook, Philips’ latest Velo, and Accent’s translation software. plus, IBM brings its “giant MR” technology to a desktop disk. Byte Magazine, 23 (2):155–??, February 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:WNW

[Ano98-114] Anonymous. What’s new — Wintel notebooks from a former Mac cloner, HP’s Kayaks, NT clustering, Web commerce tools, and more new hardware and software. Byte Magazine, 23(1):151– ??, January 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:WNH

[Ano98-115] Anonymous. What’s new: HP’s multifunction machine, Photoshop 5, graphics accelerators, doit-yourself G3 Mac upgrades, PII servers, and lots of development tools. Byte Magazine, 23(6):155– ??, June 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:WT

[Ano98-116] Anonymous. What’s up with tape? Byte Magazine, 23(4):26– ??, April 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:WPC

[Ano98-117] Anonymous. When push comes to serve. Byte Magazine, 23(7):

98–??, July 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:WWE

[Ano98-118] Anonymous. When will E-cash jingle in your E-pocket? Byte Magazine, 23(1):85–??, January 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:WWF

[Ano98-119] Anonymous. Wherefore warehouse? the future of data mining. Byte Magazine, 23(1): 88NA1–??, January 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:NN

[Ano98-120] Anonymous. Where’s NetWare headed? Eric Schmidt, CEO and chairman of the board, discusses the technology directions he’s setting for Novell. Byte Magazine, 23(1):32, January 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). First question: “What’s your greatest technology challenge at Novell?” Answer: “We need to get to the TCP/IP platform as fast as we can.”.

Anonymous:1998:WCF

[Ano98-121] Anonymous. Which compiler is fastest? our tests reveal surprising bugs and inefficiencies. Byte Magazine, 23(1):122– ??, January 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:WR

[Ano98-122] Anonymous. Win98 roadmap.

Byte Magazine, 23(6):63–??, June 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:WCG

[Ano98-123] Anonymous. Windows CE goes global — with Windows CE 2.0, you can now put Windows on color hand-helds. Byte Magazine, 23(1):40–??, January 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:WFI

[Ano98-124] Anonymous. Windows: Fast inkjet takes on lasers. Byte Magazine, 23(7):24–??, July 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:WTS

[Ano98-125] Anonymous. Windows Terminal Server gets a hand. Byte Magazine, 23(7):30–??, July 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:WPH

[Ano98-126] Anonymous. Workgroup printer — HP makes an even better impression — the new LaserJet 4000 delivers 1200 dpi, lets you manage more of the print process from your . . .. Byte Magazine, 23 (2):129–??, February 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:X

[Ano98-127] Anonymous. XML. Byte Magazine, 23(1):80–??, January 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:YBS

[Ano98-128] Anonymous. Your brain in software — Natrificial’s The Brain. Byte Magazine, 23(4):45– ??, April 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1995:OCG

[Api95a] Steve Apiki. OLE controls from the ground up: OLE controls are the technology of choice for lightweight software components under Windows. building one is easy using the Control Development Kit, but starting from scratch provides an inside look at the underlying technology. Byte Magazine, 20(3):169– ??, March 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1995:SSR

[Api95b] Steve Apiki. Simple, scalable RAID: Micropolis makes RAID both scalable and multiplatform with the Raidion LTX, a modular RAID system that works with any SCSI host. Byte Magazine, 20(2):115–??, February 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1995:VCG

[Api95c] Steve Apiki. Visual C++ goes multiplatform: With versions on Intel, mips, and Alpha; a slick, well-designed integrated environment; class-library enhancements; and a Macintosh cross-development package, Visual C++ is an impressive package. but be prepared to jump into 32-bit development with both feet. Byte Magazine, 20 (4):167–??, April 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1997:CEL

[Api97] Steve Apiki. C++ on every level — Microsoft’s Visual C++ 5.0. Byte Magazine, 22(5):47– ??, May 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Archide:1998:FCD

[Arc98] Reynaldo Archide. A flexible CPU for digital cameras — the MIPS processor architecture offers ways to extend computing power for custom applications. LSI Logic used it. Byte Magazine, 23(3):49–??, March 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Amine:1995:CIG

[ARH95] Gilbert Amine, Annette Riggio, and Ellis Hill. Caller ID goes to work. Byte Magazine, 20(1):149– ??, January 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Andrews:1996:BCF

[ATA96] Dave Andrews, Tom Thompson, and Rowland Aertker. Books and CD-ROMs: The future, according to Bill. Byte Magazine, 21(3):41–??, March 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baker:1995:PXWa

[Bak95a] Steven Baker. The power of X for Windows NT. Byte Magazine, 20(7):149–??, July 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baker:1995:PXWb

[Bak95b] Steven Baker. The power of X for Windows NT: Hummingbird’s eXceed delivers native X Window System to Windows NT. Byte Magazine, 20(7):149– ??, July 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baldazo:1995:RMP

[Bal95a]  R. Baldazo.  Red-hot 100

MHz portable Pentium (Tadpole P1000). Byte Magazine, 20(6): 165, June 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baldazo:1995:ATG

[Bal95b] Rex Baldazo. 3-D Action — three-dimensional games have changed enormously since the days of Atari Battlezone. Byte Magazine, 20(12):123–??, December 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baldazo:1995:ATL

[Bal95c] Rex Baldazo. ATM with a twist of LAN — ATM must interoperate with legacy systems, especially LANs. LAN emulation might do the trick. Byte Magazine, 20(11):255–??, November 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baldazo:1995:CPG

[Bal95d] Rex Baldazo. Color and a Pentium to go: Toshiba’s T4900CT lightweight, low-voltage Pentium powerhouse makes a great traveling companion. Byte Magazine, 20(7):153–??, July 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baldazo:1995:LAN

[Bal95e] Rex Baldazo. Local air networks: With roaming-enabled access points, good range, and high throughout, WaveLANbase wireless LANs are the ones to beat in North America. Byte Magazine, 20(6):201– ??, June 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baldazo:1995:PMS

[Bal95f] Rex Baldazo. Portable multimedia: Satellite Pro — power at the right price. Byte Magazine, 20(6):173–??, June 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baldazo:1995:RPP

[Bal95g] Rex Baldazo. Red-Hot 100-MHz portable Pentium: The Tadpole PI000 – screaming performance. Byte Magazine, 20 (6):167–??, June 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baldazo:1995:TMR

[Bal95h]  Rex Baldazo.  Tru multi-

media road warrior: IBM’s latest ThinkPad with CD-ROM, motion-video support, and mwave DSP brings multimedia to the road warrior. this new machine also moves laptops a step closer to desktops. Byte Magazine, 20(3):129–??, March 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baldazo:1995:VCL

[Bal95i] Rex Baldazo. Virtual CDs on the LAN — CD-QuickShare speeds up your shared network CD-ROM drives. Byte Magazine, 20(12):153–??, December 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baldazo:1996:BWS

[Bal96a] Rex Baldazo. Battle of the Web site builders. Byte Magazine, 21(7):157–158, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baldazo:1996:BOS

[Bal96b] Rex Baldazo. The be-all of operating systems. Byte Magazine, 21(5):163–??, May 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baldazo:1996:JA

[Bal96c] Rex Baldazo. JavaScript adventures. Byte Magazine, 21(8):117– ??, August 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baldazo:1996:MCN

[Bal96d] Rex Baldazo. Microsoft catches up with Netscape — Microsoft’s latest Web browser matches Netscape Navigator feature for feature, plus adds some tricks all its own. Byte Magazine, 21 (9):41–??, September 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baldazo:1996:NWC

[Bal96e] Rex Baldazo. Navigating with a Web compass. Byte Magazine, 21(3):97–98, March 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baldazo:1996:NT

[Bal96f] Rex Baldazo. Navigator turns 3.0. Byte Magazine, 21(8):42, August 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bannon:1997:AAA

[Ban97] Peter J. Bannon. Alpha arrives at the desktop — A low-cost version brings workstation power to desktop PCs. Byte Magazine, 22(5):59–??, May 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baran:1995:IGS

[Bar95a] Nicholas Baran. The Internet: The greatest show on earth: Ten tips for using the Internet as a strategic business tool. Byte Magazine, 20(7):69– ??, July 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baran:1995:CSL

[Bar95b] Nick Baran. Commentary: Stop, look, and listen! just because you can put it on-line doesn’t mean people want it. Byte Magazine, 20(8):218–??, August 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barker:1995:BAB

[Bar95c] Dennis Barker. 1994 BYTE awards: BYTE editors pick the 50 outstanding products or technologies of the past year. Byte Magazine, 20(1):49–??, January 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barker:1995:BPH

[Bar95d] Dennis Barker. Blasts from the past — highlights from two decades of the PC revolution. Byte Magazine, 20(8):41–??, August 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bassak:1995:BBB

[Bas95] Gil Bassak. Brainy, brawny batteries: Battery technology for mobile computing is at a crossroads. Byte Magazine, 20 (6):131–??, June 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baum:1995:AWF

[Bau95] David Baum. Assets in wonderland: Financial traders can operate almost anywhere with new data-analysis software and virtual-reality interfaces. Byte Magazine, 20(7):111–116, 118, July 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baum:1996:IWC

[Bau96] David W. Baum. Installations:

When COBOL is cool. Byte Magazine, 21(4):64–??, April 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baum:1997:IPT

[Bau97] David Baum. Intranet politics and technologies — VARs and integrators are successfully tangling with the intranet maelstrom. Byte Magazine, 22 (5):88A–??, May 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baldazo:1995:WCG

[BD95a] R. Baldazo and S. Diehl. Workgroup conferencing (groupware for Windows). Byte Magazine, 20(3):125–126, 128, March 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baldazo:1995:WCL

[BD95b] Rex Baldazo and Stanford Diehl. Workgroup conferencing: A look at two new groupware solutions for Windows: Collabra Share and Attachmate’s OpenMind. both products deliver an effective conferencing system to large and small workgroups. Open-Mind also includes document management and OLE automation features. Byte Magazine, 20(3):125–??, March 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baldazo:1996:BCB

[BDF96] Rex Baldazo, Stanford Diehl, and Rich Friedman. Books and CD-ROMs: Baby’s first Java book. Byte Magazine, 21(4):41– ??, April 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Beard:1996:DGO

[Bea96] N. Beard. Doctors go online. Byte Magazine, 21(6):7– 8, June 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Beebe:1996:LEW

[Bee96] Nelson H. F. Beebe. Letter to the editor: Web lessons learned. Byte Magazine, 21(9):20, 22, September 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Advocates use of HTML/SGML grammar checkers.

Baldazo:1995:SRI

[BEM95] Rex Baldazo, David Essex, and Stan Miastkowski. Software roundup: Industrial-strength fax servers — after some hairpulling, teeth-gnashing installations, we test network-based faxing software that can handle the heavy loads of a workgroup. Byte Magazine, 20(10):137–??, October 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Berg:1995:BOD

[Ber95] Kevin Sven Berg. Business objects done right: Using objectoriented tools to capture business processes and data is a hot trend in client/server computing. Next’s Enterprise Objects Framework is among the best at exploiting this potent application of the object paradigm. Byte Magazine, 20(4):175–176, April 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baldazo:1995:SDU

[BGR95a] R. Baldazo, R. Grehan, and D. Rowell. Sudden darkness (uninterruptible power supplies). Byte Magazine, 20(11):227–231, November 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baldazo:1995:SDW

[BGR95b] Rex Baldazo, Rick Grehan, and Dave Rowell. Sudden darkness — when the lights go out, which uninterruptible power supply is going to keep your network running? we test 31 UPS models to find the best. Byte Magazine, 20(11):227–??, November 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bender:1996:MTB

[BHKM96a] Maggi Bender, Dorothy Hudson, Jim Kane, and John Mc-Donough. 10 multifunction telephony boards — NSTL tests allin-one boards that let you dial up, dial in, send faxes, browse on- line, and do voice mail. Byte Magazine, 21(9):104–??, September 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bender:1996:ETP

[BHKM96b] Maggi Bender, Dorothy Hudson, Jim Kane, and John Mc-Donough. Eight twin-engine Pentium Pro workstations — our tests point out the peak performers and the best buys. Byte Magazine, 21(11):112–??, November 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bienvenu:1995:SDO

[Bie95] Mike Bienvenu. Systems design in object time — Objec-Time can reduce development time. Byte Magazine, 20(12): 189–??, December 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bender:1996:PPD

[BKM96] Maggi Bender, Jim Kane, and John McDonough. 19 Pentium portables do multimedia — the current crop of Pentium. Byte Magazine, 21(10):114–??, October 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Blake:1996:RGI

[Bla96] Joy-Lyn Blake. A (re)touch of genius — image manipulation gets easier and more powerful with Adobe’s Photoshop 4.0. Byte Magazine, 21(12): 48–??, December 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barker:1995:BCW

[BM95] Dennis Barker and Jeff Macclay.

Books and CD-ROMs: Working in the Code Mine: A novel about working at Microsoft. plus, a Marilyn Monroe CD-ROM. Byte Magazine, 20(10):49–??, October 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bielefeld:1997:UWM

[BMW97] Alexa Bielefeld, Rick Martin, and Brian Wachter. Using the Web to maintain legacy databases — Salvo lets you use protocols to manage mainframe data. Byte Magazine, 22(7):41– ??, July 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Borsook:1995:SC

[BN95] Paulina Borsook and Michael Nadeau. Small-office computing. Byte Magazine, 20(1):120– ??, January 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bond:1995:CBO

[Bon95a] George Bond. Commentary: Bosnia On-Line: The Internet is becoming populated with enclaves of xenophobic, crabby egotists. Byte Magazine, 20(3):250– ??, March 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bond:1995:GI

[Bon95b] George Bond. Gateways to the Internet. Byte Magazine, 20(9): 229–231, September 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Boucher:1995:DHF

[Bou95] Grant S. Boucher. Desktop Hollywood F/X — Hollywood-style computer-generated 3-D special effects are no further than your desktop computer. Byte Magazine, 20(7):103–??, July 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Boyle:1997:BMD

[Boy97] James Boyle. A blueprint for managing documents — ways to find the needle in the gigabyte haystack. Byte Magazine, 22(5): 75–??, May 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Breton:1998:RDR

[Bre98] Bob Breton. Reliable data replication. Byte Magazine, 23(4):57– ??, April 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Brickman:1997:FIA

[Bri97a] Jonathan E. Brickman. Faster Internet access — by caching name/address translations on your own system, you can speed your way to Web sites. Byte Magazine, 22(8):47–??, August 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Briggs:1997:SFI

[Bri97b] Chris Briggs. Smarter and faster I/O for servers — I2O blasts bottlenecks. Byte Magazine, 22(5): 51–??, May 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bryan:1995:BFN

[Bry95a] J. Bryan. Build a firewall (networks protection). Byte Magazine, 20(4):91–94, 96, April 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bryan:1995:FS

[Bry95b] J. Bryan. Firewalls for sale. Byte Magazine, 20(4):99–100, 102, 104, April 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bryan:1995:BFP

[Bry95c] John Bryan. Build a firewall: Protect your networks from unwanted intruders — build a firewall. Byte Magazine, 20(4):91– ??, April 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bryan:1995:CSC

[Bry95d] John Bryan. Compression scorecard: Compression’s more efficient than ever if you pick the right code. our scorecards compare compression quality and speed of seven leading choices. Byte Magazine, 20(5):107–??, May 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bryan:1995:FSH

[Bry95e] John Bryan. Firewalls for sale: Here’s a look at five different firewall products and services that you can install today. Byte Magazine, 20(4):99–??, April 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Burton:1995:SIC

[Bur95a] J. Burton. Standard issue (computer telephony integration). Byte Magazine, 20(9):201–204, 206, 208, September 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Burton:1995:BTA

[Bur95b] James Burton. Building telephony applications. Byte Magazine, 20(9):211–??, September 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Burton:1995:SI

[Bur95c] James Burton. Standard issue. Byte Magazine, 20(9): 201–??, September 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baldazo:1995:SRW

[BVN95] Rex Baldazo and Steven J.

Vaughan-Nichols. Software roundup: Web publishing made easier — HTML authoring and editing tools promise to make your documents Web-ready in no time. but few live up to their claims. Byte Magazine, 20(12): 170–??, December 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Campanale:1997:BFN

[Cam97a] Michelle Campanale. Back to the future with 9 network computers — we test the current crop of thin clients to see what they’re good for. Byte Magazine, 22(12): 124–??, December 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Campanale:1997:SPM

[Cam97b] Michelle Campanale. Smart phones make the grade — Mitsubishi’s MobileAccess and Samsung’s duette. Byte Magazine, 22 (12):47–??, December 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Campanale:1998:EHN

[Cam98a] Michelle Campanale. Eight heavy-hitting NT workstations — we stress-test 300-MHz PII workstations with the 440LX chip set. Byte Magazine, 23 (1):98–??, January 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Campanale:1998:RAS

[Cam98b] Michelle Campanale. Remote access for short attention spans: We check out Covad’s TeleSpeed ADSL service. Byte Magazine, 23(5):129–??, May 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Carls:1995:DDW

[Car95] Jim Carls. dBase does Windows: The Windows version of dBase makes all the right noises in object orientation and ease of use, and it knows how to read old dBase files. Byte Magazine, 20(1):193–194, 196, January 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Celko:1997:DZM

[CC97] Joe Celko and Jackie Celko. Double zero — 2000 means trouble for computers, including — gulp — PCs. Byte Magazine, 22(7):89–??, July 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Celko:1998:SPT

[CC98] Joe Celko and Jackie Celko.

Stored procedures: Threat or menace? — stored procedures are fine — if you want to produce slow code. Byte Magazine, 23(3):87–??, March 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Carreira:1997:FTW

[CCS97] Joao Carreira, Diamantino Costa, and Joao Gabriel Silva. Fault tolerance for Windows applications — here’s a utility that provides fault tolerance so that mission-critical 32-bit Windows applications can run reliably and continuously. Byte Magazine, 22 (2):51–??, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Campanale:1998:HHP

[CE98] Michelle Campanale and David Essex. Hand-held PCs: Computing out of your palm. Byte Magazine, 23(7):103–??, July 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cook:1995:BCI

[CFDT95] Rich Cook, Rich Friedman, Brenda Daunt, and Matt Trask. Books and CD-ROMs: Is There a God: An original work of theology, a CD-ROM to help you redesign your yard, and graphics file formats. Byte Magazine, 20(4):37–??, April 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Chappell:1997:ERW

[Cha97]  David Chappell.  Enterprise

ready 5.0 — Windows NT is gaining powerful distributed features for advanced networking. Byte Magazine, 22(5):66– ??, May 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Chinnock:1995:CGN

[Chi95]  Chris Chinnock.  Color to

go: New display technologies are coming, but LCD is still king. Byte Magazine, 20(6):115– ??, June 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Chiquoine:1996:ALD

[Chi96a] Selinda Chiquoine. Alpha learns to do Windows — Digital’s clever technology lets you run 32-bit x86 code on an Alpha machine. Byte Magazine, 21(10):56– ??, October 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Chiquoine:1996:MPP

[Chi96b] Selinda Chiquoine. Mainstreaming Pentium Pro [Bravo MS-T 6200]. Byte Magazine, 21(8):41, August 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Chiquoine:1996:NFA

[Chi96c] Selinda Chiquoine. NT flies at 333 MHz. Byte Magazine, 21(6): 65, June 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Chien:1997:VEC

[Chi97] William Chien. Video for everyone — C-Cube’s low-cost MPEG encoder can turn a desktop PC into a video-editing station. Byte Magazine, 22(1):59– ??, January 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Circello:1995:CHA

[Cir95] Joe Circello. ColdFire: A hot architecture: A new architecture for 68000 processing means low cost and high power for embedded applications. Byte Magazine, 20(5):173–??, May 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ckappell:1997:CNG

[Cka97] David Ckappell. COM+: The next generation — Microsoft’s next component object model will make it easier to create and use objects. Byte Magazine, 22 (12):99–??, December 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Chappell:1997:ADI

[CL97] David Chappell and David S.

Linthicum. ActiveX demystified: It’s invasive. it’s ubiquitous. but what, exactly, is ActiveX? Byte Magazine, 22(9):56–62, 64, September 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Contains a comparison of Java and ActiveX.

Clarkson:1995:MEH

[Cla95a] M. Clarkson. Moody’s evolving help desk (financial services agency). Byte Magazine, 20(2): 76–78, 80, February 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Clarkson:1995:EEB

[Cla95b] Mark Clarkson. Eyes, ears, and brains on a chip: Economical pattern-recognition horsepower comes to the desktop thanks to the latest generations of DSPs (digital signal processors) and dedicated processors. here’s how systems developers are using this power in commercial products. Byte Magazine, 20 (2):91–??, February 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Clarkson:1995:SFM

[Cla95c] Mark Clarkson. Solutions focus: Moody’s evolving help desk: A distinctly different approach to departmental scheduling uses genetic algorithms to maximize productivity. Byte Magazine, 20 (2):76–??, February 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Clip:1998:DME

[Cli98a] Paul Clip. DCOM: Microsoft enhances DCE — the Distributed Component Object Model is an object-oriented RPC architecture that’s built on standards, yet it’s extensible. Byte Magazine, 23(3):47–??, March 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Clip:1998:INH

[Cli98b] Paul Clip. IIOP: The next HTTP? TCP/IP version of a messaging protocol enables client/server communications over the Net. Byte Magazine, 23(1):47–??, January 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Clip:1998:NJO

[Cli98c] Paul Clip. Networks — Java’s object-oriented communications — remote method invocation is a powerful mechanism that enables flexible and scalable distributed communications among Java programs. Byte Magazine, 23(2):53–54, February 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Clip:1998:SCJ

[Cli98d] Paul Clip. Servlets: CGI the Java way: These server-oriented Java classes offer better performance than CGI programs. they bring system-independence, too. Byte Magazine, 23(5):55– 56, May 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cobb:1995:IFI

[Cob95] Stephen Cobb. Internet firewalls — an industry group is attempting to bring order to firewalls, demand for which is spurred by intrusions, building real-time error. Byte Magazine, 20(10):179– ??, October 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Colwell:1996:BPL

[Col96] Susan Colwell. Big picture, little box [NEC’s LCD300 monitor]. Byte Magazine, 21(6):66, June 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cook:1995:MML

[Coo95]  Rick Cook.  More memory

in less space: Packing silicon wafers tightly creates compact new memory devices. Byte Magazine, 20(6):197–??, June 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cook:1996:WAF

[Coo96a] Rick Cook. The way of all flash. Byte Magazine, 21(6): 99–100, 102, June 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cooper:1996:YRA

[Coo96b] Lane F. Cooper. Are you ready for ATM? Byte Magazine, 21(8):

85–86, 88, 90, August 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cooper:1996:PAC

[Coo96c] Lane F. Cooper. Playing the ATM card. Byte Magazine, 21(5):71–??, May 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Corcoran:1996:IUT

[Cor96] Cate T. Corcoran. Installations:

Unix takes off. Byte Magazine, 21(1):80–??, January 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Corcoran:1997:CHC

[Cor97] Cate T. Corcoran. Cutting-edge health care — medicine counts on IT to take the knife to paperwork and inefficiency. Byte Magazine, 22(5):88M–??, May 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

GaCote:1995:AWS

[Cot95a] Raymond Ga Cote. Apple’s Workgroup Server 9150: Apple presents its top-of-the-line PowerPC-based system as an application server. Byte Magazine, 20(1):175–176, 178, January 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

GaCote:1995:PCP

[Cot95b] Raymond Ga Cote. Prograph CPX: Purely visual: Macintosh programmers have access to a powerful visual-programming environment — and it’s not from Microsoft. Byte Magazine, 20 (1):179–??, January 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cote:1995:SCD

[Cot95c] Raymond Ga Cote. Symantec C++ differences — Symantec’s new object-oriented compiler lets you distribute compilation jobs across the network. Byte Magazine, 20(12):183–??, December 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

GaCote:1996:OSB

[Cot96a] R. Ga Cote. OpenDoc: small is beautiful. Byte Magazine, 21 (2):167–168, February 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cote:1996:NBB

[Cot96b] Raymond Ga Cote. The new OS in the blue box. Byte Magazine, 21(7):41–??, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cote:1996:OSI

[Cot96c]  Raymond Ga Cote.  Open-

Doc: Small is beautiful. Byte Magazine, 21(2):167–??, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cote:1997:MCM

[Cot97] Raymond G. A. Cote. Multiplatform CodeWarrior — Metrowerks’ CodeWarrior 10 hits the scene with development tools for some PDAs. Byte Magazine, 22 (1):47–??, January 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cox:1998:MLP

[Cox98a] Tony Cox. MkLinux: Linux for the Power Mac. Byte Magazine, 23(6):51–??, June 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cox:1998:UPD

[Cox98b] Tony Cox. Using Perl with databases: For a Unix shop, getting a Windows database connected to the Web need not be an expensive and time-consuming task. Byte Magazine, 23(5):57– ??, May 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Coyle:1998:SYE

[Coy98] Dan Coyle. Security for your eyes only: Iris-scanning security. Byte Magazine, 23(5):26, 30, May 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Chiquioine:1996:PPM

[CR96a] Selinda Chiquioine and Dave Rowell. Pentium Pro makes NT fly. Byte Magazine, 21(2):155–??, February 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Chiquoine:1996:PPM

[CR96b] S. Chiquoine and D. Rowell. Pentium Pro makes NT fly. Byte Magazine, 21(2):155–156, 158, 160, 162, February 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cranford:1996:SWR

[Cra96] Richard Cranford. The spreadsheet war, Revive — Lotus’s release of a 32-bit version of 1-2-3 for Windows revives the spreadsheet fight, but not for long. our reviewer examines why. Byte Magazine, 21(12):149–??, December 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crehan:1996:RRV

[Cre96] Rick Crehan. Real-time RAD — Verilog’s ObjectGeode does a decent job of marrying realtime control to OO-modeling standards. Byte Magazine, 21(11): 181–??, November 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cronk:1995:SFN

[Cro95] Randall D. Cronk. Solutions focus: The net that manages the mail: A new, traffic-oriented network of multimedia, multiprocessor workstations with integrated telephony gives the U.S. Postal Service the ability to handle bad weather, deal with unforeseen operational contingencies, and manage the flow of mail based on real-time information. Byte Magazine, 20(3):84– 86, 88, 90, 92, March 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cuadrado:1995:ODQa

[Cua95a] John Cuadrado. Optimize database queries: Optimizing complex queries on distributed databases. Byte Magazine, 20(7): 57–??, July 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cuadrado:1995:MSS

[Cua95b] John L. Cuadrado. Mining statistics: Statistical-pattern analysis provides the technical underpinnings to make many types of recognition systems more accurate. Byte Magazine, 20(2):97–??, February 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cuadrado:1995:ODQb

[Cua95c] John L. Cuadrado. Optimize database queries. Byte Magazine, 20(7):51–??, July 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cunningham:1995:UFE

[Cun95] Paul Cunningham. Untangling fast Ethernet cables — the fast Ethernet standard specifies a variety of cabling types used in 100-Mbps networks. Byte Magazine, 20(12):213–??, December 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dobson:1995:WNU

[DA95] R. Dobson and D. Andrews. What Notes users want (groupware requirements and prospects). Byte Magazine, 20(1):24–25, January 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Davis:1995:FFV

[Dav95] Andrew W. Davis. Face to face — videoconferencing systems from different vendors can now wor together, thanks to the H.32x standard. plus, some products that let you. Byte Magazine, 20(10):69–??, October 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Davis:1996:DS

[Dav96] Andrew W. Davis. A DSP sampler. Byte Magazine, 21(5):56– ??, May 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dawson:1995:IPN

[Daw95] J. Bruce Dawson. Intrusion protection for networks: The Internet is a two-way street that can let unauthorized users into your network. CheckPont’s FireWall-I software monitors your gateways for such security breaches. Byte Magazine, 20(4):171–172, April 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Deitel:1997:JMC

[DD97] Harvey M. Deitel and Paul J. Deitel. Java Multimedia Cyber Classroom. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, January 1997. ISBN 0-13-271974-6. ???? pp. LCCN ???? US$69.95. URL http:

//www.deitel.com;   http://
www.prenhall.com/. See review

[Die97b].

DeJesus:1995:PQI

[DeJ95a] E. X. DeJesus. Presentation quality (IBM ThinkPad 755CV). Byte Magazine, 20 (9):233, September 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dejesus:1995:GTG

[Dej95b] Edmund Dejesus. 3-D graphic technology: 3-D graphics is becoming essential in applications from financial analysis to entertainment, and new standards support the trend. Byte Magazine, 20(7):101–??, July 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dejesus:1995:GT

[Dej95c] Edmund X. Dejesus. 3-D graphics technology. Byte Magazine, 20(7):101–??, July 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dejesus:1995:DMT

[Dej95d] Edmund X. Dejesus. Data mining — there’s gold in those hills of data. Byte Magazine, 20 (10):81–??, October 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dejesus:1995:DND

[Dej95e] Edmund X. Dejesus. Data to the nth dimension — Arbor Software’s Essbase adds multidimensional analysis to familiar front ends, such as Lotus 1-2-3. Byte Magazine, 20(12):193–??, December 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dejesus:1995:DDT

[Dej95f] Edmund X. Dejesus. Dimensions of data: Tools that bring multidimensional analysis to the client/ server world. Byte Magazine, 20(4):139–??, April 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dejesus:1995:FVN

[Dej95g] Edmund X. Dejesus. Face value: New facial-recognition research is yielding fast, accurate, and commercially viable algorithms for a variety of applications. Byte Magazine, 20(2):85–??, February 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dejesus:1995:MDB

[Dej95h] Edmund X. Dejesus. Manufacturing data — big changes on the factory floor. new technology for data acquisition is changing the way that manufacturing works. Byte Magazine, 20 (12):63–??, December 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dejesus:1995:PQ

[Dej95i] Edmund X. Dejesus. Presentation quality. Byte Magazine, 20 (9):233–??, September 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dejesus:1995:SFB

[Dej95j] Edmund X. Dejesus. Solutions focus: Big OOP, no oops — moving to object-oriented development can be like diving into shark-infested waters. GTE took the plunge. here’s what happened. Byte Magazine, 20(8):74– ??, August 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

DeJesus:1996:PED

[DeJ96a] E. X. DeJesus. A place for everything [data storage]. Byte Magazine, 21(6):88–89, June 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

DeJesus:1996:CQM

[DeJ96b] Edmund X. DeJesus. Control quality — maintaining software quality when developing client/ server applications requires new tools–and new concepts. Byte Magazine, 21(9):82–??, September 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

DeJesus:1996:HTB

[DeJ96c] Edmund X. DeJesus. How to tame big networks — it’s easier when you use intelligent design, simulation, and maintenance tools. Byte Magazine, 21 (10):90–??, October 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dejesus:1996:ID

[DeJ96d] Edmund X. DeJesus. Irresistible drives. Byte Magazine, 21(6):91– 96, June 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dejesus:1996:MR

[Dej96e] Edmund X. Dejesus. The middleware riddle. Byte Magazine, 21(4):65–??, April 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

DeJesus:1996:PE

[DeJ96f] Edmund X. DeJesus. A place for everything. Byte Magazine, 21(6):88–??, June 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

DeJesus:1996:TYT

[Dej96g] Edmund X. Dejesus. Toss your TV. Byte Magazine, 21(2):50– 54, 56, 60, 62, 64, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

DeJesus:1997:PPF

[DeJ97] Edmund X. DeJesus. The pull of push — five types of “push” technology are transforming info delivery. Byte Magazine, 22(8):84– ??, August 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

DeJesus:1998:PII

[DeJ98]  Edmund X. DeJesus.  Part

II: Infinite space — holograms, molecules, and atoms will be the components of tomorrow’s storage technology. Byte Magazine, 23(2):121–??, February 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Delrossi:1996:LLM

[Del96a] R. A. Delrossi. Learn the lingo [multimedia authoring tools]. Byte Magazine, 21(3):85–86, 88, 90, March 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Delrossi:1996:LL

[Del96b] Robert A. Delrossi. Learn the lingo. Byte Magazine, 21(3):85– ??, March 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dejesus:1995:BCI

[DFT95a] Edmund X. Dejesus, Rich Friedman, and Doug Tamasanis. Books and CD-ROMs: Impossible-to-use software. Byte Magazine, 20(7):45–??, July 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dejusus:1995:BCI

[DFT95b] Edmund X. Dejusus, Rich Friedman, and Doug Tamasanis. Books and CD-ROMs: Impossible-to-use software: What’s wrong with computers, a CD-ROM for real golf fans, and — believe it or not — a Unix book that’s fun to read. Byte Magazine, 20(7):45–??, July 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Diehl:1995:WGA

[Die95a] Standford Diehl. Windows 95 Graphics Architecture: Windows 95 will eventually unify a series of graphics APIs — for video, games, and 3-D — using a new set of device-driver interfaces. Byte Magazine, 20(6):241– ??, June 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Diehl:1995:AAP

[Die95b] Stanford Diehl. Applications 95 arrive — programs that make Windows 95 worth the wait. Byte Magazine, 20(10):117–??, October 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Diehl:1995:BVW

[Die95c] Stanford Diehl. BYTE’s video workshop: BYTE sets up a digital video studio to test the viability of corporate video production on the PC desktop. using Windows tools for digital and analog video production, we address the major issues of developing and deploying video content. Byte Magazine, 20(5):97– ??, May 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Diehl:1995:WNN

[Die95d] Stanford Diehl. Windows 95:

The numbers — the new Windows architecture delivers enhanced performance, but as BYTE testing reveals, there are strings attached. Byte Magazine, 20(10):113–??, October 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Diehl:1996:FNM

[Die96a] Stanford Diehl. Forging a new medium [Strata’s MediaForge 2.0 multimedia authoring tool]. Byte Magazine, 21(7):161–162, July 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Diehl:1996:PC

[Die96b] Stanford Diehl. The paperless cubicle. Byte Magazine, 21 (3):107–108, March 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:KNK

[Die97a] Stan Diehl. Book reviews: The bible, updated. Byte Magazine, 22(11):28, November 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Review of D. E. Knuth’s third edition of Volume 1 of the Art of Computer Programming.

Diehl:1997:JCC

[Die97b] Stan Diehl. Java cyber class: Java multimedia study hall. Byte Magazine, 22(9):30, September 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Review of [DD97].

Diehl:1995:VFH

[DL95] Stanford Diehl and Greg Loveria. Video for free: High-performance desktop systems, graphics-chip innovations, and the new DCI (Display Control Interface) software layer will soon make digital video a standard feature for Windows graphics cards. Byte Magazine, 20(2):105–??, February 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dern:1998:NII

[DM98] Daniel P. Dern and Scott Mace. Network integration — the Internet reinvented — changing the way researcher work and collaborate, newly minted high-speed networks will also benefit. Byte Magazine, 23(2):89–??, February 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dobson:1995:AAD

[Dob95] Rick Dobson. Access 95 advances database design — Microsoft’s new version of Access gets more than just the Windows 95 look. replication technology helps coordinate changes. Byte Magazine, 20(12):18–??, December 1995. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Dobson:1996:MAW

[Dob96] Rick Dobson. Make Access and the Web work together — how to use Microsoft Access and ODBC drivers to produce interactive Web pages and on-line databases. notebooks are expensive, but testing shows they’re worth it. Byte Magazine, 21 (10):71–??, October 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dobson:1997:DHEa

[Dob97a] Rick Dobson. Dynamic HTML explained, part 1 — dynamic objects and client-side processing mean better Web sites and better browsing. Byte Magazine, 22 (11):53–54, November 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dobson:1997:DHEb

[Dob97b] Rick Dobson. Dynamic HTML explained, part II. Byte Magazine, 22(12):61–62, December 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dobson:1997:FFA

[Dob97c] Rick Dobson. Fast and flexible access to databases — Microsoft’s ODBCDirect gives programmers better ways of processing ODBC data sources. Byte Magazine, 22(8):53–??, August 1997. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Dobson:1997:VBW

[Dob97d] Rick Dobson. A Visual Basic for Web work — A variant of Visual Basic allows you to script and control objects on a Web page. Byte Magazine, 22(3):51– ??, March 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dobson:1998:DHE

[Dob98a] Rick Dobson. Dynamic HTML explained, part III — data binding utilizes data source objects that can cache Web pages, speeding up Internet access. Byte Magazine, 23(1):43–??, January 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dobson:1998:PEH

[Dob98b] Rick Dobson. Programming ECMAScript: The holy standard? Byte Magazine, 23(7):47– ??, July 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dobson:1998:SS

[Dob98c] Rick Dobson. Scriptlets simplified. Byte Magazine, 23(4):55– ??, April 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dobson:1998:SEY

[Dob98d] Rick Dobson. Scriptlets to energize your site — this technology could have major impact on Web developers and their clients. we look at its strengths and weaknesses. Byte Magazine, 23(4):96NA1, 96NA4, 96NA6, April 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dreystadt:1998:SJR

[Dre98] John Dreystadt. Storing Java in a relational database. Byte Magazine, 23(6):61–??, June 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dutt:1998:BQS

[Dut98] Dinesh Dutt. Bringing quality of service to IP. Byte Magazine, 23(6):49–??, June 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Diehl:1996:SGG

[DV96] Stanford Diehl and G. Armour Van Horn. Sumo graphics giants weigh in. Byte Magazine, 21(1): 129–130, 132, January 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Emmerson:1996:GEG

[EG96a] B. Emmerson and D. Greetham.

GSM’s extraordinary growth. Byte Magazine, 21(3):IS40/21– 22, IS40/24, March 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Emmerson:1996:TN

[EG96b] B. Emmerson and D. Greetham.

Teamwork on the net. Byte Magazine, 21(7):IS40/9–12, July 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Egan:1995:PL

[Ega95] David Egan. Publisher’s letter. Byte Magazine, 20(1):14–??, January 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Em:1998:LPG

[Em98] David Em. 3D Labs puts a glint in graphics world’s eye: The first single-chip geometry and lighting processor, 3D Lab’s Glint GMX 2000 screams on Windows NT. Byte Magazine, 23(5):42, May 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Emigh:1997:ACD

[Emi97] Jacqueline Emigh. All 12X CD-ROM drives are not equal — A look at the first four 12X CD-ROM drives. Byte Magazine, 22 (2):141–??, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Emmerson:1996:CWC

[Emm96] Bob Emmerson. Commentary: Wake-up call. Byte Magazine, 21(5):228–??, May 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Engelbart:1995:CDF

[Eng95] Douglas Engelbart. Commentaly: Dreaming of the future. Byte Magazine, 20(9):

330–??, September 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Engst:1997:IRG

[Eng97] Tonya Engst. The Internet robot’s guide to a Web site — how to block Netbot access. Byte Magazine, 22(5):63– ??, May 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Em:1997:ABN

[EP97] David Em and Alex Pournelle.

3-D animation blossoms on NT — three high-end programs let you create Hollywood-quality 3-D animation on an NT workstation. Byte Magazine, 22(12): 133–??, December 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Esch:1995:FM

[Esc95a] J. Esch. A fine MES. Byte Magazine, 20(12):67–68, 70, 74, 76, December 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Esch:1995:FMM

[Esc95b] Jim Esch. A fine MES — manufacturing execution systems unite factory computers from the planning level to the machine level. (A new dimension in bar codes p 68). Byte Magazine, 20 (12):67–??, December 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Esposito:1995:OSC

[Esp95] Bill Esposito. On-Line service on the cheap: Mustang Software’s popular wildcat BBS software gets increased capacity, customizable menus, V.34 modem support, and an optional utilities suite that includes a QuickBasiclike language. Byte Magazine, 20 (2):131–??, February 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Essex:1995:BBP

[Ess95] David Essex. Better business processes: Scitor links an optimized spreadsheet to flowchart objects, and — presto! — an instant simulator that most anyone can use. Byte Magazine, 20(6): 211–212, 214, June 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Evans:1998:OSE

[Eva98] Daniel Evans. The OQL standard emerges — Object Query Language offers a good solution for complex data management and distributed object computing. Byte Magazine, 23(3):55– ??, March 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Essex:1996:SIF

[EVN96a] D. Essex and S. J. Vaughan-Nichols. Self-serve information [fax-on-demand]. Byte Magazine, 21(1):133–134, 136, January 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Essex:1996:SI

[EVN96b] David Essex and Steven J.

Vaughan-Nichols. Self-serve information. Byte Magazine, 21 (1):133–??, January 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Faust:1995:DAS

[Fau95] Bruce Faust. Designing Alphabased systems: Digital’s three CPUs bring powerful possibilities to desktop and workstation systems design. Byte Magazine, 20(6):239–??, June 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fedyk:1997:LEJ

[Fed97] Roger Fedyk. Letter to the editor: Java unveiled. Byte Magazine, 22(6):19, June 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Feniello:1995:OVW

[Fen95] Mark Feniello. Open(ing) VMS to Win32 — adding Win32 support to Digital’s OS means good news for developers: They can build for both OpenVMS and NT machines. Byte Magazine, 20(11):251–??, November 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Froning:1997:MFW

[FH97] Andrew Froning and Helen Holzbaur. 13 modems: Fast for the Web — if you browse the Web, it’s time to upgrade your 14.4-Kbps modem. we test a baker’s dozen of the first 33.6 kbps modems. Byte Magazine, 22(3):120–??, March 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fillinich:1996:PFF

[Fil96] Paul Fillinich. Plan 9: Feature film to feature-rich OS. Byte Magazine, 21(3):143–144, March 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fink:1998:NIB

[Fin98] Robert Fink. Network integration — boning up on IPv6 — the 6bone global test bed will become the new Internet. Byte Magazine, 23(3):96NA–3–96NA– 8, March 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Flohr:1995:HOW

[Flo95] Udo Flohr. Hyper-G organizes the Web — this could be the next big thing on the Web: a new hypermedia architecture that makes it easier to find what you’re looking for. Byte Magazine, 20(11):59–??, November 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Flohr:1996:ECS

[Flo96a] Udo Flohr. 3D for everyone — coming soon to a screen near you: low-cost 3-D graphics. but first, some technical difficulties. Byte Magazine, 21(10):76–??, October 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Flohr:1996:EM

[Flo96b] Udo Flohr. Electric money. Byte Magazine, 21(6):74–78, 80, 82, 84, June 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Flohr:1996:PSC

[Flo96c] Udo Flohr. Put the space in cyberspace [Virtual Reality Markup Language]. Byte Magazine, 21(3):61–62, 64, March 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Flohr:1997:IIS

[Flo97] Udo Flohr. Intelligent intranets — some tips and products that will help you avoid anarchy. Byte Magazine, 22(8):69–??, August 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Flohr:1998:QL

[Flo98a] Udo Flohr. Quantum leaps.

Byte Magazine, 23(6):111–??, June 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Flohr:1998:SCD

[Flo98b] Udo Flohr. The smallest components: Deep inside IBM’s R&D labs, the future of computing looks nanomechanical and optical. Byte Magazine, 23(5):111– ??, May 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fomook:1998:DMM

[Fom98] John Fomook. Databases: Moving mission-critical apps to the Web. Byte Magazine, 23(7):49– ??, July 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Foust:1995:SFNa

[Fou95a] John Foust. 3-D step forward: New 3-D standards will enhance the applications — and even the OS — you’ll be using. Byte Magazine, 20(7):123–??, July 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Foust:1995:SFNb

[Fou95b] John Foust. 3-D steps forward — new 3-D standards will enhance the applications — and even the OS — you’ll be using. Byte Magazine, 20(7):123– ??, July 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fowles:1997:GSU

[Fow97] Ken Fowles. Globalization software: Unicode evolves: Worried about localization problems? content-enabled software, through Unicode, may be what really matters. Byte Magazine, 22(3):105, 106, 108, 110, March 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Franklin:1996:COG

[Fra96] Curtis Franklin, Jr. Commentary: Oh, grow up. Byte Magazine, 21(3):216–??, March 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Frederick:1996:WNE

[Fre96] Terry Frederick. Windows NT event logging. Byte Magazine, 21(4):183–??, April 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Friesenhahn:1995:BYO

[Fri95a] Bob Friesenhahn. Build your own WWW server: Setting up a WWW server offers real benefits for many organizations, and it’s probably easier to accomplish than you think. here’s a look at what’s involved. Byte Magazine, 20(4):83–??, April 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fritz:1995:TI

[Fri95b] Jeffrey Fritz. Tuning in to ISDN. Byte Magazine, 20(9): 273–??, September 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fritz:1995:VCI

[Fri95c] Jeffrey Fritz. Video connections:

Isochronous Ethernet and ATM race to bring video to your desktop. Byte Magazine, 20(5):113– ??, May 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fritz:1995:YCT

[Fri95d] Jeffrey Fritz. You can take it with you. Byte Magazine, 20 (9):41–??, September 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Friesenhahn:1996:FSW

[Fri96a] Bob Friesenhahn. A file system for the Web — A modified version of Sun’s NFS can deal with large amounts of Web data. Byte Magazine, 21(11):63–??, November 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fritz:1996:BB

[Fri96b] J. N. Fritz. Building bandwidth. Byte Magazine, 21(6):146–148, 150, June 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fritz:1996:CIT

[Fri96c] Jeffery N. Fritz. The complete ISDN telecommuter — A look at connection issues of using an ISDN line to support the home office. Byte Magazine, 21(12): 53–??, December 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fritz:1996:LWI

[Fri96d] Jeffrey N. Fritz. From LAN to WAN with ISDN — ISDN can extend your LAN to the Internet, remote users, and enterprise networks. an expert examines the technologies, the . . .. Byte Magazine, 21(11):104–??, November 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Friesenhahn:1997:AMP

[Fri97a]  Bob Friesenhahn.  Autoconf

makes for portable software — use of OS features and a freeware scripting utility solves application portability across various flavors of Unix. Byte Magazine, 22(11):45–46, November 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Friesenhahn:1997:EOU

[Fri97b] Bob Friesenhahn. Expect offers Unix scripting — embeddable language offers patternmatching capabilities. Byte Magazine, 22(4):63–??, April 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fritz:1997:BBS

[Fri97c] Jeffrey Fritz. Building bridges and secure connections — A look at the Point-to-Point Protocol and how it fits into the corporate network. Byte Magazine, 22 (2):55–??, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fritz:1997:BAP

[Fri97d] Jeffrey Fritz. Bulletproofing ATM, part 2 — how LAN emulation and proper switch and router configuration can establish redundancy. Byte Magazine, 22(7):55–??, July 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fritz:1997:DAA

[Fri97e] Jeffrey N. Fritz. Demystifying ATM addressing — before connecting your ATM network to the outside world, consider its addressing scheme. Byte Magazine, 22(12):57–??, December 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Friesenhahn:1998:OSR

[Fri98a] Bob Friesenhahn. Operating systems: Real time goes mainstream. new additions to the post. Byte Magazine, 23(7):39– ??, July 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Friesenhahn:1998:PWJ

[Fri98b] Bob Friesenhahn. Programming — writing JavaScript applications — some tricks for building useful applications, and some ways to get around this language’s limitations. Byte Magazine, 23(2):59–60, February 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Froning:1996:CNM

[Fro96] Andrew Froning. CD-ROM has a new MO. Byte Magazine, 21 (2):165–??, February 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fryer:1998:DWS

[Fry98] Ron Fryer. Data-warehouse scalability. Byte Magazine, 23(6):63– ??, June 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fritz:1996:MI

[FS96] Jeffrey Fritz and Salvatore Salamone. Marrying ISDN to the OS. Byte Magazine, 21(4):185– ??, April 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gaida:1996:IH

[Gai96] K. Gaida. Interactive highway.

Byte Magazine, 21(1):IS40/17– 18, IS40/20, January 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gardner:1997:QSM

[Gar97]  Stephen R. Gardner.  The

quest to standardize metadata — someday, disparate data manipulation tools will be able to exchange information. Byte Magazine, 22(11):47–??, November 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gartenberg:1998:GRW

[Gar98] Michael Gartenberg. Gartnergroup report: Win 98 or NT for the PC desktop? it depends. Byte Magazine, 23(5): 38, May 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gasteiger:1995:WGO

[Gas95] Daniel Gasteiger. Wanted: A good OS/2 spreadsheet: In the OS/2 spreadsheet duel, upstart Mesa 2 has the advanced GUI features that 1-2-3 should have. Byte Magazine, 20(6):215– ??, June 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gates:1998:VCG

[Gat98] Bill Gates. VBA and COM — Gates discusses his view of the future of application programmability. Byte Magazine, 23(3):70–??, March 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gray:1995:STM

[GE95] Jim Gray and Jeri Edwards. Scale up with TP monitors: How TP monitors enable client/ server applications to scale up. Byte Magazine, 20(4):123–??, April 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gillmor:1995:SEM

[Gil95] Steve Gillmor. Storming the enterprise — Microsoft bolsters Visual Basic 4.0 with distributedobject technology and enhanced OLE automation. but, as our reviewer explains, not all objects are alike. Byte Magazine, 20(11): 219–??, November 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gillmer:1996:NFS

[Gil96a] S. Gillmer. NetWare: fight or switch? Byte Magazine, 21(7): 127–128, 130, 132, July 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gillmor:1996:NNW

[Gil96b] S. Gillmor. Notes 4.0: now it’s Webware. Byte Magazine, 21 (4):133–136, April 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gillmor:1996:ENI

[Gil96c] Steve Gillmor. Enter the Net-Ware impostor. Byte Magazine, 21(2):163–??, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gillmor:1996:GSM

[Gil96d] Steve Gillmor. GroupWise sends a message — Novell’s Release 5 makes the choice of groupware a tougher decision. Byte Magazine, 21(11):53–??, November 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gillmor:1996:NNI

[Gil96e] Steve Gillmor. Notes 4.0: Now it’s Webware. Byte Magazine, 21(4):133–??, April 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gillmor:1996:NOW

[Gil96f] Steve Gillmor. Notes opens up to the Web — Lotus Domino delivers Notes databases to Web clients and powerful tools to Notes developers. Byte Magazine, 21(10):49–??, October 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gillmor:1996:TMP

[Gil96g] Steve Gillmor. Toward a more productive office in ’97 — Microsoft’s Office 97 sets a new standard in applications software suites. Byte Magazine, 21(12): 42–??, December 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gilmor:1996:RS

[Gil96h]  Steve Gilmor.  Reinventing

SmartSuite. Byte Magazine, 21 (5):131–??, May 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gillmor:1997:BWD

[Gil97a] Steve Gillmor. The basics of Web development — Visual Basic 5 supports easy component construction. Byte Magazine, 22(4):41–??, April 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gillmor:1997:DDA

[Gil97b] Steve Gillmor. Dial 411 for directory assistance — Novell’s latest NetWare release bets big on Java and the intranet. Byte Magazine, 22(1):137–??, January 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gillmor:1997:TBW

[Gil97c] Steve Gillmor. Toolkits for building Web applications — we look at seven sitebuilding and maintenance kits to help you choose the right one. Byte Magazine, 22(12):118–??, December 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gillmor:1998:JWM

[Gil98a] Steve Gillmor. Java for windows:

Microsoft’s Visual J++ 6.0 ties Java ever more tightly into the Windows platform. Byte Magazine, 23(5):41, May 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gillmor:1998:PSD

[Gil98b] Steve Gillmor. Productivity software — A different desktop for the workplace — Lotus’ eSuite uses JavaBeans. Byte Magazine, 23(2):41, February 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Giorgis:1995:SRN

[Gio95] Tadesse W. Giorgis. Software roundup: Networks for the enterprise: Heterogeneous enterprise networks require a robust, reliable, and secure software environment. NSTL evaluates the major server-based network operating systems, including OS/2 Lan Server, Windows NT, and two flavors of NetWare. Byte Magazine, 20(2):119–??, February 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Giorgis:1996:TNN

[Gio96a]  T. W. Giorgis.  Tame the

network [network management tools]. Byte Magazine, 21(1): 147–148, 150–151, January 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Giorgis:1996:SCP

[Gio96b] Tadesse W. Giorgis. Six comm programs that speak Windows 95. Byte Magazine, 21(7):110– 114, July 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Giorgis:1996:SRB

[Gio96c] Tadesse W. Giorgis. Software roundup: Big decision: Warp vs. Windows. Byte Magazine, 21(4):154–??, April 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Giorgis:1996:SRT

[Gio96d]  Tadesse W. Giorgis.  Soft-

ware roundup: Tame the network. Byte Magazine, 21(1):147– ??, January 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Giorgis:1997:TWY

[Gio97] Tadesse Giorgis. 8 tools for weaving your Web site — build your Web site with one of these authoring tools. Byte Magazine, 22 (1):116–??, January 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Glass:1995:HBI

[Gla95] Brett Glass. How to build an Internet app — with a bit of Visual Basic code, the author builds his own weather channel in a flash. Byte Magazine, 20(12): 211–??, December 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Glass:1996:VTA

[Gla96] Brett Glass. Visual toolkits for audio apps. Byte Magazine, 21 (4):147–??, April 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Giorgis:1995:SAL

[GM95a] T. W. Giorgis and J. Mc-Donough. 26 safeguards against LAN data loss (single-medium tape drives for backing up midrange networks). Byte Magazine, 20(3):144–148, 150, 152, 154, 156, 158, 160–161, March 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Guy:1995:PMB

[GM95b] M. Guy and J. McDonough.

29 printers that mean business (evaluation). Byte Magazine, 20 (5):154–158, 160–162, 164, 166, 168, 170–171, May 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Goodman:1995:MRW

[Goo95a] John M. Goodman. More RAM for Win 3 holdouts — some RAM doublers can help Windows 3.x manage physical and virtual RAM. and some can’t. Byte Magazine, 20(12):191–??, December 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Goodrich:1995:AMF

[Goo95b] Floyd Goodrich. AMD’s 29030 microprocessor: A flexible bus interface and the ability to sustain high performance with lowcost memory make the 29030 a good choice for the embedded market. Byte Magazine, 20 (1):225–??, January 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gould:1995:IAR

[Gou95] Lawrence Gould. If Al ran the zoo — ready to turn your operation over to a bunch of algorithms? neural networks and fuzzy logic are helping to control complex manufacturing processes. Byte Magazine, 20(12): 79–83, December 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grace:1996:BB

[Gra96] Rich Grace. Beyond benchmarketing. Byte Magazine, 21 (8):121–??, August 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1995:BNB

[Gre95a] Rick Grehan. BYTE’s new benchmarks: The benchmark picture just got a whole lot brighter: BYTE has released new cross-platform benchmarks, the BYTE Native Mode suite, for testing CPUs and FPUs. and NSTL has released its InterMark suite for testing hardware under Windows. Byte Magazine, 20(3):73–??, March 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1995:CASa

[Gre95b] Rick Grehan. Casting arrays — sometimes a wrong answer to a programming problem leads you to the right answer. Byte Magazine, 20(7):179–??, July 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1995:CASb

[Gre95c] Rick Grehan. Casting arrays:

Sometimes a wrong answer to a programming problem leads you to the right answer. Byte Magazine, 20(7):179–??, July 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1995:LGS

[Gre95d] Rick Grehan. Let’s get small — A remarkably complete embedded development system. Byte Magazine, 20(10):133–??, October 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1995:RXO

[Gre95e] Rick Grehan. Radical Xbase objects: CA-Visual objects, Computer Associate’s classy new entry into the Windows application-development field, is also a migration path for DOSbased Clipper developers. Byte Magazine, 20(4):179–??, April 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1995:SSO

[Gre95f] Rick Grehan. The software stopwatch: Optimize timeconsuming routines with highresolution software timers. Byte Magazine, 20(4):215–??, April 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1996:NMM

[Gre96a] R. Grehan. New media is the message. Byte Magazine, 21(1): 113–114, January 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1996:BBH

[Gre96b] Rick Grehan. Beyond the bouncing heads — Rogue Wave’s JFactory can get you started building Java applications quickly. Byte Magazine, 21(10):216–??, October 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1996:BB

[Gre96c] Rick Grehan. Bug-free benchmarks. Byte Magazine, 21(3):

149–150, March 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1996:CTC

[Gre96d] Rick Grehan. A C++ tool that cures VB envy — drag-and-drop C++ programming doesn’t get any better than Powersoft’s slick Optima++. Byte Magazine, 21 (10):153–??, October 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1996:CC

[Gre96e] Rick Grehan. Compiling convenience +. Byte Magazine, 21 (3):109–112, March 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1996:NMI

[Gre96f] Rick Grehan. New media is the message. Byte Magazine, 21 (1):113–??, January 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1996:NRT

[Gre96g] Rick Grehan. NT in real time — embedded-systems developers can now add Windows NT to their list of real-time OSes. Byte Magazine, 21(10):88–??, October 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1996:VAB

[Gre96h] Rick Grehan. Visual Age for BASIC… sort of — Visual Age for BASIC has more in common with Visual Basic than with the other Visual Age products. Byte Magazine, 21(11):248–??, November 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1996:VPS

[Gre96i] Rick Grehan. Visual programming for science — Rick checks out Visual Science, a mathematical programming and simulation tool. Byte Magazine, 21(12): 208–??, December 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1996:VW

[Gre96j] Rick Grehan. VxDs in Windows 95. Byte Magazine, 21(6):63– 64, June 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1996:WCD

[Gre96k] Rick Grehan. Womb-to-tomb C++ development. Byte Magazine, 21(7):159–160, July 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Greenberg:1997:MMP

[Gre97a] Ilan Greenberg. Making money with point-of-sale. Byte Magazine, 22(11):104M–??, November 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Greenberg:1997:UI

[Gre97b] Ilan Greenberg. The universal inbox. Byte Magazine, 22(11):

104S–??, November 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1997:JCV

[Gre97c] Rich Grehan. Javatalk: Convert Visual Basic to Java. Byte Magazine, 22(3):188, March 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1997:JJA

[Gre97d] Rich Grehan. Javatalk: Jetassisted Java: Optima++ developers will feel at home in Powersoft’s PowerJ for Java. Byte Magazine, 22(5):103–104, May 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1997:JRP

[Gre97e] Rich Grehan. Javatalk: Rebuilt parts: Parts for Java delivers an excellent visual programming environment for Java. Byte Magazine, 22(10):117–118, October 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1997:CSC

[Gre97f]  Rick Grehan.  Client/server

C++: Write fast, run fast — picking a visual tool to program client/server databases in C++ involves many trade-offs. we try toolkits. Byte Magazine, 22(3):139–??, March 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1997:DDC

[Gre97g]  Rick Grehan.  Delphi does

C++ — Borland’s popular RAD tool sheds its Pascal beginnings and gets a new name: C++ Builder. Byte Magazine, 22 (2):42–??, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1997:JRR

[Gre97h] Rick Grehan. Java’s RAD route to data access — developing distributed data access applications? Visual Café Pro can handle some of the ugliness for you. Byte Magazine, 22(2):192, February 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1997:JSL

[Gre97i] Rick Grehan. Javatalk: SuperCede lives up to its names. Byte Magazine, 22(4):125–126, April 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1997:JVJ

[Gre97j] Rick Grehan. Javatalk: Vibe Java: Feel that cross-platform Vibe is one thing; deploying it is something else. Byte Magazine, 22(9):89–90, September 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Praises Visix’s [Nan95g] Vibe springs-and-strut geometry management, but is severely critical of its non-standardness and lack of security.

Grehan:1997:JVA

[Gre97k] Rick Grehan. Javatalk: Visual Age for Java: IBM’s Visual Age for Java isn’t just beans. Byte Magazine, 22(7):101–102, July 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1997:JWC

[Gre97l] Rick Grehan. Javatalk: Write cosmic code: Cosmo Code, soon to be combined with SGI’s Cosmo Worlds 3-D development system, is a worthy Java IDE on its own. Byte Magazine, 22 (8):113–114, August 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1997:PGJ

[Gre97m] Rick Grehan. Poet goes Java — an OODBMS takes the plunge into Java. Byte Magazine, 22(12):113–??, December 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1997:RDM

[Gre97n] Rick Grehan. Redesigned Delphi masters networked apps — the emphasis is on Web and distributed applications. Byte Magazine, 22(5):121–??, May 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1997:RDS

[Gre97o] Rick Grehan. Review: Development software: Java Development Kit, take two. Byte Magazine, 22(4):145–146, April 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1997:UDM

[Gre97p] Rick Grehan. An uncertain discovery — Metrowerks’ Discover Programming will give you a decent introduction to Java, but it can be annoying. Byte Magazine, 22(11):101–??, November 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1997:WSA

[Gre97q] Rick Grehan. Web server activator — Microsoft’s Visual InterDev provides the means for building Active Server-based Web sites. Byte Magazine, 22 (3):41–??, March 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1997:YCS

[Gre97r] Rick Grehan. You can start developing now — Moss Micro’s programming tool guides you down the path of Visual Basic. Byte Magazine, 22(1):192– ??, January 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1998:MJP

[Gre98a] Rich Grehan. More Java persistence: A look at another Javaenable database reveals what’s in store for you. Byte Magazine, 23(1):93–94, January 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1998:OMR

[Gre98b] Rich Grehan. Object marries relational — Ardent’s Java Relational Binding turns a relational database into a Java objectoriented database management system. Byte Magazine, 23(3): 101–102, March 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1998:FDO

[Gre98c] Rick Grehan. The federated database — Objectivity/DB, a Java-based OODBMS Plus, Powersoft’s component transaction server. Byte Magazine, 23 (2):101–102, February 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1998:RJD

[Gre98d] Rick Grehan. Rapid Java development. Byte Magazine, 23 (6):109–??, June 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1998:SFW

[Gre98e] Rick Grehan. Small footprint, will travel. Byte Magazine, 23 (4):109–??, April 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1998:TBT

[Gre98f] Rick Grehan. TopLink bridges two worlds: A pure Java object-relational database system maps Java objects to relational databases. Byte Magazine, 23(5):107–108, May 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gronert:1996:GMC

[Gro96] E. Gronert. Global mobile communications. Byte Magazine, 21 (7):IS40/6–8, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Guy:1995:FRR

[Guy95] M. Guy. 16 fast, reliable RAID subsystems. Byte Magazine, 20 (9):248–252, 254, 256, 259–261, September 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gwennap:1995:WMH

[Gwe95] Linley Gwennap. Why the 615 matters — here’s our design for how IBM could build a PowerPC chip that runs x86 software. Byte Magazine, 20(11): 198–??, November 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gwenmap:1996:BCB

[Gwe96] Linley Gwenmap. Birth of a chip — A brief history of the microprocessor, from the 4004 to today’s nitro-burning rockets. Byte Magazine, 21(12):77–??, December 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hackathorn:1997:DWC

[Hac97] Richard Hackathorn. Data warehousing’s credibility crisis — “we’ve got a problem: The VP of finance is questioning our reports”. Byte Magazine, 22(8):43– ??, August 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1995:IPW

[Hal95a] Tom R. Halfhill. Intel’s P6: A worthy successor to the Pentium, the P6 furthers blurs the already fuzzy boundaries between CISC and RISC. Byte Magazine, 20(4):42–??, April 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1995:MML

[Hal95b] Tom R. Halfhill. Multimedia masterpieces — A look at the technology behind the scenes of two dazzling, movie-like adventures: Phantasmagoria and Buried in Time. Byte Magazine, 20(12):129–??, December 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1995:NPD

[Hal95c] Tom R. Halfhill. The new PC — despite faster and faster chips, today’s PC is basically yesterday’s AT. but that’s about to change. here are the technologies that will. Byte Magazine, 20 (10):52–??, October 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1995:SYA

[Hal95d] Tom R. Halfhill. See you around: A look at how the panoramicvideo technology used by Apple’s QuickTime VR and Microsoft’s Surround Video can open up a wraparound, photo-quality, virtual-reality world right on your computer screen. Byte Magazine, 20(5):85–??, May 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1995:TBP

[Hal95e] Tom R. Halfhill. The truth behind the Pentium bug: How often do the five empty cells in the Pentium’s FPU lookup table spell miscalculation? Byte Magazine, 20(3):163–??, March 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1996:AA

[Hal96a] Tom R. Halfhill. Agents and avatars. Byte Magazine, 21(2): 69–70, 72, February 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1996:AKT

[Hal96b] Tom R. Halfhill. AMD K6 takes on Intel P6. Byte Magazine, 21 (1):67–68, 70, 72, January 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1996:BBB

[Hal96c] Tom R. Halfhill. Break the bandwidth barrier — new technologies will deliver an unprecedented amount of speedy on-line bandwidth. Byte Magazine, 21 (9):68–??, September 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1996:CGE

[Hal96d] Tom R. Halfhill. CDs for the gigabyte era — thanks to DVD technology, next-generation optical discs will hold 25 times more data than current CD-ROMs. Byte Magazine, 21(10):139–??, October 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1996:IWP

[Hal96e] Tom R. Halfhill. Inside the Web PC. Byte Magazine, 21(3):44–48, 50, 52, 54, 56, March 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1996:MRF

[Hal96f] Tom R. Halfhill. Mips R5000:

Fast, affordable 3-D. Byte Magazine, 21(5):161–??, May 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1996:PRP

[Hal96g] Tom R. Halfhill. PowerPC regroups — the PowerPC may break the 300-MHz barrier before the Pentium. Byte Magazine, 21(11):101–??, November 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1996:PSD

[Hal96h] Tom R. Halfhill. PowerPC speed demon — bipolar transistors give Exponential’s 533-MHz PowerPC chip the performance of an Alpha at an affordable price. Byte Magazine, 21(12): 88NA1–??, December 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1996:WSY

[Hal96i] Tom R. Halfhill. Which side are you on? Byte Magazine, 21(5): 42–??, May 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1996:XEM

[Hal96j] Tom R. Halfhill. x86 enters the multimedia era. Byte Magazine, 21(7):59–60, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1996:XFA

[Hal96k] Tom R. Halfhill. The x86, faster with age — A guide to the PC chip choices you’ll face in the next year. Byte Magazine, 21 (11):89–??, November 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1996:GGF

[Hal96l] Tonl R. Halfhill. GUIs get a facelift. Byte Magazine, 21(7): 83–84, 86, 88, 90, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfbill:1997:GGH

[Hal97a] Tom R. Halfbill. Good-bye GUI, hello NUI — the user interface is about to undergo the biggest revolution since 1984 — how IBM’s team defeated the world champ. Byte Magazine, 22(7): 60–??, July 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1997:TWT

[Hal97b] Tom H. Halfhill. Today the Web, tomorrow the world. is Java a serious programming language? yes. will it become a Windows killer? maybe. Byte Magazine, 22(1):68–71, 74, 76, 78, 80, January 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1997:BMN

[Hal97c] Tom R. Halfhill. Beyond MMX — new 3-D power of AMD, Centaur, and Cyrix chips will challenge Intel. Byte Magazine, 22 (12):87–??, December 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1997:BPI

[Hal97d] Tom R. Halfhill. Beyond Pentium II — the first detailed look at the new breakthrough microprocessor architecture from Intel and Hewlett–Packard — and what it will mean for developers and users. Byte Magazine, 22 (12):80–??, December 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1997:SMP

[Hal97e] Tom R. Halfhill. Socket to me — the Pentium II’s proprietary slots could mean trouble for Intel’s competitors. what does it mean for users? Byte Magazine, 22(11):73–??, November 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1998:DPN

[Hal98a] Tom R. Halfhill. Disposable PCs: New technology, fragmenting standards, and a revolt against rapid obsolescence will radically change the PC landscape over the next two years. are you ready? Byte Magazine, 23 (2):62–??, February 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1998:HSJ

[Hal98b] Tom R. Halfhill. How to soup up Java: Part 1: The rap on Java is that it’s too slow. proponents are inventing ways to speed it up, including better compilers. Byte Magazine, 23(5):60–64, 66, 68, 70, 72, 74, May 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1998:IPC

[Hal98c] Tom R. Halfhill. IBM’s powerhouse chip. Byte Magazine, 23(4):51–??, April 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1998:II

[Hal98d] Tom R. Halfhill. Inside IA-64. Byte Magazine, 23(6):81– 88, June 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1998:RFB

[Hal98e] Tom R. Halfhill. RISC fights back with the Mips R12000 — SGI tweaks its chip with better bus bandwidth, wider parallelism, and stronger emphasis on FP performance. Byte Magazine, 23(1):49–??, January 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1998:SRJ

[Hal98f] Tom R. Halfhill. Survey reveals Java adoption plans. Byte Magazine, 23(3):26, 30, March 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1998:TPM

[Hal98g] Tom R. Halfhill. Today’s PCs are the most crash-prone computers ever built here’s why, and here’s how you can make yours more reliable. Byte Magazine, 23(4):60– ??, April 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hars:1996:OB

[Har96] A. Hars. Organizing Babylon.

Byte Magazine, 21(3):IS40/11– 12, IS40/14, IS40/16, IS40/18, March 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Haskin:1996:WW

[Has96] D. Haskin. Windows without walls. Byte Magazine, 21(2): 109–110, 112, 114, 116, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1996:BCP

[HCM96] Tom R. Halfhill, Rick Cook, and Jeff Macclay. Books and CD-ROMs: Pentium pin-out signals and more. Byte Magazine, 21 (1):41–??, January 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

ReeseHedberg:1995:DGR

[Hed95] Sara Reese Hedberg. The data gold rush — here’s how corporations, researchers, and scientists are using data-mining techniques to discover everything from new customers to new galaxies. Byte Magazine, 20(10):83–??, October 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Henderson:1996:PWC

[Hen96] Liza Henderson. The price of WAN connectivity. Byte Magazine, 21(5):65–??, May 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hess:1997:DRN

[Hes97] Deborah Hess. Datapro report: New Java tools cozy up to servers. Byte Magazine, 22(7):

28, July 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hettler:1996:SFE

[Het96a] M. Hettler. Smart forms for the enterprise. Byte Magazine, 21(3): 119–120, 122–123, March 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hettler:1996:LNV

[Het96b] Mark Hettler. Lotus Notes vs.

Microsoft Exchange — our testing shows that Lotus Notes remains a generation or two ahead of Microsoft Exchange, its closest groupware competitor. Byte Magazine, 21(12):112–??, December 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hettler:1996:NLC

[Het96c] Mark Hettler. New leaders of the client/server migration. Byte Magazine, 21(6):124–126, 128, 130, June 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hettler:1996:SDW

[Het96d] Mark Hettler. Serving up data on the Web — we take a close look at three Web-server programs designed to deliver database access to your browser. Byte Magazine, 21(9):112–??, September 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hettler:1996:SRS

[Het96e] Mark Hettler. Software roundup: Smart forms for the enterprise. Byte Magazine, 21(3):119–??, March 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hettler:1996:SSS

[Het96f]  Mark Hettler.  SQL server:

The sequel’s better. Byte Magazine, 21(8):131–132, August 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hettler:1996:UCT

[Het96g] Mark Hettler. Upgraded C/S tools: How much better? evaluating new versions of two client/ server toolkits: Power-Builder and Centura Team Builder (the program formerly known as SQLWindows). Byte Magazine, 21(11):177–??, November 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hettler:1997:OSY

[Het97] Mark Hettler. OLAP serves up your data — powerful on-line analytical processing servers from Arbor Software and Oracle offer serious enterprise-wide data analysis. Byte Magazine, 22 (3):130–??, March 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Heckart:1996:VGF

[HG96] Christine Heckart and Beth Gage. Voice gets framed. Byte Magazine, 21(2):89–90, 92, 94, 96, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Higgs:1995:PMW

[Hig95a] S. Higgs. Project Management for Windows (software review). Byte Magazine, 20(4):185–186, 188, 190, 192, April 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Higgs:1995:SRWa

[Hig95b] Scott Higgs. Software roundup:

Windows to the internet. Byte Magazine, 20(7):142–??, July 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Higgs:1995:SRWb

[Hig95c] Scott Higgs. Software roundup:

Windows to the Internet: Five all-in-one packages make it easy to get on the Internet, but not without some bumps along the way. Byte Magazine, 20(7):142– ??, July 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Higgs:1995:SRP

[Hig95d] Scott Higgs. Sofware roundup:

Project Management for Windows: While supporting timehonored models, such as Gantt and PERT charts, the current generation of projects managers also use newer E-mail and workgroup concepts to involve groups of people in project tracking. Byte Magazine, 20(4):185– ??, April 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hilliard:1995:OCT

[Hil95] Bill Hilliard. On-Screen color: Today’s crop of monitors are producing better color than ever. here are some tips on how you can pick the best monitor and video card to get the kind of color accuracy you need. Byte Magazine, 20(1):101–??, January 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hudson:1996:SLP

[HKM96] Dorothy Hudson, Jim Kane, and John McDonough. Six laser printers for workgroups — we test six of the latest 17- and 24-ppm network laser printers for workgroups. Byte Magazine, 21(12):100–??, December 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hudson:1997:GCB

[HKM97a] Dorothy Hudson, Jim Kane, and John McDonough. 13 graphics cards for business — we test 13 high-end graphics cards that can really speed up the Windows interface and even have some low-end 3-D features. Byte Magazine, 22(2):100–??, February 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hudson:1997:LRHa

[HKM97b] Dorothy Hudson, Jim Kane, and John McDonough. LAN report: Hardware — big screens for big jobs — we pick the best of the big-screen monitors. Byte Magazine, 22(1):104– ??, January 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1995:CFY

[HM95] Tom R. Halfhill and John Montgomery. Chip fashion — you ain’t seen nothing yet. new multimediachips, appearing in products next year, will power eyepopping graphics. Byte Magazine, 20(11):171–??, November 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hyon:1996:CIY

[HM96] Jason J. Hyon and Michael D.

Martin. CD it for yourself. Byte Magazine, 21(6):105–106, 108, 110, 112, June 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hudson:1997:LRHb

[HR97] Dorothy Hudson and Dave Rowell. Lab report: Hardware: Seven fast color printers: These midrange printers, though still pricy, give you business-quality color without trading off performance. Byte Magazine, 22(4):128–135, April 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1996:CE

[HS96] Tom R. Halfhill and Salvatore Salamone. Components everywhere. Byte Magazine, 21(1):97– 98, 102, 104, 106, January 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hubbard:1995:JAF

[Hub95] Jordan Hubbard. Not just another free Unix — FreeBSD is fast and open, it runs powerful applications, and it won’t cost you a cent. Byte Magazine, 20(12):207–??, December 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hummel:1996:UDS

[Hum96a] R. L. Hummel. UltraSCSI doubles speed. Byte Magazine, 21 (8):126–128, 130, August 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hummel:1996:AWD

[Hum96b] Rob Hummel. Affordable 3-D workstations — with dual Pentium Pros and accelerators, these PCs bring 3-D graphics down to earth. Byte Magazine, 21(12): 145–??, December 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hummel:1996:EWF

[Hum96c] Robert Hummel. Eight ways to the future — technology experts agree: Change is the only sure thing about computing in the next decade. Byte Magazine, 21 (12):85–??, December 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hummel:1996:BCL

[Hum96d] Robert L. Hummel. Browse this color laser. Byte Magazine, 21(5):119–??, May 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hummel:1996:DYD

[Hum96e] Robert L. Hummel. Ditto your data — Iomega’s Ditto 2GB external drive delivers everything you expect from a portable backup unit: convenience, low cost, ease of use. Byte Magazine, 21(11):184–??, November 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hummel:1996:DS

[Hum96f] Robert L. Hummel. Doublespeed SCSI. Byte Magazine, 21 (8):126–??, August 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hummel:1996:QPF

[Hum96g] Robert L. Hummel. Quality printing, fast and cheap — NEC’s SuperScript 860 combines great features with Adobe Print-Gear technology for good Windows performance in a low-cost laser printer. Byte Magazine, 21 (9):47–??, September 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hummel:1996:RNN

[Hum96h] Robert L. Hummel. Running on NT — NT 4.0’s Win 95 interface and leading-edge connectivity features make for a valuable upgrade — but beware those nasty installation glitches. Byte Magazine, 21(10):126–??, October 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hummel:1996:WCL

[Hum96i] Robert L. Hummel. Waterproof color that lasts — the Alps MD-4000 prints smudge-proof color images and razor-sharp text. it’s also a 600-dpi color scanner. Byte Magazine, 21(11):50–??, November 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hummel:1996:HFL

[Hum96j] Robert S. Hummel. A high five for the latest LaserJet — Hewlett–Packard’s LaserJet 5 is an impressive workgroup/ network printer. Byte Magazine, 21(9):167–??, September 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hummel:1997:MMD

[Hum97a] Robert L. Hummel. 3-D mighty mite — Dynamic Pictures’ Oxygen 202 is a single-slot PCI card that can compete with 3-D graphics accelerators. Byte Magazine, 22(2):47–??, February 1997. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Hummel:1997:MIS

[Hum97b] Robert L. Hummel. Multiprocessor intranet servers — we stresstest six SMP boxes. Byte Magazine, 22(5):106–??, May 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hummel:1997:NVN

[Hum97c] Robert L. Hummel. Notebook with a view — NEC’s Versa 6200MX is a natural for presentations. Byte Magazine, 22 (4):45–??, April 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hummel:1997:QMM

[Hum97d] Robert L. Hummel. Quickstriking mighty mite — Avatar’s Shark portable drive. Byte Magazine, 22(5):48–??, May 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hummel:1998:BBB

[Hum98a] Robert L. Hummel. Big backup on a budget, image editing tools, easy Web access, and PA-RISC notebooks. Byte Magazine, 23 (5):155–??, May 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hummel:1998:CPB

[Hum98b] Robert L. Hummel. Cheap PCs:

Bargain or blunder? they cost less than a grand: are they worth it? Byte Magazine, 23(5):125– ??, May 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hummel:1998:EWS

[Hum98c] Robert L. Hummel. Eightway servers race for the NT prize. Byte Magazine, 23(6):114– ??, June 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hurwicz:1997:PSB

[Hur97a] Mike Hurwicz. PCs strike back — armed with zero administration and NetPCs, Microsoft and Intel counterattack the NC rebels. Byte Magazine, 22(5):81– ??, May 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hurwicz:1997:RTE

[Hur97b] Mike Hurwicz. Reach out and touch everyone — global enterprise directories offer a big payoff, but the technology’s new, and the integration requires hard work. Byte Magazine, 22(12): 92NA1–??, December 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hurwicz:1997:SPS

[Hur97c] Mike Hurwicz. Superfast peripherals — SCSI is nearing its limits, and serial is going to eat its breakfast. Byte Magazine, 22(3):89–??, March 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hurwicz:1997:TYD

[Hur97d] Mike Hurwicz. Take your data to the cleaners — “dirty” data could bring your warehouse tumbling down. Byte Magazine, 22 (1):97–??, January 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hurwicz:1997:VPA

[Hur97e] Mike Hurwicz. A virtual private affair — finally, a real way to save money with the Internet: Virtual private networks. Byte Magazine, 22(7):79–??, July 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hurwicz:1998:CTM

[Hur98a] Michael Hurwicz. Cracker tracking: Tighter security with intrusion detection. Byte Magazine, 23(5):112C, 112D, 112F, 112H, 112J, May 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hurwicz:1998:SFC

[Hur98b] Michael Hurwicz. Swimming the fibre channel. Byte Magazine, 23(2):104C–??, February 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hurwicz:1998:BBC

[Hur98c]  Mike Hurwicz.  Behind the

benchmarks — can you really believe those test results the vendor gave you? Byte Magazine, 23(4):75–??, April 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Husselbaugh:1995:BYN

[Hus95] Brett Husselbaugh. Break up your network: Dividing your network into two or more parts can boost performance. Byte Magazine, 20(6):91–??, June 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hyde:1996:HMP

[Hyd96] John Hyde. How to make Pentium Pros cooperate. Byte Magazine, 21(4):177–??, April 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Irvine:1997:LEL

[IF97] Aaron Irvine and Ken Fowles. Letter to the editor: Latin-X. Byte Magazine, 22(7):16, July 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Comments and response on ISO 8859 Roman code pages, and Unicode.

Indermaur:1995:BST

[Ind95] Kurt Indermaur. Baby steps: They may not fulfill our dreams yet, but agents and smart software are beginning to help us find information and do our jobs more effectively. Byte Magazine, 20(3):97–??, March 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Jennings:1995:CWN

[Jen95] Richard Jennings. Commentary: Who needs the Internet: Heresy or a cogent view? Byte Magazine, 20(1):296–??, January 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Joch:1995:BHI

[Joc95a] Alan Joch. Beyond Hollywood:

It’s getting easier to bring digital video to your desktop. Byte Magazine, 20(5):95–??, May 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Joch:1995:HNG

[Joc95b] Alan Joch. Heartthrobs — the next generation of processors are guaranteed to get your heart, and your computer, pumping. Byte Magazine, 20(11):169–??, November 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Joch:1995:HIG

[Joc95c] Alan Joch. Herd instincts — Groupware’s technical problems are huge, but they’re being solved. Byte Magazine, 20(8):83– ??, August 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Joch:1995:HSD

[Joc95d] Alan Joch. How software doesn’t work — bad code can lead to disaster. here’s why there’s a software crisis. and here’s what you can do about it. (how to build reliable code p 50) (make quality job 1 p 54). Byte Magazine, 20 (12):48–??, December 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Joch:1995:IEK

[Joc95e] Alan Joch. Incredible expanding keyboard: IBM’s revolutionary ThinkPad 701C fits a full-size keyboard and a 10.4inch LCD screen into a subnotebook. you must see it to believe it. Byte Magazine, 20(5):119– ??, May 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Joch:1996:WPT

[Joc96a] A. Joch. What pupils teach computers. Byte Magazine, 21(7):99– 100, July 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Joch:1996:BNI

[Joc96b] Alan Joch. Brave new interface worlds. Byte Magazine, 21(7): 80–??, July 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Joch:1996:ARC

[Joc96c] Alan Joch. Is ATM ready to catch fire? Byte Magazine, 21 (8):82–83, August 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Joch:1996:KC

[Joc96d]  Alan Joch.  Killer compo-

nents. Byte Magazine, 21(1):81– ??, January 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Joch:1996:TCT

[Joc96e] Alan Joch. Tomorrow’s CPUs — they just keep getting faster. Byte Magazine, 21(11):76–??, November 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Joch:1996:VWC

[Joc96f] Alan Joch. Virtually well connected. Byte Magazine, 21 (8):93–95, August 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Jock:1996:WPT

[Joc96g] Alan Jock. What pupils teach computers. Byte Magazine, 21 (7):99–??, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Joch:1997:BAF

[Joc97a] Alan Joch. BackOffice at the forefront. Byte Magazine, 22(11): 104C–??, November 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Joch:1997:CMI

[Joc97b] Alan Joch. CTI matures — integrating computers and telephones is still tricky, but new hardware and development tools are making the job easier. Byte Magazine, 22(2):82–??, February 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Joch:1998:IGP

[Joc98a] Alan Joch. IDBC’s growing pains: Java database connectivity: Developers need more than promises. Byte Magazine, 23(5): 112M, 112N, 112P, May 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Joch:1998:SNS

[Joc98b] Alan Joch. Sales with no strings attached. Byte Magazine, 23(2): 104K–??, February 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Johnson:1996:DRM

[Joh96] Stephen P. Johnson. Direct3D revealed — Microsoft’s Direct3D API allows 3-D applications to run on a wide variety of PCs. A seasoned programmer of device drivers explains how to accomplish this task. Byte Magazine, 21(12):63–??, December 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kador:1996:UM

[Kad96] John Kador. The ultimate middleware. Byte Magazine, 21 (4):79–??, April 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kay:1995:BNF

[Kay95a] Russell Kay. Barricading the Net: Finding a safe path through the potential minefields of the Internet calls for planning. Byte Magazine, 20(4):89–??, April 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kay:1995:C

[Kay95b] Russell Kay. Collision! Byte Magazine, 20(9):199–??, September 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kay:1995:IYM

[Kay95c] Russell Kay. Improve your memory: Sorting out the dazzling variety of new memory technologies. Byte Magazine, 20(6):185– ??, June 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kay:1995:MCG

[Kay95d] Russell Kay. Matching colors:

Getting the color you want without unpleasant surprises is no simple trick. Byte Magazine, 20 (1):88–??, January 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kay:1996:AFM

[Kay96a] Russell Kay. Adobe’s finally the managing type — get unified font management on the Mac and Windows PCs with Adobe’s new ATM 4.0. Byte Magazine, 21 (9):42–??, September 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kay:1996:BBB

[Kay96b] Russell Kay. Big, bright, and beautiful. Byte Magazine, 21 (3):101–102, March 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kay:1997:CRR

[Kay97a] Russell Kay. CD-ROM review: Prevent Java attacks: Good explanations but lame advice. Byte Magazine, 22(4):38, April 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Brief review of [MF97d]. See rebuttal in [MF97b].

Kay:1997:CTS

[Kay97b] Russell Kay. Color that’s simply amazing — Epson’s nifty and thrifty color ink-jet prints super images. Byte Magazine, 22(5):124–??, May 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kay:1997:HSC

[Kay97c] Russell Kay. Hot sauce for cooking up databases — if you’ve got Salsa, you don’t need to understand database theory to build an application. Byte Magazine, 22(1):139–??, January 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kay:1998:CGW

[Kay98a] Russell Kay. Computers to go: What to expect in the future of portable computing. Byte Magazine, 23(5):80–??, May 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kay:1998:HDC

[Kay98b] Russell Kay. Hardware — 15 disks cover more data than ever — platter clatter: We test gigabyte disk drives for workstations and servers. Byte Magazine, 23 (2):112–??, February 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kay:1998:SPI

[Kay98c] Russell Kay. Storage — part I: 15 disks cover more data than ever — platter clatter: We test gigabyte disk drives for workstations and servers. Byte Magazine, 23 (2):112–??, February 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Khazam:1996:PSI

[KB96] Jonathan Khazam and Bev Bachmayer. Programming strategies for Intel’s MMX. Byte Magazine, 21(8):63–64, August 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kohn:1997:MPV

[KE97] Les Kohn and Greg Efland. Making personal video a reality — A video compressor/decompresser from C-Cube enables real-time MPEG-2 encoding for PCs. Byte Magazine, 22(12):59–??, December 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Keyes:1995:CWI

[Key95] Jessica Keyes. Commentary: So where’s the IT payoff: Flawed technology and advice can seriously affect the bottom line. Byte Magazine, 20(5):260–??, May 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kay:1995:BCR

[KFT95] Russell Kay, Rich Friedman, and Tom Thompson. Books and CD-ROMs: Risky business and private E-Mail: An on-line forum in book form, secure communications, a CD-ROM to help you travel in Europe cheaply, and a major work on Mac programming. Byte Magazine, 20(5):43– ??, May 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kane:1995:VMA

[KH95] J. Kane and H. Holzbaur. 32 V.34 modems answer the call (28.8-kbps external and PC Card modems, review). Byte Magazine, 20(6):224–228, 230–232, 234–237, June 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kane:1996:IPQ

[KH96] Jim Kane and Dorothy Hudson. 12 ink-jet printers for quality color. Byte Magazine, 21(6):114– 118, 120, 122–123, June 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Khare:1996:NBT

[Kha96] Rohit Khare. The next big thing on the Web. Byte Magazine, 21(7):42–??, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Khan:1998:TPV

[Kha98] Osman Khan. TriMedia powers a videophone. Byte Magazine, 23(6):57–??, June 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kane:1996:TLN

[KHB96a] J. Kane, D. Hudson, and M. Bender. 12 tape libraries for network backup. Byte Magazine, 21(8): 102–106, 109–111, August 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kane:1996:LRH

[KHB96b] Jim Kane, Dorthy Hudson, and Maggi Bender. LAB report: Hardware 12 tape libraries for network backup. Byte Magazine, 21(8):102–??, August 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kane:1996:RSR

[KHM^{+}96] Jim Kane, Dorothy Hudson, J. McDonough, M. Bender, and S. Colwell. 5 reliable servers for remote access. Byte Magazine, 21(7):102–108, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kinal:1996:WEA

[Kin96] George V. Kinal. Where on Earth am I? [GPS]. Byte Magazine, 21(2):175–176, 178, February 1996. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Kane:1996:NVM

[KM96a] J. Kane and J. McDonough. 13 notebooks with video muscle. Byte Magazine, 21(4):158–162, 164, 166–167, April 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kane:1996:GCQ

[KM96b] J. Kane and J. McDonough. 18 graphics cards quick on the draw. Byte Magazine, 21(2):142–145, 148, 150–151, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kane:1996:LRN

[KM96c] Jim Kane and John McDonough. Lab report: 13 notebooks with video muscle. Byte Magazine, 21(4):158–??, April 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kane:1996:LRG

[KM96d] Jim Kane and John McDonough. Lab report: 18 graphics cards quick on the draw. Byte Magazine, 21(2):142–??, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Knuth:1996:IKC

[Knu96a] Don Knuth. Interview: Knuth comments on code. Byte Magazine, 21(9):40, September 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Knutson:1996:BMN

[Knu96b] Charles D. Knutson. Building the model network — modeling and simulation provide a cure for pain. Byte Magazine, 21 (10):101–??, October 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Knutson:1996:DQA

[Knu96c] Charles D. Knutson. Developing quality — advice on developing those scary distributed applications. Byte Magazine, 21 (9):93–??, September 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Korzeniowski:1995:NHY

[Kor95] Paul Korzeniowski. Needle hunting — you can’t tell the baseball players without a scorecard. and you can’t find network users without a directory. Byte Magazine, 20(11):51–??, November 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kosiur:1996:EGI

[Kos96] Dave Kosiur. E-Mail grows up — IMAP will provide email services and options impossible with the current POP. Byte Magazine, 21(12):135–??, December 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kempf:1998:WYS

[KP98] James Kempf and Charles Perkins. What’s your service? the new Service Location Protocol enables computers to automatically locate and use resources on the network. Byte Magazine, 23(5):49– ??, May 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Krause:1998:UM

[Kra98]  Jason Krause.  Unixes for

Merced. Byte Magazine, 23(5): 25–26, May 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Krivda:1995:DDS

[Kri95] Cheryl D. Krivda. Data-mining dynamite — supercharge your data-mining projects with data cleansing, data warehouses, parallel processing, and megastorage. Byte Magazine, 20 (10):97–??, October 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Krten:1998:TSQ

[Krt98] Robert Krten. A tiny OS that scales up: QNX/Neutrino is a microkernel-based OS suitable for embedded applications. it can also be scaled up for bigger jobs. Byte Magazine, 23(5): 47–48, May 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lachal:1996:ULC

[Lac96] Laurent Lachal. Unix leads the 64-bit charge — 64-bit OSes are moving from the horizon to the desktop. Byte Magazine, 21(11): 139–??, November 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

LaPedus:1996:AIF

[LaP96a] M. LaPedus. Asian Internet fever. Byte Magazine, 21(8): IS40/16–20, August 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

LaPedus:1996:FN

[LaP96b] M. LaPedus. Future notebooks. Byte Magazine, 21(6):19– 20, June 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

LaPedus:1996:MS

[LaP96c] M. LaPedus. Motherboards smarten-up. Byte Magazine, 21 (7):IS40/13–17, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lawrence:1995:HBB

[Law95] Bill Lawrence. Hands-off backup:

Bundled with Palindrome’s automated backup software for NetWare, HP’s SureStore Tape 12000e DAT Autoloader can provide weeks of unattended network backup. Byte Magazine, 20(6):209–??, June 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lawrence:1996:NMT

[Law96] Bill Lawrence. NetWare mirror with a twist. Byte Magazine, 21 (3):103–104, March 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lazzaro:1996:WAD

[Laz96] Joseph J. Lazzaro. Windows 95: Aiding the disabled — computer users with disabilities need support from their OS to accomplish tasks independently. Byte Magazine, 21(10):59–??, October 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lazarro:1997:UHD

[Laz97a] Joseph J. Lazarro. Unix helps the disabled — long-awaited adaptive technology comes to Unix. Byte Magazine, 22(4):51– 52, April 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Discusses speech output, Braille output, screen readers, and voice recognition tools for UNIX, some of them freely available.

Lazzaro:1997:SEC

[Laz97b] Joseph J. Lazzaro. Speech enables the common desktop PC — powerful processors, cheap memory, and a hardware standard enable speech recognition and synthesis on out-of-the-box PCs. Byte Magazine, 22(12): 55–??, December 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lennon:1995:BBC

[Len95] A. J. Lennon. 62 big, beautiful color monitors. Byte Magazine, 20(1):204–206, 209–210, 212, 215–216, 218–221, January 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Leslie:1995:CAM

[Les95] Jacques Leslie. Commentary:

Ambiguity machines — computers are better at poetry than mathematics. Byte Magazine, 20 (10):250–??, October 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Levkoff:1998:RAD

[Lev98] Keith Levkoff. Remote access devices mix analog and digital. Byte Magazine, 23(1):26–27, January 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lewis:1996:LUI

[Lew96] Jamie Lewis. LDAP unites the Internet — new directory services, especially Lightweight Directory Access Protocol, extend the directory metaphor to the Internet. Byte Magazine, 21(12): 121–??, December 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Linthicum:1995:EPI

[Lin95a] David S. Linthicum. The end of programming — it’s over. finis. kaput. you’ll never have to write another line of code ever again. yeah. right. Byte Magazine, 20 (8):69–??, August 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Linthicum:1995:HBMa

[Lin95b]  David S. Linthicum.  How

best to migrate to Windows 95. Byte Magazine, 20(7):51– ??, July 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Linthicum:1995:HBMb

[Lin95c] David S. Linthicum. How best to migrate to Windows 95: Slowly, very slowly. Byte Magazine, 20(7):51–??, July 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Linthicum:1995:MBS

[Lin95d] David S. Linthicum. Make bulletproof SQL queries: Software AG’s Esperant, a front end for SQL database querying, uses an SQL “expert” to guard against syntactically flawed queries. Byte Magazine, 20(2):111–??, February 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Linthicum:1996:IEM

[Lin96a] D. S. Linthicum. 4.0 isn’t for everyone [Microsoft NT 4.0 release]. Byte Magazine, 21(7): 121–122, 124, 126, July 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Linthicum:1996:IP

[Lin96b] David S. Linthicum. Integration, not perspiration. Byte Magazine, 21(1):83–84, 86–88, 90, 92, 94, 96, January 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Linthicum:1996:MVR

[Lin96c] David S. Linthicum. Make voice response sing. Byte Magazine, 21(5):53–??, May 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Linthicum:1996:PTD

[Lin96d] David S. Linthicum. Power tools for data drilling. Byte Magazine, 21(1):143–144, 146, January 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Linthicum:1998:JES

[Lin98]  David S. Linthicum.  Java

evolves: With solid standards and tools, Java is becoming important in enterprise and embedded development. Byte Magazine, 23(1):60, January 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lennon:1995:NNR

[LM95a] A. J. Lennon and J. McDonough.

30 no-compromise notebooks (review, testing). Byte Magazine, 20(4):194–196, 199–200, 202– 204, 206–209, April 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lennon:1995:LRP

[LM95b] Anthony J. Lennon and John McDonough. Lab repoet: 16 Pentiums high on Win 95 — can’t get enough speed? we test the latest 120-and 133-MHz Pentium machines using new Windows 95-based benchmarks. (120-MHz Pentiums p 158) (32-Bit performance advantages p 158) (133-MHz Pentiums p 160) (Triton-based Pentiums p 161) (honorable mentions p 161) (how we tested p 164). Byte Magazine, 20(12):156–??, December 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lo:1997:IGE

[Lo97] Selina Lo. Inside gigabit Ethernet — Ethernet and Fibre Channel team up for 1-Gb throughput. Byte Magazine, 22(5):55– ??, May 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Logue:1996:CWI

[Log96] Bob Logue. Commentary: A word is worth a thousand pictures. Byte Magazine, 21(4):236– ??, April 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lohse:1997:RAN

[Loh97] Steen Lohse. Reliable ATM networking — LAN emulation links legacy networks and ATM segments. Byte Magazine, 22(4):55– ??, April 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Loshin:1997:CGP

[Los97a] Pete Loshin. Cryptography gets personal — six products that promise to secure your data, both over the Internet and on your desktop. Byte Magazine, 22(11):121–??, November 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Loshin:1997:NES

[Los97b] Pete Loshin. Netscape’s enterprising server upgrade — 3.0 has promising cross-platform directory support. Byte Magazine, 22(5):123–??, May 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Loshin:1997:TTW

[Los97c] Pete Loshin. Tune in, turn on the Web — marimba’s Castanet aims to solve the Web’s bandwidth problems by “broadcasting” pages and applets over a distributed transmission system. Byte Magazine, 22(2): 145–??, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Loveria:1995:GGZ

[Lov95] Greg Loveria. 3-D graphics go zoom. Byte Magazine, 20(9): 243–244, September 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Laird:1998:GGS

[LS98] Cameron Laird and Kathryn Soraiz. Get a grip on scripts. Byte Magazine, 23(6):89–??, June 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Moyer:1998:CRG

[MA98] Bill Moyer and John Arends. CPUs — RISC gets small — Motorola’s MCore embedded RISC processor has low-power modes that enable battery-juiced products to run for longer time. Byte Magazine, 23(2):55–56, February 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mackenzie:1995:DRP

[Mac95a] J. Mackenzie. Document repositories (product review). Byte Magazine, 20(4):131–132, 134, 136–138, April 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mackenzie:1995:DRN

[Mac95b] Jonathan Mackenzie. Document repositories: With Notes, users can build client/server document repositories. Byte Magazine, 20 (4):131–??, April 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mace:1998:BBB

[Mac98a] Scott Mace. Breaking bandwidth bottlenecks: Management tools can prioritize traffic, guarantee bandwidth, and cache data closer to users. Byte Magazine, 23(5): 89–??, May 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mace:1998:CCV

[Mac98b] Scott Mace. Commodity clusters — the virtual interface architecture will tie together clusters from multiple vendors. Byte Magazine, 23(4):87–??, April 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Markowitz:1995:TMS

[Mar95a] Judith Markowitz. Talking to machines — Star Trek officers talk to their computers. here’s how we can, too. (speechrecognition products p 98) (hidden Markov models p 100). Byte Magazine, 20(12):97–??, December 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Marshak:1995:CPN

[Mar95b] David Marshak. Competing platforms — now do Notes, Exchange, and GroupWise compare today? how will they compare tomorrow? Byte Magazine, 20 (8):84–??, August 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Marshall:1995:LPP

[Mar95c] T. Marshall. Lively pictures (Power Mac image editing). Byte Magazine, 20(1):171–172, 174, January 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Marshall:1995:LPH

[Mar95d] Trevor Marshall. Lively pictures: HSC Software’s Live Picture is an impressive image editor for the Power Mac. Byte Magazine, 20(1):171–??, January 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Martin:1995:CCR

[Mar95e] James Martin. Commentary: CyberDavid rocks Goliath — you don’t have to be a giant to succeed. Byte Magazine, 20(12): 268–??, December 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mauth:1996:CYP

[Mau96] R. Mauth. A consultant in your PC. Byte Magazine, 21 (1):IS40/5–6, IS40/8, IS40/10– 11, January 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mudge:1997:DVA

[MB97] Peter Mudge and Yobie Benjamin. Deja vu all over again — those holes in NT security sure look familiar. no wonder: Other OSes patched them years ago. Byte Magazine, 22(11): 81–??, November 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mirho:1997:JTC

[MC97] Charles Mirho and Tom Clements. JavaOS: Thin client, fat service — JavaOS uses a small memory footprint, yet its network-centric design lets it access large-scale services. Byte Magazine, 22(7): 53–54, July 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Monti:1998:SRN

[MC98] Gary Monti and John Clay.

SMS: Ready for NT 5.0: Microsoft SMS (Systems Management Server) 2.0 has better inventorying and distribution, but upgrading might be tricky. Byte Magazine, 23(5):46, May 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Marshall:1995:ITG

[MD95] Trevor Marshall and John Davey.

Is there a GLINT in your future?

— 3Dlabs’ chip produces fast 3-D graphics. here’s how it does those dazzling displays. Byte Magazine, 20(10):167–??, October 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Montgomery:1995:BCR

[MDM95] John Montgomery, Edmund X.

Dejesus, and Jeff Macclay. Books and CD ROMs: Batlle of the sexes, part II — the confusion of a digital culture, computer “minds,” and two CD-ROMs on getting back to nature. Byte Magazine, 20(11):49–??, November 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mehta:1995:CRFa

[Meh95a] Arun Mehta. Commentary: Radio Free Usenet. Byte Magazine, 20(7):242–??, July 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mehta:1995:CRFb

[Meh95b] Arun Mehta. Commentary: Radio Free Usenet: Broadcasting Usenet can help avoid the problems of cost and censorship. Byte Magazine, 20(7):242– ??, July 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mello:1995:TKA

[Mel95] John P. Mello, Jr. Telephony’s killer app. Byte Magazine, 20 (9):215–??, September 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

McGraw:1997:AHA

[MF97a] G. McGraw and E. Felten. Avoiding hostile applets: They’re out there. and they want to get in. Byte Magazine, 22(5):89–??, May 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

McGraw:1997:LEL

[MF97b] Gary McGraw and Ed Felton. Letter to the editor: Lame review? Byte Magazine, 22(6): 20, June 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Authors’ response to [Kay97a].

McGraw:1997:JST

[MF97c] Gary McGraw and Edward Felton. Java security and type safety — Java’s sophisticated security mechanism performs runtime safety checks before it allows a downloaded applet to execute. Byte Magazine, 22(1):63– 64, January 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

McGraw:1997:JSM

[MF97d] Gary McGraw and Edward Felton. Java Security: Managing the Risks. MindQ Publishing, Inc., 450 Springpark Place Ste. 1200, Herndon, VA 20170, USA, Tel: +1 703 708-9380, 1997. ISBN 1-57559-012-3. US$59.95.

URL http://www.mindq.com/
java/javasecu.html.  CD

ROM. See review [Kay97a].

Mace:1998:WBW

[MFDG98] Scott Mace, Udo Flohr, Rick Dobson, and Tony Graham. Weaving a better Web — the Web is falling apart due to HTML’s lack of essential features. but help is on the way. Byte Magazine, 23(3):58–??, March 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Montgomery:1996:BCT

[MG96] John Montgomery and Rick Grehan. Books and CD-ROMs: Take home — takedown. Byte Magazine, 21(2):41–??, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mann:1997:HDO

[MG97] Steve Mann and Rick Grehan. Hand-held development options — we test software for building applications for handhelds, including CE-based HPCs. Byte Magazine, 22(4):141–??, April 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Moore:1997:LEJ

[MH97a] David L. Moore and Tom R. Halfhill. Letters to the Editor: On Java. Byte Magazine, 22(4):19–20, April 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Moran:1997:LEA

[MH97b] Tom Moran and Tom R. Halfhill. Letter to the Editor: Ada speaks Java. Byte Magazine, 22(4):20, April 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1995:CWF

[Mia95a] Stan Miastkowski. Curing the Windows fax blues. Byte Magazine, 20(1):133–??, January 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1995:DEB

[Mia95b] Stan Miastkowski. Daisy-Chain Ethernet: Bring the legendary simplicity and flexibility of PhoneNet to the twisted-pair Ethernet LAN. Byte Magazine, 20(1):229–??, January 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1995:PPU

[Mia95c] Stan Miastkowski. Peer power upgrade: If your LANtastic peerto-peer network is running out of steam, Artisoft’s CorStream adds a NetWare 4.01 server wrapped in a familiar interface. Byte Magazine, 20(3):135– ??, March 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1995:PSM

[Mia95d] Stan Miastkowski. The Penthouse Suite — Microsoft Office 95 moves up to true 32-bitness, provides full OLE 2 support, and introduces the Bindery for mixing data from different applications. Byte Magazine, 20(12): 179–??, December 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1995:WW

[Mia95e]  Stan Miastkowski.  Weeding

Windows. Byte Magazine, 20 (7):133–??, July 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1995:WWS

[Mia95f]  Stan Miastkowski.  Weeding

Windows: Six new uninstaller utilities take the pain and suffering out of removing Windows applications. Byte Magazine, 20(7):133–??, July 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1997:MBH

[Mia97a] Stan Miastkowski. Maximum DAT backup — HP’s Sure-Store DAT24 offers great backup speed. Byte Magazine, 22 (5):125–??, May 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1997:CLS

[Mia97b] Stan Miastkowski. This cartridge is loaded — SyQuest reaches a new milestone in fast, cost-effective, removable-media drives. Byte Magazine, 22(3):48– ??, March 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1997:WFD

[Mia97c]  Stan Miastkowski.  World’s

fastest disk drive — Seagate makes a giant leap in hard drive speed with the 10,000-rpm Cheetah. Byte Magazine, 22(1):48– ??, January 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miller:1996:CCG

[Mil96] Justin Miller. Clean up: C++ garbage collection. Byte Magazine, 21(1):157–158, January 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Minasi:1996:UCF

[Min96] Mark Minasi. Unearthing Cairo — follow the signs to Cairo to get a glimpse of NT’s future. Byte Magazine, 21(11):145–??, November 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mysore:1996:STW

[MM96] Chandrika Mysore and Brent Melson. 6 servers tangle on the Web. Byte Magazine, 21(3):124– 128, 130, 132–133, March 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mohr:1997:LGF

[Moh97a] Jim Mohr. Linux in a gray flannel suit: It’s powerful. it’s open. it’s free. that’s why this Unix is entering corporate IS. Byte Magazine, 22(3):96NA3, 96NA6, 96NA8, March 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mohr:1997:SLU

[Moh97b] Jim Mohr. The state of Linux — an update on this 32-bit Unixlike OS, with pointers to even more Linux info. Byte Magazine, 22(1):49–50, January 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mohseni:1998:OSG

[Moh98] Piroz Mohseni. Operating systems — Glasgow enhances JavaBeans. Byte Magazine, 23(2):

51–52, February 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Montgomery:1995:GPW

[Mon95a] John Montgomery. The games people write — Microsoft’s Win32 development kit will help make designing games for Windows 95 a less-scary adventure. Byte Magazine, 20(12):135–??, December 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Montgomery:1995:OML

[Mon95b] John Montgomery. Organization made loud and clear — Net-Manage’s ECCO Pro 3.0 offers clever OLE 2.0 integration and a compelling outline metaphor for the organizationally challenged. Byte Magazine, 20(11):217–218, November 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Moncur:1996:LEA

[Mon96a] Michael Moncur. Letter to the editor: Adventures in JavaScript. Byte Magazine, 21 (12):20, December 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Montgomery:1996:SOL

[Mon96b] J. Montgomery. Sunsoft’s object lesson. Byte Magazine, 21 (2):117–120, 122, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Montgomery:1997:FSB

[Mon97] John Montgomery. Fiber in the sky — broadband satellite systems are ready to beam data around the world at multimegabit rates. so what’s the catch? Byte Magazine, 22(11): 58–??, November 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mondrus:1998:HBE

[Mon98a] Eugene Mondrus. How to build an electronic store fast. Byte Magazine, 23(5):32, 36, May 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Montgomery:1998:WR

[Mon98b] John Montgomery. Win98 roadmap. Byte Magazine, 23(6): 66–??, June 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Morgan:1997:CDC

[Mor97] Justin Morgan. Cross-platform development with Concert — here’s a way to rapidly port a Mac OS application to Windows and Unix. Byte Magazine, 22(7): 51–??, July 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Muller:1996:D

[Mul96a] Nathan Muller. Dial 1-800-Internet. Byte Magazine, 21(2): 83–84, 86, 88, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Muller:1996:DCI

[Mul96b] Nathan J. Muller. Design and conquer — intelligent tools can validate your network design and gauge its performance before you buy a single piece of equipment. Byte Magazine, 21(10):93– ??, October 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Muller:1996:MN

[Mul96c] Nathan J. Muller. Multimedia over the network. Byte Magazine, 21(3):73–74, 76–78, 80, 82, March 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Morgan:1995:DCR

[MW95] Alan Morgan and Scott Wallace. Digital cameras for real work — new digital cameras from Apple, Kodak, and Logitech find a balance between price and image quality. Byte Magazine, 20 (10):129–??, October 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mysore:1996:LRD

[Mys96a] Chandrika Mysore. Lab report: 16 drives for fast data backup. Byte Magazine, 21 (5):142–??, May 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mysore:1996:LRB

[Mys96b] Chandrika Mysore. Lab report: 20 big-picture monitors. Byte Magazine, 21(1):116–??, January 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nadeau:1995:PBB

[Nad95a] Michael Nadeau. PDAs bounce back: PDAs finally evolve into useful business tools for mobile workers. Byte Magazine, 20 (6):147–??, June 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nadeau:1995:QCD

[Nad95b] Michael Nadeau. Quad-Speed CD-ROM delivers: Four new quad-speed CD-ROM drives are affordable if you work with highspeed video or animation. Byte Magazine, 20(1):165–167, January 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nadeau:1995:SFL

[Nad95c] Michael Nadeau. Solutions focus:

Not lost in space: Xerox’s entire sales force will soon be working from virtual offices. Byte Magazine, 20(6):50–??, June 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nadeau:1997:TTA

[Nad97] Michael Nadeau. Tools for telephony apps — here’s how to find the right telephony development tool for your needs and skill level. Byte Magazine, 22 (2):91–??, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1995:BBS

[Nan95a] Barry Nance. Big Blue’s speed trip: Behind all the hype about OS/2 warp, we found a stable, fast operating system that sports an updated GUI, improved hardware support, and a Bonus Pack loaded with useful OS/2 applications. but is it good enough to displace Windows? Byte Magazine, 20(3):131–??, March 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1995:EDM

[Nan95b] Barry Nance. Enterprise database managers — the latest versions of DB2/ 2, Oracle, and SQL server are strong enough to support a business, but they differ in reliability and scalability. Byte Magazine, 20(11):213– 214, 216, November 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1995:FTS

[Nan95c] Barry Nance. File transfer on steroids: LapLink for Windows includes some additional features beyond simple file transfers, including remote access, file synchronization, a chat facility, filedelta transfers, and security features. Byte Magazine, 20(2): 129–130, February 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1995:NFA

[Nan95d] Barry Nance. NetWare 4.1 forges ahead: Can NetWare 4.1 stand up to strong competition from NT Server 3.5 and LAN server 4.O? Byte Magazine, 20(5):175– ??, May 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1995:NAW

[Nan95e] Barry Nance. Networking at warp speed. Byte Magazine, 20 (9):235–??, September 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1995:RNS

[Nan95f] Barry Nance. Roundup: Network storage economizers: Previously available only on minicomputers and mainframes, HSM (hierarchical storage management) is an increasingly popular way to economize on network storage. we took a look at three HSM products for PC LANs. Byte Magazine, 20(3):137–??, March 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1995:WOG

[Nan95g] Barry Nance. A whole other galaxy: quickly develop objectoriented applications without the shortcomings of high-level abstraction. Byte Magazine, 20 (7):155–??, July 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). See also [Gre97j].

Nance:1996:DTR

[Nan96a] B. Nance. Diagnosing token ring ailments. Byte Magazine, 21(7): 63–64, July 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1996:NSG

[Nan96b] B. Nance. NT Server’s growing pains. Byte Magazine, 21 (7):137–140, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1996:DTA

[Nan96c] Barry Nance. Diagnosing tokenring ailments. Byte Magazine, 21(7):63–??, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1996:KNS

[Nan96d] Barry Nance. Keep networks safe from viruses — networks put you at risk from viruses, an enterprise-wide strategy can help keep you safe. Byte Magazine, 21(11):167–??, November 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1996:TYA

[Nan96e] Barry Nance. Teach your apps to speak ATM. Byte Magazine, 21(8):97–101, August 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1996:WDA

[Nan96f] Barry Nance. Workgroups dock at IR port. Byte Magazine, 21(5):135–??, May 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1996:YTW

[Nan96g] Barry Nance. You talk, Warp listens — thanks to VoiceType technology, you can tell the new version of OS/2 where you want to go. Byte Magazine, 21(9): 48–??, September 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1997:BNS

[Nan97a] Barry Nance. Battle of the network superservers — Netscape’s SuiteSpot 3 vs. Microsoft’s IIS 4. Byte Magazine, 22(11):125–??, November 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1997:DYD

[Nan97b] Barry Nance. Drive your data through the middle(ware) — we test two middleware programs that take the muddle out of multiple databases. Byte Magazine, 22(12):132–??, December 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1997:YPR

[Nan97c] Barry Nance. Your PC’s ringing — answer it! — with Ether-Phone, you can turn your Ethernet LAN into a phone system easily and inexpensively. Byte Magazine, 22(2):48–??, February 1997. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1998:JCT

[Nan98a] Barry Nance. Jasmine challenges traditional databases. Byte Magazine, 23(4):129–??, April 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1998:PIR

[Nan98b] Barry Nance. Part II: Relations with your data. Byte Magazine, 23(6):100–??, June 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1998:RYD

[Nan98c] Barry Nance. Relations with your data. Byte Magazine, 23 (6):100–??, June 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1998:UGM

[Nan98d] Barry Nance. Unix gears up for Merced. SCO’s new UnixWare. Byte Magazine, 23(5):45, May 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1998:WSS

[Nan98e] Barry Nance. Workgroup server: A server with headroom. Byte Magazine, 23(7):110–??, July 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Navas:1997:LEJ

[Nav97] John Navas. Letter to the editor:

The Java race. Byte Magazine, 22(2):20, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Needleman:1995:Ec

[Nee95a] Raphael Needleman. Editorial. Byte Magazine, 20(2):10–??, February 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Needleman:1995:Ed

[Nee95b] Raphael Needleman. Editorial. Byte Magazine, 20(3):10– ??, March 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Needleman:1995:Ee

[Nee95c] Raphael Needleman. Editorial. Byte Magazine, 20(4):10– ??, April 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Needleman:1995:Ef

[Nee95d] Raphael Needleman. Editorial. Byte Magazine, 20(5):12– ??, May 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Needleman:1995:Eg

[Nee95e] Raphael Needleman. Editorial. Byte Magazine, 20(6):10– ??, June 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Needleman:1995:Eh

[Nee95f] Raphael Needleman. Editorial. Byte Magazine, 20(7):10– ??, July 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Needleman:1995:Ei

[Nee95g] Raphael Needleman. Editorial. Byte Magazine, 20(7):10– ??, ?? 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Needleman:1995:Ej

[Nee95h] Raphael Needleman. Editorial. Byte Magazine, 20(8):10–??, August 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Needleman:1995:Ek

[Nee95i] Raphael Needleman. Editorial. Byte Magazine, 20(9):10–??, September 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Needleman:1995:Ea

[Nee95j] Raphael Needleman. Editorial. Byte Magazine, 20(10):12– ??, October 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Needleman:1995:Eb

[Nee95k] Raphael Needleman. Editorial. Byte Magazine, 20(11): 12–??, November 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Needleman:1995:EBE

[Nee95l] Raphael Needleman. Editorial: The butterfly effect. Byte Magazine, 20(12):10–??, December 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Needleman:1995:TTT

[Nee95m] Raphael Needleman. Tales from the trip: Tip from a savvy world traveler on getting on-line from anywhere. Byte Magazine, 20 (6):159–??, June 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Needleman:1996:ESN

[Nee96a] Rapahel Needleman. Editorial: But sir, I have no pornograph! Byte Magazine, 21(4):10– ??, April 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Needleman:1996:BNI

[Nee96b] Raphael Needleman. BYTE’s — new interface. Byte Magazine, 21(6):14–??, June 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Needleman:1996:EKY

[Nee96c] Raphael Needleman. Editorial: Keep your eye off the ball. Byte Magazine, 21(3):12– ??, March 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Needleman:1996:ETI

[Nee96d] Raphael Needleman. Editorial: Not till it flies! Byte Magazine, 21(2):10–??, February 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Needleman:1996:EWS

[Nee96e] Raphael Needleman. Editorial: Web sites: Don’t blink. Byte Magazine, 21(1):10–??, January 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Needleman:1996:EWY

[Nee96f] Raphael Needleman. Editorial: While you’re waiting for the Web. Byte Magazine, 21(5):10– ??, May 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nelson:1995:CTC

[Nel95]  Ted Nelson.  Commentary:

Things to come — the inventor of hypertext makes his forecasts on the future of computing and the human race. Byte Magazine, 20(11):312–??, November 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Orfali:1995:CSD

[OH95] Robert Orfali and Dan Harkey. Client/server with distributed objects: CORBA 2.0 mechanisms and services pave the way for client/server computing with distributed objects. Byte Magazine, 20(4):151–??, April 1995.

CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Orfall:1995:ICS

[OHE95] Robert Orfall, Dan Harkey, and Jeri Edwards. Intergalactic client/server computing: A look at four dominant client/server paradigms. Byte Magazine, 20 (4):108–??, April 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Orfali:1997:CJO

[OHE97] Robert Orfali, Dan Harkey, and Jeri Edwards. CORBA, Java, and the Object Web: The Web is in trouble, CORBA and Java are out to save it. Byte Magazine, 22(10):95–98, 100, October 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

OMalley:1996:CCI

[O’M96]  Chris O’Malley.  CD-ROM

changers: the inside story. Byte Magazine, 21(3):113–114, March 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Orchard:1998:BPE

[Orc98] David Orchard. Better performance with exceptions in Java — you can use the language’s runtime error checking to improve a program’s performance. Byte Magazine, 23 (3):53–54, March 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Palmer:1996:RAS

[Pal96] Stewart Palmer. A RISC OS for all seasons — tired of OSes that require acres of memory and huge hard drives? Acorn’s multifaceted RISC OS fits in ROM and runs in just 4 MB of RAM. Byte Magazine, 21 (12):49–??, December 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pende:1998:MDO

[Pen98] Nigel Pende. Managing data — OLAP omnipresent — online analytical processing is popping up in spreadsheets, on the Web, and in other new locations. Byte Magazine, 23(2):88NA1–??, February 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pepper:1997:CPB

[Pep97] Jon Pepper. Corel PDA to bundle Java suite. Byte Magazine, 22(1):38, January 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pepper:1998:BFP

[Pep98] Jon Pepper. Bigger flat panels. Byte Magazine, 23(1):23, 25, January 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Preston:1997:GSG

[PF97] Holly Hubbard Preston and Udo Flohr. Globalizing software: Global from day one. Byte Magazine, 22(3):97, 98, 100, 102, 104, March 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Discusses Unicode.

Pountain:1995:CST

[PH95] Dick Pountain and Tom R.

Halfhill. CPU scorecards — technical details on nine nextgeneration chips, including three Intel challengers from AMD, Cyrix, and NexGen. Byte Magazine, 20(11):179–??, November 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Phillip:1997:ZAC

[Phi97] Mike Phillip. Zen and the art of cache maintenance — proper understanding and use of on-chip cache can improve an application’s performance. Byte Magazine, 22(3):59–??, March 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1997:WCC

[PM97] Dick Pountain and John Montgomery. Web components: Components and the Web are a match made in developer heaven. Byte Magazine, 22(8):56–60, 62, 64, 66, 68, August 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Discusses Sun’s Java Beans, Microsoft’s ActiveX, IBM’s San Francisco Project, and OMG’s CORBA.

Potter:1995:KIW

[Pot95] Caren D. Potter. Kaboom, Inc — whether it’s a factory explosion or James Dean’s fatal car crash, the computer-toting detectives from Failure Analysis find out what went wrong. Byte Magazine, 20(11):48–??, November 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1995:BLC

[Pou95a] D. Pountain. The British Library’s catalog is on-line. Byte Magazine, 20(5):62–64, 66, 68, 70, May 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1995:CLP

[Pou95b] Dick Pountain. Constraint Logic Programming: CLP’s power to tackle difficult combinatorial problems may make it the most significant commercial programming paradigm over the next five years. Byte Magazine, 20 (2):159–??, February 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1995:HSRa

[Pou95c] Dick Pountain. HP’s speedy RISC. Byte Magazine, 20 (7):175–??, July 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1995:HSRb

[Pou95d] Dick Pountain. HP’s speedy RISC: HP’s PA-8000 employs multiple function units and a radical out-of-oder execution strategy aimed at creating the fastest RISC yet. Byte Magazine, 20(7):175–??, July 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1995:JJ

[Pou95e] Dick Pountain. The joy of J. Byte Magazine, 20(9):267–??, September 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1995:OFS

[Pou95f] Dick Pountain. The Oberon/ F system: Niklaus Wirth’s Oberon/F is a lightweight, portable, object-oriented component framework for the Windows and Mac environments. Byte Magazine, 20(1):227–??, January 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1995:PSI

[Pou95g] Dick Pountain. PixelFlow: Scalable image processing — this new chip architecture takes an innovative approach to generating photo-realistic images. here’s how this rapid renderer works. Byte Magazine, 20(11):249–??, November 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1995:SFB

[Pou95h] Dick Pountain. Solutions focus: The British Library’s Catalog is On-Line: The British Library computerizes its paper catalogs. Byte Magazine, 20(5):62– ??, May 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1995:TAU

[Pou95i]  Dick Pountain.  Transport-

Triggered architectures: The ultimate expression of the RISC philosophy is the Transport-Triggered architecture. Byte Magazine, 20(2):151–??, February 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1995:IGA

[Pou95j] Jerry Pournelle. I’m game — all work and no play at Chaos Manor? not hardly. Byte Magazine, 20(12):139–??, December 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1995:PNM

[Pou95k] Jerry Pournelle. Pournelle: A new mutation — Jerry provides advice for dodging a new breed of virus and then returns to his explorations of Windows 95. Byte Magazine, 20(12):197–??, December 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1995:PCI

[Pou95l]  Jerry Pournelle.  Pournelle:

Communications issues: Jerry discusses modems and networking. Byte Magazine, 20(1):231– ??, January 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1995:PDS

[Pou95m] Jerry Pournelle. Pournelle:

Death swoops and upgrades — Jerry witnesses several test flights. first the DC/X spacecraft, and then a new Pentium, a new hard drive, and . . .. Byte Magazine, 20(10):181–??, October 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1995:PDM

[Pou95n] Jerry Pournelle. Pournelle: Digital models — Jerry returns from SIGGRAPH, explores mountains of CD-ROMs and, yes, manages to find one more problem with Windows 95. Byte Magazine, 20(11):257–??, November 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1995:PCP

[Pou95o] Jerry Pournelle. Pournelle: Of COM ports and digital frogs. Byte Magazine, 20(9):275–??, September 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1995:POO

[Pou95p] Jerry Pournelle. Pournelle: Orchids and onions: Part 1: What Jerry liked and didn’t like in 1994. Byte Magazine, 20(4):221– ??, April 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1995:PPL

[Pou95q] Jerry Pournelle. Pournelle: Privacy and liberty: AAAS’s annual meeting and Washington, DC, highlight Jerry’s month. Byte Magazine, 20(6):249–??, June 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1995:PPS

[Pou95r]  Jerry Pournelle.  Pournelle:

Prizes and surprises: A visit to Microsoft finds Jerry coming away favorably impressed. Byte Magazine, 20(5):187–??, May 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1995:PSH

[Pou95s]  Jerry Pournelle.  Pournelle:

Software-installation hell: Jerry finds installing software a pain. Byte Magazine, 20(2):165–??, February 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1995:PUA

[Pou95t] Jerry Pournelle. Pournelle: Unexpected adventures: The simplest things can be an adventure. Byte Magazine, 20(3):175– ??, March 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1995:PWAa

[Pou95u] Jerry Pournelle. Pournelle: Windows 95 arrives. Byte Magazine, 20(7):183–??, July 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1995:PWAb

[Pou95v] Jerry Pournelle. Pournelle: Windows 95 arrives: Windows 95 comes to Chaos Manor, and Jerry gets yet another lesson in networking. Byte Magazine, 20(7):183–??, July 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1995:PWP

[Pou95w] Jerry Pournelle. Pournelle: Windows 95 pastiche — Jerry continues to pry the seams of Windows 95. Byte Magazine, 20(8):159– ??, August 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1996:BCS

[Pou96a] D. Pountain. Best CASE scenario: choosing OO methods. Byte Magazine, 21(8):IS40/11– 15, August 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1996:PPB

[Pou96b] Dick Pountain. Parallel processing in bulk — A new programming model offers hope for solving problems of complexity. Byte Magazine, 21(11):71–??, November 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1996:ST

[Pou96c] Dick Pountain. StrongARM tactics. Byte Magazine, 21(1): 153–154, January 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1996:VMR

[Pou96d] Dick Pountain. VR meets reality. Byte Magazine, 21(7):93–94, 96, 98, July 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1996:WFD

[Pou96e] Dick Pountain. What’s the future of Dylan? Byte Magazine, 21(2):197–198, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1996:WV

[Pou96f] Dick Pountain. The word on VLIW. Byte Magazine, 21(4):61– ??, April 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1996:YNS

[Pou96g] Dick Pountain. Your next OS — six key issues to consider. Byte Magazine, 21(11):134–??, November 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1996:APC

[Pou96h] Jerry Pournelle. Adieu, pentafluge… hello, cyrus. Byte Magazine, 21(8): 133–??, August 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1996:DSJ

[Pou96i] Jerry Pournelle. Don’t swap– network! Jerry offers advice on upgrading and engages in another struggle over IRQ assignments, this time under Windows 3.11. Byte Magazine, 21(11): 185–??, November 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1996:HNA

[Pou96j] Jerry Pournelle. A hot night at the opera — A disk error on Cyrus’s hard drive forces Jerry to reach for several of his favorite Win 95 utilities. Byte Magazine, 21(12):153–??, December 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1996:LTC

[Pou96k] Jerry Pournelle. A little taste of crow — Jerry tracks down the real reason his new Cyrix machine crashed and finds an answer in an unexpected place. Byte Magazine, 21(9):169–??, September 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1996:ZSN

[Pou96l] Jerry Pournelle. Of Zip and Spam and NT 4.0 — Jerry finds uses that go beyond backup for Iomega’s Zip drives and takes a first look at the Windows NT 4.0 beta version. Byte Magazine, 21 (10):157–??, October 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1996:PWD

[Pou96m] Jerry Pournelle. Pournelle — when D equals E. Byte Magazine, 21(6):157–??, June 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1996:PUM

[Pou96n] Jerry Pournelle. Pournelle: An upgrade for Mrs Pournelle. Byte Magazine, 21(3):137–??, March 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1996:PDC

[Pou96o] Jerry Pournelle. Pournelle: Dead chickens and portable data. Byte Magazine, 21(1):161–??, January 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1996:PCC

[Pou96p] Jerry Pournelle. Pournelle: Of cables and cards. Byte Magazine, 21(5):153–??, May 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1996:PSB

[Pou96q] Jerry Pournelle. Pournelle: SuperCow on the beach. Byte Magazine, 21(2):185–??, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1996:PFF

[Pou96r] Jerry Pournelle. Pournelle: The fragrant and the foul. Byte Magazine, 21(4):169–??, April 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1997:APS

[Pou97a] Dick Pountain. Atomic power speeds data into the home — A new family of chips helps handle Digital Subscriber Line connections between the central office and the home. Byte Magazine, 22(8):51–??, August 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1997:CEF

[Pou97b] Dick Pountain. The component enterprise — future apps will create a more flexible client/server architecture. Byte Magazine, 22(5):93–??, May 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1997:PGP

[Pou97c] Dick Pountain. Parallel goes populist — three new technical trends could push parallel processing onto your desktop. Byte Magazine, 22(5):88NA3–??, May 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1997:FTE

[Pou97d] Jerry Pournelle. Fire three for effect! building a new dual Pentium Pro system is a snap, until you try to put it on the network. Byte Magazine, 22(12): 139–??, December 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1997:FAW

[Pou97e] Jerry Pournelle. Fooling around with the Web — new tools help build the Manor’s Web site. plus, a visit to the world’s greatest graphics show: SIGGRAPH. Byte Magazine, 22(11):129–??, November 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1997:HDH

[Pou97f] Jerry Pournelle. A hard drive and a hot Santa Ana — it’s a month of disasters at Chaos Manor: Jerry has to replace a baked hard drive, and Cyrus needs internal surgery after taking a fall. Byte Magazine, 22 (1):141–??, January 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1997:IWG

[Pou97g] Jerry Pournelle. It was a great Comdex — Comdex according to Jerry: What he liked — and a few words of warning. Byte Magazine, 22(3):143–??, March 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1997:BNF

[Pou97h] Jerry Pournelle. Of bug-hunting and a new frontier — Jerry offers his opinions on what’s best for the desktop in the new age of graphics. Byte Magazine, 22 (2):149–??, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1997:OOB

[Pou97i] Jerry Pournelle. Orchids and onions are blooming — Jerry looks back on the hits and misses of 1996. Byte Magazine, 22 (4):147–??, April 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1997:STM

[Pou97j] Jerry Pournelle. Some things make you feel stupid — Win 95 rains plagues upon Jerry. but Softimage’s graphics wizardry cheers him up. Byte Magazine, 22(8):139–??, August 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1997:THB

[Pou97k] Jerry Pournelle. Two heads are better than one — Chaos Manor gets a new multiprocessor Windows NT machine for highend graphics. Byte Magazine, 22(5):127–??, May 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1997:WSC

[Pou97l] Jerry Pournelle. A Web site for Chaos Manor — Jerry builds a Web site, installs an OverDrive chip in Pentafluge, and pays a visit to WinHEC. Byte Magazine, 22(7):125–??, July 1997.

CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1998:AML

[Pou98a]  Dick Pountain.  Amending

Moore’s law — technology, finance, and physical law may force chip transistor density to level off. Byte Magazine, 23 (3):91–??, March 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1998:TPB

[Pou98b] Dick Pountain. TeraFLOPS power — big problem to solve? first you get 9000 Pentium chips. Byte Magazine, 23(3):73– ??, March 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1998:FWM

[Pou98c] Jerry Pournelle. Four ways to more storage: Jerry tests five methods to increase storage, from adding disk drives, to RAID, to digital video. Byte Magazine, 23(5):131–??, May 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1998:WJB

[Pou98d] Jerry Pournelle. From the workbench — Jerry builds a new system that uses AMD’s K6 MMX enhanced chip. Byte Magazine, 23(1):123–??, January 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1998:GEG

[Pou98e] Jerry Pournelle. Good enough is good enough. Byte Magazine, 23(4):135–??, April 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1998:SW

[Pou98f] Jerry Pournelle. A smarter Windows 95. Byte Magazine, 23 (6):129–??, June 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Prasad:1995:WTS

[Pra95a]  Shashi Prasad.  Weaving a

thread — Solaris and Windows NT bring the power, speed, and efficiency of multithreading and symmetric multiprocessing to the desktop. Byte Magazine, 20(10):173–??, October 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Prasad:1995:WNT

[Pra95b] Shashi Prasad. Windows NT threads — A multithreaded application may actually run slower on an SMP machine than on its single-threaded equivalent. here’s how to avoid that. Byte Magazine, 20(11):253–??, November 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Prince:1995:CAA

[Pri95] Ted Prince. Commentary: ATM = After the Millennium: Don’t blindly jump on the bandwagon.

Byte Magazine, 20(6):306–??, June 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Quinn:1996:TTS

[QS96] Stephen R. Quinn and Muralidhar Sitaram. Testing, testing — simplify testing client/ server applications with shrinkwrapped and custom tools. Byte Magazine, 21(9):97–??, September 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Quinn:1997:TWF

[Qui97] Bob Quinn. “thin” Windows firms up — Citrix’s WinFrame 2.0. Byte Magazine, 22(5):41– ??, May 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Raghunathan:1997:LEJ

[Rag97] Balaji Raghunathan. Letter to the editor: Java for C++. Byte Magazine, 22(5):19–20, May 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rau:1997:CPP

[Rau97] Larry Rau. Core programming: Programming in Limbo. Byte Magazine, 22(10):53–54, October 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1995:NGV

[Rei95a] A. Reinhardt. The Newton goes vertical (phone testing equipment: automated field service). Byte Magazine, 20(1):64–66, 68, 70, 72, January 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1995:DIA

[Rei95b] Andy Reinhardt. Doing it all on one line. Byte Magazine, 20 (1):145–??, January 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1995:HMW

[Rei95c] Andy Reinhardt. From here to mobility: What you need to know to hit the road. Byte Magazine, 20(6):100–??, June 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1995:NWL

[Rei95d] Andy Reinhardt. New ways to learn: As networking, multimedia, and better software converge, corporations and schools will be offering much-improved ways of learning. Byte Magazine, 20(3):50–??, March 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1995:SFN

[Rei95e] Andy Reinhardt. Solutions focus: The Newton goes vertical: Why a maker of phone-testing equipment chose to repackage the Apple Newton MessagePad as its solution for client/serverbased automated field service. Byte Magazine, 20(1):64–??, January 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1995:YNM

[Rei95f] Andy Reinhardt. Your next mainframe: PCs are becoming as powerful as mainframes, and mainframes are being made from PC parts. but the mainframe of the future won’t be a gigantic PC; it will be software and a network. Byte Magazine, 20(5):48– 52, 54, 56, 58, May 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Renshaw:1998:BNA

[Ren98] David S. Renshaw. Building network apps — making components portable with JavaBeans — A key to integrated solutions across multiple platforms. Byte Magazine, 23(2):81–84, 86–88, February 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reynolds:1996:AWN

[Rey96] Jim Reynolds. Air-tight Windows NT. Byte Magazine, 21 (8):51–52, August 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1995:BCW

[RFD95] Andy Reinhardt, Rich Friedman, and Edmund X. Dejesus. Books and CD-ROMs: Word overload, information shortage: Four books about the Internet, CD-ROMs for learning applications and computer languages, and Nicholas Negroponte’s experiences, insights, and predictions. Byte Magazine, 20(6):45– ??, June 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Richardson:1997:CBC

[Ric97] Robert Richardson. Components battling components — software plug-ins for Windows and Macintosh can do remarkable things. but they’re also a source of conflict. Byte Magazine, 22(11): 87–??, November 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Riddick:1996:CCT

[Rid96] Wade Riddick. Commentary:

From copyrights to telerights. Byte Magazine, 21(2):248–??, February 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ranum:1998:FYF

[RM98] Marcus Ranum and Scott Mace.

Finding your firewall. Byte Magazine, 23(6):96–??, June 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Roberts:1996:GS

[Rob96] Bill Roberts. Groupware strategies. Byte Magazine, 21(7):68– 72, 74, 76, 78, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rowell:1996:R

[Row96a] D. Rowell. 3-D without RISC. Byte Magazine, 21(1):137, January 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rowell:1996:BLP

[Row96b] Dave Rowell. Best low-cost printer on the LAN. Byte Magazine, 21(1):141–??, January 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rowell:1996:NUC

[Row96c] Dave Rowell. Notebook under construction — you can reconfigure Compaq’s Armada from a 5-pound notebook to a full-featured multimedia machine. Byte Magazine, 21(10): 154–??, October 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rowell:1997:SGW

[Row97] Dave Rowell. Silicon Graphics’ Wintel killer — just when NT-based Pentium Pro PCs are starting to look like hot 3-D machines, SGI introduces the cool blue 02. Byte Magazine, 22 (1):41–??, January 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rudin:1998:WPL

[Rud98] Ken Rudin. When parallel lines meet: Parallel databases can wring the best performance out of parallel processing hardware. Byte Magazine, 23(5):81– ??, May 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rowell:1996:PPP

[RV96] Dave Rowell and Peter D.

Varhol. Powerhouse Pentium pros — we test the fastest Intel boxes on earth: 200-MHz quad-Pentium Pro servers from ALR and Compaq. Byte Magazine, 21 (9):161–??, September 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rymer:1996:MM

[Rym96] John R. Rymer. The muddle in the middle. Byte Magazine, 21(4):67–??, April 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salamon:1995:RDW

[Sal95a] Salvatore Salamon. Radio days:

Wireless links are expanding — despite obstacles. Byte Magazine, 20(6):107–??, June 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salamone:1995:ALC

[Sal95b] S. Salamone. Assets on the line (computer asset management). Byte Magazine, 20(9):37–38, 40, September 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salamone:1995:CTW

[Sal95c] Sal Salamone. Compatibility testing: Will computers replace singles bars? Byte Magazine, 20(4):286–??, April 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salamone:1995:CSC

[Sal95d] Salvatore Salamone. Control software costs: An efficient software-monitoring strategy can save organisations thousands of dollars in unnecessary licensing fees. Byte Magazine, 20(4):75– 78, 80, 82, April 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salamone:1995:IAAa

[Sal95e]  Salvatore Salamone.  ISDN

and analog access in one package. Byte Magazine, 20(7):181– ??, July 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salamone:1995:IAAb

[Sal95f] Salvatore Salamone. ISDN and analog access in one package: For users who need both an ISDN link and a traditional analog phone line, the new hybrid ISDN/analog modems may be the perfect solution. Byte Magazine, 20(7):181–??, July 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salamone:1995:MAE

[Sal95g] Salvatore Salamone. Merging ATM and Ethernet — combining ATM switches with Ethernet switching hubs economically increases network bandwidth. Byte Magazine, 20(8): 155–??, August 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salamone:1995:NBN

[Sal95h]  Salvatore Salamone.  Novell

builds a NEST. Novell’s vision of pervasive computing requires NetWare to move beyond traditional networked computers. Embedded Systems Technology brings NetWare to the broader market for controllers and embedded devices. Byte Magazine, 20(8):151–??, August 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salamone:1995:ODP

[Sal95i] Salvatore Salamone. Olympic-Size data pool — athletes aren’t the only ones in training for the Olympic Games in Atlanta. technology jocks have been running around to set. Byte Magazine, 20(10):80–??, October 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salamone:1995:SRM

[Sal95j] Salvatore Salamone. Simplifying remote management: Remote monitoring lets managers find out what’s going on at distant locations. Byte Magazine, 20(6):245–??, June 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salamone:1995:SFPa

[Sal95k] Salvatore Salamone. Solution focus: Pruning branch-office problems: The payoff is big if you replace an SNA backbone with access-routed networks. Byte Magazine, 20(7): 64NA4–??, July 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salamone:1995:SFPb

[Sal95l] Salvatore Salamone. Solutions focus: Pruning branch-office problems. Byte Magazine, 20 (7):64NA4–??, July 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salamone:1995:UWA

[Sal95m] Salvatore Salamone. Untangling wireless — advice on choosing the right technology and service for your unplugged applications. Byte Magazine, 20(12): 96NA1–??, December 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salamone:1995:VLG

[Sal95n] Salvatore Salamone. Virtual LANs get real. Byte Magazine, 20(5):181–??, May 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salamone:1996:SA

[Sal96a] S. Salamone. Stack attack. Byte Magazine, 21(2):105–106, 108, February 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salamone:1996:WSM

[Sal96b] S. Salamone. What’s the story?

[multimedia project development]. Byte Magazine, 21(3): 67–68, 70, March 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salamone:1996:HPM

[Sal96c] Salvatore Salamone. How to put mainframes on the Web. Byte Magazine, 21(6):53–54, June 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salamone:1996:IA

[Sal96d] Salvatore Salamone. Internet-Aware applications -. Byte Magazine, 21(3):151–152, March 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salamone:1996:IGG

[Sal96e] Salvatore Salamone. ISDN: Give up and go? Byte Magazine, 21 (2):75–76, 78, 80, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salamone:1996:LFL

[Sal96f] Salvatore Salamone. Life in the fast LANE [ATM application]. Byte Magazine, 21(7):53– 54, July 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salamone:1996:LU

[Sal96g] Salvatore Salamone. The lines unleashed. Byte Magazine, 21 (5):57–??, May 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salamone:1996:MMH

[Sal96h] Salvatore Salamone. Make multimedia happen. Byte Magazine, 21(3):65–??, March 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salamone:1996:MVD

[Sal96i] Salvatore Salamone. Merging voice and data. Byte Magazine, 21(2):73–??, February 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salamone:1996:MM

[Sal96j]  Salvatore Salamone.  Mid-

dle(ware) management. Byte Magazine, 21(4):71–??, April 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salamone:1996:MCF

[Sal96k] Salvatore Salamone. More control, fewer headaches. Byte Magazine, 21(1):159–160, January 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salamone:1996:NW

[Sal96l] Salvatore Salamone. The new WAN. Byte Magazine, 21(5):59– ??, May 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salamone:1996:OPD

[Sal96m] Salvatore Salamone. Opening PBX doors. Byte Magazine, 21 (3):64–??, March 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salamone:1996:WS

[Sal96n] Salvatore Salamone. What’s the story? Byte Magazine, 21(3):67– ??, March 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sartin:1998:HJO

[Sar98] Randy Sartin. HP JetSend: Offthe-cuff communications. Byte Magazine, 23(4):49–??, April 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sohal:1996:RRT

[SB96] Vik Sohal and Mitch Bunnell. A real OS for real time — LynxOS provides a good, portable environment for embedded applications. Byte Magazine, 21(9): 51–??, September 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Satchell:1996:IVS

[SC96] Stephen Satchell and H. B. J.

Clifford. Internet video: Small isn’t beautiful. Byte Magazine, 21(6):69, June 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schlack:1996:CCW

[Sch96a] Mark Schlack. Computing crossroads: Will it be Intel inside or Java everywhere? Byte Magazine, 21(11):14, November 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schlack:1996:PWP

[Sch96b] Mark Schlack. Peace on the wired planet — Flaming’s one thing, but the lack of civility online is getting out of control. Byte Magazine, 21(12):14–??, December 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schlack:1996:SFW

[Sch96c] Mark Schlack. Smart, fast, and well connected. Byte Magazine, 21(9):14–??, September 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schlack:1996:TT

[Sch96d] Mark Schlack. Triumphant technologists. Byte Magazine, 21(7): 14–??, July 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schlack:1996:WNS

[Sch96e] Mark Schlack. Wanted: New software — HTML converters for five-year-old applications: Is this the best we can hope for? Byte Magazine, 21(10):14–??, October 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schlack:1996:WWW

[Sch96f] Mark Schlack. Weird, wacky, and wonderful. Byte Magazine, 21(8):14–??, August 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schoettle:1996:IML

[Sch96g] B. Schoettle. IP-address management on LANs. Byte Magazine, 21(2):199–200, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schoettle:1996:IAM

[Sch96h]  Bob Schoettle.  IP address

management on LANs. Byte Magazine, 21(2):199–??, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schlack:1997:AOM

[Sch97a] Mark Schlack. Apple’s opening move — the IT community does indeed have an interest in seeing a revived Apple. Byte Magazine, 22(3):14–??, March 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schlack:1997:CCI

[Sch97b] Mark Schlack. Cadillacs or Cherokees? — it’s wise not to confuse newness with appropriateness. Byte Magazine, 22 (2):14–??, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schlack:1997:ERJ

[Sch97c] Mark Schlack. Editorial — raising the Java standard. Byte Magazine, 22(11):10, November 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schlack:1997:ELD

[Sch97d] Mark Schlack. Editorial: Living dangerously: The computer industry is engaged in a platform war. which side should you be on? Byte Magazine, 22(1):14, January 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Discusses Java, Network Computers, and NetPCs.

Schlack:1997:SH

[Sch97e] Mark Schlack. Serve ’em up hot. Byte Magazine, 22(5):14– ??, May 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schlack:1997:TNT

[Sch97f] Mark Schlack. Tracking new technologies. Byte Magazine, 22(4):14–??, April 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schneier:1998:CBT

[Sch98] Bruce Schneier. The crypto bomb is ticking: Cracking encryption gets easier as computers get faster and cheaper. now what? Byte Magazine, 23(5):97– ??, May 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Seachrist:1995:DIM

[Sea95a] D. Seachrist. Document image managers. Byte Magazine, 20 (5):143–144, 146, 148, 150, 152, May 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Seachrist:1995:VWD

[Sea95b] D. Seachrist. Virtual whiteboards (desktop conferencing). Byte Magazine, 20(11):233–234, 236–237, November 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Seachrist:1995:SRD

[Sea95c]  David Seachrist.  Software

roundup: Document image managers: NSTL surveyed eight document image managers for converting paper documents into electronic form and found promising programs with much room for improvement. Byte Magazine, 20(5):143–??, May 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Seachrist:1995:SRV

[Sea95d]  David Seachrist.  Software

roundup: Virtual whiteboards — desktop-conferencing — whiteboarding — software is often a more viable solution than videoconferencing. BYTE and NSTL pick the best from seven Windows programs. Byte Magazine, 20(11):233–??, November 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Seachrist:1996:DPP

[Sea96a] D. Seachrist. Desktop publishing photo finish. Byte Magazine, 21(2):179–180, 182, 184, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Seachrist:1996:HMM

[Sea96b] David Seachrist. How multimedia multitools compare — NSTL tests four authoring programs for building mixed-media applications. Byte Magazine, 21(11): 122–??, November 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Seachrist:1996:SRD

[Sea96c]  David Seachrist.  Software

roundup: Desktop publishing photo finish. Byte Magazine, 21 (2):179–??, February 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Seachrist:1996:SRH

[Sea96d]  David Seachrist.  Software

roundup: Help-desk helpers. Byte Magazine, 21(5):138–??, May 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Seachrist:1997:PBS

[Sea97] David Seachrist. Present a businesslike slide show — present your case more effectively with one of these five graphics packages. Byte Magazine, 22(2): 108–??, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Shapiro:1995:CNG

[Sha95a] Ezra Shapiro. Commentary:

Needed: A GUI revolution: GUIs have become fat — and not very intuitive. Byte Magazine, 20 (2):238–??, February 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sharp:1995:FPS

[Sha95b] Oliver Sharp. FPU precision:

A simple comparison program can help programmers who grapple with FPU imprecision. Byte Magazine, 20(1):223–??, January 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sharp:1995:GCR

[Sha95c] Oliver Sharp. The grand challenges: Researchers are beginning to tackle problems in geography, weather, and other areas that require more computing capability than today’s most powerful computers can muster. here’s a look at the biggest of these challenges and the ways in which scientists are attacking them with supercomputers. Byte Magazine, 20(2):65–??, February 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Shaw:1996:HPW

[Sha96] Andrew Shaw. How PrintGear works — Adobe’s custom processor implements a print architecture for high-performance, low-cost laser printers. Byte Magazine, 21(9):59–??, September 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Shafron:1998:SSW

[Sha98] Tom Shafron. Slip into Silverstream’s Web app services. Byte Magazine, 23(4):133–??, April 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sheldon:1995:HBP

[She95a] K. M. Sheldon. High-tech bookkeeping (personal computer applications). Byte Magazine, 20 (1):155–156, January 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sheldon:1995:HB

[She95b] Ken Sheldon. High-tech bookkeeping. Byte Magazine, 20 (1):155–??, January 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sheldon:1995:ARM

[She95c] Kenneth M. Sheldon. Au revoir, mon ami — Ami Pro gets a major upgrade, new workgroup features, and a new name, Word Pro. Byte Magazine, 20(10):127– ??, October 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Shell:1997:PDS

[She97] Jeffrey P. Shell. Python does scripts and objects — Python is a platform-independent OOP language with capabilities ranging from simple scripting to sophisticated object libraries. Byte Magazine, 22(2):63–??, February 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Shimshock:1998:CRM

[Shi98]  Susan Shimshock.  ColdFire

revisited: Motorola’s thirdgeneration embedded processor boosts performance with a RISC core. Byte Magazine, 23(5):51– ??, May 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Shoffner:1996:JBA

[Sho96] Michael Shoffner. Java’s busting out all over. Byte Magazine, 21 (5):32, May 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Silling:1997:CPP

[Sil97] John Silling. CTI, piece by piece — these basic building blocks can help you launch a successful CTI implementation. Byte Magazine, 22(2):85–??, February 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Simha:1996:SMC

[Sim96] Satya Simha. Super mario chip — inside the 64-bit RISC processor that powers the new Nintendo game machine. Byte Magazine, 21(12):59–??, December 1996. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Simon:1997:BCB

[Sim97] Alan Simon. Better clients, better decisions — the right tools can turn all that data into useful information. Byte Magazine, 22(1):91–??, January 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Singleton:1995:CWS

[Sin95a] Andrew Singleton. Cash on the wirehead: Six new pavement systems aim to move your money across the Internet. Byte Magazine, 20(6):71–??, June 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Singleton:1995:VS

[Sin95b] Andrew Singleton. The virtual storefront. Byte Magazine, 20 (1):125–??, January 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Singleton:1996:HCJ

[Sin96a] Andrew Singleton. A hot cup of Java. Byte Magazine, 21(4):129– ??, April 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Singleton:1996:WW

[Sin96b] Andrew Singleton. Wired on the Web. Byte Magazine, 21(1): 77–78, 80, January 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Discusses Java and HotJava.

Singleton:1995:BCC

[SK95] Andrew Singleton and Russell Kay. Books and CD-ROMs: Creating life in a computer: A look at books on artificial life and genetic programming and a CD-ROM on warplanes. Byte Magazine, 20(1):45–??, January 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Seachrist:1996:MSV

[SK96a] D. Seachrist and R. Kay. Microsoft’s suite victory [Office 95/Office Professional]. Byte Magazine, 21(8):112–116, August 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Seachrist:1996:LRS

[SK96b] David Seachrist and Russel Kay.

LAB report: Software — Microsoft’s suite victory. Byte Magazine, 21(8):112–??, August 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Skeen:1998:DRQ

[Ske98a] Dale Skeen. Databases — realtime queries in the enterprise — new forms of persistent queries are necessary to handle live data as it speeds through a network. Byte Magazine, 23(2): 47–??, February 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Skeen:1998:ERE

[Ske98b]  Dale Skeen.  Enabling the

real-time enterprise — A new communications infrastructure moves critical data through intranets at nearly real-time speed. Byte Magazine, 23(1):51–??, January 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stern:1997:MLE

[SM97] Morgan Stern and Gary Monti. Microsoft laces up for the enterprise race — enterprise-ready Windows NT 4.0. Byte Magazine, 22(12):45–??, December 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1995:ISN

[Smi95a] Ben Smith. Internet with style: Netscape Communication’s free Mosaic browser is optimized for 14.4-Kpbs modem connections. versions of its commercial Mosaic server offer security features to enable commerce on the Internet. Byte Magazine, 20(1):197– ??, January 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1995:SSM

[Smi95b] Ben Smith. Server with a slot: Microplex’s M204 multiprotocol print server is winning the features-for-the-price war among stand-alone print servers, and it has little to do with the unit’s PCMCIA network adapter design. Byte Magazine, 20(4):181– ??, April 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1995:SOI

[Smi95c] Ben Smith. Short order Internet access. Byte Magazine, 20(7):139–??, July 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1995:SIA

[Smi95d] Ben Smith. Short-Order Internet access: Sun’s Netra and Performance Technology’s Instant Internet can both get your LAN onto the Internet fast. Byte Magazine, 20(7):139–??, July 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1996:HNN

[Smi96a] Ben Smith. Holes in the neural network. Byte Magazine, 21(4):145–??, April 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1996:UU

[Smi96b] Ben Smith. Ultrafast Ultra-SPARCs. Byte Magazine, 21(1): 139–140, January 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1996:VAW

[Smi96c] Ben Smith. VMS: Alive and well — A young horse is fast, but an old horse knows what’s going on. Byte Magazine, 21 (11):59–??, November 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sontag:1995:HHP

[Son95] John Sontag. HP-UX 10.0: Hewlett–Packard’s new HP-UX 10.0 will be faster, more reliable, and easier to use. Byte Magazine, 20(4):213–??, April 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Seachrist:1997:WAA

[SP97] David Seachrist and Stephen Platt. Web applications at your service — enliven your Web site with one of these application servers. Byte Magazine, 22(7):114–??, July 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stallings:1995:EI

[Sta95a] William Stallings. Endian issues. Byte Magazine, 20(9): 263–??, September 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stallings:1995:PWT

[Sta95b] William Stallings. The PGP Web of trust: Managing public keys with the PGP (Pretty Good Privacy) web of trust. Byte Magazine, 20(2):161–??, February 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stallings:1996:PSS

[Sta96a] W. Stallings. Patching the stacks in SNMP. Byte Magazine, 21 (8):55–56, August 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stallings:1996:BWH

[Sta96b] William Stallings. The backbone of the Web — HTTP 1.1 is a much-needed standard for the growing Internet. here’s what it does and how it works. Byte Magazine, 21(10):63–??, October 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stallings:1996:IAS

[Sta96c] William Stallings. Internet armor — secure IP lets you use the open Internet with privacy and confidence. Byte Magazine, 21(12):127–??, December 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stallings:1996:NII

[Sta96d] William Stallings. The new and improved Internet Protocol — additions to the Internet Protocol ensure future growth and new functions. Byte Magazine, 21 (9):55–??, September 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stallings:1996:PCS

[Sta96e] William Stallings. Patching the cracks in SNMP. Byte Magazine, 21(8):55–??, August 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stallings:1997:ACR

[Sta97a] William Stallings. Adjustable cell rates for ATM networks — A new ATM service compensates for network congestion. Byte Magazine, 22(11):43–??, November 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stark:1997:ESP

[Sta97b] Thom Stark. Encryption for a small planet — U.S. restrictions on encryption exports are cramping development of secure international applications. what are your options for competing in the world market? Byte Magazine, 22(3):111–??, March 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stallings:1998:NMM

[Sta98] William Stallings. Networks: S/MIME E-mail gets secure. Byte Magazine, 23(7):41–??, July 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stepanov:1995:STL

[Ste95] Alexander Stepanov. The Standard Template Library — how do you build an algorithm that is both generic and efficient? Byte Magazine, 20(10):177–??, October 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stewart:1998:PSB

[Ste98] George A. Stewart. Prophetable software — business forecasting software is using new ideas in statistics to improve the bottom line. Byte Magazine, 23(4):97– ??, April 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stonebraker:1998:BOI

[Sto98]  Michael Stonebraker.  Bet-

ting on ORDBMS — Informix’s CTO analyzes the capabilities of object-relational databases in line-of-business applications. Byte Magazine, 23(4):91–??, April 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sugihara:1995:CCN

[Sug95] Michael Sugihara. Consistent color: New software standards help manage color for consistent and accurate reproduction. Byte Magazine, 20(1):93– ??, January 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Supp:1997:HHP

[Sup97] Lyle Supp. The heart of handheld PCs — Hitachi’s SH-3 addresses the demands of handheld PCs. Byte Magazine, 22(7): 45–??, July 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Seachrist:1997:PYE

[SY97] David Seachrist and Tom Yager.

Put your E-mail on the IMAP — these e-mail servers make for handier management. Byte Magazine, 22(5):116–??, May 1997.

CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Tamaanis:1995:MMW

[Tam95a] Doug Tamaanis. Mathematica meets Warp: It’s no surprise that the new OS/2 version of Mathematica runs faster than the Windows version on IBM’s 32-bit OS. Byte Magazine, 20(5):137– ??, May 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Tamasanis:1995:LWWa

[Tam95b] Doug Tamasanis. A less wobbly Wabi. Byte Magazine, 20 (7):159–??, July 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Tamasanis:1995:LWWb

[Tam95c] Doug Tamasanis. A less wobbly Wabi: Wabi 2.0 makes Unix more Windows-friendly. Byte Magazine, 20(7):159–??, July 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Tamasanis:1995:RRP

[Tam95d] Doug Tamasanis. RISC for the road: The SparcBook 3XP — A workstation to go. Byte Magazine, 20(6):171–??, June 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Tamasanis:1995:SAS

[Tam95e] Doug Tamasanis. Springtime at Sun. Byte Magazine, 20(9):

271–??, September 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Tanner:1998:PIS

[Tan98] Dan Tanner. 333-MHz Pentium IIs: Slow-bus swan song: We test 19 NT workstations powered by Intel’s Deschutes processor. Byte Magazine, 23(5):114– ??, May 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Taylor:1997:ENP

[TC97] Andrew Taylor and Raymond G. A. Cote. Enhancing Netscape with plug-ins — A how-to guide to the tricks and traps of writing a Netscape plug-in on the Macintosh. Byte Magazine, 22(3):63– ??, March 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Tomkowiak:1996:CYW

[TH96] S. Tomkowiak and P. Hofland. A computer in your wallet. Byte Magazine, 21(6):11–12, 14, 16, June 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1995:PSE

[Tho95a] Kevin Thompson. PostScript sins — errors in PostScript files are becoming painfully plentiful. here are some of the most common ones. Byte Magazine, 20 (8):153–??, August 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1995:PRC

[Tho95b] T. Thompson. To print a rainbow (colour laser printers). Byte Magazine, 20(9):239–240, September 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1995:AAEa

[Tho95c] Tom Thompson. Adding Apple events to your Mac application: Part 2: This month, we write the event handlers and the dispatch table. Byte Magazine, 20(6):243– ??, June 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1995:AAEb

[Tho95d] Tom Thompson. Adding Apple events to your Mac applications: Part 1: How to make your application a good neighbor to other applications — and thus of value to your customer. Byte Magazine, 20(5):177–??, May 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1995:ACR

[Tho95e] Tom Thompson. Apple completely redesigned its Mac OS to provide a growth path to the future while continuing to support the existing software base. Byte Magazine, 20(6):59– ??, June 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1995:AIM

[Tho95f] Tom Thompson. Apple’s improved MessagePad: With bug fixes, more memory, and thirdparty support, Apple’s new MessagePad 120 strengthens its position in the PDA marathon. Byte Magazine, 20(5):131–??, May 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1995:BVP

[Tho95g] Tom Thompson. The better virtual PC — SoftWindows 2.0 puts a 33-MHz 486 inside your Power Macintosh. Byte Magazine, 20(12):195–??, December 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1995:BBV

[Tho95h] Tom Thompson. Building the better virtual CPU — how Apple revved up its 680x0 emulator. Byte Magazine, 20(8):149– ??, August 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1995:CAMa

[Tho95i] Tom Thompson. Copland: The abstract Mac OS. Byte Magazine, 20(7):177–??, July 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1995:CAMb

[Tho95j] Tom Thompson. Copland: The abstract Mac OS: The next generation of Macintosh OS has been delayed until early 1996 so that Apple can implement a Hardware Abstraction Layer — which should help support a Mac clone market. Byte Magazine, 20(7):177–??, July 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1995:NPN

[Tho95k] Tom Thompson. New Power-PCs for notebooks and PDAs: The 602, 603, and 603e PowerPC processors zero in on both camps in the low-power market: the low-cost, single-user segment and the long-battery-life, highperformance group. Byte Magazine, 20(4):211–??, April 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1995:OBT

[Tho95l] Tom Thompson. One box, two computers: If you want to run both PC and Mac applications, Apple’s DOS-Compatible Power Macintosh offers the best of both worlds. Byte Magazine, 20 (4):165–166, April 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1995:PMG

[Tho95m] Tom Thompson. Power Mac gets PCI — Apple’s Power Mac 9500 delivers great performance and the promise of fast, low-cost PCI peripherals. Byte Magazine, 20 (10):123–??, October 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1995:PR

[Tho95n] Tom Thompson. To print a rainbow. Byte Magazine, 20(9): 239–??, September 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1995:TTP

[Tho95o] Tom Thompson. Two turbocharged PowerPCs — IBM and Motorola rev up the 603 and 604 and reduce their chips’ hunger for power. Byte Magazine, 20(12):209–??, December 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1996:ME

[Tho96a] T. Thompson. Must-see 3-D engines. Byte Magazine, 21(6):137– 138, 140, 142, 144, June 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1996:PGC

[Tho96b] T. Thompson. The PowerPC goes consumer. Byte Magazine, 21(8):59–60, August 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1996:RCA

[Tho96c] Tom Thompson. The 1996 Readers’ Choice Awards. Byte Magazine, 21(7):145–??, July 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1996:APC

[Tho96d] Tom Thompson. An Alpha in PC clothing. Byte Magazine, 21 (2):195–196, February 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1996:CWF

[Tho96e] Tom Thompson. Componentized WAV of the future — Digital Harhor’s WAV word processor is among the first OpenDoc components for the Macintosh. Byte Magazine, 21(11):49–??, November 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1996:CPR

[Tho96f] Tom Thompson. The consumer PowerPC revisited — Motorola’s slimmed-down PowerPC could be the brains of your next handheld PC. Byte Magazine, 21 (11):67–??, November 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1996:CRL

[Tho96g] Tom Thompson. Copland, revisited — A look at the current state of Apple’s next OS. Byte Magazine, 21(11):151–??, November 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1996:ED

[Tho96h] Tom Thompson. The elements of design. Byte Magazine, 21(8):80– ??, August 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1996:HCC

[Tho96i] Tom Thompson. How Copland communicates. Byte Magazine, 21(4):179–??, April 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1996:MAM

[Tho96j] Tom Thompson. Move aside, Mac. Byte Magazine, 21(8):48, August 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1996:NS

[Tho96k] Tom Thompson. The network in the server. Byte Magazine, 21(7):151–154, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1996:NR

[Tho96l] Tom Thompson. Newton reborn. Byte Magazine, 21(5):123– ??, May 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1996:PPG

[Tho96m] Tom Thompson. The Power PC goes consumer. Byte Magazine, 21(8):59–??, August 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1996:SHI

[Tho96n] Tom Thompson. The server’s helper — Intel’s new i960 RP intelligent VO processor simplifies a server’s architecture while also improving performance. Byte Magazine, 21(10):67–??, October 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1996:TTC

[Tho96o] Tom Thompson. Two, two chips in one! Byte Magazine, 21(6):59– 60, June 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1996:WSH

[Tho96p] Tom Thompson. When silicon hits its limits — what’s next? Byte Magazine, 21(4):44– ??, April 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1996:WFC

[Tho96q] Tom Thompson. The world’s fastest computers. Byte Magazine, 21(1):45–47, 50, 52, 54, 58, 62, 64, January 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thomson:1996:ME

[Tho96r] Tom Thomson. Must-see 3-D engines. Byte Magazine, 21(6):137– ??, June 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1997:FLA

[Tho97a] Tom Thompson. First look at PowerPC G3 — A new generation offers speed and low fuel consumption. Byte Magazine, 22(4):59–??, April 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1997:HMW

[Tho97b] Tom Thompson. How to make the world’s fastest CPUs: Enter a world where dust is your enemy, day and night run together, and blazingly fast CPUs are born. Byte Magazine, 22 (2):20, February 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). This article offers a very understandable account of the complete fabrication process for the DEC Alpha CPU chips.

Thompson:1997:IBI

[Tho97c] Tom Thompson. I2O beats I/O bottlenecks — A new architecture promises a big boost in server performance but not a big boost in cost. Byte Magazine, 22(8):85–??, August 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1997:MGM

[Tho97d] Tom Thompson. The Mac goes multiprocessor — A new library and API enable preemptive multitasking and multiprocessing on Mac OS machines. Byte Magazine, 22(2):59–??, February 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1997:MMN

[Tho97e] Tom Thompson. The MessagePad makeover — the newest Newton puts a big screen into a smallish package and a fast StrongARM CPU inside. it’s the best one yet. Byte Magazine, 22(3):44–??, March 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1998:AMG

[Tho98a] Tom Thompson. Apple Macs get faster and more affordable. Byte Magazine, 23(1):28– 29, January 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1998:CPD

[Tho98b] Tom Thompson. CPUs: PowerPC does blazing vector. Byte Magazine, 23(7):43–??, July 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1998:FLA

[Tho98c] Tom Thompson. A first look at Rhapsody — Apple’s Rhapsody offers modern OS features, but the UI needs some work, and certain Mac technologies are missing. Byte Magazine, 23(1):45–??, January 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1998:MNM

[Tho98d] Tom Thompson. Mac notebook: The most powerful notebook. Byte Magazine, 23(7):107– ??, July 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1998:WRA

[Tho98e] Tom Thompson. Writing for Rhapsody — Apple’s latest OS offers a rich set of object-oriented frameworks for rapidly building GUI applications. Byte Magazine, 23(3):79–??, March 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1995:BCH

[TJF95] Tom Thompson, Alan Joch, and Rich Friedman. Books and CD-ROMs: How to optimize your PowerPC code. Byte Magazine, 20(9):33–??, September 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Tropiano:1995:HII

[TM95] Lenny Tropiano and Dinah Mc-Nutt. How to implement ISDN: Knowing the fundamentals of ISDN costs and service options can help you save money and avoid frustration. Byte Magazine, 20(4):67–68, 70, 72, 74, April 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Trammell:1995:UCP

[Tra95] Kelly Trammell. Under construction — have pity on groupware developers. they’ve got some serious problems to deal with. doin’ the LN:DI. telephony’s wake-up call. Byte Magazine, 20(8):93–??, August 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Trammell:1996:PMA

[Tra96a]  Kelly Trammell.  Prospero’s

magic application integrator. Byte Magazine, 21(2):173–174, February 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Trammell:1996:WFF

[Tra96b] Kelly Trammell. Work flow without fear. Byte Magazine, 21(4):55–??, April 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Traut:1997:BVP

[Tra97] Eric Traut. Building the virtual PC — A software emulator transforms the PowerPC into a Pentium PC. Byte Magazine, 22 (11):51–??, November 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Tristram:1995:KDM

[Tri95] Claire Tristram. Keep the data moving — with ruggedized computers and PC cards you can computerize collection of data in the factory or field. (buying rugged Hand-Held computers p 90). Byte Magazine, 20 (12):87–88, 90, 92, 94, 96, December 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thomsen:1996:YNS

[TS96a] D. Thomsen and W. Schwartau. Is your network secure? Byte Magazine, 21(1):155–156, January 1996. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Thomsen:1996:IYN

[TS96b] Dan Thomsen and Winn Schwartau.

Is your network secure? Byte Magazine, 21(1):155–??, January 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Turk:1997:BBI

[Tur97] Andy Turk. Building a better interface with Java: Internet Foundation Classes lets you build Java applets with useful features such as drag and drop. Byte Magazine, 22(8):45–46, August 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1995:BHW

[Ude95a] Jon Udell. BOMB’s away — how we built a Web-based survey to get feedback from readers. and how you can turn your Web server into an application server that connects you to customers, suppliers, and business partners. Byte Magazine, 20 (10):105–??, October 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1995:GGB

[Ude95b] Jon Udell. Global groupware — BYTE implements the virtual press room, a Web archive where companies can publish press releases and users can find out about the latest products. Byte Magazine, 20(11):203–??, November 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1995:HWJ

[Ude95c] Jon Udell. Hello, world: Jon Udell, executive editor for new media, launches a monthly column, the BYTE Network Project, to help you mine the opportunities of business-to-business networking. Byte Magazine, 20(7):87–??, July 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1995:LWB

[Ude95d] Jon Udell. Live wire — BYTE’s new 56-Kbps link to the Internet sparks concern over public versus private IP addressing. Byte Magazine, 20(8):103– ??, August 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1995:NCN

[Ude95e] Jon Udell. Novell’s campaign: Novell is embarking on a grand strategy to get 1 billion users and devices connected to NetWare by the year 2000. Byte Magazine, 20 (2):42–??, February 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1995:PMO

[Ude95f] Jon Udell. Perl magic — our Webmeister explains the Perl and HTML programming tricks under the hood of the new Virtual Press Room. Byte Magazine, 20(12):115–??, December 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1995:WPI

[Ude95g] Jon Udell. A warped perspective: IBM says Warp is for end users. wrong. it’s an integrator’s dream. Byte Magazine, 20(3):165–??, March 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1995:WS

[Ude95h] Jon Udell. Web search. Byte Magazine, 20(9):223–??, September 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1995:YCT

[Ude95i] Jon Udell. You can take it with you: What does Windows 95 offer users on the go? Byte Magazine, 20(6):145–??, June 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1996:DL

[Ude96a] Jon Udell. Damn lies. Byte Magazine, 21(2):137–??, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1996:DCB

[Ude96b] Jon Udell. Dual-mode conferencing — the BYTE Webmaster explains how to bring both news-style conferencing and Web-style conferencing to your site. Byte Magazine, 21(12): 141–??, December 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1996:EA

[Ude96c]  Jon Udell.  E-Mail adven-

tures. Byte Magazine, 21(4):85– ??, April 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1996:HWC

[Ude96d] Jon Udell. Hybrid Web/CD-ROM: Do it yourself — what bandwidth problem? WebCD lets you publish your Web site on CD-ROMs that retain live Internet links. Byte Magazine, 21 (11):54–??, November 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1996:LL

[Ude96e] Jon Udell. Lessons learned. Byte Magazine, 21(7):115–??, July 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1996:LT

[Ude96f] Jon Udell. Let’s talk. Byte Magazine, 21(5):77–??, May 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1996:OCH

[Ude96g] Jon Udell. On-line componentware — here’s how you can turn your Web site into a flexible software component to build new applications quickly. Byte Magazine, 21(11):129–??, November 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1996:PMP

[Ude96h] Jon Udell. Push me, pull you — how to combine e-mail push and Web pull with Next’s WebObjects. Byte Magazine, 21(9): 117–??, September 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1996:SM

[Ude96i] Jon Udell. Server management. Byte Magazine, 21(1):107– 108, 110, January 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1996:WC

[Ude96j]  Jon Udell.  Web conferenc-

ing. Byte Magazine, 21(6):131– ??, June 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1996:WD

[Ude96k] Jon Udell. Web design. Byte Magazine, 21(3):91–??, March 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1996:WSP

[Ude96l] Jon Udell. Web surveys — with properly designed forms, the Web becomes a valuable resource for data collection. Byte Magazine, 21(10):133–??, October 1996. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1996:YBN

[Ude96m] Jon Udell. Your business needs the Web. Byte Magazine, 21 (8):68–70, 72, 74, 76, 78, 80, August 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Discusses NASA’s distributed data visualization using Java.

Udell:1997:WPJa

[Ude97a] John Udell. Web project: JavaScript revisited. Byte Magazine, 22(5):99, 100, 102, May 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1997:WPP

[Ude97b] John Udell. Web project: Persistent Java: A servlet-based group calendar becomes a surprising success and prompts an exploration of ways to bind Java programs to persistent storage. Byte Magazine, 22(8):109, 110, 112, August 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1997:CSS

[Ude97c] Jon Udell. Cascading style sheets — CSS will change how Web designers work. here are some ways to learn and use it. Byte Magazine, 22(11):97–??, November 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1997:DIT

[Ude97d] Jon Udell. Digital IDs — today’s digital identification technologies promise long-term benefits but currently inflict shortterm pain. Byte Magazine, 22 (3):115–??, March 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1997:NAW

[Ude97e] Jon Udell. Net applications: Will Netscape set the standard? Byte Magazine, 22(3):66–69, 72, 74, 76, 80, 82, 84, March 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Discusses Java, Javascript, and the Java Runtime Environment.

Udell:1997:NSH

[Ude97f] Jon Udell. NT security — how to make Windows NT a secure Internet platform. Byte Magazine, 22(2):117–??, February 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1997:PFH

[Ude97g] Jon Udell. Perl’s future — how to benefit from Perl and its challengers. Byte Magazine, 22(4):121–??, April 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1997:SAO

[Ude97h] Jon Udell. Search again — our Webmaster explains how to move beyond basic search-engine capabilities to make your site more informative. Byte Magazine, 22 (1):123–??, January 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1997:TTS

[Ude97i] Jon Udell. Textbase tricks — some tips on making good use of all that information on the Web. Byte Magazine, 22(7):97– ??, July 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1997:VFS

[Ude97j] Jon Udell. The value of free software — freeware gems like Apache, Linux, and Perl keep the Web spinning. Byte Magazine, 22(12):109–??, December 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1997:WPJb

[Ude97k] Jon Udell. Web project: Java servlets: Servlets, the Java equivalent of CGI applications, can deliver on many of Java’s promises while dodging some of its worst limitations. Byte Magazine, 22(6):115, 116, 118, June 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1998:DEH

[Ude98a] Jon Udell. Deploying effective HTML forms: 10 tips for creating the best Web forms you can. Byte Magazine, 23(5):103– ??, May 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1998:HAW

[Ude98b] Jon Udell. HTTP authentication — worried that anyone can get into your site? authentication is the answer, but not all Web servers do it the same. Byte Magazine, 23(1):89– ??, January 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1998:SSS

[Ude98c] Jon Udell. In search of SSL spidering — building tools that fetch and process secure pages teaches some painful lessons. Byte Magazine, 23(2):97–??, February 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1998:LWS

[Ude98d] Jon Udell. Local Web servers.

Byte Magazine, 23(6):105–??, June 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1998:SMN

[Ude98e] Jon Udell. Securing mail and news. Byte Magazine, 23(4):105– ??, April 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1995:BCS

[UGB95] Jon Udell, Rick Grehan, and Rex Baldazo. Books and CD-ROMs: A savvy guide to client/server computing: Programming CD-ROMs, LANs, and the development of Windows NT. Byte Magazine, 20(2):37–??, February 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1995:BCM

[UT95] Jon Udell and Tom Thompson. Books and CD-ROMs: A model for future AI research: An AI model from Douglas Hofstadter, a Star Trek technical manual, and an unauthorized look at Windows 95. Byte Magazine, 20(3):45–??, March 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vacca:1995:MBN

[Vac95a] John Vacca. Mosaic: Beyond net surfing: This GUI navigator is also a tool for doing business on the World Wide Web. Byte Magazine, 20(1):75– ??, January 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vacca:1995:FIT

[Vac95b] John R. Vacca. Faking it, then making it — simulation software helped design production flow in Siemens Solar’s new clean room. Byte Magazine, 20(11): 66–??, November 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanHorn:1995:ASM

[Van95a] G. Armour Van Horn. Audience share: As more people incorporate sound and video in their presentations, projection panels must keep up. Proxima’s new Ovation+ active-matrix LCD projection panels support 24-bit color and stereo sound. Byte Magazine, 20(1):191–??, January 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

vanHorn:1995:ASO

[van95b] G. Armour van Horn. Audience share (ovation+ projection panels). Byte Magazine, 20(1): 191–192, January 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanHorn:1996:ADR

[Van96a] G. Armour Van Horn. Apple’s desktop replacement. Byte Magazine, 21(6):70, June 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanHorn:1996:OGG

[Van96b] G. Armour Van Horn. One gig to go. Byte Magazine, 21(4):151– ??, April 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanHorn:1996:RSR

[Van96c] G. Armour Van Horn. Roadshow romance — Proxima’s Desktop Projector 4100Z sets a new standard for direct digital projection. Byte Magazine, 21 (10):55–??, October 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanderVeer:1996:OSO

[Van96d] Emily R. Vander Veer. Open-Doc says OLE to developers [operating systems]. Byte Magazine, 21(7):49–50, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanHorn:1997:PMC

[Van97] G. Armour Van Horn. Power Macs: Cookin’ at 225 MHz — should you buy a clone or stick with the original brand? we test the latest 200- and 225-MHz Power Macs to find the best. Byte Magazine, 22(3):135– ??, March 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Varhol:1995:WCF

[Var95] Peter Varhol. Windows for control freaks: QNX’s Photon kernel gives developers welcome control over windowing programming. Byte Magazine, 20(6):83– ??, June 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Varhol:1996:OGW

[Var96] Peter D. Varhol. OOP from the guy who wrote the book. Byte Magazine, 21(2):169–170, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vaughan-Nichols:1996:EES

[Vau96] Steven J. Vaughan- Nichols. Exchange: E-mail on steroids. Byte Magazine, 21(6):153–??, June 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VaughanNichols:1996:EES

[VN96a] S. J. Vaughan-Nichols. Exchange: e-mail on steroids. Byte Magazine, 21(6):153–154, June 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vaughan-Nichols:1996:NST

[VN96b] Steven J. Vaughan-Nichols. Net saviors — two backup products take very different approaches to saving network data. Byte Magazine, 21(10):145–??, October 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vaughan-Nichols:1996:SCW

[VN96c] Steven J. Vaughan-Nichols.

Spinning the corporate Web. Byte Magazine, 21(5):113–??, May 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vaughan-Nichols:1997:EWM

[VN97a] Steven J. Vaughan-Nichols. Eudorable Web Mail — better POP and MIME support, plus new filtering plugins, make the new Windows 95 version of Eudora well worth paying for. Byte Magazine, 22(2):140–??, February 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vaughan-Nichols:1997:LAS

[VN97b] Steven J. Vaughan-Nichols. Local AltaVista searching — AltaVista now lets you aim its popular Internet search engine at your PC and LAN. but you need big metal for good performance. Byte Magazine, 22 (1):42–??, January 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vogt:1995:SRS

[Vog95] Charles Vogt. Software roundup: Simple SQL: SQL front-end tools empower nontechnical managers to make informed, time-critical decisions by giving them direct access to client/server databases. Byte Magazine, 20(6):217–??, June 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Watterson:1995:DMT

[Wat95]  Karen Watterson.  A data

miner’s tools — you can dig it with query-and-reporting programs, multidimensional analysis software, and intelligent agents. Byte Magazine, 20 (10):91–??, October 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Watterson:1997:ADS

[Wat97a] Karen Watterson. Attention, data-mart shoppers — top-down or bottom-up, building a data mart gives you the information you value most. Byte Magazine, 22(7):73–??, July 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Watterson:1997:CES

[Wat97b] Karen Watterson. Commanding the enterprise — soon you’ll be able to manage the network and all those devices from one console. Byte Magazine, 22(12): 93–??, December 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Watterson:1998:PED

[Wat98a] Karen Watterson. Part I: Enterprise databases battle on. Byte Magazine, 23(6):97–??, June 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Watterson:1998:RUR

[Wat98b] Karen Watterson. The road to a universal repository: Components technology heats up an old idea. Byte Magazine, 23 (5):112S–??, May 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1995:ACS

[Way95a] Peter Wayner. Agent-Enhanced communicator: Sony’s Magic Link provides phone, modem, sendfax, E-mail pager, and infrared communications abilities. Byte Magazine, 20(2):103–104, February 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1995:FAN

[Way95b] Peter Wayner. Free agents: A new generation of light-weight, multithreaded operating environments provide security and interoperability for agent developers. Byte Magazine, 20 (3):105–??, March 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1995:MLG

[Way95c] Peter Wayner. Machine learning grows up — what’s the difference between Elwood and Jake? the right algorithm can tell you. Byte Magazine, 20(8):63–??, August 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1995:NCF

[Way95d] Peter Wayner. Network-Ready computer: Forget the CPU and clock speed; network I/O capacity may be the new measure of a desktop system’s performance. Byte Magazine, 20 (3):171–??, March 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1995:NCN

[Way95e] Peter Wayner. New chips and new technologies compete for adoption. Byte Magazine, 20 (6):187–??, June 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1995:OHR

[Way95f] Peter Wayner. Oracle hits the road: Oracle In Motion’s agents take over network housekeeping chores, virtually eliminating the wireless performance penalty. Byte Magazine, 20 (6):207–??, June 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1995:PCL

[Way95g] Peter Wayner. Picking the crypto locks — with today’s hardware and a new technique called differential cryptanalysis. it’s getting easier for data burglars to crack your safe. what can you do to keep them out? Byte Magazine, 20(10):77–??, October 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1995:SFD

[Way95h] Peter Wayner. Solutions focus:

DragNET — G-men get a new weapon. the FBI’s DNA database is helping police nab suspects. (DNA by the numbers p 110). Byte Magazine, 20(12): 106–??, December 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1996:NPM

[Way96a] P. Wayner. Net programming for the masses. Byte Magazine, 21(2):101–102, 104, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1996:BJP

[Way96b] Peter Wayner. Better Java programming: Knowing how Java’s dynamic linking works can help you improve a program’s performance. Byte Magazine, 21(9): 63–64, September 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1996:CR

[Way96c] Peter Wayner. Crave the RAVE. Byte Magazine, 21(6):49–50, June 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1996:DLY

[Way96d] Peter Wayner. Don’t lose your crypto keys. Byte Magazine, 21(5):137–??, May 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1996:EW

[Way96e] Peter Wayner. E-Mail without wires. Byte Magazine, 21(4):138– 139, April 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1996:HBC

[Way96f] Peter Wayner. Hey baby, call me at my IP address. Byte Magazine, 21(4):142–??, April 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1996:HCC

[Way96g] Peter Wayner. How chips changed the world — 10 ways in which that silver of silicon altered the course of Planet Earth. Byte Magazine, 21(12):

68–??, December 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1996:INN

[Way96h] Peter Wayner. Inside the NC: Are network computers just stripped-down terminals? No way. The official NC platform covers everything from a set-top box to a Cray. Byte Magazine, 21(11):105, 106, 108, 110, November 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Discusses the importance of Java in a Network Computer.

Wayner:1996:NVS

[Way96i] Peter Wayner. New videophones starved for bandwidth. Byte Magazine, 21(5):125–??, May 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1996:OWO

[Way96j] Peter Wayner. An open window for OpenDoc — the promise of platform-independent reusable components comes closer to reality with the second beta of the OpenDoc toolkit for Windows 95 and NT. Byte Magazine, 21 (12):41–??, December 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1996:PPP

[Way96k] Peter Wayner. PDA packs power in your pocket. Byte Magazine, 21(7):48–??, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1996:SGJ

[Way96l] Peter Wayner. Sun gambles on Java chips: Are Java chips better than general purpose CPUs? or will new compilers make them obsolete? Byte Magazine, 21 (11):79, 80, 82, 84, 86, November 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1996:XBJ

[Way96m] Peter Wayner. XTerminal + browser + Java = Web PC — HDS’s (a) workStation is a lowcost box for Web access, Java applets, and Windows. Byte Magazine, 21(10):50–??, October 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1997:ALP

[Way97a] Peter Wayner. At last: Pocket PCs that run Windows — new hand-held PCs bring Windows to the palmtop. we test three HPCs built around Windows CE. Byte Magazine, 22(1):133– ??, January 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1997:BNO

[Way97b] Peter Wayner. The be-all-that-you-can-be OS — A new OS offers users a new set of options. Byte Magazine, 22(12): 137–??, December 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1997:NCE

[Way97c] Peter Wayner. Novera composes Epic for Java. Byte Magazine, 22(7):122–123, July 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1997:TSM

[Way97d] Peter Wayner. On TV: The set-top morphing PowerPC — Bandai Digital’s set-top box, a nice implementation of apple’s Pippin technology, is a Mac-based Web browser. Byte Magazine, 22(2):40A–??, February 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1997:RRC

[Way97e] Peter Wayner. Real remote control from your pocket PC — Symantec’s pcAnywhere CE. Byte Magazine, 22(5):42–??, May 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1997:HSV

[Way97f] Peter Wayner. This hand-held stands out — the Velo 1 is the best little PC. Byte Magazine, 22(4):46–??, April 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1997:TML

[Way97g] Peter Wayner. Throw me a LifeBook — Fujitsu squeezes all the essentials into a 4-pound package. Byte Magazine, 22 (5):126–??, May 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1998:HSJ

[Way98a] Peter Wayner. How to soup up Java, part II: Nine recipes for east, easy Java: A maturing Java brings a wide variety of programming solutions to developers. Byte Magazine, 23(5):76– 80, May 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1998:MJD

[Way98b] Peter Wayner. Making Java development JSafe — RSA’s cryptographic toolkit targets Java as the platform of choice for secured Internet applications. Byte Magazine, 23(1):117, January 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Webster:1996:CRS

[Web96] Bruce F. Webster. Commentary: The real software crisis. Byte Magazine, 21(1):218– ??, January 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Weiss:1995:FOC

[Wei95] Michael Weiss. Final output: Color printers, more affordable than ever, hold out the promise of better color. Byte Magazine, 20(1):109–??, January 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Weldon:1997:WCE

[Wel97] Jay-Louise Weldon. Warehouse cornerstones — an expert in database management explains how a well-designed data warehouse leads to smarter decisions. Byte Magazine, 22(1):85– ??, January 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Westerfeld:1996:EOU

[Wes96a] K. Westerfeld. Extending OS/2’s user interface. Byte Magazine, 21 (2):193–194, February 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Westerfeld:1996:EOI

[Wes96b] Kurt Westerfeld. Extending the OS/2 interface. Byte Magazine, 21(2):193–??, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

White:1995:OOO

[Whi95a] C. White. On-road, on-time, and on-line (trucking company: global positioning system). Byte Magazine, 20(4):60–62, 64–66, April 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

White:1995:SFO

[Whi95b] Christine White. Solutions focus:

On-Road, On-Time, and On-Line: Satellites and computerbased communications helps a trucking firm cut costs. Byte Magazine, 20(4):60–??, April 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

White:1996:IDH

[Whi96] Christine White. Installations: DMV hits pothole. Byte Magazine, 21(2):72–??, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

White:1997:MDT

[Whi97] Colin White. Managing data transformations — data requires careful control and cleansing to make it useful for business decisions. Byte Magazine, 22(12): 53–??, December 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Whittaker:1998:IAS

[Whi98] Tom Whittaker. An OS for information appliances — Sun’s Software Suite is a scalable set of OS and application components for hand-held and embedded devices. Byte Magazine, 23(3):45– ??, March 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wilansky:1995:BRT

[Wil95] Ethan Wilansky. The benefits of recentralization: There’s a move to consolidate network management tools. Microsoft fields a powerful entry in Systems Management Server, which rounds up and controls network resources multiple vendors. Byte Magazine, 20(5):121–??, May 1995.

CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wilson:1996:SBP

[Wil96] Tim Wilson. See the big picture — software that helps manage mega-size networks. Byte Magazine, 21(10):107–??, October 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wilner:1998:E

[Wil98] David Wilner. 120’s OS evolves.

Byte Magazine, 23(4):47–??, April 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Winston:1995:CMI

[Win95] Tim Winston. Create more IP addresses: Network managers facing IP address change have many options. Byte Magazine, 20(4):217–??, April 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wong:1998:NFP

[Won98a] William Wong. Nine fax programs to serve the Internet — we wring out nine fax servers for workgroups and the enterprise. Byte Magazine, 23(1):108– ??, January 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wong:1998:RRK

[Won98b] William Wong. Remote-control registry: Key Vision removes a lot of the hassle and lets you manage Windows 95 and NT registries from afar. Byte Magazine, 23(5):128–??, May 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wong:1998:STF

[Won98c] William Wong. Software — team-building on the fly — we test “teamware”- programs for building fast little workgroups across the Internet and across organizational boundaries. Byte Magazine, 23(2):106–??, February 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Weber:1996:EVB

[WS96] Volker Weber and Hans-Jochen Schmitt. Enterprise Visual Basic? almost — ComUnity OT Framework, Siemens Nixdorf’s software for augmenting Visual Basic with business logic and client/server tools, falls short of its goal. Byte Magazine, 21 (10):149–??, October 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1995:SST

[Yag95a] T. Yager. Small scale telephony (digital signal processors). Byte Magazine, 20(5):125–128, May 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1995:GLF

[Yag95b] Tom Yager. The great little file system: Flexible and secure, Veritas sets the standard for Unix file systems. Byte Magazine, 20 (2):155–??, February 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1995:STM

[Yag95c] Tom Yager. Small-Scale telephony: Multifunction telephony cards can cost-effectively make your small business sound big. we test six. Byte Magazine, 20(5):125–??, May 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1996:LM

[Yag96a] T. Yager. Linux matters. Byte Magazine, 21(2):123–124, 126– 128, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1996:NN

[Yag96b] T. Yager. NT and the Net.

Byte Magazine, 21(7):133–134, 136, July 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1996:BBS

[Yag96c] Tom Yager. Bringing benchmarks up to SPEC. Byte Magazine, 21(3):145–146, March 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1996:PFN

[Yag96d] Tom Yager. The power of fusion — need to manage a mixed Unix/Windows environment? Ross’s SPARCplug helps by combining a SparcStation and a Pentium PC. Byte Magazine, 21(12):47–??, December 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1996:WUP

[Yag96e] Tom Yager. Windows — built for Unix power. Byte Magazine, 21(3):115–116, March 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1997:CSM

[Yag97a] Tom Yager. Client/server magic — remote procedure calls provide seamless data exchange among applications across a network, regardless of the platform. Byte Magazine, 22(1):53– ??, January 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1997:UWD

[Yag97b] Tom Yager. Unix and Windows dance the Samba — the Samba suite of utilities enables Windows clients to access Unix servers. Byte Magazine, 22 (3):55–??, March 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1998:FTF

[Yag98a]  Tom Yager.  Four Tillam-

ooks face off — new 233-MHz Pentium notebooks from Compaq, Hewlett–Packard, NEC, and Toshiba. Byte Magazine, 23 (1):33–34, January 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1998:SMD

[Yag98b] Tom Yager. Sun makes a desktop gamble. Byte Magazine, 23 (4):131–??, April 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yares:1995:BMP

[Yar95] Evan Yares. Bringing 3-D modeling to PCs: Autodesk releases AutoCAD Designer 1.0, a 3-D solid-modeling program. Byte Magazine, 20(1):201–??, January 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yavin:1995:RHT

[Yav95] David Yavin. Replication’s hst (??) track — large-scale data replication becomes less daunting with the right scheduling and systems management. light at the end of the tunnel, or an oncoming train? strained relations. how exchange handles replication. Byte Magazine, 20(8):88– ??, August 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yeaton:1995:PPW

[Yea95a] S. Yeaton. Powerful presentations for Windows. Byte Magazine, 20(1):183–184, 186– 188, 190, January 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yeaton:1995:SRP

[Yea95b] Susan Yeaton. Software roundup:

Powerful presentations for Windows: The traditional slideshow leaders — Charisma, Harvard Graphics, Lotus Freelance Graphics, and Microsoft Power-Point — concentrate on power and usability. Byte Magazine, 20 (1):183–??, January 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yourdon:1996:WGE

[You96] Edward Yourdon. When good enough is best — when designing client/server systems, “good enough” is best, says the software guru. Byte Magazine, 21 (9):85–??, September 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Zhao:1996:SCP

[Zha96] William Zhao. Solving the Chinese puzzle. Byte Magazine, 21(7):47–??, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).